A Complete Bibliography of Publications in Linear Algebra and its Applications: 2015–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.81

Title word cross-reference

(0, 1) [BD19, CR18a, CLT15, CMQV16, GT18, NS15a, NA17, ZM16, dCFF17]. (2, 1, 0) [WG15]. (75, 32, 10, 16) [AM18]. (76, 30, 8, 14) [BPR17]. (A, m) [BSZ18]. (b, c) [Rak17]. (d, k, δ) [FJ19]. (δ, ϵ) [Zha18b]. (e, f) [KC16]. (I − A) [CC18]. (m, m − 1, 0) [HZ17]. (m, q) [BMN16]. (P^{1})^{n} [LM16a]. (N − 4) [HAM16]. (p, n − p) [PR18a]. (p, r) [AdF19, EJ18]. (v, b, r, k, λ) = (10, 15, 6, 4, 2) [Mic16]. ∗ [DI16]. −(1 + \sqrt{17})/2 [LM16b]. −1 [BK17a]. {−1, 0, 1} [JMPS18]. −1 ∉ σ(S^{−T} S) [GMP16c]. −2 [CRS15]. −3 [KLY18]. 0 [BBH16, HL19b, Shi18b]. {0, 1} [XLQC16, XCQ18]. 0 < p ≤ 1 [LC19]. 1 [BBH16, BR17, Hir18, HL19b, KY18, KLW16, MHL15, OP16a, PCI15, Shi18b]. 14 [Jel18]. 2 [AKR15, AK16a, AA17, BLdS16, BC15a, BH16a, CP16a, CG19b, CW15a, CL18, CJ17d, DFK^{+}15, DHS18, EV15, HPSS19, Hoo19, KLP15, KY15, KY18, KIS18, LY16, LWX17, MMS16, Mic16, ME15, NP15a, OQ18, QKP16, Sam17, STW19, Wan15b, YLT16, ZXW19, ZGW16]. 2(n − 1)/3 [LS15a]. 2n [dlCMP17a, dlCdR17]. 2 × 2 [GHK17, GLRT18, KLP18, KM16, She17, Ulu19, UG18]. 2 × 2 × 3 [LS15b]. 2 × N [Sar19]. 3 [BK17a, BKLP16, CLZ16a, DFK^{+}15, HN18, Kal16, Lee15c, LWX17, Sam17]. 3 × 3 [BBdH18, HK18b]. 4 [BDVRT15, KH17, LWX17, SS17a]. 5 [AMR^{+}16, BDVRT15, JMP17, Jør15]. 5 × 5 [LS17b, SW18a]. 7 × 7 [HLPS19]. 9 [Kar16].

[S] [BGCMASS19]. [x_{1}, y_{1}][x_{2}, y_{2}] ⋯ [x_{q}, y_{q}] = 0 [vWZ17]. ^{∗} [BCP16]. ^{⋆} [CP16a]. _{0} [LQL15]. A [BDG15, CLST18]. A^{s+1}R = RA^{∗} [CLST18]. A_{α} [CLM19b, LHX18, LXS18, LL18b, LWCL19, NPRS17, XLLS18]. A_{α}(G) [CPR18, NR17]. A_{q}(N, M) [Liu16b]. AB [GSW16]. α [CPR18, FKM17a, FKM17b, GX19, NR18].

∗ [AG15b, AG20, BBEE18]. A → Tr(P f(A)) [Gir19]. AW^{∗} [LCLW18]. AX + Y B = C [MEMM17]. AX − X̄B = C [FKS16]. AXA^{∗} + BY B^{∗} = C [MEMM17]. AXB − C = 0 [CGM17]. B [LL15a, NA15]. B(H) [AR15b, NP15a]. BA [GSW16]. □_{q} [Yan18]. BπR [OP19a]. c [Cha15, CCCNT19]. C^{∗} [EP18a, LMX19, Mol18b, MEMM17, XF17]. C_{1}^{∨}, C_{1} [NT17]. c_{2} = 1 [MMP16]. B(H) [Tur17]. B_{\} [Sim18]. C_{\} [Sim18]. D_{α} [DPR19]. E_{}[Sim18]. E_{}[Sim18]. E_{∀} [Sim18]. F_{△} [Sim18]. G_{∈} [Sim18]. D [Lin15a, BF15, FMM16, KZ17]. D_{ν} [GGMPC15]. Δ_{2} = 0 [MMP16]. DL(P) [BBFF17]. E [Sma15, BL18a]. ℓ [DDV16]. ℓ^{p} [EE15, EM15]. ℓ^{p}(I) [Lju15]. ℓ^{p}(I)^{+} [LD16b]. ℓ_{p} [HT18a, Jia15]. f [Nie15b, Vir16]. F_{2} [CKT19]. G [BBM19, zSqST15]. G_{1} [KMS17a]. Γ [LLL17]. GL(V) [Gua18]. GP(n, k) [KMM17]. H [CDR18, CS19a, FuIKT16, HQX15, HWXC18, LZM16, SWB19, WYLW17, WLWZ19, wXL14, wXZ19, YQS16, ZLLL18, MMV19, MP16b]. H_{∞} [Ste15]. H_{n}^{∗} [Gao18]. I [LD17b]. ∞ [CGL16]. J [BFY18, HR17, dlCdR17]. k [ACT16, Che19a, DMS19, HHL15, HW15c, HSS16a, KNY15, KLW15, KLW18, MM17c, Nik16b, NLWJ18, ÖK18, QH15, Stu17, YYQ19, YG15, ZL15a, ZL15b, vDS16]. K_{2} [Sin17]. K_{2,t} [Nik17c]. K_{5} [SUY16]. K_{s,t} [dFNP16]. L [Dah15a, BR16a, DPV19]. l^{1}(I)^{+} [LD17b]. L^{2} [Fre18, Li15b]. L^{p} [LNW17]. l^{p}(I) [LD17c]. L_{∞} [BBMFM17]. l_{p} [BMN16, LC19]. λ [CM17, EP18b]. LDL^{T} [LMS15]. LQ [GH16b]. LR [Nom15c, Nom16]. LU [DEH16, SP16]. M [BC16b, GL16, HWXC18, JS16, KLS19a, MY16, SH15b, VB18, ZSWB16, ANAPSR17, BBMP19]. M[[α, ∞); (s_{j})_{j=0}^{m}, ≤] [FKMS18]. m_{0}λ(p) [EF19]. M_{3}(C) [YLT16]. E^{n+1} [Kim15a]. F_{ℓ} [SSS15]. F_{q}^{3} [MC19]. H_{n} [GLZ16a, GLZ16b, GL17]. R_{∞}^{4} [HM16]. R^{n} [ÁVAK^{+}19]. Z_{2} [CR15]. Z_{ℓ} [TZK17]. p [Sab19]. B(H) [CIMW15]. H [KSMB15]. Z [SWB19]. ∫↕(∈, F) [CCD19]. N [HAM16, AMR^{+}16, CILL16, DHS18, FJM18, LS15a, TWM16, VGG17]. n + 1 [dlCdR17]. n − 3 [LHH17, MQTW18]. n > 1 [AMR^{+}16]. n ≡ 2 (mod 4) [Sma15]. n ≥ 3 [IL17b]. n ≥ 4 [GLZ16b, Gao18]. n ≥= 5 [GL17]. n ≤ 6 [BCS15]. n × n [AOK15, Tan16b, Ter18, ZSQZ19]. P [DLQ18, DdF15, FdC18, NS15b, NS15c, BSS17, CL15b, CM16, DV18, EG15, Guo19, IL17b, JM16, JLM18, KNY15, LG18, Nik17a, WQH19, Wol18]. p = 2, 3 [PR18a]. P^{2} [Kus16]. P_{0} [DLQ18]. Φ_{n} [Ter18]. ϕ_{S} [GMP15, GMP16b, GMP16c]. p ∈ (1, ∞) [LD16b]. p ≠ 2 [IL17b]. Q [FZwCW16, GK15, NY15, PS19a, Yua14, Yua15, ZLS17, ZG18, dFNP16, BT17b, CJ17a, CJ15, CJ16a, Ern18, NS17, Nag19, NT17, Ter17a, Ter17b, dS18a, dS19a]. Q^{2} [Kus15]. q ↦ q̂ [FKS16]. QR [RBKA16, Saw16, FZwCW16]. R [SS18, ÖK18, XFZD17]. R^{k} = I [CLST18]. R^{p,q} [RAAGA16]. R_{II} [IR19]. S [Bou16a, CCiT18, GMP15, GMP16b, LJL16, dlCMP15]. S^{T} S [GMP16c]. Σ [WLL19]. σ(p) [BM17c]. Σ_{1} [WLL19]. sl_{2} [AC15]. SR [FR16]. ⋆ [NP15a]. U(p, p) [MY19]. U_{q}( ˆ∫↕_{2}) [Ter17b]. U_{q}(sl_{2}) [SGH16, Ter15]. UT_{2}(E) [CR15]. φ [dlCMP17b]. Vir(a, b) [HCS17]. W [CF18]. X [PS16, LM17]. X + A^{∗}X^{−1}A = Q [Chu16]. X − AX̄B = C [FKS16]. y^{∗}f(A)x [FM16]. Z [Bal15a, DN16, LLV15, LZM16, OP19b, RT16]. Z_{2} [Ste18].

Z_{4} [Mor16b]. |V (G)| − 2m + 2c(G) [zSqST15].

* [DD16b]. *-rings [DD16b].

-adic [BSS17, EG15]. -algebra [CILL16]. -algebras [BCP16, CP16a, EP18a, LCLW18, Mol18b]. -Aluthge [CM17, EP18b]. -approximately [Zha18b]. -atomic [KY18]. -Bernoulli [Ern18]. -Bernstein [MMV19]. -Bilocal [CL18]. -by- [AMR^{+}16, DFK^{+}15, HN18, Lee15c, dlCMP17a, dlCdR17]. -characteristic [LL18b]. -chromatic [KNY15]. -coherent [GGMPC15]. -colored [Kal16]. -commuting [QH15]. -cones [IL17b]. -connected [HSS16a, NLWJ18, SS17a]. -connections [Nie15b]. -contractive [LCNZ19]. -cospectral [ZLS17]. -cut [Nik16b, vDS16]. -cyclic [MP16b]. -density [BDG15]. -derivations [NP15a]. -designs [BH16a]. -digraphs [FJ19]. -dimensional [BC15a, FMM16, KLP15, HAM16]. -divergence [Vir16]. -eigenpair [LLV15]. -eigenvalue [FuIKT16, HWXC18, Lin15a, SWB19, ZG18]. -eigenvalues [CLM19b, HQX15, YQS16]. -element [DHS18]. -Euler [Ern18]. -gons [VGG17]. -graded [CR15]. -graphs [BBM19]. -ideals [HR17]. -ifications [BR16a, DPV19, DDV16]. -independence [ACT16]. -index [NR18, Yua15, NY15, Yua14, dFNP16]. -integral [PS19a]. -inverses [KC16, PCI15]. -isometries [BMN16, BBMP19, CGL16, MMS16, BSZ18]. -isomorphisms [STW19]. -join [CDR18]. -Laplacians [NS17, Nag19]. -less [FZwCW16]. -linear [ANAPSR17, RT16, BGCMASS19]. -Local [AKR15, AK16a, AA17, CW15a, CP16a, NP15a]. -majorization [NA15]. -majorizing [Sab19]. -matrices [AMG19, BD19, CR18a, GT18, NA17, wXZ19, ZM16, dCFF17, Bal15a, CS19a, DN16, GL16, JS16, KLS19a, Kus16, MY16, OP19b, SH15b, SS18, VB18, wXL14, ZLLL18]. -matrix [CLT15, CMQV16, NS15a, BC16b, DPR19]. -minimization [LC19]. -modified [JM16, JLM18]. -module [Gua18, LMX19]. -modules [MEMM17, XF17, Yan18]. -negative [Wol18]. -norm [Hoo19, Nik17a, OP16a]. -norms [CL15b]. -nuclear [BM17c]. -numerical [CJ17a, CCCNT19, Zam19, Cha15]. -Onsager [Ter17b]. -optimal [Sma15]. -order [DV18]. -orthogonality [KZ17]. -partite [KNY15]. -permanent [dS19a]. -permanental [dS18a]. -point [KLW16, KLW18]. -polynomial [GK15]. -positive [YLT16]. -potence [Stu17]. -potent [Stu17]. -potents [XFZD17]. -problem [CG19b]. -Racah [NT17, Ter17a]. -rank [WYLW17]. -ranks [WLWZ19]. -rays [Dah15a]. -Regular [CLZ16a, MM17c]. -resolvable [Mic16]. -rings [BBEE18]. -Riordan [BT17b, CJ15]. -scalar [GX19]. -scalings [Kus15]. -Schatten [CM16]. -Schur [FKM17a, FKM17b]. -SDD [WLL19]. -semiclassical [GGMPC15]. -semigroups [AG15b, AG20]. -set [DdF15, NS15b]. -sets [NS15c]. -Sheffer [CJ16a]. -simple [PS16]. -space [LNW17]. -special [Ern18]. -spectra [JMP17, LXS18, NPRS17]. -spectral [KNY15, LZM16, LHX18, XLLS18]. -SSD [WLL19]. -stability [Fre18]. -Stieltjes [LM17]. -subsets [DHS18]. -Sylvester [DI16]. -symmetric [BFY18]. -symmetries [dlCdR17]. -tensors [LQL15, DLQ18, HWXC18, KSMB15, LL15a, OP19a, SWB19, ZSWB16]. -tessellable [HPSS19, KIS18]. -th [Guo19]. -Toeplitz [AdF19, EJ18]. -trees [HW15c, ZL15a, ZL15b]. -triples [BCP16]. -type [LJL16]. -uniform [YYQ19]. -unitary [CCiT18]. -universal [Bou16a]. -variable [KY15, KY18, LY16]. -vertex [TWM16]. -vertices [FdC18]. -Walk-regular [QKP16].

0 [BL18b, Zha15a].

1 [BL18b]. 1-tensor [BL18b]. 18th [GPW15]. 19th [CFH^{+}16].

2013 [GPW15]. 2014 [CFH^{+}16]. 2016 [FLVV18]. 2017 [GHPS19]. 20th [FLVV18].

436 [Duk15, LT16a]. 437 [Kis15]. 438 [LMO16]. 458 [wXZ19, Yua15]. 463 [EKSV18]. 467 [ZL15a]. 479 [BKSP23]. 486 [AG20].

501 [GP18]. 502 [Ano16a]. 505 [KHI16a]. 509 [SM18]. 510 [BF17b]. 512 [Bal18]. 513 [LZSD18]. 520 [FJS21]. 521 [LLS17a]. 523 [ASMN17a]. 532 [Hou20, dCDFK18]. 548 [GWL^{+}18]. 558 [RB19]. 561 [PW20]. 577 [Ali20]. 582 [KAPS20].

6 [Bar15, Bru16a].

70th [BBF19].

8 [Tsa16b].

978 [Bar15, Bru16a, Tsa16b, Zha15a]. 978-0-521-54823-6 [Bar15]. 978-0-8218-9491-0 [Zha15a]. 978-981-4667-96-8 [Tsa16b]. 978-981-4749-46-6 [Bru16a].

= [SM16, SM18].

A. [CW18a]. A6 [Jel18]. ABC [Che18c, Che19b, GS19a]. Abel [HJW19]. abelian [AJ19, TFC19]. Aberth [RV18]. abs [GBRS15]. abs-normal [GBRS15]. abscissa [HvdD16]. absolute [Eve18]. absolutely [AB19, Ege16]. absorption [BFM19, JT18a]. abstract [CW17, CFM16].

Accelerating [Pan15b]. Accessible [Lax16]. according [BDS15]. accretive [GHK17]. accretive-dissipative [GHK17]. accuracy [KR17, MH15, OZ18, TLL15]. Accurate [MMP17, MMV19, OP19b, SSB15b, BP16b, CLM19a]. Achieving [BH16a]. ACI [BC17a]. ACI-matrices [BC17a]. acoustically [BKLP16]. acting [Vir16]. Action [AM19, AMH11, Fis17]. Acute [Wei17]. acyclic [DdF15, Fer15, FKD15, kJ16]. Adaptive [CDTZ15]. Addendum [BF17b, RB19]. adding [GWL^{+}18, GTL^{+}18, JST19]. Additive [BFMK18, LL17, LY17, PZ17a, QH15]. additivity [ACM15, DFM17]. ADI [RW16]. ADI-like [RW16]. adic [BSS17, EG15]. Adjacency [CR17b, Ore16a, Ore16b, ARS16, AGH17, BK15, CLR^{+}18, LL15c, Moh16, Pea15, SY15, Shi16a]. adjoint [AG18, GG18, KNR18, MY19, Nag15, Nie15a, Pan19b, PM15b, Shi16f, WHC17]. adjointable [AG15b, AG20, XF17]. Adjusted [UM16]. admit [JT18b, Stu17]. admitting [Hua15]. AE [Hla15, LX17, Li19, RW17]. Affine [DA19a, dSP16a, Zue18, LT19a, Ma16, Pan15b, PB16, STZ15]. after [Xu17b]. again [Pan15a]. ahead [UZ19]. AI [Nie19b]. Al [Fis17]. Al-Mohy [Fis17]. Albert [VF17]. Aleksandrov [Geh15]. Alexander [KL18a]. Alexandrov [Ost15]. Algebra [ASMN17a, Ali20, Ano16a, AG20, Bal18, BMVW16, BF17b, BKSP23, Duk15, EKSV18, FJS21, GP18, GWL^{+}18, GHPS19, Hou20, KAPS20, Kis15, KHI16a, LMO16, LT16a, LLS17a, LZSD18, PW20, RB19, SM18, Tsa16b, wXZ19, Yua15, ZL15a, dCDFK18, AHV19, Bal15b, BS15a, CCiT18, CILL16, CK15a, Chi18, CW16a, CD16d, CLLS19, Dog18, DR17, DK18, FJP15a, FJP15b, GM16b, HCS17, Hol16, KS17, KY17, KMS17b, LMS16, Li19, LHG15, MMP16, Mor16b, Mos17, MP15a, MP19, Niv15, Olk15, Pas19, PS18b, RLP19, SGH16, SB15, Ser15, Ter17b, UG18, Wan19, dlP18]. Algebraic [BJ15, CCGV^{+}19, EHK16, Oi19, PS19e, Sch17, AAMM18, Alf18b, BD15a, BBEE19, BB16a, BGCMPB15, BGdCCMSS17, BS18b, CN17b, Das18, DP16, DPJ17, DD17c, GMW18, GL16, KS19, Kol15, NT19, NR16a, PMB15, PRS18, RRV15, UM16, Zas16, Zha17, DP15b]. Algebraically [KQZ16, DB19a, LS15b]. Algebras [CRVSS19, AM15a, AEV16, ALH15, AEV15, AKR15, AK16a, AK16b, BZ16a, BKRE18, BY19, Bal17a, BES16, BGV17, BC15a, BdlCL17, BCS15, BCS16, BLdSV19, BP19b, BFD18, BBS16b, Bre17, BS19c, BCP16, Bur18, CP16a, CMV16, CMV17, CKM19, CLOK13, CPS17, Cha15, CW15a, CL15a, CR18b, CDTZ15, CME18, Cos19a, Cos19b, CDH18, DIR16, Der19, DKNS16, DOR16, EL16, EP18a, EF17, EF19, FKPS18, GGH15, Gol16, Gon15, Gow17, GJ18, HA15a, HAM16, HA16, Hou15, HZ19, JG16, JG17, KKO15, KLP15, KM15b, KSTX15, KLL17, LMO16, LL19, LT16b, LZ16c, LZ17c, LGZ18, LCLW18, LZ15e, MR17a, MI16, MI17, Mol18b, Mon19, MX16, NINS16, OW17, Oi19, Pel17, PR19, PR18b, Rai19, RE16, SBB15, SY15, Shi16a, TW16, Tom15, Wan16c, WZ17b, WFS19, Xu17a, ZK17]. algebras [Zha17, Zus17a, Zus17b, vWZ17]. algorithm [AAG^{+}18, Ais18a, Ais18b, BDK^{+}17b, CLM19a, CKM16, DD17c, Fay18, Fis17, FKM17a, FKM17b, GCC18, Kwo16, LLL17, LC19, Miz16, Pan15b, RKT15, SZ17, Yin19, CA18]. Algorithms [Ano16a, BMVW16, KLS17, Per18, SdJY^{+}16]. Aliasing [BMS15]. ALIF [CGSC19]. allow [LP15]. Almost [SW16, VQ19, CS15b, Pol19, WHG17]. almost-equidistant [Pol19]. Almost-invariant [SW16]. Almost-nonsingular [VQ19]. along [Nag19]. alternate [GL15a, NAS18].

Alternating [BD19, FGH^{+}15, HS16, Kob17, Pan15b, dSP16a, dSP16d, dSP16e, Ter18]. alternative [Koz16]. Aluthge [CM17, EP18b]. amenable [NINS16]. American [Zha15a]. among [BS18b]. analog [MM19b, Nie19a]. analogs [HT18a]. analogue [GKR16]. Analogues [MP17, AJ19, LN17]. Analysis [Ano16a, BMVW16, Tsa16b, AL16, BZ16b, CL19, CN19a, CA18, CGSC19, GZB19, GCC18, GCQX15, JLT17, Kir15, KIS18, KLS17, MM16a, MP17, Miz16, OZ15b, SNP16, See18, SLY17, Bar15, Bru16a]. analytic [JLM18]. analytical [GS18]. Analyticity [DvSW18]. ancestral [ADW19]. Ando [Ric19, Wad18]. angle [LMX19]. Angles [CPZ16a]. angular [CS15a]. anisotropic [AL16]. annihilating [NN19b]. annihilator [HAM16]. anomaly [Zim15]. ansatz [FS18a]. antecedents [Bar19]. anti [AyLPS18, AT15, BF19a, DDSC^{+}16, FT18, GJK18, MV16, YWD16]. anti-banded [BF19a]. anti-circulant [AT15]. anti-community [FT18]. anti-negative [GJK18]. anti-reflective [DDSC^{+}16]. anti-regular [AyLPS18]. anti-triangular [MV16, YWD16]. anticommuting [Hru16]. antinegative [DO16]. antiperiodic [GM15]. antipodal [Fil18]. Anzahl [GL15b, GL15a]. Apartments [Pan17, Pan16b]. Appl [ASMN17a, Ali20, Ano16a, AG20, Bal18, BF17b, BKSP23, Duk15, EKSV18, FJS21, GP18, GWL^{+}18, Hou20, KAPS20, Kis15, KHI16a, LMO16, LT16a, LLS17a, LZSD18, PW20, RB19, SM18, wXZ19, Yua15, ZL15a, dCDFK18]. Application [AMH11, Abi19, AO18, ACM17, BM16a, BFMK18, CRX15, DGGP18, DHW18, FBH19, FL15, GK15, KVP19, PHW16, RGPH16, Ste15, YYSX19, BTN^{+}18]. Applications [APR^{+}17, Ano16a, BMVW16, DH15, He16a, PCI15, AAG^{+}18, AM16a, ANP16, AL19a, BBM19, BKNS17, BJKR17, BDDO16, BBP19, BR16b, Bri17, Bur18, CDR18, CDG16, CPZ16b, CIY15, CLS18, CJ15, CJB18, CSW17, Chu16, CJ16d, DR16, DLQ18, FSSW17, GZB19, GHT16, GL15b, HM17a, He16b, Hua19, HC19, KC16, KKL^{+}17, KK18, LCL15, LCWZ19, LLSX15, Lin15c, Lop17, MVPST19, MM16a, MMP17, MMV19, Mel16a, MNT19, MO18, NSCV16, PS19d, SSCS16, SSB15b, Wan16a, Xu17a, Zou17, Tsa16b]. Applied [Bru16a, LNT16c, Ser16, War17]. approach [ACPR15, BKLP16, BF19a, BBT15, BDP^{+}18b, CC17a, CLLS19, FKM17a, FKM17b, GS18, HW16, JS18a, KÖ16, Mar15b, MR19, NN19b, Ota15, Per18, SZW16, XvdBvdLS15, YDHX18]. approaches [LMM^{+}16]. approximability [KU16]. Approximate [CSW17, CKM16, PSM18, SLY17, AEV16, ACM17, FG17, dMGC19, KMOR19, LUC19, WZL17]. Approximately [Man19, Zha18b]. Approximating [Alo15, GKS19, BDK^{+}17b]. Approximation [FK18, Bob16, CCO15, Gar17, GH17, GS19b, GCC18, JMRV16, KLL16, LL16b, PQY15, PW18, PHW16, RRT16]. approximations [BHK16, LS18b, OZ18, SZ19]. Araki [Hia16b]. Arbitrary [MI16, BCS16, BNST17, BC17a, CMV16, DDCY17, GLAdS19, GM19, HA15a, KSVW17, MM17b, OvdDV17, Wan15b]. arc [XSW19a]. Archimedean [ABG16a]. arcs [ZD16a, ZD16b]. arising [HvdD16, NT18]. arithmetic [IKW16, Zou17, Zou19]. arithmetic-geometric [IKW16, Zou17, Zou19]. Arithmetical [CV18]. Arnold [Lin16a, SS17a, BSS17]. Arnoldi [BMS16a, MP18]. Aron [KVP19]. arrangement [AB17, BS19a]. array [Bar19, YDYY18]. arrays [Bar16a, CW19, CJ17b, CS18a, He15, HS16, He18, KZ17, MMW17, WZ19, YDHX18, Yan16]. arrowhead [SSB15b]. artifacts [BL15]. Artin [SB15]. aspect [CS18a]. aspects [BC16b]. assignment [Bil19, RKT15, WZ17a]. Associated [XvdBvdLS15, ABP15, AG15b, AG20, Bal15c, BM19b, CCiT18, CEM15a, CLT15, CKS16, CJ16a, CN17b, FWB^{+}19, FBH19, FX19, GL16, IHM08, LG18, Mon19, PR19, Pop16a, RMP18]. Association [WG15, BM19a, CHJM18, CGGV15, Han19a, Mor16b, MX16]. associative [AA17, Pel17, Zus17a]. assortativity [EB17]. assuming [LMS16]. asymmetric [CLHL15, FI16]. Asymptotic [CLR16, iT16, VG18, BÓ15, CUPW15, De 16b, Gar17, KLS17]. asymptotically [Dal17a]. Asymptotics [BGS15, BGM^{+}19, MNT19]. atomic [KY18]. attain [CR19a]. attainment [SPH15]. attenuated [Liu16b]. attractive [SH15b]. Audenaert [Lin16c]. augmentation [PZ17a, PS18a]. augmented [Mas17, Ste15]. autocovariance [CIY15]. Automatic [DA19b]. automorphic [GPI18]. automorphism [FKS16, HA15b, TZ19]. Automorphisms [CD16d, AK16a, CW15a, Cos19a, CDH18, Wan15a]. autonomous [SSF16]. autoregression [FK18]. average [Das18, KLL17]. AVMM [BT17a]. Avoiding [KN16]. Avram [BEF16]. aware [Tyg19].

B [AMG19, DGP^{+}15, MR19]. B-splines [DGP^{+}15]. backtracking [AGHN18]. Backward [BCMASS15, Ma15, MP18]. Baer [BBEE18]. balance [CRRY15]. balanced [Mic16]. balancing [Zha15c]. Ball [MM16a, Dym16]. balls [Jia15]. Banach [AHV19, BM17c, BBS16b, CG19b, CDFK19, DK18, JC18, KVP19, KL19b, Mos17, NINS16, NP15c, PMW19, Pop16b, SPBB19, SB15, SBB15, XS17]. band [BGS15, CG15]. banded [BF19a, LDL18]. bandwidth [LDL18]. Bannai [WHG17]. Bannai/Itô [WHG17]. Bapat [Lin15c]. bar [Alf15, KS15b]. bar-joint [KS15b].

Barabanov [GMW18]. Bartholdi [TH16a]. barycenter [HL17a, HL18, KLL16]. barycentric [EG19]. base [BS16e, CDR^{+}19]. base-orderability [BS16e]. based [Ais18a, AM16b, BL15, BB18, BKLP16, CIY15, DGM^{+}18, DV19a, FT18, GCC18, MHA18, RD15, RMKJ18, RBKA16, RW17, Som17, SZW16, VT18, WG15, WL16, XXZ16, wXL14, wXZ19, ZD15, ZZR15]. Bases [Shi15, BDP^{+}18b, DA19b, DA19a, DA19c, VD18, DPV19, DV19b, FS17a, HHR16, LS17a, MP16a, MB15, PM18, Rai19, Szá16, DDMV16, DDV16, WZW19, War17]. Basic [CKM19, SSS15, Ste15]. basis [AC15, DMQ19, Hür15, JMPS18, Lax16, MI16, You15, YMZ19]. Baxter [SS18]. Bayesian [RH15]. BCCB [Kar16]. be [Nom16, dCdSA15]. behavior [AB15a, Tab16, VG18]. behaviors [BMO15, Obe16]. behaviour [BCF^{+}18]. Belgium [FLVV18]. Belitskii [CXLF16, CNX17]. Bell [ZSQZ19]. benefits [LMS15]. bent [ÇMP18]. Bergman [Lee19a]. Berner [KVP19]. Bernoulli [DTZ16, Ern18, He16a]. Bernstein [MP16a, MMV19, YMZ19]. Bessel [JC18]. Best [KRZ^{+}17, BaHOS15, GH17, dMGC19, GCC18, Reb19]. Bethe [KSTY18, SdJY^{+}16]. better [OZ18]. between [AJO15, BBM19, Bar17b, Bün17, CCW16, CDM19, CVV19, DM16a, DS17a, DMG17, DdFR16, EG19, HS19b, LL16a, LS19, LHL18b, MSSZ17, Min17, Mol18b, Obo19c, PMW19, RS17a, TW18a, WDFS17, Zui17]. Beurling [AS17]. Beyond [Nik16a, Fie15a, KL18a]. Bezout [BS19b, ACDT^{+}15, Kwo16]. Bezoutian [Ota15]. Bhatia [DDF17a, Kim18, Lin17b]. bi [LMR16]. bi-infinite [LMR16]. biased [RBKA16]. BIB [HMS17]. bicolored [CD16c]. Bicyclic [BSP18, BBC18, LWX17, OQY17].

Biderivations [LGZ18]. Bidiagonal [FN17, MM16a, AAG^{+}18, Hua15, MMP17, MMV19]. bidiagonal-type [Hua15]. bidual [KVP19]. bijection [Ter18]. Bilinear [CG19b, HSTW17, AAdFS19, DPS19, FM16, GLW17, Mar18, MdlP16, dFFRS17, CA18]. Billiard [Yan16]. Bilocal [CL18]. Binary [HI16, MBB^{+}18, TVD15, Alo15, BS18b, CD16c, GCQX15, KM18, PS18a, Sei18, Tok17, Yor17]. biography [Olk15]. bipartite [BN17, DMS19, FuIKT16, GSZ16, HLSP17, HLW15, JKS15, uIKF15, LN17, aLwW15, MMP16, MM17a, MHN18, PS15, SZ17, Sim18, Sim19, VSV18, WHG17, WZL15, Zaj19, ZG18]. bipartiteness [RMS16]. bipartitive [BmBCC17]. bipolar [GM17]. Birkhoff [AR15b, CQCZ19, DDCY17, DD16a, DD19, DU16b, DKPU18, Fer15, kJ16, KL19b, PSG16, PSM18, PMW19, SPH15, VB19]. Birkhoff-James [KL19b]. birth [CLM19a]. birthday [BBF19]. bisexual [DOR16]. Bishop [DGK^{+}17]. bisolvents [CP17]. bisymmetric [JS15, SW18b, SW18a]. biunimodular [FR15]. bivariate [Liu17, Ple17]. Blaschke [DGSV17]. blind [Win16]. Block [AMNR18, DPV19, ELN18, FS18a, APT17, AL19b, BJ16, BDLM16, BFD18, BDFR15, BDP^{+}18b, CL19, CGMSR15, Cho17b, dMGC19, GHK17, GLRT18, Had17, Hay19, KS15a, KLL19, KL17, KM16, Lee15a, LZ17b, LGZ18, LSM16, MV16, Mas17, MN19a, Mic16, NZZ15, Nie19c, NK18, PCL15, PQY15, RRT16, SNDM17, SP16, Sot17, Ulu19, You15, YWD16, Zha19a, SK17a]. block-augmented [Mas17]. block-matrix [Hay19]. block-monotone [Mas17]. block-shift [APT17]. block-symmetric [BDFR15]. block-term [dMGC19]. block-Toeplitz [BDLM16, NK18]. block-triangular [BDLM16]. blocked [SY16b]. blocks [AMNR18, Hir18, Kal16, Sot17, Ulu19]. blow [LM16a]. blow-up [LM16a]. Board [Ano18-27, Ano18-28, Ano18-29, Ano18z, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18a, Ano18-30, Ano18-31, Ano18-32, Ano18-33, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano18-40, Ano18-41, Ano18-42]. Board [Ano18-43, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18y, Ano19a, Ano19-27, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano19-36, Ano19-37, Ano19-38, Ano19-39, Ano19-40, Ano19-41, Ano19-42, Ano19-43, Ano19-44, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z]. boards [War17]. Bohnenblust [CJRMF^{+}15]. Bohnenblust-Hille [CJRMF^{+}15]. Bollobás [DGK^{+}17]. Book [Bru16a, Tsa16b, Zha15a]. Boolean [ABC^{+}16b, Kim15b, PS18a, PRS19]. border [Zui17]. bordered [DE15]. borderenergetic [DLG16]. Bordering [OvdDV17, CN17a, CN17b, CN18]. borderline [FM19b]. Bott [KC16]. bottleneck [Mol18a]. bound [ARR18, AL15, ARS17, BL18a, Bal15c, Ber18, BH16a, CLF18, GRS15, HW15b, IL17a, LCWZ19, LwW16, Mol18a, Obo19b, Vec16, WF19, Xu17b, ZWS18]. boundaries [AGV17, BS18b]. Boundary [CN19b, BL15, CN15, DDSC^{+}16, EHK16, EH18, GM16a, HJW19, Piv19, YLB15]. Bounded [Pan19a, AB19, CM19, DDL17, DD17b, GSP16, LT12, LT16a, LZ16c, PSG16, PSM18, dSP16a, dSP16c, dSP16d, Pep17, RMS16]. Bounding [Sta19, AO18]. Bounds [ALMZ^{+}19, DdSJdFDV15, FW17, Gil16, HHL19, HMNP15, Lia17, LL18a, Mel16a, OP16a, Pep17, Rum19, VV18, WM17a, dlPJG18, ACT16, AR16, ACRR15, AGRR16, AdFRR18, ADLR19, AP15a, BG15, BaHOS15, CR19b, bCR18, CLL15, CIY15, CL17b, CLS18, CJ16c, CJ16d, DHHZI19, DV19a, DR18, DM18, Drn19, EA17, GM17, IGW17, Ken16, Lan15, LLV15, LWY16, LZM16, Lot15, MR17b, MS17b, MHS15, OVW18, Rod19, SZ23, SK17b, Sku17, TLL15, TWM16, WZL17, WLL19, WZW19, WWC18, YZL15, vDS16, ELN18]. Braess [HK19b]. Brauer [BJLD17, DLB18, LL16a]. Brauer-type [BJLD17, DLB18, LL16a]. Bregman [MPV16]. bridge [CDM19]. bridging [LS17a]. Brill [Gua18]. Brouwer [GP18, Che18b, Che19a, GAP16, HT17]. Brown [BŚS18]. Brownian [MMS16]. Brualdi [BWSZ15b, LL16a]. Brualdi-type [BWSZ15b, LL16a]. Bruhat [dCFF17]. Brunn [Liu16a]. bundles [DFK^{+}15, EM16, Gal17]. Bures [Mol18b]. Butson [EC19]. butterfly [LZSD17, LZSD18]. butterfly-like [LZSD17, LZSD18].

C [COS18, BCP16, CP16a]. cactoid [HC15]. cactus [Gal17]. cages [Fil18]. cake [CA19]. calculation [PW18]. calculus [Nev18]. CAM [Bru16a]. Cambridge [Bar15]. Can [dCdSA15, Nom16]. CANDECOMP [KL18b]. CANDECOMP/PARAFAC [KL18b]. Canonical [BBdH18, DD17c, LS15b, AEKP16, BTW16, BGSS18, CXLF16, CNX17, CW17, DFK^{+}15, FdC15, GLAdS19, MN19b, Rad17]. Carathéodory [BFMK18, IL17a]. cardinality [CRVSS19, MMR17, Yor17]. Carlen [Bek16]. Carleson [DPS19]. Carlitz [KT19]. Carlo [WC15]. Cartan [HL17a, HL18, KLL16, LL19, MM19a, PR19, WZ17b]. Cartesian [BK19, KMY15]. case [ACS19, Bar17a, BDK^{+}17a, DV19b, EHHL17, lHS16, Mat15a, Pan15b, ZHC16, dFRS16]. cases [MP17]. Catalan [HS17, YDHX18]. category [MVPST19]. Cauchon [AAG^{+}18]. Cauchy [JM16, JKP16, Kal18, LL16b, Lyn15, MMP17, Mel16b]. Cauchy-like [Mel16b]. causal [FKD15]. Cayley [AT15, BBM19, BBS16b, CFH19, KM15b, KLW15, LZ15d, RB16, SLY17, TFC19]. Cayley-type [RB16]. cell [KT15]. cells [GMLT19]. Central [Yin18, BDDO16, EF17, EF19, Lék16]. centrality [AM16b, BFM19, HKS19]. centralizer [AGP^{+}17]. centralizers [Han19a, dCDFK17, dCDFK18]. centralizing [Wan16c]. centrally [Mig16]. Centrosymmetric [JRS19, Bur15]. certain [AKYD16, Ege16, EEV17, FKPS18, Hia16a, Joe16, JMP17, Kol15, KLW18, Lin17a, Liu18, Lot15, Mol19, NP15c, PP18, Sch17, She17, Tab16, VHB18, vWZ17]. certificate [CJ17c, Dic19]. Cesàro [AB19]. chain [AAS16, BK17c, Cho19, WC15]. chains [BLPT15, BCF^{+}18, Hun16, Hun18a, Hun18b, JLT17, LZ15a, Mas17, MP16b, NP15b, SY16a, Sku17, Ter18, VB18, Vas15]. Change [DFK^{+}15, JST19, VSV18]. changing [GM15, SV17]. channels [FL16, HJ19, LPPZ19, Lev18]. chaos [BCMASS15]. character [Che15b]. Characteristic [AW15a, BLdS16, Ben18, BCS16, BBFF17, BCF^{+}16, CCD19, CK15b, EV15, FOvdD16, FMR18, HY18, HZ19, HZ17, KMS17b, LL18b, MW15, MMR16, MMR17, MMR18, Niv15, Shi16f, WW18, WLWZ19, ZGW16]. characteristics [AKPS19, JKKL15, KAPS20, Koz16, MNTX16]. Characterization [BJKR17, DP16, GLZ16a, GSSvdD16, HT18b, JT18b, Kim15b, LD16a, MWT16a, QH15, RW17, Wój19, ZXW19, Abi19, CRRY15, CLR^{+}18, DGCC16, DP18, DP15b, DPJ17, lFwW16, GGMPC15, GH16a, GX17, HvD18, HJW19, Hoa15, KY18, LZSD17, LZSD18, LWCL19, MHL15, MM16b, NS18, PP15, SW18a, zSqST15, TSH16, WM17b, XvdBvdLS15, YCL17]. Characterizations [HM17a, Mol19, RS17b, AyLPS18, ABDN18, BP15, DDF18, He15, He18, LL15b, LWC18, SBB15, Sol19, UYY17, WZ15, ZSWB16]. Characterizing [BDMB19, CJ16c, FMR19, LZ15b, HK19a]. characters [AR15a]. Charles [Bar15]. Chebyshev [Guo19, KMM17, Yam15]. Chevalley [MMR18]. Chief [Ano18-44, Bru15, Bru16b, Bru17]. Chip [BMZB18]. Chip-firing [BMZB18]. Choi [GKR16]. Cholesky [LWY16]. Cholesky-like [LWY16]. Christoffel [AM15b]. Chromatic [KM19, AL15, FCL^{+}16, KNY15, Tom15, vDS16]. circle [Cas19, IR19, MS15a]. circles [Rum19]. circuits [MT15]. circulant [AT15, AMNR18, Hil17b, Joe16, KK17, Lee19a, NW19, Nie19c, PSZ15, PS18b, QWC15, RR18b, San15, San18, SB18, TG19]. circulants [DM16b, Sbu17]. class [ASMN17a, ASMN17b, AMR18, AMG19, BG15, BPP16, BR16a, BHP17, CX17, Chu19, CM16, FPGPV17, Hou15, HSTW17, JLM18, uIKF15, KY18, LDL18, MI17, Min17, Mir18, Nob16, ÖK18, Ris16, RBKA16, Sam17, XXZ16, ZD15, Zhu19]. classes [AG15a, CHJM18, DGS18, EC19, Egg15, ETF18, Far19, FMR19, Gre18, GJLS16, LCF19, MR17a, Mar15a, Pan19b, Pel17, Rob19, Sha16, SS18, VLGS15, Ye17, ZM16, dSRST17]. classical [HLG15, OVW18, Ota15, Saw16]. Classification [AMC18, BKRE18, BCS15, CMV17, CLOK13, DdFR16, EEV17, HA16, LMO16, LZ16c, MSSZ17, RE16, SWZ18, GSZ15, GSZ16, Gas19, HAM16, Sim18, YdB19, dFRS16, dFFRS17]. classifying [BP19b, FKPS18]. clean [BM16b]. Clifford [CK15a]. clique [FHJT19, HFS15, Yus17]. clique-width [FHJT19]. close [BCF^{+}18, BaHOS15]. Closed [DKSV19, AYK17, Bie15, LS15b, LMX19, XS17]. closest [KT15]. Cluster [TLL15]. Cluster-robust [TLL15]. Clustering [BCF^{+}18, EB17, Bai18, Est16]. clusters [CR17b]. clutter [MFdM15]. CMV [MS15a]. Co [Bru16a, Tsa16b, Alf18b]. co-rank [Alf18b]. coadjoint [BC15a]. coarse [KN16]. cocycle [Ong18]. code [LA16]. codes [AGP^{+}17, ANP16, BDMB19, CG16, CG19a, CLHL15, CMPZ18, DGGP18, GZ19, GR16, La 16, LCF19, Ouy17, Pud16, SSS15, TZK17, TZ19]. codimension [MS19, RZ17]. coefficient [BDDSC17, PRS18]. coefficients [BS15c, BF17a, Ben18, CLR^{+}18, JLM18, LG18, Sug17, TH16a, WZZ19]. Coffey [SSCS16]. cognitive [Dog18]. Coherence [CKL17, RGPH16]. Coherent [DD16b, GGMPC15, Xu17a]. Cohomologically [BCS16]. Collatz [GH16a]. collections [HJ16]. Collocation [JKP16]. Collocation-quadrature [JKP16]. color [AT15]. colored [Kal16]. Colorings [KL18a]. colourful [BFMK18]. column [SS17b]. combination [Kis15, XX12, Ziz16]. combinations [KP17]. Combinatorial [ACD19, Nik17a, BR18, CJ15, CJS16, EJ19, MR19, Ser15]. Combinatorics [AMPT16, LMP^{+}17, AM16b]. Combined [BGGS16, Mat15c]. cominuscule [MM15a]. comments [AdF19]. common [CKM16, Gór17, GX16, GX19]. communicability [DHHZI19]. community [FT18]. commutation [Nie18]. Commutative [BS19b, AM15a, AA17, GPI18, GMM16, KM19, LZ15d, OW17, SNDM17, Ser16]. Commutativity [Pan19b, Pie18, SB15]. Commutator [GH15, CL17b]. Commutators [SB15, CL15b, GW17a, Hou19, XCJ18, dlCdR17]. commute [BFFN16, Kis15, XX12]. Commuting [Bou16b, CKT19, Shi18a, Wan16c, BF18a, BS19c, BKS15, BKPS17, BKSP23, CM19, CJW19, DBK17, FKPS18, GMM19, IKVZ18, KS17, LdlP15, LY17, Mig16, MT16, QH15, Sha16, aHRT15]. Compact [ABM16, CN19a, JLM18, LWC18, PS19c]. compactly [AKPS19, KAPS20]. Companion [EKSV14, EKSV18, BM16b, GSSvdD16, Lee15c, VV18]. compared [QKP16]. comparison [ABD19, DU16a]. Compatible [BKS15, BKSP23, LT19a, dSP15, dSP16e]. competition [FGS^{+}16, Kim15b]. complement [AG16, AAMM18, CGMSR15, Row18, WYL19]. Complementarity [JS16, See18, Bal15a, Bal15b, BP16a, CFNP17, DN16, DN18, WLL19, XXZ16, wXL14, wXZ19, ZZR15, ZLQ16]. Complementary [FJT17, Szá16, CCdFV18]. complements [CMM19, LD17a, Row15, YCL17, YZLC18]. completable [SMBDK16]. Complete [BSMSZ19, HK19b, BK18, BES19, CI15a, CGSZ16, DP18, DD17b, LNW17, NS18, Obo19a, Obo19c, Shi19, VSV18, WYL19]. Completely [Gow16, LT12, LT16a, Sal16, Zha18a, BDS15, GdLL17, SSV17, DDMV16, XLQC16]. Completion [Ber17, Dod16, MFdM15, BA15, CMQV16, DH15, Dod17]. completions [Cal16, CJ16b, JLD17, PCI15, VQ17]. Complex [HP15, LWX17, AM19, AGQ19, AS17, AP15a, BFY18, BH16a, BSS17, Bün17, CDM19, DL16, ET18, Fik18, FM19b, FKM17a, GT16, Gór17, Hut17, Kar16, Mag16, Mat15a, Nev18, NW19, PM15a, PR19, QWW18, RGPH16, SK17a, Tan16b, TCD15, dlCMP17a, dlCdR17]. complexes [LT18]. complexity [BC16a, FMY16, HI16, KMM17, RR18a, VGG17]. component [FK18]. components [BFFN16, CLT15, DDL17, DBK17, ÖK18]. componentwise [DS18b]. Composition [AL19a, BMN16, HA15b]. compound [AGN18, APR^{+}17, BSS16, FK18, RS17b]. comprehensive [BBT15]. Compressed [KKLP17, BÓ15, Fli16]. compressing [BE15]. Compressions [PS19c, He18]. computation [BFM19, CRX15, Hun16, Hun18a, Miy19, Sim19]. computations [BP16b, Win16]. compute [BV18a, Pas19]. Computing [AMH11, BDLM16, DU16a, Eld15, FI19a, Fer15, HR17, Hir18, KL18b, MMR17, Saw16, Fis17, SK17a, SSH15, SZW16]. Concatenated [DGGP18]. Concavity [Hia16a, FS17b]. Concentration [Bar16b, Ngu18, FL15]. concerning [AKYD16, Hil15]. Concini [FH19]. Concise [OBRA15]. concrete [Uch15]. Condition [ME15, Van17, WYL15, ABD19, Ben16, CC17a, DR16, DS18b, Fli16, LS17b, Nie19b, Sch17, Zim15]. conditioned [Bar16b, NSCV16]. conditioning [FR16]. conditions [Bar17a, CYD19, DDSC^{+}16, DD17c, ES16, EH18, FZL^{+}17, GM16a, Gu16, Li15a, LX17, LSX15, MPS17, PM15a, ZLG15]. cone [Bal15a, BP16a, CCCNT19, CFNP17, FG17, GR15, Gow16, HC19, KOST15, Lee19b, Ser15]. cones [BFN15, BMZB18, GP19b, GJ18, IL17b, JG16, JL16b, KLL17, Kut18, Sol18a, Sol19, Zha18a]. Conference [CFH^{+}16, GPW15, GHPS19, ET18, TG19]. configurations [BF19b, FGG^{+}18, Xu17a]. congruence [BGSS18, BHSS18, DFK^{+}15, Dmy16, Dmy19, Egg15, YZZ18, dlCF16]. Congruences [Sbu17, Sim19]. coninvolutory [AMR^{+}16]. conjecture [AB15a, AKYD16, Bah19, Che18b, Che19a, CL15d, Das18, DDF17a, DK16, eFLYL15, GAP16, GP18, Gho16, GO18, GLMS18, HT17, KM19, Koz16, LOS19, Lin15a, LMP^{+}17, Riz19, SY16a, Shc16, XS18, YQS16, AA19, BS16f]. conjectured [AL15]. conjectures [BCJ^{+}16, Das15, TWR15, Zha17]. conjugacy [Pan19b, VLGS15]. conjugate [DMG17, KR17, Lin15b, Nev18]. conjugate-normal [Lin15b]. conjugation [Bik19]. Connected [LM16c, CLT15, COvdD16, DPR19, DS19c, GWL^{+}18, GTL^{+}18, HSS16a, JMM16, LLL15, NLWJ18, PS15, SS17a, See18, VK19]. connectedness [GJ18]. Connecting [ES16]. connection [CHB15, DGSV17, RS17a]. Connections [GHPS19, CJ17a, Nie15b]. Connectivity [ZLG17, AAMM18, Das18, Gu16, HXL19, Kol15, NLW18, O16, Row15, XSW19a]. consensus [VK19]. Conservative [KLP15, Geh15]. considerations [AAC^{+}15]. consistency [Shi16b, XvdBvdLS15]. consistent [BMS15]. constacyclic [CLHL15, La 16]. Constant [PRS18, BBH16, BC17a, EM16, MMM19, PV17, WDV17]. Constant-coefficient [PRS18]. constants [BCKL17, Nun17, Reb19]. Constrained [CMQV16, DOR16, OS17, Rak17, TD15]. constraint [LSM18, MT15]. constraints [BT17a, DAG16, LPPZ19]. construct [LMPM17]. constructed [Mir18]. Constructing [CR18a, CJW19, CJ17d, DDV16, WC15, ATM18, BH16a, MLW15]. Construction [HHL15, JCW19, LS15c, MM15b, BBM16, CW17, DGGP18, FdC15, Gór17, Hey17, MW15, MR17a, PS17]. constructions [FJMP15, GF17, LA16, HL17c, Pud16]. contain [FJKM18]. containing [AMG19, FJ19, Gre18]. Contemporary [Bru16a]. Continuity [CG19c, Gol17, AB15b, BP16a, GMW18]. Continuous [BS19c, Cos16a, Mas17, ZH17, Zhu19]. Continuous-time [Mas17, ZH17]. Contour [Wim16, Van16a]. contracted [CEM19, XCJ18]. contraction [McK18]. Contractions [BC15a, LPW16]. Contractive [AKS17a, AKS17b, LNW17, LCNZ19, Lin16b]. contribution [Geh15]. control [Bel16, BTN^{+}18, Ste15, BCMASS15]. controllability [AG15a, TD15]. Controlling [Yas16]. convection [BDDSC17]. convection-diffusion [BDDSC17]. Convergence [CN19a, GCC18, JMRV16, KR17, OZ15b, SSF16, BM17a, BW18, BMS16a, CGSC19, DHW18, GKS18, Guo19, KH17, Sch16, SLY17, ZZR15]. convergent [Ais18a, Ais18b, EE16, JCW19, QWW18, Sku17]. converse [Lev15]. Conversions [EG19]. Convex [BKNS17, EE15, Han19b, BGCMASS19, BU18, CW18b, EM15, Eve18, KP17, Naj17, RW17, SXD16, Vir19, WH19b, ZC15]. Convexity [Dym16, JK15, LNT16a, Sab16, Bek16, CFL16, FNX19, Kum19, Nor18, Sab18]. convolution [ABC^{+}16b, Ziz16]. convolutional [ANP16, CLHL15, DGGP18]. convolutions [RLP19]. coordinate [KRZ^{+}18]. coordinatization [Lax16]. Copositive [Hil17b, SM16, SM18]. copositivity [Dic19]. coprime [AMZ16, Lie18]. corank [GSZ15, GSZ16, SZ17]. corank-two [GSZ16]. Core [EB17, CMV19, HPZ18, HPZ19, JS19, KY18, NW19, WL16, Wan16a]. Core-EP [Wan16a]. core-nilpotent [JS19, WL16]. Core-satellite [EB17]. cores [GT17]. corner [bCR18, ZMT19]. corona [BS17, LJK16]. coronas [ABC^{+}16a]. correction [WP17, Zas16]. correlated [Sma15]. correlation [CPZ16a, Tab16, iT16, Zim15]. correspondence [GPI18]. Corrigendum [ASMN17a, Ali20, Ano16a, AG20, Bal18, BKSP23, DIPR18a, Duk15, EKSV18, FJS21, GP18, GWL^{+}18, Hou20, Kis15, KHI16a, LMO16, LT16a, LLS17a, LZSD18, PW20, SM18, wXZ19, Yua15, ZL15a, dCDFK18]. cosets [LA16]. Cospectral [CHJM18, DF16, WQH19, AHMM18, BH16b, Hey17, LMT19, ZLS17]. cosquare [GMP15, GMP16b]. cost [CYD19]. Coulomb [BS18a]. Coulson [ALOR19]. countable [JLT17]. counterexample [Bah19, Shi16c, Shi17a]. Counterexamples [SY16a, FJS17, FJS21]. Counting [FMM16, Mat15b, PD17, Ram16, VLGS15, GRV16, MT15]. counts [KS15b]. coupled [HAWM16, Lee19b]. coupling [CRC17]. Covariance [CLZ^{+}16b, GH16b, KRZ^{+}17, LHC15, MNT19, iT19, Wan16b, Yin18]. cover [PF18, WM17a]. Covering [WH19b]. coverings [FHH15, MC19]. covers [CGSZ16]. Cox [Sim19]. Cox-regular [Sim19]. Coxeter [GSZ15, Gas19, Mró16, MdlP16, Sim18, dlPJG18, dlP18]. Coxeter-Dynkin [GSZ15]. CP [BDS15, XCQ18]. CP-plus-rank [BDS15]. CP-rank [BDS15]. CRAIG [HKP17, TW16]. criteria [DIPR18a, DIPR18b, DFKS17, FKS16, SFW18]. criterion [FHS17b, SS15a, Sam17]. Critical [BBT15, Gae16, GP19b, DHS18, FCL^{+}16, HMPT18]. crossings [dS19a]. Crouzeix [GO18]. CS [CRX15]. cubature [BBM16]. cubic [CLR19, Jel18]. cumulant [OBRA15]. curvature [NR16a, Opo16]. curve [CN15, LHC15]. Curves [Ber18, ÁVAK^{+}19, BGCMPB15, CS15a, CN17b, CN19b, LP18]. cut [Alo15, CA19, DV19a, Nik16b, WWT18, vDS16]. cut-norm [Alo15]. cycle [DF15, DF18, DR19, HFS15, Liu15, PF18, WZW19, dS19a]. cycle-clique [HFS15]. cycles [ABDN18, Ash19, GH19, MWT16a, Mac16, NY15, WZZ19, Yua14, Yua15, ZLG15]. Cyclic [BGV17, BM18, BMR17, CLUP17, Dal17b, Far16, MP16b, Ziz16]. cyclomatic [HHL19, TH17]. cyclotomic [CFW15, LA16].

D [BKLP16, Sam17]. DAHA [NT17]. Darboux [AM15b, CG15]. Dashnic [LCWZ19, ZLLL18]. Data [SUBG18, BCJ^{+}16, CS15a, CJ16c, HPZ19, KRZ^{+}17, KRZ^{+}18, MM15b, Wei15]. Data-driven [SUBG18]. Davidson [BM17a, WP17]. Davis [ABR18, CJ17a, LSZ18]. DCT [BP19a]. DDVV [GXYZ17]. DDVV-type [GXYZ17]. Dealing [BL15]. death [CLM19a]. deblurring [DDSC^{+}16]. Decay [MR17b, BMR17, PPZ19]. decomposability [LL17]. decomposable [CW18b, YLT16]. Decomposition [FM19a, GT16, Hou19, ÁVAK^{+}19, BC19b, CRX15, CQCZ19, CP16b, DD17c, DU16b, DKPU18, Duk12, Duk15, FZwCW16, FR16, FX19, GH16b, dMGC19, GMP15, GMP16b, GM17, HKM18, HAWM16, Hir18, JS19, KMY15, KL18b, Lop15, MVPST19, MM16a, MMP17, MMV19, MFdM15, MV16, MMR18, Nie18, Nie19a, RBKA16, Saw16, SP16, SSB15b, SSB15a, SG17b, Van17, Wan16a, WL16]. Decompositions [KK18, Mar18, BBEE18, BFW17, CMV16, CCGO17, FSSW17, FSW19, Lee16, Nom15c, Ota15, Szá16, Tyg19, Urs18, Ye17, dlCMP17b]. deconvolution [BL15, Win16]. decreases [JDS17]. dedicated [BBF19]. defined [He16b, Hon18, HC19, PMW19, Ram17, Zho17]. defining [CR18a]. definite [ACM17, BS15d, BJL19, BC19a, BDK^{+}17b, CJ16b, CLS17, Fie15b, GLRT18, HL17a, HK17, KL15, KLL16, LZ17b, LNT16c, Min17, MS15b, MPV16, PR19, Riv15b, Sch16, SH15b, STZ15]. definiteness [DLB18, HT18a]. deflated [KR17]. deflating [WZL17]. Deflation [LNT16c]. deformations [BSS17, Dmy16, Dmy19]. degenerate [Bob16, KK17]. Degree [DGM^{+}18, Chu16, CW16a, Das18, De 16b, DD17b, DK16, Fer15, NLW18, Sku17, Wan15b, XWL17, YHY15, ZLS17, ZWS18]. Degree-based [DGM^{+}18]. degrees [BD15a, DMG17, DV19b, LLS17a, LLS17b, PS19a]. delay [BTN^{+}18, FM17, PMB15, SNP16, WZ17a]. deletion [VSV18]. Dellac [BF19b]. Demmel [AGV18]. dense [Dal17a, Duk12, Duk15, FGH^{+}15]. density [BDG15, Lie18, Mol18b, SCD17]. dependent [DHLT19, GM16a, PCL15]. depending [EH18]. derangement [MR19]. derivable [WMC15]. derivation [LZ16c]. Derivations [Bre17, Hou15, AKR15, AK16b, AA17, BCP16, CL15a, CL18, EP18a, LZ17c, NP15a, ZK17]. derivative [UUG15]. derivatives [Eld15, Gir19]. derived [JLD17, LMR16]. Desarguesian [Van16b]. descent [VK16]. descent-like [VK16]. descents [AR15a]. description [BmBCC17, DDL17, Ong18, PS19e, Ser15]. descriptions [DSCD17, May15]. descriptor [Mar15b]. design [AKPS19, Bel16, KAPS20, Mic16, RGPH16]. Designing [Van16a]. designs [BMS15, BH16a, BL16, HMS17, Ma16, Sma15]. detection [FT18]. determinant [BDK^{+}17b, BR17, Cho17b, Cos19b, DdC16, GSS15, Hil18, Hol18, HLS^{+}16, KMS16, Rum18, RB19, Rum19, Shi18b, Yur15, Min17]. Determinantal [LP18, CNY18, CN18, DP18, HI16, LS18a, MT15, Ple17, dCdSA15]. Determinants [HZ19, Riz19, Sot17, ABO15, BS18a, BK17b, BaHOS15, BR18, CLLS19, HM16, Sbu17, Sei18, Zhe19]. determination [AAG^{+}18]. determined [ALH15, BCC^{+}18, BSMSZ19, HNR16, KHI16a, KHI16b, MW15, MLW15]. detrended [CIY15]. Deveci [CL15d].

Developed [DKS15]. deviation [DHLT19]. DGL [BBMFM17]. DGMRES [GKS18]. Diagonal [AM16a, AMV17, APS18, BBH16, Bik19, BR16b, BMR17, ELN18, Far19, GX16, GX17, HN18, HK18b, Lee15a, Nom15b, PM15a, PPZ19, SP16, SV17, SSB15a, ZH17]. Diagonality [ACM17]. diagonalizability [dlCF16]. diagonalizable [JLDS18, JKP^{+}17, wXjMZL19]. diagonalization [ACM17, CL19, LUC19, cSfCfX15]. Diagonally [Roh18, DM18, HW15b]. diagram [Bal15c, LCF19]. diagrams [Sch15]. diameter [BK17a, CC17b, CNM19, DBK17, HSS16a, LS15a, Mig16, XWD18, ZLG17]. Diaz [MF16]. dichotomy [Dra18]. dicyclic [CFH19]. difference [Dra18, EJ19, PHW16, SNP16, SJD16, Sug17, UWYY15, UZ19]. differences [Cos16b]. different [ALMZ^{+}19, SY16b]. Differential [Gon15, BMO15, HLM17, LMS15, Mat15a, Obe16, PMB15, PRS18, RRV15]. differential-algebraic [PMB15, PRS18]. Differentiating [Hol18]. Differing [Dog18]. Diffuse [GHS16]. Diffusion [VK19, BDDSC17, CS15a, RW16]. diffusively [Lee19b]. digit [JT18b]. digraph [ARR18, BDMB19, HC15]. digraphs [AR16, ARR18, AGH17, DF16, Dal17a, Dal17b, FGS^{+}16, FJ19, GLMPC18, GMO19, JL16b, Kal16, LWCL19, LMP^{+}17, MS17b, Ost15, SG16, XSW19a, XSW19b, ZSWB16, dS19a]. Dilations [BEKS17]. dilogarithm [BV18b]. dimension [BCS15, BBT15, CMV16, DD16a, DD19, FI19b, Hol16, IL17b, JKKL15, KS17, MB15, Pas19, SB15, VB19, YT16]. dimensional [Alf15, AKPS19, AKR15, AK16a, AK16b, BC15a, BB16a, Ben18, BGdCCMSS17, CMV17, CDFK19, CW15a, CL16, CJ17b, CG19c, Cos19a, EHHL17, EHLP18, ET18, EF17, EF19, FJM18, FMM16, FKPS18, Gol17, HAM16, HA16, lHS16, KAPS20, KLP15, KL19b, KU16, LT16b, LZ15e, MS17a, Min17, NINS16, OW17, Pan17, PMW19, Pel17, QWC15, RE16, Sab19, Vir16, WH19a, Yan18, Yin18]. Dimensionality [BF19c]. dimensions [BNST17, CLUP17, bCR18, DDCY17, Fre18, Geh15, PP15]. dimers [NS15a]. Dirac [FKRS17]. Direct [KP19, DSCD17, GF17, KS15a, Xu17a]. Directed [Jør15, ASMN17a, ASMN17b, AM16b, BK17c, CLN15, DL16, Kob17, VK19]. directional [AKPS19, Gir19, KAPS20]. directions [JZ16]. Dirichlet [LN18]. discontinuity [CG19c]. discrepancy [CLZ^{+}16b]. Discrete [ACE15, EH18, Lev15, BB18, BPT15, CEM15b, DPJ17, EE15, EM15, FCLP18, GM15, GM16a, GKS19, HM17b, LZQ19, LS18b, LN18, SL16, SH15a, Shi16b, TNK16, ZD15, Zhu19, CGSC19]. discrete-time [GKS19, SL16]. discriminant [Bar16b, HS18]. discriminants [CP16b]. Disentangling [ZS17]. disjoint [WM17b]. Disjointness [LCLW18]. displacement [TW18b]. dispositions [LMP^{+}17]. dissipative [GHK17]. Distance [AJO15, AAKS17, BK17a, BF15, CDR18, Che15a, DS17a, LZ16a, LZW17, Obo19a, VSV18, AAB^{+}16, Alf18a, AP15b, AB18, BD15a, BG15, BS15b, BS16a, BK19, BGOY15, BF18b, BMS16a, BGT^{+}19, CG16, CVV19, Cho19, CGGV15, CHT19, DdSJdFDV15, Das15, DR18, DL15, FdFdSDV18, GK15, Geh15, GLW17, Hey17, HC15, HFS15, KPL18, KHI16a, KHI16b, KT15, KB15, LM16b, LZ15c, LLL15, LD16a, LWCS16, LZ16b, LD17a, LDZ18, LHG15, LHH17, LHL18a, Ma16, MQTW18, MMP16, MM17a, MS15b, Som17, TWR15, TWM16, TW17, TW18a, XSW19a, XSW19b, XS18, ZLG17, ZD16a, ZD16b, Zho17]. distance- [BF15]. distance-based [Som17]. Distance-regular [AAKS17, BK17a, BF15, CGGV15, GK15, GLW17, LHG15, MMP16, MM17a]. distances [CEM19, Ngu18]. distinct [AAKS17, BBH16, BDVRT15, BF15, HZ15, HH17, JDS17, LOS19, Mon17, Row16b, Xu17b]. Distinguishing [DM15]. Distribution [BB16b, DMT16, HMSC19, HLLZ17]. Distributional [BCMASS15]. Distributions [Jia15, ABvN19, SSF16]. diverge [Est16]. divergence [Vir16]. divergences [Hia19, Min17, MPV16, Vir19]. divisibility [AYK17, BC16a]. divisible [VB18]. division [BZ16a, BKRE18, CKM16, LT16b, Mig16, dSP16b, RE16]. divisor [CQ15a, Wan15a]. divisors [AMZ17, SJS15]. DM [Hir18]. DM-decomposition [Hir18]. do [CR19a, GL17, Gao18]. does [MHL15, Wad18]. domain [HR17, Kwo16]. domains [BGCMLSSS16, BS19b, CQ15a, Cal16, Ris16, Vol17]. dominance [ELN18, Far19, HT17]. dominant [DM18, HW15b]. domination [Lin19]. Double [FM19b, Han19a, LL15a, LMR16, AAC^{+}15, He18, PF18]. doubling [CLM19a, KLS17]. doubly [ATM18, BM18, Ben18, CC19, CQCZ19, Chu15, Dah15a, DU16b, DKPU18, Dya17, Gow17, KRZ^{+}17, KRZ^{+}18, Lju15, LD17c, NMBA19, Nat19, dCDFK17, dCDFK18]. Dowling [LM17]. down [LT18]. Drazin [QWW18, Rak17, YWD16]. Drinfel’d [Yan18]. driven [SUBG18]. Drury [Lin17b, Nie19a]. DST [Per18]. dual [BMS15, De 16a, DN18, Fay18, GR15, Gow16, KK18, NAS18, RR18a, Shi16a, DDMV16, DDV16]. Duality [NP15b, STW18, WO17, NR16a, RT16]. duals [BM17c, ÇMP18, CJ17c, JLD17]. Ducci [OZ15a, SB18]. due [JST19, Lin15c, Nun17, VSV18]. Duffin [KC16]. Dye [BL19]. dynamical [BBMP19, CNX17, CR18b, DOR16, Hoo19, Kur19]. dynamics [BTN^{+}18]. Dynkin [GSZ15, Sim18]. Dyukarev [Riv15a].

Each [AMR^{+}16, dlC15, dlCdR17, GLZ16b]. Eaton [Nie19b]. Edge [CR17b, O16, OQ18, FCL^{+}16, GSZ16, Gu16, NT18, NLW18, PS19a, Row15, SZ17, Sim18, Sim19, VSV18, Yus17, Zaj19]. edge-bipartite [GSZ16, SZ17, Sim18, Sim19, Zaj19]. Edge-connectivity [O16, Row15]. edge-degrees [PS19a]. Edge-minimal [OQ18]. edges [BL18a, GWL^{+}18, GTL^{+}18, HK19b, JZ18, JST19]. edition [Bar15]. Editor [Bru15, Bru16b, Bru17]. Editor-in-Chief [Bru15, Bru16b, Bru17]. Editorial [Ano18-27, Ano18-28, Ano18-29, Ano18z, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano18a, Ano18-30, Ano18-31, Ano18-32, Ano18-33, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano18-40, Ano18-41, Ano18-42]. Editorial [Ano18-43, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18y, Ano19a, Ano19-27, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano19-36, Ano19-37, Ano19-38, Ano19-39, Ano19-40, Ano19-41, Ano19-42, Ano19-43, Ano19-44, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z]. Editors [Ano18-44]. Editors-in-Chief [Ano18-44]. effect [BLPT15, BTN^{+}18, FJP15a, GWL^{+}18, GTL^{+}18, LMS16, LZ16b, Yus17]. effective [KLS19b, WDV17]. effects [Ruk15, Wan16b]. efficiencies [SSH15]. Efficient [RV18, Sas18, CP16b, Per18, RKT15]. Ehrhart [BJS18]. Ehrlich [RV18]. eigenanalyses [UZ19]. Eigenbasis [HPSS19, MS16c]. eigencones [PS16]. eigendata [DU16a]. eigenpair [LLV15]. eigenparameter [GM16a]. eigenproblem [BFY18]. Eigenschemes [AEKP16]. eigenspaces [BSS16, KT17, MP19, PS18b]. Eigenstructure [BCU15, BLdS16]. eigenstructures [DD17b]. Eigenvalue [AAS16, CR17a, ES16, FHJT19, Gho19, Mel19, Nag19, Row16a, SSCS16, ATM18, Ais18a, Ais18b, ABH^{+}19, ABD19, BM16a, BK17a, BPP16, BPP17, BZ16c, BBC18, BMS16a, BC17b, BWSZ15b, BJLD17, CRX15, CPR18, CLQW17, CRS15, DLB18, DHW18, Duk12, Duk15, ELN18, EH18, FuIKT16, FWB^{+}19, FI16, GM15, Gil16, GRS17, HL17b, HWXC18, HL19a, IK18, IR19, JP16, JMP17, JST19, JP18, JKS15, JS15, Kal16, KKY16, KLY18, KU16, LCL15, LM16b, LL16a, LJL16, LLL17, LDL18, Lin15a, Ma15, MHL15, MQTW18, MPS17, McK18, MNT19, NN19a, Nik16b, Obo16, PP17b, RMKJ18, See18, SdJY^{+}16, SLY17, SWB19, SAdFZ15, SW18b, Sta19, SSB15b, SSB15a, TW18a, Van16a, VT18, WZL17, WZ17a, Wei15, WZL15, XL18, Yas16, Yin19, ZG18]. Eigenvalue-free [Gho19]. Eigenvalues [BŚS18, GM15, GM16a, JTT15, YSZ17, ASMN17a, ASMN17b, Ais18b, ABG16a, AGS17, AAKS17, APS18, AB15a, AKYD16, ARSB18, ABvN19, BK15, BK16, BM18, BGS15, BGM^{+}19, BM19b, BDVRT15, BCF^{+}18, BB16b, BF15, CR17b, Che18c, Che18b, CLF18, Che19a, CLM19b, CJS18, Das15, DMT16, DMS19, EA17, ET16, EG15, lFwW16, FJT17, FdFdSDV18, FM19b, GAP16, GP18, Gho16, GWL^{+}18, GTL^{+}18, HKLQ16, HL18, HXL19, HQX15, HHL15, HH17, HK18b, IHM08, Jia15, JDS17, JLDS18, JMM16, KM15a, KPL18, Kri15, KLW16, KLW18, LOS19, LD16a, LWCS16, LDZ18, LLT19, Mat15c, Mel16a, Mon17, MST15, NN19a, Nik15, O16, OS17, PM15b, PG15, Piv19, PP18, Ref16, Row16b, SW15, Sam17, Sod18, TY15, TS18, TWR15, TW17, WLP16, WH18, Xu17b, wXjMZL19, YYMK17, YG15, YZL15, YQS16, Yua16, Zhu19]. Eigenvariety [FBH19]. eigenvector [AM16b, BMS16a, DdSJdFDV15, HW16, Ken16, Rin17, SSCS16]. eigenvector-based [AM16b]. Eigenvectors [BBGM16, DT15, FWB^{+}19, AB16, BM19b, BS18c, Chu19, DA19b, DA19a, DV19a, GHT16, LZB18, LKY16]. eight [CF18]. element [DHS18]. Elementary [CGL16, AHV19, AMZ17, Bou16a, BR17, CQ15a, JLM18, KMS17a, Lee16, MB18, Pan19a, Pas19, PR18b, SJS15]. elements [Bal17a, Bre18, CDG16, DD16b, GLAdS19, HM17a, HC19, KLW15, MY19, Rai19, Van16b]. elimination [PQY15, PZ17b]. Ellipsoids [Kim15a, KKP19]. elliptic [CEM15b, HA15b, Kim15a, Yur15]. Elliptical [AAS18]. embedding [DPS19, VB18]. embeddings [De 16a, Han18, Pan16a]. empirical [GJLS16]. enclosure [Miy15]. encountered [BDLM16]. end [Nom15a]. end-entries [Nom15a]. endomorphism [CD16d, MMR18]. endomorphisms [KOST15, SR18, Shi16g]. endpoint [LS18b, MM17a]. energies [DMG17, DGM^{+}18, DR18, LL15c]. Energy [AHJR18, Ash19, AAG^{+}19, AR16, APR^{+}17, ARS17, ALOR19, BSP18, CLL15, CLZ16a, Che18c, Dah15b, DMG16, DM16a, DG18, DMS19, DL15, GCP18, GS19a, GW17b, GRS15, HLSP17, HT17, Ito17, JTT15, LMZRZ18, LLL15, MMM19, Mró16, Nik16a, Nik16c, NA17, Obo19b, PG15, VSV18, WM17a, WZZ19, WF19, WWC18]. Engel [BRN18]. enhanced [BCF^{+}16, FOvdD16]. entangled [GW19a, Poo15]. entanglement [Car17, HK19a, JP18]. entries [Alf18a, APS18, DdSJdFDV15, GP19a, HLPS19, Hil17a, HN18, HK18b, Nom15a]. Entringer [FHS17a]. entropies [DHL^{+}17]. Entropy [HYH15, FS18b, KP17, KKS18, SCD17]. Entropy-preserving [HYH15]. entry [AG17, HZ15, VQ17, VQ19]. entry-wise [AG17]. Enumerating [OL16]. enumeration [KS19]. envelope [APT17]. envelopes [Ber18]. EP [Wan16a]. Equality [GW18, GWW19b, GWW19a, BT17a, LL15b, SR19]. equals [MWT16a]. equation [CLM19a, CGM17, DI16, IT18, LMS15, LM19, SS18, WYL15, Wój19, WP17, DP15b]. equations [Ais18a, AHV19, BB16a, BDDSC17, BHP17, Bou15, CN19a, CFNP17, DIPR18a, DIPR18b, DP16, DPJ17, DFKS17, Dra18, EJ19, EM17, FI19a, FL15, FKS16, Gal17, Gua18, GL16, HK18a, HAWM16, HS19b, JKP16, KU16, MEMM17, Mys16, PMB15, RW16, SR18, SNP16, Shi16b, UZ19, WT18, Wim16, Zue18, dS18a]. equi [ET18]. equi-isoclinic [ET18]. Equiangular [CGSZ16, FJKM18, Gre18, BFG^{+}16, BF17b, RNC17]. Equidistant [GR16, Pol19]. equilibrium [SSF16]. equimodular [CK19]. Equitable [BFW17, JZZ18, AC15, FSSW17, FSW19, YYSX19]. Equivalence [Bar17b, BBdH16, BBdH18, BS16f, BS18c, CJS16, Chu15, Had17, Jin15, Moh16, NSCV16, dlCF16]. equivalent [BR16a, NS15a]. equivalently [dCdSA15]. erasure [FJMP15, LS15c]. erasure-robust [FJMP15]. erasures [CLS18]. ERES [CKM16]. Ergodic [BBMP19]. Ergodicity [Sab19, AB19, RH15]. Erratum [KAPS20]. Error [MM16a, WZL17, CW17, CJ16c, CJ16d, Ema18, Fis17, KL19a, Ma15, WLL19]. Essential [NSCV16, SS15c]. essentially [ÁKM17, SW16]. essentially-invariant [SW16]. Estimability [BFFN16]. estimate [Fre18, Nun17]. Estimates [Rum18, RB19, Duk12, Duk15, Hua19, KRZ^{+}17, OZ18, wXjMZL19]. Estimating [PSZ15, Wol18, FL15]. Estimation [CS15a, FM16, LHC15, SH15b, BBP19, BSS17, DR16, DLM15, Ema18, Fis17, KRZ^{+}18, LLH17, PCL15, WYL15]. estimator [LLH17, iT19]. estimators [RBKA16]. Estrada [APR^{+}17, CR19b, HLW15, HW15c, LZ15b, Rod19, ZW16]. etc [Bar15, JS18b]. ETFs [RGPH16]. Euclidean [Alf18a, BD15a, BK19, ET18, Fie15a, Gow17, GJ18, JG16, JG17, Jia15, Kim15a, KSTX15, KT15, KB15, Kwo16, PR18a, SSS15, TZK17, TW16]. Euler [Ern18, He16a]. evaluated [AEV15, CW16a, MO16]. evaluating [Fas19]. evaluation [Sas18]. even [DS19b, Fil18, FKM17a, FKM17b, uIKF15, NY15, WZZ19]. eventually [SA17]. Every [dlCMP17a, TXZ18]. Evolution [CMV16, CMV17, CKM19, DOR16, EL16, HA15a, HPSS19]. exact [SSS15]. exactly [HKLQ16, HHL15, MM17a]. Examples [Gol16, MM17c]. exceed [MHL15]. exceptional [Zus17b]. exchangeable [KRZ^{+}17]. excluded [BS16e]. exist [GLZ16b]. Existence [Mon17, BPR17, BÓ15, DS17b, Fil18, FJ19, dMGC19, HHR16, NS15a, Sam17, ZLG15, ZMT19]. Exotic [KOST15]. expander [Lev15]. expansions [BCKL17, LG18, dS19a]. expected [DHLT19]. Explicit [Ban16, BG15, EJ18, Guo19, SR18, DDL17, DSCD17, FKRS17, Nat19, Pud16]. exponent [BS16b, HYY15, Iop18, OQ18]. Exponential [AMH11, HM17b, Tan19, AGHN18, BDLM16, BMO15, CJB18, Dra18, Fis17, FL15, Obe16, PPZ19]. exponents [ANAPSR17]. exposed [Zha18a]. expressed [DMQ19]. Expressing [HLZ17, Hou20, Ste18]. Expressions [CGMSR15]. Ext [CG19b]. extendability [BGdCCMSS17]. extendable [CRVSS19]. Extended [Nie18, AG16, BW19, CEM19, LHC15, Nom16, Sch16, WHG17]. Extending [JZC19, PP15, Shi16d]. extensible [dCdSA15]. Extension [Zha15b, ZH17, Ali20, CS17, DP15a, HAD19, LS18a, Nie19b, PQY15, Ric19, Row18, VGG17]. Extensions [CTZ19, FSSW17, SST15, BBM16, EF17, EF19, KVP19, KSVW17, MM17b, NT19]. exterior [Mag16]. extinction [GCQX15]. Extracting [You15]. extraordinary [FM19b]. Extrema [Nik15]. Extremal [AdlPR16, Kri15, RMS16, SUY16, BZ16c, BM19b, BDK^{+}17a, GP19a, Hut17, LD16a, WZ15]. Extremality [GW16a, GW17a, GL19, Che19b]. Extremals [Ser15]. Extreme [HLM17, NAS18, Eve18, FL16, PP18]. Extremizing [Kal16].

facets [AK17]. facially [Zha18a]. factor [LZ17c, WFS19, Yor17]. factoring [FR18]. factorisations [ABP15]. Factorization [CP17, Kur19, Mar16, AAG^{+}18, Bot16b, Bur15, Chu16, DEH16, LMS15, LWY16, LSM16, LSM18, MV16, SP16, ZSQZ19, ZZC15]. Factorizations [BJ16, Lev18, SG17a, AAC17, Hua15, LW15, SSV17, TVD15]. factors [CQ15a, DS15, FR16, Ram17]. Failed [AJPS16]. Failure [DPS19]. false [Shc16]. families [AN17, ABGJR^{+}18, BmBCC17, BGCMASS19, CLHL15, CM19, DGGP18, DR19, FJ19, FL15, GMO19, Hru16, Kol15, LD17c]. family [APR^{+}17, BCC^{+}18, BS16e, BH16b, CPS17, COS18, CMPZ18, ET18, GSSvdD16, PRS19, SSH15, UWYY15]. Fan [HRM18, Pet15, SWB19, TY15]. Faria [ACPR15]. Farkas’ [CS15b, Alf15, CJ17c, DP15a, LX17]. Farkas-type [LX17]. farthest [GMLT19]. Fast [LL16b, Miy15, BP19a, JKP16, KP19]. fat [KLY18]. FEAST [Yin19]. Feedback [DS19c, BR19, Bel16, NN19b]. Fejér [Chu16]. Ferrers [Bal15c, LCF19]. few [JZ18, OQY17]. fewer [BF15]. Fibonacci [HHJM15]. fidelity [LPP15]. Fiedler [AB16, ARSB18, BDP^{+}18b, DA19b, DA19c, GP19a, LZ17b, PD17, RRS16]. Fiedler-like [AB16, BDP^{+}18b, DA19c]. Field [HK19c, BZ16a, Bal17b, Bal18, Bie15, Cox15, FKS16, JSST16, JLDS18, LS15b, MM17b, Pel17, Ram16, Sha16, Sin17, UG18, Van16b]. fields [Bal19, BLdS16, BP19b, BC17a, Bre18, CM18, CFW15, CJK^{+}16, HA15a, JKKL15, LS15b, Lie18, NT19, PW19, PW20, RFPS18, Shi18a, SH15b, WG15]. fifteen [HAWM16]. filiform [CLOK13, EF19, LMO16]. filling [BGCMPB15]. Fillmore [Bor17]. filter [CMV19, Van16a]. filtering [DLM15]. filtration [FH19]. finding [RV18]. finitary [CDH18]. Finite [BGOY15, LZ15e, MS19, NINS16, Yan18, AMZ17, AKR15, AK16a, AK16b, Bal17b, Bal18, Bal19, BLPT15, Bre18, CM18, CDFK19, CW15a, Che15b, CJK^{+}16, Cos19a, DHLT19, Far16, Gol17, Gór17, HA15b, IM19, JT18b, KKLP17, KT19, LS15b, LT16b, LH17, Lie18, LZ15d, Man19, MS17a, Mig16, OW17, Pan19b, PMB15, PS19e, PW19, PW20, QWW18, Ram16, RFPS18, RE16, Row15, Sab19, SR18, Sha16, UZ19, UG18, VB18, Vir16, WG15, Wol18]. Finite-dimensional [LZ15e, Yan18, AKR15, AK16a, AK16b, CDFK19, CW15a, Gol17, LT16b, OW17, RE16]. finite-time [PMB15, QWW18]. finite-valued [KKLP17]. Finiteness [Egg15, Koz16, NW19]. Finsler [Cim15]. firing [BMZB18]. First [PM15b, BK17c, Chu16, GHPS19, Hun18a, Kal16, KT17, LDZ18, Lom17, OQY17, XWD18]. Fischer [CTZ19]. fitting [HPZ19, UM16]. five [HA16]. five-dimensional [HA16]. fixed [CMQV16, DD18, FG17, JKS15, JKP16, KLL16, KKY16, NLW18, PS15, RR18a, SJD16, You15]. fixed-point [FG17]. fixing [KLW15, KLW16, KLW18]. flag [BF19b]. flags [Lax16]. Flanders [dSP16b]. Flat [KY15, BBM16, JSST16, SS17a, WH19a]. flattenings [Hu15]. flow [DvSW18, Li15b, MGM17, Per18]. flows [Ban16, HL17a, VG18]. fluctuations [MNT19]. fluid [BDLM16]. forbidden [FdFdSDV18, Yua14, Yua15, NY15]. forcing [AJPS16, FMY16, KL19a, Lin16a, Lin19, TD15]. forest [JMPS18]. form [AEKP16, BBdH18, BLdS16, BTW16, BGSS18, BD19, CS18b, Cus17, DEH16, DFK^{+}15, EG15, FM16, GLAdS19, Gir19, GBRS15, IW15, IGW17, Joe16, Li15a, MN15, Miz16, Nom15b, PRS18, Rad17, SZ23, Wój15]. formal [CJS16, GPI18]. formalism [MdlP16]. formats [YLB15]. forms [AAdFS19, BS18b, CR19a, CG19b, CXLF16, CNX17, CN17a, CFP18, CQS18, De 16b, GKR16, GLW17, HSTW17, IKR16, LS15b, MN19b, Ota15, Tok17, dFRS16, dFFRS17]. formula [AM15b, ACM15, ALOR19, BLdS16, BBM16, CDMP15, CUPW15, DKSV19, Hil18, LMPM17]. formulae [Ban16, OBRA15, XW16]. formulas [CM17, CJB18]. formulation [HPZ18, HPZ19]. Forward [SSB15a, MP18, Per18, BCMASS15]. forward-backward [MP18]. found [BH16b]. foundations [Gar17]. four [Ben18, BP19b, HH17, Pel17, TWR15, dlC15]. fourfolds [Jel18]. Fourier [BLdS16, KP19]. Fox [KL18a]. Fractional [Mat15a, SKC18, RW16, ZH17]. Frame [CC17a, CDFK19, LH17, BCKL17, LC19]. framelets [AKPS19, KAPS20]. Frames [CLS18, AW15a, BGOY15, BFG^{+}16, BF17b, CH19, CFW15, CW17, DV18, FJMP15, FJKM18, HJ19, HL17c, LS15c, MRS15, MBB^{+}18, NAS18, RNC17, RGPH16, WO17, Zas16]. framework [FM17, JLT17, OZ15b]. frameworks [Alf15, KS15b]. Frechet [Eld15]. Free [BdlCL17, HNS15, KRZ^{+}18, Tan16a, ABM16, AL19b, CLR16, CD16d, GSZ16, Gho19, IKR16, JMRV16, Row16a, Shi15, Shi16g, ySpW17, SZ17, Zaj19]. Free-coordinate [KRZ^{+}18]. Friedland [Tsa16b]. Friedrichs [LMX19]. Frobenius [CLZ^{+}16b, FK16, GH16a, LT19b, Mon19, Tar18]. Frobenius-norm [CLZ^{+}16b]. Fucík [LN18]. Fueter [ACS19]. full [SZW16]. function [AB15b, AMZ16, AGHN18, BS18a, BGCMASS19, CFL16, CPW17, DS18b, DSX15, Hun18b, JCW19, Kli19, KMS16, KMMS19, Som17, TH16a]. Functional [Ere16, Wan15b, Nev18, Wój19]. functionals [Gir19, dSP19a, Pop16b, SST15, WH19b]. Functions [MN19b, AB16, AL18, AL19a, ACS19, BKNS17, Bar17b, BBSS18, BGCMLSSS16, BEKS17, Bil19, BS19c, BU18, CPZ16a, ÇMP18, CR18b, CW18b, CRVSS19, CLP16, DR16, DDF18, DM15, FI19a, Fas19, Fay18, FNX19, FKM17b, FKRS17, GS16, HM17a, He16a, HNS15, Hia16a, Hil15, Hoa15, Hon18, IHM08, JMRV16, JG17, JS18a, JLM18, KM16, LT19a, MY16, MR17b, Mon19, NW15, Naj17, Nie19b, Pas18, RRT16, RV18, SR19, Sch16, Shi18a, dSN15, Skr17, SZ19, SS15b, UUG15, Van16a, Vol17, Yam15]. fundamental [AHV19, GZB19, lHS16, ZD15]. Further [AAG^{+}18, BTN^{+}18, DKPU18, LW15, YHY15, ZM16]. Fuss [HS17]. fuzzy [MP15b].

Gabidulin [TZ19]. Gabor [DV18]. Gaddum [EA17, HL19a, LOS19]. gain [LWX17, Ref16]. Gale [RT16]. Galois [GL15b, GL15a]. game [GR15]. game-theoretic [GR15]. games [GC19, Gow16, XvdBvdLS15]. gap [BGT^{+}19, CCW16, Car17, Wol18, Zui17]. gaps [FCLP18]. Gau [CDR^{+}19, Lee15b]. Gauss [BV18a]. Gaussian [FL15, FI16, Hoo19, PCL15, PQY15, PZ17b, SH15b, Zim15]. gcd [AYK17, AKYD16, AYK17, GW19b, HT18a]. gcd-closed [AYK17]. Gelfand [FH19]. General [DSCD17, FSW19, ABP15, BCM17, Bar17a, BBM16, CDR18, DL16, EH18, Hla15, KSMB15, KM18, Li15a, LX17, Miz16, RH15, SK17a, SZB19, Uch15, wXL14, wXZ19, ZLKB17]. Generalization [DFKS17, Roo19, AOK19, BEF16, Hia16b, HP15, KPRvdO16, Rao18, TW16, TG19, Vec16]. Generalizations [Liu16a, HS19a]. Generalized [BS18c, FT16, IK18, MT15, Zem16, ZZC15, AMPT16, AHV19, Bai18, BLdS16, Bar16a, BBS16b, CEM15a, CCW18, CK15a, CJ17c, CHT19, DA19b, DI16, DIPR18a, DIPR18b, Dra16a, DHW18, DGP^{+}15, DM15, FuIKT16, FGS^{+}16, FM19a, FX19, GSSvdD16, GH16a, HAWM16, HPZ19, Hua19, IR19, JT18a, JZZ18, KP17, Kim15b, Koo19, KSTY18, KMM17, LJK16, LNT16b, LLPS19, LZ15e, Ma15, MW15, MLW15, MS17a, MS15b, NN19a, Nie19b, Pep17, PF18, RD15, SMH19, SR18, SR19, SB15, TZ19, WM17b, XSW19b, XSW17, wXjMZL19, Yin19, YMZ19, ZH15, dCdSA15, BS16f]. Generalizing [SP16]. generated [FM19a, Jia15, JLM18, Kan16, KS17, KLW15, MS19, Pas19]. generates [RS18]. Generating [Mon19, SÖ16, AGHN18, BN17, CN15, CN19b, Vij15]. generation [RAAGA16]. generator [LMS16]. Generators [CS18b, Chi18, FH19]. Generic [DD17b, DD18, Mor17, SS15a, BC15a, Bou15]. genericity [BGdCCMSS17]. geodesic [BHML15, STZ15]. geodesic-affine [STZ15]. geodesics [WO17]. Geometric [BHML15, BC16b, HL17a, JS18b, KLL19, RH15, Tsa16a, DGSV17, EHN16, FS17b, FS19b, GT17, IKW16, KLL18, LPP15, MHA18, Nob16, NK18, Pep17, Zou17, Zou19]. geometrical [Fli16]. geometries [ABP15, CD16a, Hor17, IM19]. Geometry [ÁVAK^{+}19, BPZ17, CRC17, Che15a, DvSW18, Fie15a, lHS16, Li15b, LM19, MHS15, Sar19]. geometry/trigonometry [CRC17]. gerechte [BL16]. Gersgorin [HW16]. Gershgorin [DeV19, Rum19]. Gerstenhaber [HO15, dSP19b, dSP19c]. Ghorbani [Riz19]. ghost [ME15]. Gibbs [RH15]. Gilmour [KÖ16]. girth [Fil18, LLT19]. give [Gal17]. given [BR16a, CJ17d, De 16b, DV19b, FdC15, HLW15, HSS16a, IKVZ18, JL16a, Kal16, KT15, LdlP15, LLT19, MS16a, Mon17, Pas19, Ram17, SÖ16, SUY16, TH17, WHG17, WZL15, XSW19a, XWL17, XWD18]. glN [FH19]. Global [LM19, CJ16d, Est16, KH17, NT19]. globally [JCW19]. GLT [Bar17b]. GM [WQH19]. GM-switching [WQH19]. GMRES [DHW18]. Gohberg [XW16]. gons [VGG17]. Gorenstein [Jel18]. Grad [Zha15a]. grade [DD18]. Graded [BY19, BFD18, BZ16a, BKRE18, CC16, CLOK13, CR15, LMO16, RZ17]. graded-division [BKRE18]. gradient [KR17]. grading [Zue18]. Gradings [CME18, KY17, RZ17, RE16, Zus17a]. graft [LZ16b]. grafting [HLSP17]. Gram [CA19, GSZ16, Sei18, Sim19, Yas16]. Graph [BK16, dSN15, YYMK17, ACT16, ARS16, AAMM18, AL15, AJPS16, AvdHS15, ALOR19, AM16b, AM18, AB18, BDMB19, Ben16, BPR17, BF15, CLT15, CD16b, Che18b, CLF18, CD16c, CGSZ16, COvdD16, DHL^{+}17, Dal19, DM16a, DMG17, DR18, DT15, DHS18, EA17, FW17, GAP16, GP18, GCP18, GK15, Gho16, GRV16, HHL19, HFS15, HZ17, JT17, JLDS18, Ken16, Kob17, KMS16, KIS18, KMMS19, KLW15, KLW16, KLW18, KMM17, LOS19, LLL15, LL18b, LWZ18, LHL18b, MWT16b, Mes17, Mig16, MMS18, MM15b, MS16c, Mon17, NS15c, Nik15, Nik16a, OS17, OP16b, Per18, PG15, Piv19, SW15, SAdFZ15, STW18, STW19, TNK16, TWR15, TW18a, VSV18, Wan15a, WM17a, WDV17, WH18, WYL19, WF19, WWC18, XLLS18, ZG18, Zho17, ZWS18, vDS16]. Graphical [SFW18, FZL^{+}17]. Graphs [Alf18b, BFM19, CRS15, LS15a, MQTW18, TW17, dFNP16, AAB^{+}16, AAG^{+}19, AJO15, Abi19, AG15a, AyLPS18, ASMN17a, ASMN17b, AHMM18, ABH^{+}19, AAS16, AAKS17, AT15, AAC^{+}15, ACRR15, AGRR16, APR^{+}17, ARS17, ALMZ^{+}19, AB17, Ash19, AH16, AP15b, ABK19, BBM19, BK15, Bah19, BM17b, BK17a, BPP16, BPP17, BK18, BFW17, BS15c, BP15, BSS16, BZ16c, BBC18, BCC^{+}18, BSMSZ19, BSP18, BK17c, BF15, BH16b, CR17a, CR17b, CDR18, CPR18, CCdFV18, CR19b, CLL15, CLZ16a, CL17a, Che19a, CLM19b, CFH19, CJKM19b, CJKM19a, CV18, CW16b, CGGV15, CJ17d, Dah15b, Das15, DMT16, DMG16, DS17a, DGM^{+}18, DG18, Das18, DHHZI19, DMS19, DLG16, DPR19, DBK17, DL16, DL15, DM15, ET16, EB17, EHHL17, EHLP18, ETF18, lFwW16, FCL^{+}16, FJT17, Fio16, FKD15, FSSW17, FSW19, FHJT19, GGH15, GH19, GSZ16]. graphs [GK15, Gho19, GHS16, GLW17, Gre18, GJLS16, GRS17, GWL^{+}18, GTL^{+}18, GRS15, HKLQ16, HLSP17, HS19a, Her16, Hey17, HXL19, HMT19, HLLZ17, HK19b, HHL15, HLW15, HSS16a, HSTW17, HH17, IM19, JTT15, JT18a, JZ18, JKP^{+}17, Jør15, JMM16, JKS15, KM15b, KKY16, KM19, Kol15, Koo19, KLY18, KSTY19, LJK16, LMT19, LMZRZ18, LZ15b, LM16b, LN17, LS19, LZ15c, LwW16, LD16a, LWCS16, LDZ18, LHX18, LXS18, LL15c, LHG15, LZ15d, aLwW15, LSX15, Liu17, LLS17a, LLS17b, LZSD17, LLZ17, LL18a, LZSD18, LLT19, LWX17, LHH17, LHL18a, MHL15, MWT16a, Ma16, MG18a, Mac16, MMP16, MM17a, Man19, MW15, MLW15, Mir18, Moh16, Mol18a, MM16b, MK18, Mró16, MST15, Nik16a, Nik16c, Nik17c, NR18, NLW18, NLWJ18, OQ18, Obo16, Obo19a, Obo19b, Obo19c, OL16, PP17a, PP17b, Pan16a, PS19a, PS19b, PS17, PF18, PS15]. graphs [QKP16, RFPS18, RFPS19, Ref16, RS17a, RB16, RR18b, RMS16, RJK18, Row15, Row16a, Row16b, Row18, San15, San18, SS17a, See18, SMBDK16, SM16, SM18, SY15, SSF16, SS16, SZ17, Sim18, Sim19, SUY16, zSqST15, Sta19, SV17, TH17, TFC19, TWM16, Tom15, TSH16, Urs18, VK19, WYLW17, WM17b, WW18, WWT18, WQH19, WLWZ19, WZW19, WYL19, WZZ19, WZL15, WZ15, XS18, YCL17, YZLC18, Zaj19, ZLG15, ZW16, ZLG17, ZLS17, ZXW19, ZGW16, dS18a, dlPJG18]. Grassmann [BA15, CG16, CG19a, CLLS19, GGH15, Jen17, Pan16a, SY15]. Grassmannians [BHML15, Pan16b, Pan17]. Green [CEGM16, CEM15a, GJK18]. Grid [CDRT17, KN16, OL16]. grids [Sar19]. ground [LM19]. Group [FJMP15, KY17, AMC18, ABP15, Bie15, BS19a, BJ15, CS19a, CFH19, CH19, CJ17b, CS18b, DHS18, Gae16, GH15, He16b, HS17, KLW15, KMM17, LH17, LMM^{+}16, MBB^{+}18, MY19, PB16, Pop16a, RB16, Sri17, Van16b, Wei17, WLH18, XW16, ZMT19]. group-subgroup [RB16]. Group-theoretic [FJMP15]. groups [AJ19, BD15a, BRN18, BMZB18, CLUP17, Che15b, CHK17, CW17, DF15, DF18, DH19, Far16, GG18, Gae16, GP19b, KMOR19, LMPM17, MS19, MOR16a, NT19, Oi19, Saw16, Sin17, TZ19, Zem16]. Grover [KSTY18, LT18]. Growth [AB19, CR15, bCR18, LHC15, RZ17]. Grundy [Lin19]. GTS [NS19]. guarantee [Lin16a]. guarantees [ZC15]. Guo [AB15b, AMNR18, Rob19]. gyroscopic [CL15c].

H [CS19b]. H-matrices [CS19b]. Haagerup [Ali20, HAD19]. Hadamard [FJS21, AMV17, BS15a, BGKP19, CM18, DHW16, FJS17, GW16b, GX17, HL17b, HW15b, Jai17, JKP^{+}17, Kar16, LK16, NS17, NW19, Pep17, RS17a, Roo19, RWH17, TG19, Zha18c]. Hadamard-type [HW15b]. Hadamards [MB15]. Hajós [Jen17]. Hajós-type [Jen17]. Half [EV15, YDYY18, AM19, CD16a, HT18b]. half-radial [HT18b]. half-spin [CD16a]. halfspaces [SW16]. Halley [LG18]. Halmos [LMX19]. halves [Bar19]. Hamburger [CJ16b, HZC15]. Hamilton [BBS16b, LN17]. Hamiltonian [LSX15, LS18b, SS15c, ZG18, Zhu19]. Hamiltonicity [Ben16]. hand [HPZ18]. Hankel [Al’17, BS18a, CJ16b, CI15a, CI15b, DS19b, FM19a, HKM18, Xu16, Yur15]. Hans [BMS16b]. hard [BC17b, FS19a, RSS17]. HARDI [CS15a]. Hardy [AP15a, ANAPSR17, HJW19, Nun17]. harmonic [Vec16, Yam15, Yin19]. Harpe [Ali20, HAD19]. Haukkanen [AKYD16]. having [BK17a, BBH16, ET16, HLG15, IKVZ18, Nom15a, Nom15b]. Hecke [Rai19]. Hedetniemi [KM19]. Heisenberg [LZ15e, Pop16a]. Helly [BJC15]. hemispaces [EHN16]. Herglotz [ACS19]. Hermite [BDM18, ACDT^{+}15, Joe16, PP18, Reb19]. Hermitian [ACM17, BŚS18, BBGM16, BBFF17, CG19a, CRC17, CLZ16a, CS18b, CQS18, FM16, GXYZ17, GMO19, Lau18, LL15c, MNTX16, Moh16, Ore16a, Ore16b, Sch16, SWZ18, TC15, Vec16, Yin19]. Hermitian-adjacency [LL15c]. Hermitianity [HK19c]. Hermiticity [BdP18a]. Heron [Hoa17]. Hessenberg [BR16b, DHW18, Mar16]. Hessenberg-upper [DHW18]. Hessian [CPW17]. heterogeneous [Ruk15, Zha15c]. Hiai [Wad18]. hidden [DN16, LL16b]. Hierarchical [DH15]. hierarchy [Ong18]. high [CS15a, GdLL17, KU16, Lom17, MH15, SSH15, TW18b, WZ17a]. high-dimensional [KU16]. Higham [Fis17]. Higher [BBMFM17, AAS18, AMC18, BGV17, BBT15, CLUP17, CL16, Fre18, GH16b, GS15, HMNP15, KKP19, LZ15a, SL16]. higher-dimensional [CL16]. higher-order [GH16b, LZ15a, SL16]. Highly [CLM19a]. Higman [LMP^{+}17]. Hilbert [AK19, BF18a, BCKL17, BSZ18, Bil19, CR19a, CGL16, DY15, GKR16, Jel18, LMX19, MS17a, MEMM17, OZ15b, Pan16b, PCI15, Pop16a, Vir16, XF17]. Hilbertian [BFA18, FG16, Zam19]. Hille [Bob16, CJRMF^{+}15]. Hille-type [Bob16]. Hilton [GLW17]. Hlawka [BS15d]. Hoffman [KLY18, KSTY19]. hold [DV18, Wad18]. Hölder [Kal18, Zou19]. Holomorphic [BGCMLSSS16, Bal17a, BGdCCMSS17, Bil19, FKM17b, Nev18]. homogeneous [BB16b, IL17b, PC15, Vas15, Zue18]. homology [BBMFM17, CCW18]. homomorphism [Jen17]. homomorphisms [GRV16, OP16b, dSP15, dSP16e]. homotopies [BV18a]. homotopy [KL18b]. Hook [NS17, Sri17]. Hopf [BBdH16, BBdH18, MG18b]. Horizontal [CJB18]. Horn [Bar15, CW18a, DP15a]. Hou [Bah19]. Hourglass [Koz16]. Householder [RMP18]. HSS [Bai18]. Hua [lHS16, LZH16, Lin16b]. Hua-type [LZH16]. hulls [AMR19, KS15a]. Hurwitz [BS18a, Dya17, PM15a, Riv15b]. Hurwitz-type [BS18a]. Hyper [BT15]. Hyper-ideals [BT15]. hyperbolic [GT16, LP18, MHS15]. hyperbolical [SÖ16]. hyperboloid [SÖ16]. hyperboloids [Kim15a]. hypercube [Mir18]. Hypercubes [KHI16a, KHI16b]. hyperdeterminant [YLB15]. hypergraph [HQX15, Pea15]. hypergraphic [CRRY15, CLR^{+}18]. Hypergraphs [Nik17b, BL18a, BCM17, CLN15, Duk12, Duk15, DR19, FuIKT16, FWB^{+}19, FBH19, JZZ18, KNY15, uIKF15, LZB18, LZ16a, LZM16, LZW17, LZ17a, LKY16, LM16c, OQY17, RS17a, SY17, SZB19, XLQC16, YYQ19, YZL15, YQS16, ZLKB17]. Hyperinvariant [AW15b, AW15a, FMM16]. hypermatrices [Nik17a, Nik17b]. hypermatrix [GF17, Pea15]. hyperplane [FGG^{+}18]. hyperplanes [Pan15b]. hyperpower [SSH15]. hypersurfaces [UM16]. hypertrees [GZ18, ZKSB17]. Hyponormality [Lee19a].

ideal [HR17, Pop16b, Ris16]. ideals [AKS17a, AKS17b, BT15, CKM19, HR17, JLM18, KVV17, Ris16]. idempotence [LP15]. idempotent [Bot16a, GT17, LP15, Ste18, TXZ18]. idempotents [AJL16, BES16]. Identifiability [Wan16b]. Identification [LPPZ19, Hoo19, NAS18]. identifications [SBB15]. Identifiers [GZ19]. identifying [BDMB19]. Identities [UG18, AGN18, BY19, BKNS18, BLdSV19, BV18b, CR15, Ere16, HHJM15, LMR16, TZK17, Wan15b]. identity [AMPT16, BaHOS15, LMS16, LLSX15, Rum18, RB19, Zha18b, dCdSA15, vWZ17]. ifications [BR16a, DPV19, DDV16]. Ihara [KMMS19, Som17]. II [dSP19c, AKS17a, BS16c, CJKM19b, CSJ16, DV19b, EHLP18, FKM17b, GH15, Hia16a, KLY18, Ore16b, PT19, She17]. III [AKS17b]. ILAS [CFH^{+}16, FLVV18, GPW15, GHPS19]. ill [BB18, NSCV16]. ill-posed [BB18]. Ilmonen [AKYD16]. image [DDSC^{+}16, PS16, Win16]. images [AEV15, Fag19, LNT16a, Lau18, LT16b, MO16]. imaging [CS15a]. immanantal [NS17, NS19]. immanants [Tab16]. Implementing [CG19a]. Implications [BdP18a]. implicit [CMV19, SNP16]. imprecise [Sku17]. imprecision [Sku17]. imprimitive [MP16c]. Imprimitivity [AGH17]. improve [AM16b]. Improved [Che18b, DM18, XL18, LLH17, Xu17b]. Improving [OVW18]. incentive [LT19a]. incidence [AB18, CR18b, CDH18, IHM08, Ma16, MMM19, WZZ19, ZK17]. including [Ern18, JLD17]. inclusion [BWSZ15b, BJLD17, DLB18, ELN18, LCL15, LL16a, XF17, YZL^{+}19]. incomplete [LSM18, Mic16, iT16, iT19]. inconsistent [BW19]. increasing [CR19b, Gar17]. Indecomposable [CPS17]. indefinite [MV16, MSSZ17, ME15, NN19a, SG17a, VK16]. independence [ACT16, JL16a, LL18a, ZLG17]. independent [bCR18, MC19, OL16, Vij15]. indeterminates [HZ15]. index [AGH17, AMNR18, BOT19, CR19b, DG18, DS19b, FW17, JDS17, Kim15b, LZ15b, LS19, LLZ17, LL18a, MMS18, NY15, NR18, PS15, Rob19, SPBB19, SG16, Yua14, Yua15, ZW16, dFNP16, Rod19]. indexes [WWT18]. indices [APR^{+}17, DA19c, FGS^{+}16, HLW15, HW15c, Pal16, Som17, SS17b]. induced [KSTY19, Mar18]. inductive [MHA18]. Inequalities [BM19b, BJL19, CS19a, Cho17a, Cho17b, HvdD16, AOK15, AO18, AKS17a, AKS17b, AN17, AK19, ANAPSR17, Bar17a, BS15d, BLY16, BBP19, CPZ16a, CM16, DKS15, DS15, GW18, GHK17, GLRT18, HRM18, Hoa17, HWXC18, JM16, JLM18, Kal18, KMY15, KL17, KMS17a, Lin16b, Lék16, MN15, MRS15, MF16, NS17, Nie15a, Nie15b, Reb19, Ric19, Sab15, Sab18, SMH19, SDK17, She17, SWB19, TNK16, Uch15, Ulu19, WDFS17, XCJ18, Zam19, Zha18c, Zha19a, Zou19]. inequality [Ali20, ACPR15, AP15a, BNST17, CJRMF^{+}15, CPZ16b, CTZ19, CS17, CW15b, CJ16c, CJ16d, FS19b, GXYZ17, HAD19, Hay19, HL17b, IKW16, JZC19, LS18a, Lin15c, Liu16a, Nun17, RRV15, Roo19, RWH17, Wad18, Zha19b, Zou17]. Inertia [HT18a, NN19a, AvdHS15, AvdHS16, BS16a, FW17, LS19, LLL15, LOvdD18, WWT18]. inertias [BOvdD17, COvdD16]. inexact [BM17a]. Infimum [WLL19]. Infinite [BGdCCMSS17, DE15, ET18, MS17a, Min17, ABO15, AMZ17, Bie15, BS16e, CG19c, CJ17c, Dya17, Egg15, EGT15, HLZ17, Hou19, Hou20, KL19b, LD17b, LMR16, MK18, NT19, Pan17, PMW19, PP15, PPZ19, SR18]. infinite-by-infinite [Egg15]. Infinite-dimensional [Min17, CG19c, Pan17]. Infinitely [VB18]. infinity [GS19b, LCWZ19]. Inflation [SZ17]. inflations [MM19a]. influence [BK17c]. information [DKS17, Mel19]. initial [MGM17]. injectivity [FMR19]. inner [AGQ19, AL19b, BEKS17, FZwCW16, FS19b, Mag16, MSSZ17, PP18, SG17a]. inputs [LPPZ19]. insight [CQ15b]. Instances [DK16]. instructional [Dog18]. integer [Bor17, Mac16]. integers [San15]. Integral [BK18, CFH19, San15, ALOR19, Bal17a, Cal16, CDMP15, CS15a, CX17, CW16b, DN18, FdC15, HK19a, JKP16, Mir18, Ota15, PS19a, San18, Wim16]. integrality [DN18]. Integrally [MS16a, PS17]. integration [BPT15, Van16a]. Integrators [AMH11]. intercyclic [VV18]. interlaced [MM15b]. Interlacing [GMO19, DS15, Tya17]. International [GHPS19]. interplay [Pal16]. interpolants [ACDT^{+}15, EG19]. Interpolated [Sab15]. Interpolating [Zha19b, RV18]. Interpolation [BDM18, AL18, HZC15, MM17c, Ste15]. interpolational [UYY17]. interpretation [FKD15]. interpretations [LPP15]. interrelations [Riv15b]. intersection [KKL^{+}17, PRS19, SXD16]. intersections [RV18]. intertwiners [Ter17a]. intertwining [DSCD17, MMS16]. Interval [Mys16, RS18, WT18, Gho19, Hla15, HS19b, Li15a, LX17, Li19, MP15b, MP19, PM15a, PS18b, Rub19, wXjMZL19]. intervals [XL18]. intriguing [BM19a]. Introduction [Bru16a]. Invariance [AG17, BJ15]. Invariant [CQ15a, ACRR15, AG15b, AG20, BBEE19, BB16a, CKL17, CFP18, CJW19, DG18, DH19, DS15, JLM18, KT17, KLL18, KMS17a, Lee19b, LP18, LNW17, MRS18, MOR16a, NINS16, NP15c, Ouy17, Ram17, Sab15, SK17a, SW16, WL15, Zha19b, Zha15c, Zou19]. invariants [DS19c, GRV16, KLP18, Obe16]. Inverse [BS16a, CS19b, CIY15, CNY18, ES16, HJ16, HC15, HFS15, Wei15, ATM18, Ais18a, Ais18b, AB18, BW16, BPP17, BFM19, Bik19, BC17b, BC16b, CJ17a, CGMSR15, CGDM16, CL15c, CCO15, DHW18, EJ18, GMLdS16, Hoo19, JT18a, JMRV16, JP16, JMP17, JP18, JS15, KB15, LDL18, LCWZ19, LL16c, LMM^{+}16, Ma15, MN15, MPS17, Miy15, MK18, NMBA19, Nor18, QWW18, Rak17, SLY17, SW18b, Wei17, XSW17, YWD16, ZC15, Zho17]. inverse-free [JMRV16]. Inverses [LL19, PP17a, WZ17b, BJKR17, BKNS18, BBH16, CFM16, DM18, Dra16a, ELN18, HWZ15, KC16, Niv15, OP16a, OP19b, PSZ15, PCI15, PHW16, RD15, SR18, Shi17a, SSH15, SZW16, XW16, ZMT19, ZZC15]. inversion [EGT15, KP19]. Invertibility [PPZ19, DS19b, YYMK17]. invertible [CS19a, CKT19, KKA15, LY17, Ore16a, Ore16b, PS17]. investigation [GO18]. involution [BLdSV19, UG18, ZZC15]. involutions [AR18, BKRE18, GT16, HLZ17, Hou19, Hou20, LMPM17, dlC15]. involutive [FKS16]. involutory [Kis15, XX12]. involving [CGMSR15, Das18, FI19a, GHK17, HL17b, KMS17a, Lin15a, SS17b, UWYY15]. Iowa [GHPS19]. irrationals [OZ15a]. irreducibility [GJ18]. irreducible [CUPW15, DDL17, DLB18, FBH19, KLS19a, Kim15b, LLV15, LZ16c, MM17a, MP16c, WFS19, Yan18]. irregular [NLWJ18]. ISBN [Bar15, Bru16a, Tsa16b, Zha15a]. isoclinic [ET18]. Isometries [Hor17, Nag15, BMN16, BSZ18, BBMP19, Bil19, BJ15, BKS15, BKPS17, BKSP23, Cha15, CGL16, FG16, GS17, GT16, GS15, MMS16, Mol18b, aHRT15]. isometry [CCW16, FX19, GG18, Oi19]. Isomorphism [Pel17, ET16]. isomorphisms [BFD18, GJK18, PR18a, STW19, Wan16c]. Isospectral [MGM17, BS18c, DT15]. Isotonic [EE16, NN16]. Issue [DLW15, BBF19]. Itô [WHG17]. Iterated [Dal17a, BMZB18, Col19, GP19b]. iteration [BB18, Lia17, RW16, RV18, XXZ16, ZZR15, SK17a]. iterations [SSH15]. Iterative [CA18, FS19a, LC19, AL19b, Had17, RSS17]. IV [HW16]. Iwasawa [Saw16].

J [Dra16b, LCNZ19]. J. [Dra16b]. Jacobi [AM15b, BW16, BM17a, BKNS18, HP15, KH17, SS15b, Wei15, WP17]. Jacobian [KMM17, dB19b]. Jacobians [YT16, YdB19]. James [AR15b, KL19b, LLH17, PSG16, PSM18, PMW19, SPH15]. JB [BCP16, KLL17]. JB-algebras [KLL17]. Jensen [MPV16, Nie15a, Vir19]. Jersey [Bru16a, Tsa16b]. Jin [Bru16a]. Johnson [Bar15, CHJM18]. join [CDR18, IHM08, IK18, Mat15c]. Joining [SNDM17]. Joint [BFA18, CM19, cSfCfX15, AMC18, ACM17, BF18a, CL19, GMW18, KS15b, Lau18, Pep17, SWZ18]. Jointly [Vir19, Alf18a]. Jordan [AEKP16, ALH15, AA17, Bal15b, BLdS16, BLPT15, Ben15, BM15, GLAdS19, GW17a, Gow17, GJ18, HA16, IKVZ18, JG16, JG17, KS15a, KVP19, KSTX15, KY17, LNW17, MO16, MP16b, MN19b, MMR18, Pet15, Rad17, TW16, YL17]. July [GHPS19]. just [Row16b].

Kac [Fik18, FM19b, FH19, Sin17]. Kaczmarz [BW18, BW19, NZZ15, OZ15b]. Kaczmarz-type [OZ15b]. Kadison [Hou15]. Kalman [RRV15]. Kaplansky [DK16, RY15]. Karaduman [CL15d]. Karcher [FS18b, LK16, SMH19]. Karlsson [MB15]. Karpelevic [JP17]. Kautz [Dal17b]. Keller [PC15]. Kemeny [Hun18b, MMM19, WDV17]. kernel [Pep17, Ram16, San18]. kernels [AG15b, AG20, BGKP19, CJW19, GW19b, Pan19a]. Kerov [GH15]. key [Hun16]. kind [KT17, YMZ19]. Kirchhoff [DG18, MMS18, RFPS18, RFPS19]. Kirchhoffian [Pal16]. kissing [Che15a]. Kittaneh [Lin17b]. Kneading [ABO15]. Kneser [DHS18]. Knill [MB18]. Kogbetliantz [MH15]. Kolchin [KLS19b]. Korea [CFH^{+}16]. Korovkin [KNR18]. Korovkin-type [KNR18]. Koteljanskii [JZC19]. Krawtchouk [HK19a]. Krein [CMM19, GMP16a]. Kreiss [Rao18]. Kronecker [BTW16, FS18a, FK18, GF17, HF18, Hil18, Lot15, PRS18]. Krylov [CA18, BMS16a, CMV19, CS18a, Dax17, Sch16]. kurtosis [Hür15, Lop17]. Kutta [BDDSC17]. Ky [HRM18, Pet15, TY15].

Lagrange [CPW17, VT18]. Lam [GKR16]. Laman [KS15b]. Laplace [Reb19, Sch16, SY17]. Laplacian [ASMN17a, GP18, LLS17a, ZL15a, ABC^{+}16a, AA19, ARR18, AG15a, ASMN17b, ACPR15, ACRR15, AGRR16, ADLR19, ARSB18, AH16, BK16, BS17, BS15c, BZ16c, BBC18, BCC^{+}18, BDVRT15, BH16b, CR17b, CCdFV18, CD16b, CX17, CLR^{+}18, Che18b, CLF18, Che19a, Dah15b, DdSJdFDV15, Das15, DMT16, DM16a, DS17a, DG18, DV19a, DL16, EHHL17, EHLP18, lFwW16, FuIKT16, FWB^{+}19, FdFdSDV18, GAP16, GCP18, Gho16, GRS17, HvD18, HT17, HQX15, HL19a, JMM16, LZ15b, LZ15c, LwW16, LD16a, LWCS16, LZ16b, LZW17, LDZ18, LLS17b, LLZ17, LL18a, LHH17, MQTW18, MMM19, NS19, NLWJ18, Obo19c, Ost15, PG15, SW15, STW18, STW19, SV17, TWR15, TWM16, TW18a, WH18, WZZ19, WZL15, XSW19a, XS18, YG15, YQS16, Yua16, ZL15b, ZW16]. Laplacian-energy-like [DG18]. Laplacians [FCLP18, MM19b, NS17, Nag19, SCD17]. Large [BDFR15, dSP16c, dSP16d, BW19, BGM^{+}19, BB16a, BDLM16, CR18a, EM17, HK18a, Jia15, LMS15, MNT19, PS19d, VHB18, Yin18]. large-scale [BB16a, EM17, HK18a, LMS15]. larger [YSS16]. largest [ABH^{+}19, AB15a, BK17c, Che19a, CLM19b, DMS19, lFwW16, FuIKT16, HQX15, Kim15b, Lin15a, LDZ18, MHL15, McK18, MNT19, Obo16, OQY17, SAdFZ15, Sta19, TW18a, XWD18, YQS16]. lasso [WLH18]. latin [War17, Shi18c]. lattice [EHLP18, Joe16, MMR16, MMR17, MMR18, WDFS17, YDYY18]. Lattices [BFG^{+}16, BF17b, Bie15, LM17, MdlP16, Str18, Yam15]. Laurent [Che18a]. law [PCI15, XSW17, dCdSA15]. laws [NN19a]. Lax [AS17]. Layout [RR18b]. LCM [AYK17, HT18a]. leading [De 16b, HJ16]. learning [SH15b]. least [BZ16c, BBC18, CGDM16, CKM16, CGM17, CRS15, DS18b, Eld15, GRS17, HK18a, JKS15, KL15, KLY18, LM16b, LC19, LHC15, NZZ15, SZ17, UM16, WZL15, Yas16, ZG18]. Lee [Mor16b, SSS15, TZK17]. Left [Dra16a, HW16, Lie18]. Legendre [Reb19]. Leibniz [LMO16, CC16, CLOK13, CILL16, KKO15]. lemma [Alf15, CS15b, CJ17c, DP15a, Lev15, Cim15]. Length [GMM19, WZW19, DF15, DF18, LMS16]. Lengths [GMM16]. Leonard [AC15, HLG15, LHG15, Nom15a, Nom15b, NT17, SGH16, Ter17a, WHG17]. Leslie [Ben18]. less [FZwCW16]. Leuven [FLVV18]. level [KRZ^{+}17, MP18, WQH19, Zas16]. lexicographic [WW18]. LFED [Zha17]. Lidskii [MRS15, MRS18]. Lie [ABM18, AEV15, AKR15, AK16a, AK16b, BC15a, BdlCL17, BCS15, BCS16, BP19b, CCiT18, CPS17, CCW18, CW15a, CL15a, Chi18, Cos19a, CLP16, Der19, DR17, EF17, EF19, FKPS18, GPI18, HAM16, Hol16, HZ19, HK17, KLS19b, LL19, LZ16c, LZ17c, LGZ18, PR19, RZ17, Saw16, Wan16c, WZ17b, ZK17, Zus17b, vWZ17]. Lieb [Bek16, FS17b, Hua19]. lifts [CV17]. like [AB16, BDP^{+}18b, CCGO17, DG18, DA19c, DF15, DF18, LWY16, LZSD17, LZSD18, Mel16b, Nie15b, RW16, VK16, Zhe19]. likelihood [Ruk15]. Lim [DDF17a]. limit [Ban16, LMM^{+}16, Yin18]. limited [EM17]. limited-memory [EM17]. Limiting [Tab16, CG15]. limits [Bik19, Ege16, Nat19, PR18b]. Line [Her16, ALMZ^{+}19, CG19a, DF16, Dal17a, DMG16, DM16a, FdC15, GCP18, Gre18, HS19a, LMZRZ18, MST15, Ref16]. line-Hermitian [CG19a]. Linear [AMR19, ASMN17a, Ali20, Ano16a, AG20, Bal15a, Bal15c, Bal18, BR19, BTW16, BMVW16, BF17b, BO17, BKSP23, CC17b, CR18b, Dog18, Duk15, EKSV18, EP18a, FJS21, FL15, GP18, GWL^{+}18, GJK18, GHPS19, HF18, Hou20, HSS16b, KAPS20, KS19, KHI16a, LMO16, LM16a, Lau18, LT16a, LLS17a, LZSD18, LD16b, LD17b, MRV15, Nom15c, PW20, Pud16, RB19, SM18, XFZD17, Yua15, ZL15a, dCDFK18, dS19a, ABP15, AAdFS19, AJ19, ÁKM17, ABO15, ANAPSR17, ABGJR^{+}18, AW15b, BW19, Bal15b, BP16a, BLPT15, Bel16, BF17a, BGCMASS19, BS16c, BRN18, BM17c, BL19, BMO15, CEM15a, CVV19, CNX17, CJ16c, CJ16d, CJ17c, COS18, CFNP17, Cus17, DGCC16, DGGP18, DS18b, DdFR16, DU16a, DN16, DN18, DdC16, EE16, Ema18, EJ19, FM17, FHS17b, FKPS18, GM15, GSP16, GC19, GR15, Gow16]. linear [Grc18, GBRS15, GHT16, Hla15, HR16, HS19b, Hoo19, HM17b, KL18a, KKA15, Kis15, KÖ16, LNT16a, LPPZ19, Lee19b, Li15a, LX17, LLL17, Li19, LS18b, Lom17, Mar18, MSSZ17, MMR16, NA15, NN19b, NINS16, NT19, NP15a, Nie15b, Obe16, Olk15, ÖK18, PSG16, PSM18, dSP19a, PMB15, Ram17, RJK18, RH15, RT16, RW17, SFW18, SR18, SNP16, SDK17, SK17b, SSS15, Shi16b, Shi16c, Sug17, SS15c, TZK17, TXZ18, VGG17, VF17, VG18, Wan16b, WZL17, WT18, WLL19, Wan19, Wój19, XX12, wXL14, XvdBvdLS15, wXZ19, YLT16, Yin18, ZD15, ZZR15, ZC15, ZH17, ZLQ16, Zhu19]. linearization [BŚS18, VT18]. Linearizations [MP16a, PM18, AB16, BDFR15, BBFF17, CCGV^{+}19, DA19a, DMQ19, FS17a, NP15b]. linearly [Mar15a, MC19, Vij15]. lines [CGSZ16]. Liouville [GM16a, YSZ17]. Lipschitz [AAdFS19, BP16a, BZ16b, GMW18, Oi19]. list [HMPT18]. lists [AMR18, BCJ^{+}16, JZ18, MAR17, MN19a]. Littlewood [AP15a, ANAPSR17, Nun17]. LNED [Zha17]. load [Zha15c]. Local [AKR15, AK16a, AK16b, AA17, CW15a, CL15a, Cos19a, CDH18, LZ17c, MRS18, AAM15, ABM18, ÁVAK^{+}19, BM17a, BE15, Ben15, BM15, BL18c, BCP16, CP16a, CJ16c, Cos16a, CS18b, CQS18, EJ17, EG15, Est16, HLM17, NT19, NP15a]. local-global [NT19]. Localization [BM18, BM16a, LJL16, Mel19, NN19a]. localizations [SKC18]. locally [AW15b, Beh17, Bik19, Che18a, DF16, Man19, TXZ18]. location [AAS16, FHJT19]. loci [BBT15]. locus [Jel18]. Log [AGS17, HL18, Hia19, BDK^{+}17b, Hia16b, LS18a, Min17]. Log-Determinant [Min17]. Log-majorization [AGS17, Hia19, Hia16b]. Log-majorizations [HL18, LS18a]. logarithm [Miy19]. logarithmic [BES16]. logarithms [BNST17]. logistic [KM18]. lollipop [BP15]. long [GH19]. look [UZ19]. look-ahead [UZ19]. loop [BGM^{+}19, BBGM16, GSZ16, SZ17, Zaj19]. loop-free [GSZ16, SZ17, Zaj19]. loops [AvdHS15, AvdHS16, GPI18, Sim19]. loose [Xu17a]. loose-coherent [Xu17a]. Lorentz [Bal15a, HC19]. losing [BdP18a]. Lovász [OP15]. Low [BA15, GS19b, RSS17, BB16a, BR16b, CCO15, FS19a, GH17, HKM18, lHS16, KU16, MS19, Mor17, OVW18, PQY15, RGPH16]. low-dimensional [BB16a]. Low-rank [BA15, GS19b, FS19a, GH17, KU16, Mor17, PQY15]. Lower [AR16, AP15a, CLL15, DTZ16, TWM16, WWC18, AL15, ARS17, AdFRR18, Bal15c, CR19b, HW15b, Mor17, Obo19b, SP16, ACRR15]. lower-upper-lower [SP16]. Lowering [Ter15]. Lowering-raising [Ter15]. Löwner [Pas18]. LS [MVPST19]. LS-category [MVPST19]. LSMR [HKP17]. LSQR [CN19a, HKP17]. Lur’e [SL16]. Lvov [DK16]. Lyapunov [eFLYL15, Gór17, GX16, GX17, GX19].

MacWilliams [TZK17]. made [Kal16]. magic [JKKL15, War17]. magnetic [FCLP18]. main [HKLQ16, HHL15]. maintaining [MGM17]. Majorization [DGS18, Fie15a, Nie15a, AGS17, Dah15b, EE15, EM15, Gow17, Hia16b, Hia19, Lin15c, Lju15, LD16b, LD17b, NA15, Obo19a, Obo19c, PP15, XF17]. majorizations [HL18, LS18a]. majorizing [Sab19]. Malcev [HAM16]. manifold [BA15, DH15, DSX15]. manifolds [BC19a]. map [BP16a, Dym16, MPS17]. mapping [Cus17, GMLdS16, Tsa16a]. mappings [AAdFS19, BMGP15, CC17b, DGK^{+}17, DKS15, DdFR16, FJ17, FJM18, eFLYL15, FHS17b, MSSZ17, Wój19, WMC15, Zha18b]. Maps [AAM15, BE15, BM15, KP17, MPV16, Pet15, ZH15, AKS17a, AKS17b, BS16c, Bou16b, BL18c, BL19, CP16a, CHK19, CKL17, CKT19, COS18, Cos16a, Cos16b, Cos19b, DdC16, EJ17, EP18a, FMR19, Gow17, HF18, HYH15, HLS^{+}16, KVP19, LT12, LT16a, LT16b, LY17, MR17a, Mar18, MS16b, MM17c, Nie15b, Oi19, Ong18, PC15, QH15, SDK17, SK17b, STZ15, Vir16, XFZD17, YT16, YdB19, YLT16, YL17, dB19b]. Mardia [Hür15]. marked [CFP18]. Markham [LZ17b]. Markov [BPZ17, BK17c, BCF^{+}18, CLR19, Cho19, GZB19, Hun16, Hun18a, Hun18b, JLT17, LZ15a, Mas17, Reb19, SY16a, Sab19, Sku17, SH15b, VB18, Vas15, WC15]. Markovian [GCQX15]. mass [HZC15]. matchable [AJ19]. matching [Ash19, HHL19, MWT16b, TH17, TWM16, TW18a, WWC18]. matchings [Mes17, MM16b, TD15, YG15]. mates [CHJM18]. Math. [Zha15a]. Mathematical [Zha15a]. Mathematics [Bru16a]. Matic [Zha15b]. Matrices [AR15a, Bre18, CLST18, CSJ16, GS17, GdLL17, Hil17a, LdlP15, RFPS18, AMR^{+}16, AOK15, AM19, AAS18, ATM18, AG17, AMR19, AJL16, AK19, Alf18a, APS18, ANP16, Alo15, AKYD16, AYK17, ACM17, AMZ16, AdF19, ARS17, AMNR18, AMR18, AMG19, APT17, AGV18, ABK19, AL19b, BJ16, BW16, Bai18, BFFN16, BM18, Bal15a, Bal15c, Bal17b, Bal18, BM16a, BS15a, BK19, BBdH16, BBdH18, BDDO16, BP16b, BBH16, Bar16b, BGS15, BGM^{+}19, BF18b, BF19a, BŚS18, BM19b, BE15, BN17, Ben18, Ber17, BS16c, BJL19, BC19a, Bie15, BDLM16, BR16b, BBGM16, BBC^{+}15, BDS15, BC17a, Bor17, BC19b, Bot16a, Bot16b, Bou16b, BM15, BC15b, BSS17, BGSS18, BHSS18, BS16f, BM16b, BaHOS15, Bri17, BCU15, BGGS16, BD19, BR18, BÓ15, BO17, Bur15, BV18b, CCiT18, CRX15, CU18]. matrices [CC19, CDR18, CEM15a, CR19b, CLR19, CPZ16a, CGMSR15, CM18, CHK19, CR18a, Cha19, bCR18, CRC17, CS19a, CS19b, CLW15, CXLF16, CIY15, CL17b, CKS16, CJB18, CS18a, CK19, CN17a, CN17b, CN18, Cho17b, CKT19, CS15b, CHB15, Chu19, CI15a, CI15b, Chu15, CM19, CD16c, CW16a, Cos16a, Cos16b, CJW19, CLP16, CLS17, DD17a, Dah15a, DA19a, DA19c, DB19a, Dax17, DDCY17, DeV19, DD16a, DD19, DAG16, DEH16, DDF17b, DFK^{+}15, Dmy16, Dmy19, DO16, DL16, DMQ19, DM18, DdF15, DE15, DN16, DdC16, DS17b, DU16b, DKPU18, DHW16, Dya17, ELN18, Ege16, Egg15, EHK16, EM16, Ern18, EGT15, ET18, EW19, Fag19, FOvdD16, Far19, FT16, FMR19, FdC15, FM19a, Fie15b, FGH^{+}15, FM16, Fik18, FM19b, FI16, FKM17a, FR15, FR18, GG18, GSW16, GSSvdD16]. matrices [Gar17, GW16b, GWW16, GL19, GXYZ17, GH17, GIM16, GM16b, GT17, GMP16c, GW19b, GHT16, GHK17, GLRT18, GL15b, GL15a, GL16, GMM16, GS16, GJK18, GT18, GMM19, HS18, HKM18, HvdD16, HT18a, HAWM16, HS17, HJW19, HL17a, HJ16, Hil17b, HW15b, HP15, HT18b, Hon18, HLZ17, Hou19, Hou20, Hru16, HWZ15, Hua15, HSS16b, lHS16, HNR16, HLS^{+}16, Hua19, HL19b, Hür15, HK18b, IKVZ18, IW15, IHM08, IK18, IKW16, Jai17, JT18b, JS16, Jia15, Jin15, JSST16, JZ16, JT17, JZ18, KLS19a, Kan16, KS17, Kar16, KLP18, KKL^{+}17, Kim15b, KL15, KLL16, KK18, KLL19, KQZ16, Kis15, KL17, KM16, Kob17, KLS19b, KK17, KY17, KH17, Koz16, Kur19, Kus15, Kus16, LL17, Lau18, LLPS19, Lee15b, LP15, Lee15c, LL19, LT16b, LZH16, LDL18, LCWZ19, Lia17, Lie18, Lin15b]. matrices [Lin16b, Lin17a, LZ17b, LL15c, Liu15, LY17, LT19b, Lék16, LS17b, LSM16, LSM18, LMR16, Lyn15, MO16, MY16, MM19a, MS16a, MAR17, MS15a, MM16a, MMP17, MMV19, Mar16, Mar15a, MV16, Mat15b, Mat15c, MP16b, MP16c, MN19a, MNT19, MN19b, Mol18a, MS15b, MPV16, MM15b, Mor17, ME15, MMW17, MHN18, MP15b, MP19, NMBA19, NS15c, Ngu18, NW19, Nie19c, Nik16a, NA17, Niv15, Nob16, NK18, NSCV16, OW17, OVW18, OS17, OP16a, Ore16a, Ore16b, OP19b, Ost15, OBRA15, PSZ15, PT19, PS18a, Pas19, PM15a, PCI15, dSP16a, dSP16c, dSP16d, dSP16e, PM15b, PS19d, PR19, PS16, PS18b, PP18, PHW16, PW19, PW20, RS17a, RS17b, Ris16, Riz19, Rob19, RAAGA16, Roh18, RS18, Rub19, Sal16, SSCS16, SR19, Sch16, SK17a, Ser16, SA17, Sha16, SK17b]. matrices [She17, Shi16e, Shi18a, SSV17, SCD17, SH15b, SS18, SB18, SUY16, Sol18b, SW18a, Sot17, SJS15, SS15b, SV17, Ste18, SSB15b, SSB15a, SG17a, STZ15, SWZ18, TH16b, Tan16b, Tan19, Tar18, TS18, DDMV16, Ter18, TW18b, Tsa16a, TW19, TCD15, TC15, TG19, Ulu19, UG18, VB18, VQ17, VQ19, VV18, Vec16, VHB18, VLGS15, VB19, Wan15a, WLP16, WQH19, WLL19, WZ17b, WMC15, WHC17, XW16, XX12, wXL14, XFZD17, XSW17, wXZ19, Yan16, YDHX18, Yin18, Yor17, YZZ18, ZS17, Zha18c, ZMT19, Zha19a, ZLLL18, Zhe19, ZM16, ZF15, Zim15, dCFF17, dlCMP15, dlCMP17a, dlCG18, Tsa16b]. matricial [FKM17a, FKM17b, FKMS18, HZC15, JP17, ZHC16]. Matrix [AMH11, BLdSV19, Bru16a, BPT15, Che18d, CP17, DKS17, Eve18, He15, KM15a, MN15, MHA18, MP16c, Nat19, PW18, PV17, Rao18, ZSQZ19, Zhe19, AMR^{+}16, ARS16, AG17, ARR18, ASMN17a, ASMN17b, Ais18a, AGH17, ABG16a, AGS17, AGN18, AB16, Al’17, AB15a, AMZ17, ADW19, ABD19, AL16, ACDT^{+}15, ABK19, AA17, AB18, BR15, BmBCC17, BS15b, BS16a, BES16, BDG15, BB18, Beh17, BGKP19, Ben15, BGT^{+}19, Ber18, BLY16, BR16a, BHK16, BFD18, BHP17, BA15, Bou15, BDK^{+}17b, BC16b, BaHOS15, BS19c, BDFR15, Bün17, BR17, CL19, Cal16, CJ17a, CG15, Cha15, CCGV^{+}19, CLT15, CMQV16, CPW17, CLR^{+}18, CLM19a, CJ15, CA19, CJ16b, CS17, CYD19, Cim15, CDTZ15, Cos19b, CHT19, DGS18, DdSJdFDV15, DA19b, DI16, DDL17, DDMP19, DR16, DKS15, DHLT19, DF15, DF18, DPR19]. matrix [DD17b, DFKS17, Dmy17, DD18, Dod16, Dod17, DKNS16, DPS19, DU16a, DPV19, DE15, DvSW18, EKSV14, EKSV18, EEV17, EH15, EJ18, Ere16, FI19a, Fas19, FS17a, FS18a, FS17b, FdC18, FMM16, Fis17, FHS17a, FK16, FKM17b, FKRS17, FX19, FKS16, GGMPC15, GW16a, GW17a, GWW19b, GWW19a, Gil16, GS19b, Gol16, dMGC19, Guo19, GLMS18, HLPS19, Hay19, He16a, HAWM16, HNS15, HR17, Hia16a, Hir18, Hoa17, HC15, HFS15, Hu15, HNR16, HN18, Huh15, Hut17, JMRV16, JLD17, JLDS18, JCW19, Kal18, KS15a, KM15b, KK18, Kis15, Kli19, KPL18, KIS18, KT15, KB15, KMOR19, LMS16, LMS15, Lee15a, Lee19a, Lev15, Li15b, LWY16, LM19, LLSX15, LLL15, Lin16b, LdlP15, LW15, LLH17, Lop17, LG18, LMPM17, Mac16, Mac18, MP16a, MR17a, MI17, MR17b, MOR16a, Mat15a, MP17, MNTX16, Mel15]. matrix [Mel16a, Mes17, MR19, MRV15, Mig16, Miy15, Miy19, MGM17, Mon17, Mys16, NS15a, Nev18, NP15b, NR16b, OZ15a, Pan15a, PRS19, Pas18, dSP15, dSP19a, PM18, PPZ19, QWW18, RRT16, Rin17, Riv15a, Riv15b, RMP18, Rum18, RB19, Sab16, SR19, Sas18, Shi16f, Shi18b, SÖ16, SZ19, STW18, STW19, SS19, Tab16, TY15, TW17, Tom15, TVD15, iT16, iT19, UZ19, WYL15, Wei15, Wim16, WC15, XSW19b, XXZ16, XX12, wXL14, XvdBvdLS15, Xu17b, wXZ19, Yas16, Ye17, YYSX19, You15, YWD16, ZD16a, ZD16b, Zho17, dCDFK17, dCDFK18, dlC15, dlCF16, dlCMP17a, dlCdR17, vWZ17, MM19b, Bar15, Zha15a]. matrix-products [BR15]. Matrix-tree [BPT15]. matrix-valued [FKM17b, Lee19a]. matrix-variate [Mat15a, MP17]. Matroid [GRV16]. matroids [MFdM15, Pea15]. Max [Hoo15, Hoo19, Nik16b, EHN16, GMW18, Li19, MP15a, Mys16, MP19, PS18b, RLP19, Ser15, Shi16b, WT18, ZD15, vDS16]. max- [vDS16]. max-algebra [PS18b]. max-min [MP19]. Max-plus [Hoo15, Hoo19, EHN16, Li19, Mys16, RLP19, WT18, ZD15]. Maxima [NY15, Yua14, Yua15, dFNP16]. Maximal [BR18, DO16, Mes17, NS15c, dCDFK17, dCDFK18, CRVSS19, DMS19, DF15, DF18, EEV17, GH16a, LZ15b, LCNZ19, PS15, Ter18, ZHC16]. maximality [HT17]. maximally [GW19a, Poo15]. maximized [dSN15]. maximizers [LT19a]. Maximizing [EH15, GRS17, HSS16a, Kol15, NLW18]. maximum [BWSZ15a, CU18, CC18, De 16b, DdF15, FdC18, HZC15, HLW15, HW15c, HZ15, IKVZ18, LMZRZ18, LLS17a, LLS17b, MO18, OS17, VQ17, XWL17, ZW16, ZWS18]. maximum-volume [MO18]. Maxwell [KS15b]. Mazur [WH19a]. MB [LQL15]. MB-tensors [LQL15]. MCMC [BL15]. MCMC-based [BL15]. mean [BHML15, BJL19, BK17c, FS18b, FS19b, HL17a, Hoa17, Hun18a, IKW16, KL15, KLL16, KLL19, Nob16, NK18, Pep17, SMH19, Zou17, Zou19]. means [BLY16, DKS17, DDF17b, DU16a, FS17b, Han19b, HK17, KK18, KLL18, LK16, MHA18, Mol19, Sab16, UWYY15, UYY17]. measurable [Bar17b, CRVSS19]. measure [BŚS18, DP15a]. measure-extension [DP15a]. measurement [Ema18]. measures [ACM17, AM16b, BFM19, HZC15, KL15, MS15a, MS15b, dlPJG18]. mechanical [BTN^{+}18]. mechanics [BdP18a]. median [YYMK17]. meet [IHM08, IK18, Mat15c]. meeting [FLVV18]. Melbourne [Bar15]. memory [EM17]. Mercer [Nie15a]. Merging [MR17a]. Merikoski [AKYD16]. Mertens [CK19, Kli19]. Metcalf [MF16]. method [ATM18, BM17a, BW18, BW19, BFN15, BB18, BKLP16, BŚS18, BMS16a, Bel16, BV18a, CVV19, CGSC19, CP16b, Cox15, CFNP17, Fas19, FT18, FG17, GH16a, Guo19, HK18a, Her16, KR17, KH17, KPRvdO16, KL18b, MLW15, MH15, MP18, NZZ15, Pas19, RW16, Sch16, WP17, XXZ16, wXL14, wXZ19, ZLQ16, SK17a]. methods [AL19b, BT17a, BFM19, BHK16, CDRT17, DU16a, Gór17, Had17, HS19b, JKP16, LG18, LNT16c, Nik17a, OZ15b, PS19d, PW18, RR18a, SLY17, VK16, Zas16, ZZR15, Zhe19]. metric [CPZ16a, LCF19, Wol18]. metric-preserving [CPZ16a]. metrics [Ber17]. Metzler [Bri17]. Milner [GLW17]. min [MP19]. Minimal [Dod17, IKR16, Ple17, dCFF17, BR19, BSP18, BDP^{+}18b, CR18a, DA19b, DA19a, DA19c, VD18, DPV19, DV19b, DL15, GLMPC18, dMGC19, HM17a, LWC18, LCNZ19, MM15a, OQ18, OS17, SP16, SS17b, DDMV16, DDV16]. minimally [FGS^{+}16, SG16]. Minimising [BK17c]. minimization [CDTZ15, LC19]. minimize [JKS15]. minimizing [BBC18, DSX15, LS15a]. Minimum [BWSZ15a, CG16, FGG^{+}18, HLPS19, PF18, AAG^{+}19, Cos16b, DM16b, GS19a, HWXC18, JLDS18, LOS19, MS16b, NS18, NA17, NLW18, SWB19, ZM16]. Miniversal [Dmy16, Dmy19, BSS17]. Minkowski [Hor17, JM16, Liu16a]. minor [Bil19, Nik17c, RKT15]. minors [BS16e, BC15b, SUY16, dCdSA15]. minus [DFM17]. minuscule [MM15a]. Miroslav [RRS16]. MISO [BPZ17]. missing [Dal19, JZ18]. Mitra [Rak17]. Mixed [DS18b, MP18, WYLW17, ARS16, AM15b, Bar16b, CLZ16a, CP16b, De 16a, FKD15, Gow16, HLLZ17, KÖ16, LL15c, Moh16, ÖK18, RH15, Wan16b, WLWZ19, Wei15]. mixing [Lev15]. Mn [COS18]. Mn2 [COS18]. Möbius [Bur18, CR18b, DGSV17, HM16]. mock [Zus17b]. mock-Lie [Zus17b]. modalities [Dog18]. Model [WLH18, BFFN16, BCMASS15, HvdD16, KM18, LHC15, Mar15b]. models [DV18, Ema18, FL15, GH16b, HPZ19, KÖ16, ÖK18, QWW18, RH15, Ruk15, SNDM17, Wan16b, Zha15c]. moderate [ET16]. modifications [SSB15a]. modified [AL19b, CLLS19, JM16, JLM18, wXL14, wXZ19]. Modular [HMS17, SY15, Shi16a, TC15]. modularity [BBC^{+}15, FT16, FT18]. module [CD16d, Gua18, IKR16, LT12, LT16a, LMX19, Shi16g]. Modules [HCS17, CPS17, Gon15, LZ16c, LZ15e, MM17a, MEMM17, XF17, Yan18, dSRST17]. moduli [AGS17, Pie18]. modulus [Cos16b, MS16b, XXZ16, wXL14, wXZ19, ZZR15, ZLQ16]. modulus-based [XXZ16, wXL14, wXZ19, ZZR15]. Mohy [Fis17]. moment [CIY15, CJ16b, DE15, EGT15, FKM17a, FKM17b, FKMS18, HZC15, LM17, Lop15, PS19e, Riv15a, ZHC16]. moments [Lék16]. monic [CK15a, FK16]. monoid [CDG16, CLR16]. monomers [NS15a]. monomial [Lax16]. Monotone [Sch16, BU18, DDF18, Hoa15, JS18a, Mas17, Naj17, Nob16, Pas18, iT16, iT19]. Monotonic [Grc18, AB15a]. Monotonically [ZF15]. Monotonicity [LG18, WL15, Cas19, DKS17, DDF17b, GMW18, Nag19, NW15]. Monte [WC15]. Montgomery [Cox15]. Moody [Sin17]. Moon [LN17]. Moore [AB18, CGMSR15, CGDM16, Dal19, HWZ15, KB15, Nor18, XW16, XSW17]. Moreau [Sol19]. Moreau-type [Sol19]. Morphisms [Cha19, EC19, CLR16, DKNS16]. Morrison [ACM15]. Moser [LN17]. most [Alf18b, CJ17d, EHK16, PS19a, TW17]. Mostow [GM17]. MRD [CMPZ18, GZ19]. MRD-codes [CMPZ18, GZ19]. MUBs [Kar16]. Muir [dCdSA15]. multi [AKPS19, CJ17b, KAPS20, XLQC16, Zas16]. multi-dimensional [AKPS19, CJ17b, KAPS20]. multi-hypergraphs [XLQC16]. multi-level [Zas16]. multigraph [McK18]. multigraphs [Gu16, O16]. Multigrid [CDRT17, BDDSC17, BHK16]. Multilinear [Ano16a, BMVW16, CW16a, BT15, BM17c, CR19a, DGK^{+}17, DKS15, DK16, Fag19, LT16b, Pop16b]. multiparameter [HY18]. multipartite [BK18, BSMSZ19, HK19b, Obo19a, Obo19c, WYL19]. multiple [Ais18b, AM15b, BBT15, BDK^{+}17a, BTN^{+}18, CDMP15, CKM16, DKSV19, LPPZ19, NSCV16, PM15b]. multiplication [Beh17, CYD19, DK18]. Multiplicative [CGDM16, CL16, MRS15, XSW17, CJZ18, Hil15, Hil17a, MI16]. Multiplicatively [BL18c]. multiplicativity [San15]. Multiplicities [FdFdSDV18, Piv19]. Multiplicity [JZ18, ARSB18, BK15, BK16, BCJ^{+}16, CR17a, CPR18, ET16, JDS17, JLDS18, JST19, LHH17, MQTW18, OS17, Row16a]. Multipliers [JC18, LH17, PQY15]. multisplitting [ZZR15]. Multivariate [CHK17, FKD15, Han19b, HKM18, KRZ^{+}17, KRZ^{+}18, KPRvdO16, LW15, LLH17, Lop15, Ruk15]. Murdock [Fik18, FM19b]. mutation [CKS16]. Mutually [BL16, Kis15, Lie18, MB15, Szá16, XX12].

Narayana [AMPT16, Bar16a]. natural [Lie18]. naturally [CLOK13, LMO16]. Near [MOR16a, DL15]. Near-invariant [MOR16a]. nearest [GKS19]. nearly [MOR16a]. nearness [HNR16, NR16b]. necessarily [Mon17]. Necessary [Bar17a, Li15a, LS17b]. negative [FW17, GSZ15, GSZ16, GJK18, JLM18, Koz16, LS19, MM19a, SZ17, TW17, Wol18, Zaj19]. Neighborhood [BSS17]. Nekrasov [OP19b]. nest [CL15a]. nested [AAC^{+}15]. network [GZB19, QWW18, Sar19, SZW16]. networks [AM16b, Ban16, BS18c, CEGM16, CEM19, Est16, KKL^{+}17, Kir15, Zha15c, UZ19]. Neumann [JC18, CHASSSG15, CQCZ19, CW15b, DHL^{+}17, DGCC16, DU16b, DKPU18, LZ17c, PR18b, SCD17, WFS19]. Neural [SZW16, UZ19, QWW18]. Newton [SK17a, BV18a, CFNP17, EG19, EM17, Guo19, LG18, PM18]. Nice [CW17]. NIEP [Ben18]. Nikiforov [AR16]. Nil [BM16b]. Nil-clean [BM16b]. Nilpotent [HA15a, IKVZ18, LS17a, VF17, AW15b, BdlCL17, BP19b, Bre18, BO17, Che18a, CXLF16, Chi18, DO16, EL16, GLAdS19, GS15, HA16, JS19, Pel17, Ste18, TXZ18, WL16, YT16, YdB19]. nilradical [BCS15, BCS16, CLOK13, LMO16]. nilradicals [KKO15]. NMF [Miz16]. No [LMT19, AM18, Nik17c, PZ17b, dFNP16]. Nodal [Urs18]. nodes [PCL15]. Noetherian [Kwo16]. Noetherianity [Bik19]. Noise [HKP17, Hoo19]. Non [ABG16a, AYK17, Bal17a, BPR17, JKS15, Ser16, Zus17a, AGHN18, BŚS18, BGCMLSSS16, BGdCCMSS17, DP16, FuIKT16, Fil18, FJ19, GSZ15, GSZ16, HLPS19, HAM16, JLM18, uIKF15, KM19, Koz16, KNR18, Kwo16, MM19a, Nev18, PP18, SSF16, SZ17, Vas15, WZL15, Yin19, Zaj19, ZG18, ZLQ16, Ziz16, cSfCfX15]. Non-Archimedean [ABG16a]. non-autonomous [SSF16]. non-backtracking [AGHN18]. Non-bipartite [JKS15, WZL15, ZG18]. Non-commutative [Ser16, KM19]. Non-divisibility [AYK17]. Non-existence [BPR17, Fil18, FJ19]. non-extendability [BGdCCMSS17]. non-Hermitian [BŚS18, Yin19]. non-holomorphic [Nev18]. non-homogeneous [Vas15]. Non-integral [Bal17a]. non-Lie [HAM16]. non-modulus [ZLQ16]. non-negative [GSZ15, GSZ16, JLM18, Koz16, MM19a, SZ17, Zaj19]. non-Noetherian [Kwo16]. non-normal [BŚS18]. non-odd-bipartite [FuIKT16, uIKF15]. Non-Orthogonal [cSfCfX15]. non-Runge [BGCMLSSS16]. non-self-adjoint [KNR18]. Non-semigroup [Zus17a]. non-standard [PP18]. non-stationary [Ziz16]. non-symmetric [DP16]. non-zero-entries [HLPS19]. nonassociative [Bre17]. nonautonomous [Dra18]. Noncommutative [DH19, DvSW18, LNW17, Pas18, Vol17]. noncommuting [Sot17]. noncompact [Saw16]. Noncrossing [dS18a]. nondegenerate [ZHC16]. nonderogatory [GLMS18]. nonequispaced [KP19]. nonexistence [FHH15, Mic16]. nonincreasing [DdC16]. Nonlinear [BB17, Cos19b, MS16b, WMC15, BHP17, Cos16b, Fay18, LM19, Mar15b, Sab19, WYL15, XXZ16]. nonlinearly [EH18]. Nonnegative [ES16, SA17, SJS15, TH16b, AG17, APS18, ARSZ15, AMR18, BF17a, BC17b, BS16f, CU18, CCCNT19, CHB15, EHK16, FBH19, FdC15, GH16a, HYY15, HK18b, JP16, JMP17, JS15, JRS19, Lee15b, LLV15, Lia17, LS17b, MPS17, MP16c, Shi16e, SW18b, SW18a, TCD15, VB18, WLP16, YHY15, Zha18c]. nonnegativity [AG16, AG17, GW19b]. Nonoscillation [Sug17]. nonoscillatory [SH15a]. nonreal [Sam17]. Nonscalar [Mel15]. nonseparable [HL17c]. nonsingular [AMG19, BBdH18, BHSS18, BWSZ15a, HWXC18, MY16, Sol18b, VQ19, ZLLL18, dFRS16]. nonsingularity [Liu15, Liu18]. nonstandard [AR18]. Nonsymmetric [Bou15, HZ15, Had17]. nonuniform [Col19, Dra18, Fre18]. nonzero [BOvdD17, FGG^{+}18, Gao18, MM17b, RS17b, Shi17b]. Nordhaus [EA17, HL19a, LOS19]. Norm [CM16, GHK17, JLM18, AKS17a, AKS17b, AdlPR16, ARR18, AN17, Alo15, BLY16, CR19a, CLZ^{+}16b, eFLYL15, GH16a, GS19b, GMO19, GLRT18, HT18a, Hay19, Hia16a, Hoo19, HMT19, Hu15, JM16, KMS17a, Li15b, LL15b, LCWZ19, Nik17a, NA17, OP16a, Pet15, Reb19, SPH15, SS15a, Tan16b, WDFS17, XCJ18, Zha18c]. norm/energy [NA17]. Normal [JSST16, JT17, BŚS18, BD19, Cus17, DDL17, EG15, GSW16, GMP16a, GMP15, GMP16b, GMP16c, GBRS15, HZ15, HN18, IW15, IGW17, Joe16, KS15a, LWC18, Lin15b, LdlP15, LLH17, Nie18, Nie19a, SZ23, SG17a, YZZ18, dlCG18]. normality [BF18b, FKD15]. normalizable [MS16a]. normalized [BM17b, BCC^{+}18, BDVRT15, BH16b, DS17a, LSZ18, MMM19, SY17, cSfCfX15]. normally [ARR18]. normed [BJC15, CSW17, KS15b]. norms [CDMP15, CL15b, ELN18, GMW18, JM16, JLM18, Ken16, Lev15, LT12, LT16a, MRS18, OP15, PSZ15, Sab15, WL15, Zha19b, Zou19, Nik16a]. Northcott [LZQ19]. Note [Ano18-44, BaHOS15, CLF18, DHL^{+}17, ySpW17, Sri17, YG15, Bün17, Bur18, Dra18, FR18, GHT16, HKLQ16, HMSC19, Kum19, LHX18, LLS17a, LLS17b, NMBA19, NR17, Nor18, PS15, Rak17, Rod19, SZ23, SDK17, SSS15, Shi16e, Sin17, STW19, Str18, Wan15a, XF17, Zhu19, Zui17, dSRST17]. Notes [CW19, DU16b, SAdFZ15, DKPU18]. novel [DAG16]. nowhere [MS16c]. nowhere-zero [MS16c]. NP [BC17b]. NP-hard [BC17b]. Nuclear [Pop16b, BM17c, Hu15]. Nuclei [TZ19]. Null [KVV17, Ris16, CCW16, JMPS18, Kul15]. Null- [KVV17]. Nullities [TH17]. nullity [FdC18, zSqST15, WZ15, ZWS18]. number [ACT16, AAKS17, AB15a, AL15, Ash19, BL18b, CC17a, CLT15, CUPW15, CW16b, Cox15, DR16, FMY16, FdC18, HLPS19, HHL19, HZ15, IL17a, JL16a, JDS17, JLDS18, LOS19, Lin15a, Lin19, LL18a, LLT19, MWT16a, MWT16b, PF18, QKP16, Ram17, RMS16, TH17, TWM16, TW18a, Tom15, TD15, WM17a, WZL15, WWC18, Xu17b, ZLG17, Zim15, vDS16]. numbers [ABD19, BZ16a, CDR^{+}19, Che18d, DTZ16, DS18b, KM19, Lee15b, ME15, Van17, WYL15]. Numerical [AOK15, Ano16a, Bal19, BBP19, BMVW16, CJK^{+}16, GW16b, GWW16, GO18, KLN16, PW18, Sab18, She17, AOK19, AAS18, AMC18, AR18, AMR19, AK19, BFA18, Bal17b, Bal18, BFM19, BD15b, CJ17a, CDM19, CC17b, CCCNT19, CL16, CN15, CNY18, CN19b, CNM19, CJ17c, Dax17, DGP^{+}15, FK16, GW16a, GW17a, GW18, GWW19b, GL19, GWW19a, HA15b, HMNP15, KS15a, KKP19, KMY15, Kum19, LNT16b, LLPS19, Lee15c, LTWW15, LSZ18, MR17b, SPBB19, SWZ18, Zam19, Cha15]. Numerically [PZ17b, PQY15].

objective [Fay18]. oblique [BB17, BMS15]. observations [Sma15]. occasion [BBF19]. odd [Ash19, DF15, DD18, ET18, FuIKT16, FKM17a, FKM17b, HZ17, uIKF15, LDL18, MWT16a, Yua14, Yua15]. off [BMR17, Lee15a, PPZ19, SP16]. off-diagonal [BMR17, Lee15a, PPZ19, SP16]. often [Sol18b]. Old [Fie15b]. Oliver [MB18]. one [BM19b, BCF^{+}18, BCU15, EF17, EF19, FK18, FS19a, Gas19, HSS16b, JLDS18, KOST15, LS18b, SSB15a, TXZ18, EHHL17]. one-dimensional [EF17, EF19, EHHL17]. one-parameter [KOST15]. one-peak [Gas19]. ones [BL18b, TVD15]. only [FJ19, SS17b]. Onsager [BGV17, Ter17b]. onto [BFN15, FG17]. open [Beh17, DK18, YHY15]. operations [BN17]. Operator [HRM18, MEMM17, MHN18, NW15, SPH15, Uch15, WDFS17, AOK15, AKS17a, AKS17b, AEV16, AN17, AK19, ACM15, BBS16a, BU18, CFL16, CN19a, DDF18, Drn19, FS18b, FS19b, GS15, HJ19, Han19b, HPSS19, Hoa15, JS18a, JM16, KU16, LZH16, LN18, MF16, Naj17, OP15, OP16b, Pas18, PCI15, Reb19, SS15a, She17, Tan16b, UUG15, UWYY15, UYY17, WHC17, Zam16]. operator-valued [FS19b, HJ19, Zam16]. Operators [AR15b, DIR16, MMS16, Wój15, AB19, AAM15, AO18, And15, AAC17, AG18, ANAPSR17, ABGJR^{+}18, AG15b, AG20, BFA18, BF18a, BG15, BLPT15, BŚS18, BBSS18, BDRC16, BMN16, Bob16, BT15, BM17c, Bou16a, CDMP15, CEM15b, CEGM16, CPZ16b, CDRT17, CMM19, DS19b, DSCD17, EE16, EJ17, EM15, EHHL17, EHLP18, FPGPV17, FKPS18, GSP16, He16b, HA15b, HK17, JLM18, Kim18, KL19b, KMS17a, KNR18, Lee19a, LTWW15, LWC18, Lju15, LD17c, LMX19, MS16b, Mat15a, Min17, Mol19, Nag15, NINS16, Nie15a, Nie15b, Pan19a, Pan19b, PSG16, PSM18, PMW19, PCI15, Pep17, Pie18, PPZ19, PR18b, Pop16b, PRS18, Rao17, Sab19, Sam17, SST15, Tur17, Vir16, WFS19, XF17, Zam19, Zem16, ZH15]. Optimal [ANAPSR17, CCO15, CFM16, GMP16a, Hon18, DDSC^{+}16, Had17, lHS16, HK19c, JMPS18, La 16, LCF19, MRS15, NAS18, Sma15, WC15]. Optimality [Fas19, Li15a]. optimisation [CRC17]. Optimization [DH15, BA15, FS17b, FM17, Hil15, Kri15, Kur19]. optimization-reusing [Kur19]. Optimizing [BDDSC17, DeV19]. Optimum [GW17b, CLZ16a]. OPUC [MS15a]. orbit [IKVZ18]. orbits [BC15a, LNT16a, Lau18, MM15a, MBB^{+}18]. order [ABO15, BP16a, BGV17, BBMFM17, CFNP17, DV18, DY15, DFM17, DD17c, GM15, GL17, Gao18, GH16b, GS15, HA15b, JKS15, Kar16, KH17, LS15a, LZ15a, LUC19, Lom17, PCI15, PM15b, PS15, PPZ19, SY16a, SL16, Shi19, Sug17, TH17, WZ17a, XSW17, ZWS18, dCFF17]. orderability [BS16e]. Ordered [HW15a, CKS16]. Ordering [YSS16, BOT19, LLS17a, LLS17b]. orderings [IKR16]. orders [AB19, BMGP15, CIMW15, ET18, Mos17, RD15, WL16]. Orientation [HM16]. orientations [COvdD16]. Oriented [Ref16, AdlPR16, CRRY15, CLL15, CL17a, CLR^{+}18, DR19, HMT19, LHL18b, MWT16b, RS17a, WF19, ZXW19]. original [BS18c]. Orthogonal [Lin15b, Pan16b, War17, AM15b, BGV17, BF19b, BL16, CRX15, CG15, CJ16a, DEH16, DMQ19, DSCD17, DH19, FS17a, GGMPC15, GMP16c, HZ17, IR19, Jin15, LNT16a, LUC19, Mac18, MP18, Ota15, Riv15a, Riv15b, RAAGA16, RMP18, SNDM17, SJD16, Shi16d, Sim17, WQH19, ZS17, ZGW16, dlCMP15, cSfCfX15]. orthogonalities [Wój19]. Orthogonality [GSP16, AR15b, ABR18, CSW17, HHJM15, KL19b, KZ17, PSG16, PSM18, PMW19, SPH15, Tur17, Zha18b]. orthogonally [CW18b]. orthoplex [BH16a]. oscillation [EH18]. Osserman [NR16a]. Ostrowski [Nie19b]. other [CFW15, EA17, GSW16, KL18a]. outcome [LT19a]. Outer [BJKR17, BKNS18, LHC15, Shi17a, SSH15, SZW16]. output [Bel16, NN19b]. outputs [LPPZ19]. overdetermined [BW19].

packet [Far16]. packing [Gu16, LLT19]. Padé [PW18]. pair [BMS15, BGSS18, FHH15, Nom16, WHG17]. Pairs [SJD16, BGV17, BHSS18, BKS15, BKSP23, DdFR16, Dmy16, Dmy19, GMM16, GMM19, Lyn15, Mor17, NS15b, Nom15a, Nom15b, NT17, Ram16, RB16, SGH16, wXjMZL19]. pairwise [MC19]. palindromic [DDMP19]. PARAFAC [KL18b]. parallel [KH17, SdJY^{+}16]. parallelepipeds [Bil19]. parallelism [Zam16]. Parallelogramic [LMP^{+}17]. Parameter [AL19b, Zha15c, EH18, Fik18, FM19b, Gil16, KOST15, LMR16]. Parameter-invariant [Zha15c]. parameters [ACRR15, ACD19, BPZ17, HLW15, Ken16, KRZ^{+}17, LHC15, Mic16, Wan16b]. parametric [HS19b, RW17]. Parametrizing [And15]. paraorthogonal [Cas19]. Parking [MY16]. Parseval [KAPS20, AKPS19, MBB^{+}18]. part [DU16a, DMQ19, KLY18, LNW17, Ter17b]. Parter [BEF16, JS18b]. Partial [CIMW15, FG16, RD15, WL16, WZ17a, BS15a, BC19b, Cho17a, Eld15, FX19, GS17, Mel19, Shi19, aHRT15]. partially [BW19]. partite [KNY15]. partition [GK15, JZZ18, NS18, WDFS17]. partitioned [Hir18]. Partitions [CC19, CDG16, RNC17, Abi19, ARSB18, CD16a, Col19, dS18a]. parts [MMS16]. Pascal [KT17, Zhe19]. Pascal-like [Zhe19]. passage [BK17c, Hun18a]. Paterson [Fas19]. Path [EHHL17, EHLP18, GL17, KKL^{+}17, AHMM18, BPT15, DR19, PF18, dSN15]. paths [LN17, NR18, PS19b, YDYY18, ZLG15]. pattern [BES16, BWSZ15a, KSVW17, LP15, LT19b, OS17, SUY16, VQ17, VQ19]. patterns [BOvdD17, DB19a, EKSV14, EKSV18, FGG^{+}18, FGH^{+}15, GLZ16a, GLZ16b, GL17, Gao18, GM19, HZ15, LOvdD18, Liu18, MM17b, OvdDV17, Shi17b, Stu17]. Pauli [AMC18, RZ17, Sam17]. payoff [EW19]. Paz [GLMS18]. PDE [PCL15]. PDEs [Lom17]. peak [Gas19]. Peirce [ZMT19]. Pellet [Mel16b]. Pellet-like [Mel16b]. pencil [AHV19, Rin17]. pencils [BGT^{+}19, BR16a, BDP^{+}18b, DA19b, DA19c, DDL17, Dod16, Dod17, DHW18, dMGC19, HY18, NP15b, PD17, Sch17]. pendant [CPR18, PS19b]. pendent [NR18]. penetrable [BKLP16]. Penrose [AB18, CGMSR15, CGDM16, HWZ15, KB15, Nor18, XW16, XSW17]. Penttila [BM19a]. per-nullity [WZ15]. Per-spectral [WZ15]. Perfect [CGGV15, JKP^{+}17, TFC19, CGK^{+}17, Kir15, YG15]. Perforated [AGV18]. period [Cha19].

Periodic [Yam15, Ban16, Bre18, Fre18, GM15, SSF16, YSZ17]. Periodicity [KSTY18, MdlP16, KSTY19, Vas15]. peripheral [ZH15]. periplectic [CCW18]. permanent [CC18, GT18, Hut17, Liu17, Shc16, Shi16b, dS19a]. permanent-on-top [Shc16]. permanental [dS18a]. permanents [CP16b, CLLS19, KK17, Roo19, Sbu17]. Permutability [CI15b, CI15a]. Permutation [DF15, DF18, Ouy17, AJL16, Dah15a, TC15]. Permutation-invariant [Ouy17]. Permutation-like [DF15, DF18]. permutative [PT19]. permuting [Sim17]. Perron [AG16, ACD19, FK16, GH16a, JP16, Lia17, LT19b, Tar18]. Persymmetric [JS15, SJS15]. Perturbation [CL19, GM17, GCQX15, IGW17, SZ23, Sku17, AG17, CR17b, CGDM16, JLT17, JZ16, KR17, LWY16, MM16a, OVW18, PM15b, SK17b, cSfCfX15, Wei17, XL18, Xu17b]. Perturbations [CEM15b, XS17, BLPT15, BM19b, DFK^{+}15, VD18, DV19b, Hun16, MS15a, MRS15, SV17, WHC17, XSW17, wXjMZL19]. Perturbed [CFP18, GHT16, Rum18, RB19]. Perturbing [WLP16]. Petersen [COvdD16, KMM17]. Pfaffians [EW19]. Phase [HJ19, BZ16b]. Phase-retrievable [HJ19]. Phelps [DGK^{+}17]. phenomena [KY15]. phenomenon [Gho16, SW15]. Picard [BFN15]. piecewise [GBRS15]. pineapple [TSH16]. pivoting [PZ17b]. pivots [Kur19]. planar [VG18]. plane [BJC15, CRC17, FJ17, FJM18, LP18]. planes [ET18]. planted [Col19]. plateaux [Gho16, SW15]. player [LT19a]. pluripotential [Wan19]. plus [BR16b, BDS15, DS19b, EHN16, Hoo15, Hoo19, Li19, Mys16, RLP19, Shi16b, WT18, ZD15]. Poincaré [AL16]. point [Bai18, EP18a, FGG^{+}18, FG17, Had17, KLL16, KLW16, KLW18, Kur19, LSM16, LSM18, RR18a, WMC15]. point-hyperplane [FGG^{+}18]. points [BDG15, BDM18, CN17b, Eve18, FL16, HLM17, HMPT18, Jel18, KLW15, NAS18, Rao17, WWT18]. polar [De 16a, FX19, GMP15, GMP16b, Pan16a, Shi16a, Sol19, SG17b, dlCMP17b]. Polarity [Sol18a]. poles [PW18]. poly [RT16]. Pólya [MF16]. polyadic [DD17c]. polygons [HMNP15, VG18]. polyhedral [SPBB19]. Polynomial [KS15a, RLP19, SS16, DDMV16, VT18, Win16, YT16, AMR19, AGS17, AP15a, BY19, BBdH16, BBdH18, Ben18, BLdSV19, BR16a, BV18a, CJRMF^{+}15, CK15a, CLQW17, Che18a, CCD19, Cox15, DMQ19, FJ17, FJM18, Fio16, GK15, HWZ15, KPL18, KMS17b, Lie18, LW15, LL18b, Mel15, MMW17, NN19b, Shi16f, VV18, WW18, WZ19, YdB19, dB19b, dlP18]. polynomials [AO18, ABG16a, AMZ17, ABD19, AEV15, AM15b, BGV17, BJS18, BBP19, BB16b, BDFR15, CDR^{+}19, CW18a, Cas19, CG15, CM18, CCGO17, CCGV^{+}19, CK15b, CJ16a, CKM16, Cim15, CP17, CW16a, DA19b, DDMP19, DP18, DD17b, Dmy17, DD18, DPV19, DK16, Fag19, Fas19, FS17a, FS18a, FM19a, FK16, GGMPC15, GIM16, He16a, HK19a, HY18, HZ19, IR19, KM15a, KPL18, KMM17, Lee16, Liu17, LM17, LMR16, MO16, MP16a, MW15, MNTX16, Mel15, Mel16a, Mel16b, NS19, Niv15, Ota15, Ouy17, PM18, PP18, Ple17, Reb19, Riv15a, Riv15b, RV18, Sas18, Ser16, DDMV16, Tya17, WLWZ19, Yam15, Yan18, Zem16]. polyominoes [CEM15a]. polytope [CC19, HLM17]. polytopes [AK17, Fer15, GSS15, JK15, kJ16, NT18, SS19, War17, WH19b]. Poncelet [CN15]. Popov [RRV15]. Popoviciu [BS15d]. population [DOR16, HvdD16]. portions [JSST16]. posed [BB18]. poset [FI19b, Liu16b, Nag19, NS19, Ter18]. posets [CX17, GSZ15, Gas19]. Positive [BDRC16, BS16c, CJ16b, Gow17, GLRT18, JLNR15, SK17b, AL18, ACM17, BP16b, BS15d, BJL19, BC19a, BDS15, BL19, BDK^{+}17b, BV18b, CPZ16b, Cho17b, Chu15, CJS18, CSJ16, CLS17, DB19a, DKS15, DLB18, Drn19, EA17, FMY16, FW17, Fie15b, FG17, GS18, GdLL17, GHT16, HT18a, Hay19, HJW19, HL17a, HW15b, HK17, Jai17, JZ16, Kan16, KS17, KL15, KLL16, KQZ16, KL17, KM16, Kus16, LNW17, LPW16, LS19, Lin17a, LZ17b, LNT16c, MR17a, MM16a, MMP17, MMV19, Min17, MS15b, Mol19, MPV16, Nie15b, NR17, Pep17, PR19, Riv15b, SFW18, Sal16, Sch16, SST15, SA17, SDK17, SSV17, Sim18, Sim19, SH15b, SUY16, STZ15, Ter17b, Ulu19, Vir16, WWT18, XLQC16, YLT16, ZH17, Zha18a, Zha19a, ZF15, dS19a]. positive-definite [ACM17]. positivities [Xu16]. Positivity [SNP16, CLW15, CW19, FJS17, FJS21, LG18, MGM17, Skr17, YMZ19]. Positivstellensätze [KVV17]. possessing [SSH15]. possible [BaHOS15]. posteriori [Fis17]. potence [Stu17]. potent [SR18, Stu17]. potential [CDFK19, Fay18]. Potentials [DMS16]. potents [XFZD17]. Poupard [FHS17a]. Power [Lee16, LY17, BV18a, CJS16, CM19, DD19, DKS17, DF15, DF18, DDF17b, FuIKT16, GH16a, HNS15, JZZ18, KL15, LG18, MG18a, TCD15, UWYY15, VB19, aHRT15]. powers [BGKP19, Ege16, FJS17, FJS21, GWW19b, GWW19a, Jai17, Mag16]. PPT [Car17, Lee15a]. pre [Mos17]. pre-orders [Mos17]. precoherent [DD16b]. Preconditioned [VK16, BT17a, BA15, Miz16]. preconditioner [Bai18]. preconditioners [Hon18, LSM18]. preconditioning [AL19b, DDSC^{+}16, NSCV16, RMKJ18]. predictions [KÖ16, ÖK18]. Preface [BFH^{+}17, BDIM15, BBF19, CFH^{+}16, DLW15, FLVV18, GPW15, GHPS19]. preorders [BES19]. preprocessing [PZ17a]. prescribed [APS18, BL18b, Bal15c, DS17b, HK18b, KPL18, SJS15, WYL19, YZLC18]. presence [GLMS18]. preservation [CYD19, MR17b]. Preservers [Ben15, BBS16a, BBSS18, CLP16, EP18b, Poo15, ABM18, AMR19, AEV16, BS16b, CCCNT19, CL16, GW19a, HSS16b, LL17, LCLW18, LD16b, LD17b, NA15, Oik19, Ore16a, Ore16b, dSP19a, dS19a]. Preserving [STZ15, AAM15, AR15b, BN17, BM15, BL18c, BMGP15, Bur15, CPZ16a, CHK19, CC17b, CV17, Cos16a, Cos16b, Cos19b, Dmy17, DdC16, EJ17, EM15, GJK18, HF18, HYH15, HM16, HLS^{+}16, KP17, KM16, KLS17, Li15b, LLL17, MS16b, MS15b, MPV16, Pan17, Pan19b, Pet15, SdJY^{+}16, Vir16, Wój15, Wój19, XFZD17, ZH15, Zha18b]. Press [Bar15]. pressing [CD16c]. Primal [Fay18, RR18a]. Primal-dual [Fay18, RR18a]. primary [FI19a]. prime [DD16a, DD19, DF15, VB19]. prime-power [DD19, VB19]. Primitive [BS16b, CLN15, FGS^{+}16, HYY15, SG16, YHY15]. Principal [Bil19, LZB18, ÖK18, BC15b, BCF^{+}16, DdSJdFDV15, FOvdD16, FMR18, Gas19, HR17, HJ16, LKY16, Miy19, RKT15, Ris16]. principle [NR16a]. principles [NT19, Nie18]. priori [KL19a]. priors [RH15]. prisms [CCdFV18]. privacy [HLM17]. probabilistic [FM17, WO17]. probability [GCQX15, KL15, LZ15a, Lie18, SY16a]. problem [AO18, ATM18, AA16, ACM17, ACE15, BZ16b, BR19, Bel16, BF17a, BP19b, BC17b, BC16b, CW18a, CJ17a, CG19b, CGDM16, CL17a, CL15c, CI15a, CQ15b, DM16b, DS18b, Dod17, DE15, DN16, DN18, FPGPV17, FKM17a, FKM17b, FKMS18, Geh15, Gil16, GS18, Had17, Hil15, IR19, JP16, JMP17, JP18, JS15, Lee15c, LLH17, Ma15, MPS17, Miz16, MS16c, MK18, NMBA19, NN19b, dSP19b, dSP19c, Piv19, PS19e, Pud16, RMKJ18, RKT15, Riv15a, SdJY^{+}16, cSfCfX15, SW18b, Ste15, SS17b, Tan16b, VB18, Wei15, XL18, Yua16, ZLQ16, BDM18, ES16, VF17]. problems [Ais18a, Ais18b, BT17a, Bal15b, BP16a, BB18, BMS16a, BGCMLSSS16, BC19a, CLQW17, CFNP17, DHW18, EH18, Fay18, FKPS18, GM15, GM16a, HJ16, HPZ18, Hoo19, HZC15, HNR16, HL19a, IK18, JLD17, KU16, Kri15, LZQ19, LLL17, LDL18, Mat15b, MP18, MT15, NN19a, NS15a, NR16b, SLY17, STZ15, Van16a, WZL17, WLL19, XXZ16, wXL14, wXZ19, Yin19, YHY15, YSZ17, ZHC16, ZD15, ZZR15, ZC15]. procedure [FKRS17]. Proceedings [CFH^{+}16, GPW15, FLVV18, GHPS19]. process [CG15, CLM19a, DHLT19, Lin16a]. processes [CLR19, EHHL17, EHLP18, Hun18b]. Procrust [AA16]. Procrustes [BC19a, GS18]. Product [BS15b, HN18, LPW16, AAM15, ABM18, AGQ19, AMV17, ALH15, AL19b, BK19, BM15, CPZ16b, CF18, CJZ18, DD17a, DdC16, EH15, FZwCW16, FK18, FS19b, GW16b, HJW19, HL17b, Huh15, LK16, LTWW15, LHC15, Mag16, MSSZ17, Pet15, SWB19, Sod18, SG17a, SG17b, WW18, YLB15, ZD16b, dlC15, dlCdR17]. Products [AG18, Bot16a, CLS17, GS15, dlCG18, AJL16, BR15, BBMFM17, Ben15, BBS16a, BBSS18, Bre17, CQ15a, CHK19, CKL17, CLP16, DGSV17, FJS17, FJS21, FI16, GW16a, GW17a, GL19, GF17, GX17, Hil17a, Hil18, HLZ17, Hou19, Hou20, HSS16b, IKW16, KKA15, LL17, Lot15, MX16, PP18, XCJ18, XFZD17, Xu17a, ZH15, Zha18c]. Profiles [CKS16]. program [Li15a]. programming [BT17a, Fay18]. programs [CJ17c]. Projection [BFN15, CVV19, SXD16, BKLP16, CYD19, Col19, CFNP17, DY15, FJP15a, FG17, LNT16c, Miz16, NZZ15, NN16, Pan15b]. projection-cost [CYD19]. Projections [KM18, PR18b, And15, AG18, BB17, DSCD17, FS19a, FJP15b, GMP16a, LCNZ19, LMX19, SJD16, Sim17]. Projective [CLUP17, De 16a, BBT15, BH16a, Che15b]. Prony [KPRvdO16]. Proof [AL15, Das15, Das18, Gho16, Lin15a, TWR15, XS18, Yua16, AB15a, CS15b, Lax16, Mic16, MB18, YQS16]. Proofs [PQY15, BR17, ZZR15]. propagation [KL19a, LY16]. proper [RH15]. Properties [RGPH16, BBMP19, BBC^{+}15, BJ15, Bri17, CLR16, CJKM19b, CJKM19a, CGL16, DHW16, DvSW18, Egg15, EM15, EB17, FZL^{+}17, Fik18, GLMPC18, Hua15, Hun16, JS16, KSMB15, Kri15, Lin17a, LT19b, MB15, Mor17, PP17b, PS17, SS15a, SY16b, SY17, SZ19, SB18, SZB19, Tsa16a, iT16, UYY17, ZLKB17]. property [AAdFS19, BPP16, BPP17, CCW16, CV17, Far19, Kim18, PS18a, Tar18, WH19a, Lin16a, SS17a]. Providence [GPW15]. Pseudo [CLQW17, OZ18, PHW16, Xu17a, BBS16a, GK15, Hol18, MY19, Niv15, PR18a, Ter17a]. Pseudo-direct [Xu17a]. pseudo-Euclidean [PR18a]. Pseudo-inverses [PHW16, Niv15]. pseudo-partition [GK15]. Pseudo-skeleton [OZ18]. Pseudo-spectra [CLQW17]. pseudo-unitary [MY19]. pseudographs [XCQ18]. pseudoinverses [FG16]. pseudospectra [AGV17, AGV18, CJ17a, SKC18]. Publishing [Bru16a, Tsa16b]. PWS [RT16].

QBD [DHLT19]. Qing [Bru16a]. QR [Bur15, LLL17]. QRD [GM16b]. Quadratic [DDMP19, LZ15d, PC15, dB19b, BM17a, BT17a, BdlCL17, CM18, CLM19a, DP18, DSX15, Dym16, FJ17, FJM18, FNX19, IKR16, MP18, OZ15a, CLR19]. quadratically [Ais18a, Ais18b]. quadrature [JKP16, PCL15, RRT16]. quadrifocal [Oed17]. quandles [KL18a].

Quantity [Kim18]. Quantum

[OP16b, Vir16, BS15a, BdP18a, BN17, CLHL15, FL16, FH19, HJ19, HYH15, HK19a, HPSS19, JKP^{+}17, Kir15, KIS18, KSTY19, Lev18, LPP15, LZ16c, SGH16, SCD17, Vir19]. quartic [CR17a]. Quasi [AJL16, Hil18, CLM19a, Dod16, Dod17, EM17, FMR18, FNX19, GMM16, GMM19, KLS19b, LL15a, MM19a, PR19]. quasi-birth-and-death [CLM19a]. quasi-Cartan [MM19a, PR19]. quasi-commutative [GMM16]. quasi-commuting [GMM19]. quasi-convexity [FNX19]. quasi-double [LL15a]. quasi-Newton [EM17]. Quasi-permutation [AJL16]. quasi-regular [Dod16, Dod17]. quasi-unipotent [KLS19b]. quasiconvex [BKNS17]. Quasiseparable [BR16b, DEH16, MR17b]. quaternion [HWZ15, Yin18]. quaternionic [AR18, ABC^{+}16b, AS17, CDM19, CN19b, HM16, KMS16, Kum19, Wan19]. quaternions [FKS16]. qubit [CF18]. qudit [Ouy17]. question [Lin17b]. Questions [BCJ^{+}16]. queues [BDLM16]. quiver [Gon15]. Quotient [Fio16, SK17a, DdFR16, YYSX19]. Quotient-polynomial [Fio16]. quotients [HF18, HK19c].

R [Bar15]. Racah [LHG15, NT17, Ter17a]. radial [BBSS18, HT18b]. Radii [HR16, BFMK18, CLL15, CL17a, GW16a, GW17a, GL19, LZB18, LZM16, OQY17, TWM16, XWL17, XWD18, YSS16]. radius [AOK15, AO18, AOK19, AK19, BL18a, BL18b, BF18a, BBS16a, BBP19, BSS17, BJLD17, Bün17, CM17, Cha15, Che19b, CJ17d, Drn19, EJ17, EH15, FuIKT16, FBH19, FCL^{+}16, GH19, GW16b, GW18, Gol16, GMW18, GZ18, HSS16a, JL16a, KNY15, uIKF15, KMY15, LS15a, Lia17, LZ15c, LZ16a, LZ16b, LZW17, LZ17a, LHX18, aLwW15, LLZ17, LWCL19, LM16c, LHH17, MMM19, Mol18a, Mor17, Nik17c, NLW18, NLWJ18, Obo19a, Obo19c, PS19b, Pep17, PV17, Sab18, She17, XSW19a, XS18, XLLS18, Zam19, ZLG17, ZLKB17]. raising [Ter15]. Ramanujan [CV17, DTZ16, San15]. Randić [AdFRR18, BM16a, GRS15]. Random [OVW18, PQY15, BŚS18, BB16b, DMS16, DE15, GJLS16, HLLZ17, Hua19, LLH17, Liu18, OBRA15, PSZ15, RR18b, RJK18, Ruk15, SH15b, XL18, Yas16]. Randomized [NZZ15, BW18, BW19, BDK^{+}17b, CYD19, PZ17a]. Range [ACM15, dSP15, dSP16e, AAS18, AMC18, Bal17b, Bal18, Bal19, BS19b, CJ17a, CC17b, CCCNT19, CN15, CNY18, DFM17, FK16, HA15b, Kum19, LNT16b, LLPS19, LSZ18, Rub19, XF17]. Range-compatible [dSP15, dSP16e]. ranges [AR18, BFA18, BD15b, CDM19, CL16, CN19b, CNM19, CJK^{+}16, GWW16, GWW19b, GWW19a, HMNP15, KKP19, KLN16, Lee15c, LTWW15, SWZ18]. Rank [BBS16b, DdC16, GS16, Hua15, MV16, AAS18, AAG^{+}18, AMC18, Alf18b, ARSZ15, Al’17, Bal15c, BES19, BLPT15, BM19b, Ber17, BR16b, BS16d, BDS15, BC17a, BA15, BCU15, BWSZ15a, BCF^{+}16, CU18, CF18, CHB15, CJZ18, CCO15, Dax17, De 16b, DDL17, DM16b, DAG16, DPJ17, DD17b, DD18, DR17, EHK16, EM16, FOvdD16, FMR18, FS19a, GH17, GS19b, GdLL17, GCC18, HKM18, HHL19, Hir18, HMNP15, HSS16b, Joe16, Jør15, KKP19, KU16, LCF19, LWX17, LWZ18, LHL18b, MWT16b, Mes17, Mor17, ME15, OVW18, PQY15, Pan19b, PS18a, PRS19, PC15, dSP16a, dSP16c, dSP16d, PF18, RSS17, Rub19, Shi18c, Shi19, SZW16, SSB15a, Tok17, TW18b, VQ17, Van17, WYLW17, WF19, ZXW19, ZM16, Zui17, dB19b]. rank- [Hir18]. rank-1 [GCC18, Joe16]. rank-metric [LCF19]. rank-one [FS19a, SSB15a]. Rank-revealing [MV16]. rank/trace [BES19, Shi19]. Ranks [FGH^{+}15, BS18b, FGG^{+}18, dMGC19, Mat15b, MHN18, SP16, WLWZ19]. Rao [Rak17]. rapid [PCL15]. rate [BW18, GKS18]. rates [bCR18, JMRV16]. ratio [Bün17]. Rational [CA18, Duk12, Duk15, SSV17, AB16, AL18, AL19a, AMZ16, CMV19, CFW15, DA19a, DA19c, DMQ19, EG19, Fas19, FKRS17, JMRV16, JT18b, RV18, RAAGA16, Van16a, Vol17, WQH19, BDM18]. rationally [KVV17]. Ray [Liu15, LT19b, Liu18, Stu17]. Rayleigh [SK17a]. rays [Dah15a]. re [LC19]. re-weighted [LC19]. Reachability [MP19, PS18b]. Real [BP19a, BS16d, CW18a, DGP^{+}15, FI16, KKP19, Alf18b, AL18, BD15a, BZ16a, BKRE18, BCS15, BR16b, BS19a, BC17b, BS18b, Bün17, Bur15, CJRMF^{+}15, Che18d, CL17b, CHB15, DAG16, EE16, GM16b, JP16, LNT16a, LS15b, LDL18, MS19, MM15b, NMBA19, NN19a, RE16, RGPH16, Sei18, SSB15b, TY15]. realizability [CJS18, DDMP19, GMM19, HMPT18, JMPS19]. Realizable [AMR18, MAR17, MN19a, BPZ17, HMPT18]. realization [AL19a, BR19, JRS19, SUBG18, DDMV16]. Realizing [PT19]. receptance [WZ17a]. Recessive [SH15a]. reciprocal [BPP16, BPP17, PP17b]. reconstruction [BKLP16, BBT15, BS18c, SS16]. recoverability [WLH18]. Recovering [Col19, RJK18]. Recovery [AB16, DA19c, DA19b, DA19a, FS19a, FKRS17, KSTX15, LL16b, LS15c, RSS17]. Rectangular [MO18, BL16, CLT15, Hua15, MM16a, YZL^{+}19]. recurrence [IR19]. recurrences [ABO15, EJ19, Grc18, LMR16]. recursions [BF17a]. recursive [ATM18, CLW15, HK18a]. Recursiveness [LY16]. Redheffer [CK19]. Reduced [Pie18, CFP18, Cos16b, MS16b, TH16a]. Reducibility [CN17a, KMOR19]. reducing [BF19a]. Reduction [BGSS18, dFRS16, BK16, BŚS18, BR16b, BMR17, BF19c, CQ15b, Fay18, GP19a, HPZ18, HPZ19, JLNR15, Lom17, Mar15b, Van16b]. Refined [COvdD16, BBdH16, BBdH18, BOvdD17, Nie15a]. refinement [AL19b, JZC19]. Refinements [JM16]. reflection [BS19a]. reflective [DDSC^{+}16]. Reflexive [NT18, SR18]. reflexivity [Oi19, Pan19a]. reformulation [Ste15]. regions [Guo19]. registers [BR19, CUPW15]. Regression [Tyg19, BHML15, Eld15, Hoo19, KM18, NN16, ÖK18, RR18a]. Regression-aware [Tyg19]. Regular [CD16a, CLZ16a, LS18b, Rai19, WYL19, AAG^{+}19, Abi19, ARR18, AyLPS18, AAKS17, ÁKM17, AP15b, AM18, AB18, BK17a, BS18a, BPR17, BF15, BGGS16, CHASSSG15, CGGV15, Dod16, Dod17, lFwW16, FJKM18, GK15, GLW17, Gre18, GJLS16, GL16, HH17, IM19, Jør15, LwW16, LHG15, MMP16, MM17a, MM17c, Nik16c, O16, PB16, QKP16, Row18, Sim19, Tom15, VGG17, WQH19, ZM16]. Regularity [RS18, HKS19, KKS18]. Regularization [HNR16, CLZ^{+}16b, HKP17, NR16b]. regularized [CCO15, ZC15]. regulator [ZD15]. related [AKYD16, BGCMLSSS16, BD19, Cho17a, Cho17b, CM16, Duk12, Duk15, Hia19, HZC15, JM16, JLM18, LL17, Lin16b, LM17, LMP^{+}17, STZ15, WZ19, Zha19a, dlPJG18]. Relating [EG15]. Relation [DM16a, TW18a, BBdH16, BBdH18, CJS16, LS19, Obo19c]. Relations [DMG17, HS19b, Hu15, ÁKM17, ACRR15, BTW16, CS18b, DY15, GJK18, PR19]. relationship [BBM19, LHL18b]. Relationships [LL16a]. Relative [BD15b, MVPST19, CSJ16, FS18b, GR15, MH15]. Relaxed [DD17c, BT17a, Had17]. relevance [OW17]. reliability [MP17]. remark [FX19, Tur17, WM17b]. Remarks [Lin16c, Nik16c, BFY18, BTN^{+}18, Gol16, kJ16, JMM16, LW15, Shi17b]. removing [JST19]. rencontres [FM19a]. renewal [Hun18b]. Rényi [DHL^{+}17, Hia19]. Repetition [KLL18, DA19b]. repetitions [PD17]. representable [MFdM15]. representation [BT17b, CNY18, CN18, DAG16, HMS17, HNS15, HKP17, Kli19, Mel15, Ruk15, SZW16]. Representations [AG15b, AG20, Che18d, Che15b, CHK17, EG19, Gae16, HJ19, HJW19, JT18b, KMS17b, LP18, LH17, Ple17, Sch15, Sri17]. Representing [CK15b]. require [GLZ16a, GLZ16b, GL17, Gao18]. requiring [LOvdD18, LT19b]. rescaling [Bil19]. Residual [Lan15]. residuals [HKP17]. residue [HSTW17, Ris16]. residues [BES16, PW18]. Resistance [ABK19, CEM19, Cho19, Som17, WDV17]. resolution [CS15a, GLMS18]. resolvable [KVV17, Ma16, Mic16]. resolvent [AHV19, KLL17, Riv15a]. respect [AR18, BS16d, CFNP17, MS16a, Nom15c, Pop16b]. response [LLL17, WZL17]. Restricted [DKNS16, EF17, EF19, Ruk15, CCW16, KÖ16, YDYY18]. Restriction [Bal19]. result [Bal15b, BÓ15, CDR18, Dra16b, LZ17b, NW19, Nie19a, She16, TY15, Vij15]. Results [FJS21, AA19, Bal15a, BR15, BU18, BFMK18, CD16b, Che18b, DV18, DPR19, FHH15, GS15, KMOR19, LL17, LZH16, Lin16c, Mel16b, cSfCfX15, YHY15, Zha15b, ZM16, FJS17]. retrievable [HJ19]. retrieval [BZ16b]. reusing [Kur19]. revealing [MV16]. reverse [Kal18, PCI15, XSW17]. reversion [Bar16a]. Review [Bar15, Bru16a, Tsa16b, Zha15a]. revisited [Al’17, ELN18, KLS19a, Mac18, dSP16c, RY15, TH17, Wan16c]. Revisiting [BF17a]. reward [DHLT19]. Riccati [DP15b, AM16a, AMV17, Bar17a, BB16a, DP16, DPJ17, GL16]. Ricci [DvSW18]. Rickart [DD16b]. Ridge [Ema18, KÖ16, RR18a]. Riemannian [BC19a]. Riesz [Chu16]. right [Dra16a, HPZ18, LT12, LT16a]. right-hand [HPZ18]. rigid [BCS15, BCS16]. rigidity [Alf15]. rigorous [LWY16]. ring [Bou16b, CS18b, GLAdS19, Mig16, ZZC15]. rings [AA17, BBEE18, CHASSSG15, Che18a, CQS18, DGCC16, DGGP18, DD16b, Ere16, GM19, GL15b, GL15a, HLZ17, Hou20, HWZ15, HSTW17, KT19, LZ15d, dSP16b, QH15, RD15, Ris16, Wan15b]. Riordan [BT17b, Bar16a, Bar19, CW19, CJ15, CKS16, CJ17b, CHK17, CJB18, CS18a, CJKM19b, CJKM19a, He15, He16a, HS16, He16b, HS17, He18, KZ17, LMR16, LMM^{+}16, LMPM17, MN19b, MMW17, WZ19, YDYY18, YDHX18, Zem16]. Ritz [TLL15, Vec16]. RKFUN [EG19]. Roberts [ABR18]. Robust [CJ16d, FM17, PMB15, WLH18, CJ16c, FJMP15, LS15c, TLL15]. Robustness [VD18, DV19b, Miz16, MP15b]. Roger [Bar15]. role [BBEE19]. ROM [Hür15]. Root [MM19a, Guo19, Lia17, LG18, Miy15, Vij15]. rooted [ADW19, Sal16]. roots [AGS17, BDK^{+}17a, DKSV19, MP16c, Shi16e, TH16b, VV18]. Rosenbrock [CHASSSG15]. roses [HvD18]. rotation [BS19a, SÖ16]. rotations [DSCD17]. Rotfel’d [Zha19b]. Roth [DFKS17, FKS16]. Rough [Gal19]. roundoff [Fis17]. Row [HS16, MMW17, DV19b, Ege16, GWW16, RS17b]. row-stochastic [Ege16]. rules [Cho19, Joe16, RRT16]. Ruling [JMP17]. Runge [BGCMLSSS16, BDDSC17]. Russo [BL19].

Saad [Vec16]. Sabidussi [KM19]. saddle [Bai18, Had17, Kur19, LSM16, LSM18]. saddle-point [Bai18, Kur19, LSM18]. safe [PZ17b]. Said [MM16a]. Sakamoto [TW16]. Salpeter [SdJY^{+}16]. same [AMZ17, ZLS17]. sample [CIY15, MNT19, iT16, iT19, Yin18]. samplers [RH15]. Sampling [BKLP16, BMS15, Fre18, KL19a]. satellite [EB17]. satisfaction [MT15]. satisfy [PP17b]. scalar [GX19, Mel15, Mel16a, SG17b]. scale [BB16a, EM17, HK18a, LMS15]. scaled [VT18]. scaling [Mac16, Nat19]. scalings [CC17a, Hut17, Kus15]. scatterers [BKLP16]. scattering [GHS16]. Schatten [CL15b, CM16, JC18]. scheme [BM19a, CHJM18, Jel18]. schemes [CDRT17, CGGV15, Han19a, JCW19, Mor16b, MX16, Shi16a, WG15]. Schmidt [AK19, MHN18]. Schneider [BMS16b]. Schoenmakers [SSCS16]. Schrödinger [LM19]. Schubert [Lax16]. Schultz [BB18]. Schur [CGMSR15, CMM19, FKM17a, FKM17b, GP19a, KK17, Nie19b, PM15a, dSP19a]. Schwarz [FS19b, JM16, Kal18]. Scientific [Bru16a, Tsa16b]. scrambling [SG16]. SDD [WLL19]. SDP [CJ17c]. SDPs [BF19c]. Seak [Bru16a]. Seak-Weng [Bru16a]. secant [Han18]. Secants [MM15a]. Second [CCW18, BP16a, Bar15, BSP18, CLM19b, CFNP17, GM15, KT17, KMS16, Lin15a, MHL15, Mar15a, SY16a, SAdFZ15, Sug17, TW18a]. second-order [BP16a, GM15, SY16a, Sug17]. sectional [Opo16]. sections [Kum19]. Seidel [BSMSZ19, Riz19]. selection [Cox15]. Self [BPP17, MY19, Tya17, WHC17, AG18, GG18, GR15, Gow16, KNR18, Nag15, Nie15a, Pan19b]. Self-adjoint [MY19, WHC17, AG18, GG18, Nag15, Nie15a, Pan19b]. self-dual [GR15, Gow16]. Self-interlacing [Tya17]. Self-inverse [BPP17]. selfadjoint [And15, CMM19, FKRS17]. selfrepeat [FJ19]. Semencul [XW16]. Semi [CHB15, ÁKM17, BFA18, CJ17c, CFNP17, CLS17, FG16, FK16, GS18, GLRT18, KS17, LNT16c, Zam19]. semi-analytical [GS18]. semi-commuting [KS17]. semi-definite [CLS17, GLRT18, LNT16c]. semi-Hilbertian [BFA18, FG16, Zam19]. semi-infinite [CJ17c]. semi-monic [FK16]. Semi-nonnegative [CHB15]. semi-smooth [CFNP17]. semiclassical [GGMPC15]. semidefinite [Cho17b, FMY16, FS17b, FG17, GS18, GdLL17, HNS15, KL17, Lin17a, SUY16, Ulu19, Vir16]. semidefiniteness [KM16, NR17]. semifields [GJK18]. semigroup [Zus17a]. semigroups [AG15b, AG20, Bob16, KOST15, KMOR19, PV17, Shi17a]. semilattices [IK18]. semilinear [Van16b]. semimodule [Shi15, Tan16a]. semimodules [HM17a, Shi16d, ySpW17]. Semimonotone [TW19]. seminorms [CG19c, Gol17]. semiparametric [Ema18]. semipositive [Tsa16a]. semiring [IKR16, YZZ18]. semirings [DO16]. semisimple [CME18, PR19]. Semrl [Kim18]. sensing [BÓ15, Fli16, KKLP17]. Sensitivity [Kir15]. Seoul [CFH^{+}16]. separable [GH16b, JC18, Miz16]. Separating [KLP18]. separation [Sol18a]. Sequence [He18, BCS16, BCF^{+}16, CR19b, FOvdD16, FMR18, FHS17a, NP15c, SSS15, Yur15]. Sequences [Drn19, Bar17b, BDG15, CJ16a, CD16c, EE16, EJ19, FKM17a, FKM17b, Gar17, HMSC19, JC18, KT17, LM17, MHA18, OZ15a, PW19, PW20, Riv15b, SB18, WZ19, XWL17, ZLS17]. sequential [CR18b]. Sergeichuk [BBF19]. Series [Bru16a, Bar16a, BV18a, CIY15, CJS16, DS19c, Dya17, LG18]. Serre [CQ15b, PR19]. Serrin [ACE15]. sesqui [Bil19]. sesqui-holomorphic [Bil19]. sesquilinear [Wój15, dFRS16, dFFRS17]. Set [SS18, AvdHS15, AvdHS16, BDM18, BGT^{+}19, BGCMPB15, DDL17, DdF15, DHS18, Gór17, HYY15, Hil17b, LLPS19, LCL15, LL16a, LJL16, LD17b, MS16a, May15, NS15b, SR19, SG16, Tar18, Yor17]. Set-theoretic [SS18]. Sets [BOvdD17, Kan16, AYK17, ABM16, BM19a, BH16a, BWSZ15b, BJLD17, CR18a, bCR18, DLB18, DD17b, DO16, ELN18, Eve18, GJ18, JG17, Jin15, KL19a, Koz16, Lax16, LL16a, Lin15b, MC19, MB15, MRV15, NS15c, OL16, Pep17, Pol19, RW17, ySpW17, SXD16, Tan16a, YZL^{+}19]. setting [ABC^{+}16b, AS17]. Several [GL15b, GL15a, LCF19, CKM16, Ern18, Ken16, Naj17]. shadow [DGP^{+}15]. Shannon [Nie15b]. shape [VG18]. shaped [HP15]. shapedness [LNT16a, LNT16b]. Sharp [DR18, Ken16, BMGP15, CL17b]. Sheffer [CJ16a]. shell [ABR18, CJ17a, LSZ18]. Sherman [ACM15]. Shift [CV17, He16b, APT17, BR19, BMN16, BS16f, Chu15, CUPW15, Skr17, ZSQZ19]. shifts [KY15, KY18, LY16]. Shmuel [Tsa16b]. Shoda [SB15]. short [BP19a, Bün17, MC19, May15, Mic16, SS17b]. shorted [ACM15]. shuffled [MHA18]. SICs [CW17]. side [AL18]. sided [HAWM16]. sides [HPZ18]. Siegel [AM19]. sieve [Cox15]. Sign [Bri17, GLZ16b, LP15, LOvdD18, SS19, SV17, Stu17, BDDO16, BGGS16, BD19, BWSZ15a, BBFF17, Bün17, DB19a, FGG^{+}18, FGH^{+}15, GM15, GLZ16a, GL17, JCW19, Kob17, MNTX16, OvdDV17, Ter18]. sign-central [BDDO16]. Sign-changing [SV17, GM15]. sign-complex [Bün17]. sign-real [Bün17]. Signal [Per18, PHW16]. signals [KKLP17]. Signature [MG18a, HS19a, MWT16a, PR18a]. signatures [Ma16]. Signed [AHMM18, BZ16c, BBC18, WWT18, YYQ19, ASMN17a, ASMN17b, ABDN18, ABH^{+}19, AvdHS15, AvdHS16, AH16, Bah19, BS15c, BP15, BSS16, BSP18, CRRY15, HHL19, LWZ18, RS17a, RS17b, SS16, STW18, STW19, Sta19, WH18]. Signless [GCP18, HL19a, ACPR15, ADLR19, CR17b, CCdFV18, DdSJdFDV15, Das15, DM16a, lFwW16, FWB^{+}19, GRS17, HQX15, LD16a, LZ16b, LZW17, LLS17a, LLS17b, LLZ17, LL18a, MM19b, NLWJ18, Obo19c, TWR15, WZZ19, WZL15, XSW19a, YQS16, ZL15a, ZL15b]. similar [GS17]. similarities [Wój19]. Similarity [BP16b, BBSS18, CXLF16, CN18, CM19, CLP16, GSW16, Sha16, dlCF16]. SIMO [BPZ17]. Simple [BZ16a, Lot15, BKRE18, BGM^{+}19, BBGM16, CW15a, Cos19a, EF17, Hol16, LZ15e, MI17, MB18, PS16, RE16, Sam17]. simple-loop [BGM^{+}19, BBGM16]. simplices [FJKM18]. simplicial [BFN15, LT18]. simplification [SW15]. simplified [BM17a, BDP^{+}18b]. simulation [Hür15]. Simultaneous [BGKP19, Sha16, FT18, FKM17a, FKM17b, GX17, HK18a, HAWM16, She16]. simultaneously [DO16]. Singer [BH16a, Hou15]. single [LPPZ19]. singleton [Kan16]. Singular [CN17b, KLS19a, Lop15, YZL^{+}19, AJL16, ÁVAK^{+}19, BGT^{+}19, BMR17, GH16a, GL16, Had17, Hoo15, JS16, JKP16, KN16, Lan15, LY17, LS18b, MVPST19, Nie19c, SS15c, TY15, DDMV16, TW18b, WG15, Zhu19, Zou17]. Singularities [CDR^{+}19, Han18, JKP16]. singularity [RS18, Sch17]. singularizable [Roh18]. Sinkhorn [IW15]. six [MB15, PS19a]. size [ABGJR^{+}18, DV19a, DdF15, Gar17, JKS15, MS17b, NS18, NP15c, PS15, SWZ18]. skeleton [Fer15, OZ18]. sketching [BF19c]. Skew [BC15b, MWT16b, dlCMP17b, AAM15, AJPS16, BRN18, BGSS18, BHSS18, Che18a, CLL15, CL17a, CS15b, CS18b, CQS18, Dmy16, Dmy17, DD18, DKNS16, FOvdD16, FKRS17, FKS16, GG18, GW17b, Ito17, Joe16, LWY16, LHL18b, May15, MM15b, PM15b, WF19]. skew-adjoint [PM15b]. skew-circulant [Joe16]. skew-morphisms [DKNS16]. skew-product [AAM15]. Skew-rank [MWT16b, LHL18b]. skew-selfadjoint [FKRS17]. Skew-symmetric [BC15b, BGSS18, BHSS18, CS15b, Dmy16, Dmy17, DD18, FOvdD16, GG18, LWY16, May15, MM15b]. skewness [Hür15]. sln [Chi18]. Small [ABG16b, AAKS17, BOT19, CW16a, DHW16, Dym16, FHJT19, LM16c, QKP16, Shi15]. smallest [AB15a, AKYD16, KKY16, KLY18, KLW16, MST15, Nik16b]. Smith [BD19, EG15]. Smooth [Rao17, BGM^{+}19, BBGM16, CFNP17]. smoother [BDDSC17]. smoothness [PSG16]. snowflake [Piv19]. Sobolev [TNK16]. Society [GHPS19, Zha15a]. socle [SB15, SBB15]. Solution [BHP17, CL17a, PCL15, BP16a, BB18, Bel16, BB16a, CI15a, DI16, DTZ16, DS18b, Gór17, Lin17b, May15, RW17, SS17b, Tan16b, Yur15, ZD15, ZC15]. solutions [BV18a, CGM17, DP15b, DP16, DPJ17, DN18, FI19a, GX16, GX19, Hla15, KU16, Kri15, Li19, LM19, SFW18, SR18, SH15a, SS18, WT18, Wim16, ZHC16, Zue18]. Solvability [DIPR18a, DIPR18b, Bar17a, DFKS17, FKS16, Hla15, HR16, JLD17, LX17, vWZ17]. Solvable [KKO15, BCS15, BCS16, CLOK13, CPS17, LMO16]. solver [VT18]. solvers [LMS15]. Solving [GBRS15, BW19, EM17, HAWM16, HS19b, NZZ15, SdJY^{+}16, Van16a, ZLQ16]. Some [AA19, AN17, AAC^{+}15, AdF19, BR15, BFY18, BLY16, BU18, Cal16, CFL16, CLHL15, CD16b, CL17b, CJ15, DHHZI19, DY15, EM15, Gór17, GS15, Hoa17, HO15, HWXC18, kJ16, JMM16, KSMB15, KM16, LS18a, LL15b, LPP15, LZH16, LWC18, Lin15c, LM17, Lék16, LMM^{+}16, NR16b, Pal16, SY17, SWB19, cSfCfX15, SB18, UYY17, XCJ18, YZL15, ZLKB17, Zha18c, ZSWB16, AAS18, AAC17, BBM19, BS17, BSS16, CHJM18, DB19a, Duk12, Duk15, ET16, FZL^{+}17, FM19a, GRS15, HLW15, Jai17, Kal18, KPL18, Lan15, LHG15, LWCL19, LMR16, MQTW18, Mar15a, Mel15, Mor17, MST15, NW15, NT19, Oik19, OP16a, Ota15, PP17b, Pas19, Rob19, SY16b, YdB19, YHY15, YSS16, Zha19a]. somewhere [WH19a]. somewhere-flat [WH19a]. Soules [Hür15]. South [CFH^{+}16]. space [AM19, AL19a, BCKL17, BGCMPB15, CCW16, CGL16, CMM19, DY15, DdFR16, EE16, FJM18, HJW19, HL17a, JMPS18, Kim15a, KL15, Kul15, Lee19a, LNW17, Liu16b, Mag16, OZ15b, PSM18, Pop16a, Sab19, SG17b, WG15, Zam19]. space-filling [BGCMPB15]. Spaces [BM17c, EM16, AGQ19, BFA18, BF18a, BSZ18, Bil19, BHSS18, BDFR15, BO17, CR19a, CG19b, CDFK19, CFW15, CSW17, CG19c, DA19a, De 16a, DdFR16, EE15, EM15, ET18, FZwCW16, FS17a, FS18a, FPGPV17, FG16, FKPS18, GMP16a, Gol17, HW15a, HHR16, JC18, KL19b, KS15b, Mar18, MSSZ17, Mes17, MM17c, MRV15, Mol18b, Nag15, NP15c, Pan17, PMW19, PR18a, PCI15, dSP15, dSP16a, dSP16c, dSP16d, dSP16e, Rao17, RT16, SPBB19, Shi16c, SG17a, VF17, Vir16, WH19a, Wol18, XS17]. spanned [RMP18]. Spanning [LLT19, Gu16, KS19, MRV15]. Sparse [BHK16, KSTX15, LSM16, Alo15, BW19, BF19a, BP19a, PHW16, Shi18b]. sparser [Kli19]. sparsest [JMPS18]. sparsity [SUY16]. Specht [FHS17b, dSRST17]. Special [DLW15, SS15b, Zus17b, ABGJR^{+}18, BBF19, Ern18, LNT16a]. specified [KM15a, Nom15a, ZLG15]. Spectra [BCM17, CCdFV18, DR19, LJK16, VHB18, AAB^{+}16, AJO15, ACPR15, AB17, AH16, BW16, BR15, BS17, BCC^{+}18, CLQW17, CJS18, Cos16a, Dal17b, DS17a, DM15, GF17, JLNR15, JMP17, JRS19, JMPS19, Kan16, KHI16a, KHI16b, LXS18, LS18b, LHL18a, MAR17, MN19a, NPRS17, PT19, SY17, Shi17b, SS15c, WHC17, XS17, ZL15a, ZL15b, ZKSB17]. Spectral [ACT16, AO18, AyLPS18, ABDN18, AEV16, AL16, BL18b, BP15, BOT19, Ben16, BBC^{+}15, BS16f, CGSC19, DLM15, DHW16, FCLP18, FZL^{+}17, Fik18, Gu16, GJLS16, HT17, HMT19, JG16, JG17, LN17, LZ17a, aLwW15, MHL15, MM16b, MHS15, Pea15, SZB19, ZLG15, BL18a, Bai18, BF18a, BG15, BBS16a, BTN^{+}18, BJLD17, Bün17, BS18c, CM17, CLL15, CL17a, Che19b, CJ17d, DV19a, Drn19, DvSW18, EJ17, EH15, EH18, EB17, FuIKT16, FBH19, FT18, FCL^{+}16, GH19, GJ18, GMO19, GMW18, GZ18, HvdD16, HvD18, HJ16, HMSC19, HLLZ17, HSS16a, JL16a, KNY15, Ken16, uIKF15, KIS18, Koz16, LS15a, LZB18, Lia17, LZ15c, LZ16a, LZ16b, LZM16, LZW17, LHX18, LSX15, LLS17a, LLS17b, LZSD17, LZSD18, LWCL19, Lot15, LM16c, LHH17, Mel19, MMM19, Mol18a, MM15b, Mor17, NMBA19, Nik17a, Nik17c, NLW18, NLWJ18]. spectral [NSCV16, Obo19a, Obo19c, OQY17, PS19b, PS17, Pep17, PV17, Sim18, Skr17, TWM16, TSH16, Tsu15, WM17b, WZ15, XSW19a, XWL17, XWD18, XS18, XLLS18, Yin18, YSS16, ZLG17, ZLKB17, Zha18c, dlPJG18, CRC17, CJKM19b]. Spectrally [GM19, KSVW17, MM17b, OvdDV17]. spectratopes [JP16]. Spectrum [Chu19, AAM15, ABM18, AP15b, BDMB19, BM17b, BE15, Ben15, Ber18, BSMSZ19, BM15, BL18c, CM18, CX17, CW18b, CJ17d, Gil16, Hua19, JZZ18, KLW18, LTWW15, LD17a, LS17b, LN18, MLW15, Moh16, MK18, MP15a, Nik17b, SS15b, YYSX19, ZH15]. spectrum-preserving [BL18c]. sphere [Tan16b]. spheres [Che15a, Jia15]. spherical [FNX19]. spin [Bal15b, CD16a, Kir15]. Spiral [BK17b]. spliced [BDG15]. splines [CFM16, DGP^{+}15]. splitting [XXZ16, wXL14, wXZ19]. SPN [SM18, SMBDK16, SM16]. Spread [AA19, ACRR15, AGRR16, AdFRR18, ADLR19, CD16b, LwW16]. spreads [ALMZ^{+}19, BS16c, Van16b]. Square [Bot16b, Bot16a, Bün17, lHS16, Miy15, Shi16e, Shi16g, TH16b]. Square-zero [Bot16b, Bot16a, Shi16g]. Squared [BS16a, BNST17]. squares [CGDM16, CGM17, DS18b, Eld15, EA17, HK18a, HL19b, JKKL15, KL15, LC19, LHC15, NZZ15, UM16, War17]. Squaring [MF16]. SSD [WLL19]. Stability [CA18, DV18, FKRS17, Lee19b, SL16, SZ19, AM16a, AMV17, BMO15, Dra18, FM17, Fli16, Fre18, GX17, HM17b, Kus16, MP18, Obe16, Oik19, PM15a, PMB15, SNP16, WZL15, XW16, ZH17]. stabilization [ZH17]. stabilize [PQY15]. stabilizer [HS17]. stable [BB18, BDRC16, BS19b, GKS19, HLPS19, JCW19, Per18, SSB15a]. stage [HK18a, HvdD16]. stage-structured [HvdD16]. staggered [KIS18]. staircase [CMQV16, MGM17]. Standard [GIM16, BLdSV19, Lax16, PP18, Ser16]. Star [Row15, BMGP15, DY15, GLZ16a, LNT16a, LNT16b, PF18, Row18, WYL19, YCL17, YZLC18, aHRT15]. star-commuting [aHRT15]. star-shapedness [LNT16a, LNT16b]. State [AL19a, GHPS19, ABC^{+}16a, CGK^{+}17, CGGV15, JKP^{+}17, Kir15, TFC19, VB18]. State-space [AL19a]. states [CF18, GW19a, HYH15, KP17, LM19, Poo15, Sar19]. static [Bel16, NN19b]. Stationary [LZ15a, SY16a, VB18, Ziz16]. statistical [Lop17, Opo16, WYL15]. Statistics [DLW15, FI16, Yin18]. Staudt [KT19]. steepest [VK16]. Stein [LLH17]. Steklov [AL16]. step [NSCV16, iT16, iT19, XXZ16]. Stiefel [DSX15]. Stieltjes [CK15b, FKM17a, FKM17b, FKMS18, LM17, PS19e, Riv15a, Riv15b, Sch16]. Stieltjes-type [FKM17a, FKM17b]. stochastic [ATM18, BM16a, Ben18, CLR19, CQCZ19, Chu15, Dah15a, DLB18, DU16b, DKPU18, Ege16, GWW16, GT17, Gow17, KÖ16, NMBA19, Nat19, dCDFK17, dCDFK18]. stochastic/irreducible [DLB18]. Stockmeyer [Fas19]. strategies [KH17]. stratification [Dmy17]. strict [AGV17, AGV18, Wol18]. strictly [Bou16b, Fag19]. Strong [BPP16, BS16f, Chu15, DMQ19, AR15b, BPP17, BS16e, BDFR15, DA19a, FGS^{+}16, GLMPC18, KSMB15, SG16, Sim19, DDV16, TD15, WT18, Lin16a, SS17a]. strongly [AM18, BPR17, BU18, COvdD16, lFwW16, IM19, Jør15, LwW16, QWC15]. Structural [CYD19, GLMPC18, TD15, CJKM19a]. Structure [Dmy17, GSZ15, SdJY^{+}16, aHRT15, BBEE19, BR15, BGCMPB15, BDS15, Bur15, CC16, CILL16, CLZ^{+}16b, Der19, EHN16, FK18, Hua15, KKY16, Kob17, KRZ^{+}17, KLS17, LLL17, LL16b, MGM17, RJK18, SSCS16, SNDM17, SH15b, Tar18, Zaj19]. structure-preserving [Bur15, KLS17, LLL17]. Structured [AA16, Ano16a, BF18b, BMVW16, HKM18, HvdD16, HPZ18, MAR17, Mar15a, MK18, dSP19b, dSP19c, PM15b, SUBG18]. structures [BBMFM17, BGCMASS19, CV18, Dog18, OW17, Opo16, Pan15a]. Stud [Zha15a]. studies [PZ17a]. study [DHW18, GT17, KMS16]. Sturm [GM16a, YSZ17]. subadditive [CPZ16a]. subalgebra [BN17]. subalgebra-preserving [BN17]. subalgebras [DR17, EEV17]. subclass [ZLLL18]. Subconstituents [HZ17, ZGW16]. subdifferential [CW18b]. subdirect [AG17]. subdivision [CDRT17]. subframe [CLS18]. Subgraph [HKS19, HLSP17]. subgraphs [FdFdSDV18, KLW18]. subgroup [GH15, RB16]. subgroups [BRN18, Bie15, HS17, PB16]. subKautz [Dal17b]. submajorization [Bek16]. submatrices [HJ16, MO18, PRS19]. submatrix [DS17b]. submodules [LMX19]. subnormal [BRN18]. subnormality [KY18]. subordinate [JSST16, JT17]. subordinated [DS19b]. suborthogonal [Str18]. subpencil [SS17b]. Subresultants [BDK^{+}17a, DKSV19]. subsemimodules [ySpW17, Tan16a]. subset [ABG16b]. subsets [AJ19, DHS18, NP15c, Shi16d, Vij15]. Subspace [Zas16, BBEE19, BMS16a, BS19a, DdFR16, GR16, MMR16, NS18, Ram17, Sch16]. subspaces [AW15a, AW15b, Bal15c, Bal19, BB16a, BEKS17, CFP18, CJW19, FMM16, Huh15, MOR16a, MMR17, MMR18, NINS16, Pan15b, RMP18, SK17a, TLL15, WG15, WZL17, XS17]. substochastic [CC19, Lju15]. subtle [RWH17]. subtournaments [MS17b]. Successful [CD16c]. successive [Miz16]. such [CLST18]. Sudoku [War17]. suffice [NS15b]. Sufficient [LSX15, PM15a, Bar17a, ES16, Li15a, MPS17]. Sugeno [Mar15b]. suggested [NR16b]. Suleimanova [PT19]. sum [AMR^{+}16, AG17, AT15, BJC15, BDDO16, BNST17, Che18b, CLF18, Che19a, Cho19, DMS19, EA17, GAP16, GP18, GF17, GMP16c, Hua19, KS15a, Lee16, Sol18b, Ste18, TXZ18, WH18, dlCMP17a]. summability [BDG15]. summation [JKP16]. summing [CDMP15]. Sums [BES16, Shi16g, AO18, ABG16b, Bre18, FJS17, FJS21, FdC15, GS15, HS16, HS17, Lot15, Xu17a]. Sunder [Lin15c]. Superadditivity [UUG15]. superalgebras [CCW18, Iop18, LL19, RZ17, WZ17b]. superdiffusive [EHHL17, EHLP18]. supereigenvector [Ser15]. supereigenvectors [But16]. superinvolution [Iop18]. superinvolutions [GIM16]. supermajorization [LD17c]. Superregular [ANP16]. superstability [IT18]. superstochastic [LD17c]. supertail [NS18]. supertrees [XWL17, XWD18, YSS16]. supertropical [IKR16, Niv15, Shi16f]. support [BP19a, Hil17b]. supported [AKPS19, KAPS20, MS15a]. supporting [Pan15b]. Surjective [YL17]. survey [Dal19]. SVD [GM16b, GCC18, PS19d]. SVD-based [GCC18]. switching [BSMSZ19, ETF18, Gre18, IM19, Moh16, WQH19]. Sylvester [DIPR18a, DI16, DIPR18b, Dra16b, HAWM16, RMKJ18, Wim16, dCdSA15]. Sylvester-based [RMKJ18]. Sylvester-type [DIPR18a, DIPR18b, Wim16]. Symbolic [Sim19]. symbols [BGM^{+}19, BBGM16, Lee19a]. Symmetric [BBH16, BF19b, CR19a, SCD17, ARS17, BGS15, BGM^{+}19, BM19b, BFY18, BLdSV19, BC15b, BDK^{+}17b, BGSS18, BHSS18, BC16b, BDFR15, BO17, CP16a, CM18, CS15b, CW18b, DeV19, DP16, Dmy16, Dmy17, DD18, Dmy19, ET18, FOvdD16, FBH19, Fay18, FHH15, GG18, GKR16, GC19, HW15b, JMP17, JZ18, KH17, KLW15, KLL17, KOST15, Lee16, LWY16, LLL17, LDL18, LUC19, LS17b, LNT16c, Ma15, MO16, MPS17, MV16, May15, MM15b, Mon17, Nik17b, OS17, dSP16a, dSP16d, dSP16e, QWC15, RKT15, SR19, Sod18, Sri17, SSB15b, Tar18, VK16, XL18, dCFF17, ES16]. symmetrically [BF19a]. symmetries [BFW17, FSSW17, FSW19, GS15, LCNZ19, dlCMP15, dlCdR17]. symmetrization [Rad17]. symmetrized [HMSC19]. symmetrizers [DU16a]. Symmetry [PMW19, AH16, FK18, TC15, Tur17, dSRST17]. Symplectic [BHSS18, AM19, BF19b, BR17, CG16, DR17, DH19, EH18, HL18, IGW17, RMP18, SH15a, SZ23, WG15, dlC15, dlCF16, dlCMP17a, dlCdR17, dlCG18]. symplectic/orthogonal [BF19b]. synaptic [FJP15a, FJP15b]. synchronous [ZZR15]. synthesis [BZ16b]. system [BTN^{+}18, BB16b, GKS19, Hoo19, WZ17a]. systems [AAdFS19, BW19, BPZ17, Bel16, BTN^{+}18, CC16, CHASSSG15, CDG16, CNX17, CR18b, CL15c, CJ16c, CJ16d, DGCC16, DGGP18, DTZ16, DFKS17, DS19c, DOR16, FM17, FKRS17, FHS17b, Gór17, Gre18, GBRS15, GHT16, HK18a, Hla15, HR16, Hon18, Hoo19, HM17b, KN16, KVP19, Kur19, LM16a, Lee19b, LX17, Li19, LS18b, LNT16c, MM19a, Nie18, Nie19a, OP15, OP16b, Pas18, RRV15, RW17, SFW18, SR18, SL16, SH15a, Sug17, SS15c, Vas15, VK16, Vij15, WT18, ZH17, Zhu19, dFFRS17]. Szego [Fik18, FM19b, MF16].

T [MM17a]. T-modules [MM17a]. tableaux [SS19]. Takagi [Mar15b]. Tangent [Kut18]. Tarski [DP15a]. Taylor [DR16, JLM18, SZ19]. TD [Nom15b]. technique [Row18]. techniques [BK16, KS19, RW17]. technology [SSCS16]. Ten [CRS15]. Tensor [CJZ18, GH16a, KKA15, BL18b, BJS18, Bre17, CPZ16b, CLQW17, CF18, CKL17, DH15, DdC16, FuIKT16, GL19, GZB19, dMGC19, GCC18, Hil17a, Hu15, HSS16b, Kut18, LL17, LLV15, LL16c, RSS17, Sar19, Sch15, Shi18c, Sod18, Van17, WDFS17, XCJ18, XFZD17]. Tensor-tensor [KKA15]. Tensors [ARSZ15, YLB15, BS15d, BWSZ15b, BWSZ15a, BJLD17, CLQW17, CQCZ19, CW15b, CW18b, CLN15, DLB18, DLQ18, DD17c, FBH19, FGS19, GH16a, HYY15, HWXC18, JZZ18, KSMB15, KLN16, KL18b, LS15b, LL15a, LQL15, LCL15, LL16a, LJL16, LUC19, MS17a, NR16a, OP19a, QWC15, Sei18, SY16b, SWB19, Tsu15, XLQC16, Xu16, XCQ18, YZL^{+}19, YYQ19, YHY15, ZSWB16, dSRST17]. term [De 16b, dMGC19]. Terminating [Kwo16]. terms [DGCC16, lFwW16, HHL19, LWZ18, MWT16b, TWM16, WM17a, WF19, WWC18, ZWS18]. ternary [CN17a, De 16b]. Terwilliger [GGH15, GK15, MMP16, Mor16b, MX16]. tessellable [HPSS19, KIS18]. Test [CILL16]. th [Guo19, LG18, YG15]. their [ABP15, AJ19, Bal15c, BKNS17, Bar19, BCC^{+}18, BSMSZ19, BDS15, BM17c, BC15b, CR19a, CMV16, CR15, ÇMP18, CIY15, CHK17, CJB18, DGGP18, DLQ18, DdFR16, DHW18, FGH^{+}15, GL15b, HT18a, He16b, HS17, He18, HC19, KC16, KHI16a, KHI16b, LTWW15, Lin19, MW15, MLW15, Mas17, MO18, Mor16b, Niv15, PSZ15, PS17, PHW16, RFPS18, Riv15b, RGPH16, SKC18, SCD17, Ter17a, Ulu19, WYL15, Xu16, Yan18]. Theorem [MB18, Al’17, BW16, BM16a, BS18a, Bor17, BBS16b, CHASSSG15, CQCZ19, CS15b, DGK^{+}17, DDCY17, DP15a, DD16a, DD19, DR16, DPS19, FS17b, GH16a, GKR16, GLW17, HO15, lHS16, Hua19, JZZ18, KT19, Kul15, LN17, LMX19, Mar15a, McK18, NS15a, Pas18, dSP16b, RY15, SB15, TW16, Uch15, VB19, Yin18, BEF16, BL19, JP17, KM19, KLS19b, MG18b, MM19b, Rao18]. theorems [AS17, BJC15, BPT15, CFL16, CLLS19, EH18, FK16, GL15b, GL15a, HLSP17, Jen17, MRS18, aHRT15]. theoretic [AB15a, FJMP15, GR15, SS18]. theoretical [BBT15]. theory [BBEE19, CDG16, CGDM16, CLQW17, Che15b, DH19, DE15, Gar17, HMS17, JS18b, KKY16, KNR18, MM16a, Pea15, PM15b, RD15, Tsu15, Wan19, Zha15a]. There [AM18]. theta [OP15]. third [DD17c, Han18, LUC19, Lop15, Obo16]. third-order [DD17c]. Thompson [CS17, LZ17b]. thoughts [HO15]. Three [May15, QWC15, Shi17b, BW16, CMV17, CF18, DK16, EHK16, Far19, GW18, HK18a, IKW16, JKKL15, KRZ^{+}17, Nom15c, Pal16, Row16b, SY16b, SWZ18, TW17, iT19, YT16, dlCMP17a]. three-dimensional [CMV17]. three-level [KRZ^{+}17]. three-qubit [CF18]. three-stage [HK18a]. three-step [iT19]. Threshold [LSM18, AG15a, BM17b, Dah15b, Gho19, HT17, JTT15, LMT19, LHL18a]. thresholding [FS19a, RSS17]. Tight [CFW15, BGOY15, BFG^{+}16, BF17b, FJKM18, HL17c, LC19, LwW16, Mol18a, RNC17]. Tighter [DV19a]. Tikhonov [NR16b]. tiling [NS15a]. Time [UZ19, BMO15, BTN^{+}18, BK17c, CG15, CIY15, DHLT19, GKS19, HM17b, JMPS18, KL19a, Mas17, Obe16, PCL15, PMB15, QWW18, SL16, SSF16, WZ17a, ZD15, ZH17]. time-and-band [CG15]. time-delay [BTN^{+}18]. time-dependent [DHLT19, PCL15]. Time-varying [UZ19, BMO15, HM17b, Obe16, QWW18]. times [Hun18a]. Tingley [FPGPV17, Tan16b]. TLS [HPZ18, HPZ19]. TN [JZ16]. Toda [Ong18, Zue18]. Toeplitz [PW20, AAS18, AdF19, APT17, BGS15, BGM^{+}19, BDLM16, BBGM16, CI15a, CI15b, DTZ16, DS19b, DE15, EJ18, EW19, Hon18, HMSC19, KNR18, Lee19a, Nob16, NK18, NSCV16, PSZ15, PW19, SSCS16, Sch17, TNK16, TS18, XW16]. Toeplitz-block [NK18]. toggling [BH16b]. top [Shc16]. topical [HM17a]. topics [STZ15]. Topological [Gar17, dFRS16, dFFRS17, Shi18a, BGCMPB15]. topologies [Ost15]. topology [ABM16]. tori [LZ16c]. Toric [YMZ19]. Toric-Bernstein [YMZ19]. Total [AG16, CLW15, DN18, FJS17, FJS21, GW19b, YMZ19, AG17, ALMZ^{+}19, CW19, DHHZI19, DS18b, MGM17]. Totally [BV18b, AG17, BP16b, CU18, CSJ16, JZ16, MM16a, MMP17, MMV19, WHG17]. tournament [GW17b]. tournaments [Ito17]. TP [JZ16]. Trace [HLS^{+}16, KL17, SBB15, AdlPR16, ARR18, BG15, BES19, Ben18, CFL16, Cho17a, Cho17b, CW15b, FPGPV17, Gir19, Hia16a, HL19b, LL15b, Min17, NA17, Ota15, Shi19, SW18a, dlP18]. traceable [LSX15]. traceless [GG18, Nag15]. traces [Sch15, Wan16c]. traffic [BCMASS15]. train [Kut18]. transfer [ABC^{+}16a, bCR18, CDRT17, CGK^{+}17, CGGV15, JKP^{+}17, Kir15, TFC19]. transform [BLdS16, CM17, CK15b, KP19, SLY17]. transformation [CUPW15, GR15, MMR16, TXZ18, dlCF16]. Transformations [MS15b, Pan15a, AW15b, BS18c, CMV19, DGSV17, DEH16, DT15, HM16, LUC19, LZ16b, Nom15c, Pan17, Pan19b, Ram17]. transforms [EP18b, Far16, KKA15]. transition [WC15]. transitive [MS17b]. transplantation [Her16]. transportation [GSS15, War17]. transpose [Cho17a]. Tree [MM19b, SW15, ADW19, AvdHS16, BS15b, BS16a, Ber17, BPT15, Che19b, CP16b, CSJ16, Gon15, GCQX15, JSST16, JLDS18, KS19, Liu15, LLT19, NS15c, Pud16, ZD16a, ZD16b]. Trees [BDVRT15, AdlPR16, ACD19, BKNS17, BOT19, BCJ^{+}16, CKS16, DMS16, GS19a, GLMPC18, GP19a, Gu16, HW15c, JS19, JL16a, KSTY18, LD17a, MST15, NS17, Nag19, NS19, NPRS17, YG15, Yua16, ZL15a, ZL15b]. Trench [XW16]. trend [DLM15]. triangle [CPZ16a, LL15b, Row16a]. triangle-free [Row16a]. triangles [Bar16a, Lin15a]. Triangular [EJ19, BJ16, Bie15, BDLM16, BFD18, Bou16b, Cha15, CXLF16, CKT19, DIR16, DTZ16, DHW18, Ere16, Fag19, KKO15, KY17, MV16, OP16a, SP16, SY16b, UG18, Wan15b, Wan15a, Wan16c, WMC15, WHC17, Yan16, YWD16]. triangularization [She16]. tridiagonal [AdF19, APT17, BGV17, BF18b, Chu19, DM18, DE15, EJ18, HJ16, kJ16, Nom15c, TS18, VHB18]. Trigonometry [AGQ19, CRC17]. Triple [ABP15, Ben15, BCP16, CC16, EP18a, KVP19, Nom15c, Nom16, Wan16c, XSW17, YL17]. triples [AC15, BCP16, FN17, HLG15, LHG15, Nie19b, Ter15, Ter17a, WHG17]. Tripotency [Kis15, XX12]. tripotent [Kis15, XX12]. Tropical [AGN18, GSS15, Tsu15, AGS17, But16, GS16, JK15, JL16b, Kri15, LdlP15, MT16, PS16, Shi16c, TS18, YZZ18]. Tropicalization [AK17]. tropically [VT18].

Trotter [HK17]. truncated [FKM17a, FKM17b, FKMS18, HZC15, Riv15a]. truncation [BDRC16]. truncations [Mas17]. Tsetlin [FH19]. TU [XvdBvdLS15]. Tucker [CS15b, DH15]. tuples [PS19c]. turns [EV15]. Twenty [GHPS19]. Twenty-first [GHPS19]. Twisted [AGP^{+}17, CH19, TZ19]. Two [FJP15b, HAWM16, HLSP17, LT19a, Obe16, QWW18, TNK16, XXZ16, Alf18b, ARSZ15, AL19b, BG15, BGOY15, BBH16, CVV19, bCR18, CF18, CJS18, DS17a, DF18, DdFR16, DSCD17, DR17, DHW18, FKPS18, GSZ15, GSZ16, Geh15, Gil16, GMP16c, HKLQ16, HS19a, HN18, JKKL15, KS17, Kis15, LS17a, LPW16, Lin16c, LDZ18, LMX19, MM17a, MP18, PC15, Sim17, SZ17, iT16, WWT18, WH19a, XWD18, XX12, XvdBvdLS15, ZHC16, dB19b, EHLP18]. two-by-two [AL19b]. two-dimensional [FKPS18, WH19a, EHLP18]. two-distance [BGOY15]. two-level [MP18]. Two-player [LT19a]. Two-sided [HAWM16]. Two-step [XXZ16, iT16]. type [AM15a, AS17, AM15b, BBEE18, BKNS18, BJC15, BS18a, BLdS16, Bob16, BFMK18, BWSZ15b, BJLD17, CJ16a, CK19, De 16a, DIPR18a, DIPR18b, DLB18, FKM17a, FKM17b, GSZ15, Gas19, GXYZ17, HRM18, HW15b, HLG15, Hua15, HL19a, HZ17, IKVZ18, IR19, Jen17, KT19, KNR18, LL16a, LZH16, LJL16, LX17, LCWZ19, LHG15, Mar15a, MR19, MF16, Nie15a, NT17, OP15, OZ15b, PR19, PPZ19, RB16, Riv15b, Roo19, Saw16, Sol19, TZK17, Ter17a, WG15, WHG17, Wim16, Wol18, XW16, ZLLL18]. types [EEV17, SY16b, Sim18, TNK16]. typical [BS18b].

Ulam [WH19a]. Ultra [Shi16b]. ultrafriable [San15]. unattainable [BDM18]. unbalanced [KL18b]. Unbiased [iT19, KRZ^{+}17, MB15, Szá16]. Unbounded [MNT19]. uncertain [CJ16d]. uncertainty [CJ16c]. unconditional [BCKL17]. unconstrained [CDTZ15]. uncountable [Hol16]. underlying [Der19, LWZ18, LHL18b]. unextendible [Szá16]. unfoldings [WDFS17]. unicyclic [ABH^{+}19, BPP17, GRS17, LDZ18, MHL15, OQY17]. Unification [Zou19]. unified [ACPR15, JLT17, JS18a, Li15a, RD15, YDHX18]. Uniform [Lie18, RFPS19, CLS18, Cus17, EM16, GH17, GZ18, HQX15, KNY15, uIKF15, LZB18, LZ16a, LZM16, LZW17, LZ17a, LKY16, PRS19, SY17, XWL17, XWD18, YYQ19, YZL15, YSS16, YQS16, ZKSB17]. uniformly [Liu18, Man19]. Unifying [LZ17b]. unimodular [DS17b]. Unimodularity [MB18]. union [CHJM18, WM17b]. Unions [BES19, LS17a]. unipotent [KLS19b]. unique [KL18b, LOvdD18]. Uniquely [Sma15, Sku17]. Uniqueness [DI16, Lee15c, Bal15b, DIPR18a, DIPR18b, DD17c, SZ19, ZC15]. unit [Cas19, IR19, LWX17, MS15a, Tan16b]. unital [AHV19, CW16a]. Unitaries [Sim17]. Unitarily [KMS17a, JLM18, MRS18, Sab15, WL15, Zha19b, Zou19]. Unitary [CN18, Jin15, BBSS18, BN17, CCiT18, CN17a, CN17b, CS18b, CLP16, DDCY17, DD16a, DD19, FR15, FR18, HT18a, HC19, IW15, KM15b, KIS18, Lau18, LZ15d, MHN18, MY19, TC15, VB19]. unitriangular [VLGS15]. unity [GM19]. univariate [DKSV19]. Universal [CJS18, BY19, Bou16a, JMPS19, NT17]. Universality [CGK^{+}17]. University [Bar15, GHPS19]. unsolvability [HR16]. update [PS19d]. updated [BCU15]. Upper [ARR18, AGRR16, bCR18, GRS15, LZM16, MS17b, WF19, Yan16, AM19, BG15, BFD18, Bou16b, CXLF16, CLF18, DR18, DHW18, Ere16, Fag19, KY17, Mol18a, SP16, UG18, Wan15a, WMC15, WHC17, Xu17b, YZL15, ZWS18]. Using [Lin16a, BB16a, CMV19, CKM16, DU16a, HAWM16, JKP^{+}17, Ken16, Ong18, Pan15b, Saw16, VV18, WZ17a, YDHX18]. usual [CHB15].

valency [QKP16]. valuations [ABG16a]. value [ÁVAK^{+}19, GR15, Lop15, MVPST19, Piv19, YZL^{+}19]. valued [FKM17b, FS19b, HJ19, KKLP17, Lee19a, Zam16]. values [AdlPR16, AG15b, AG20, BBH16, BMR17, BTN^{+}18, GT18, Hoo15, JSST16, Lan15, Nie19c, TY15, TW18b, XvdBvdLS15, Zou17]. Vandermonde [BLdS16, MM16a, MMP17, MMV19, Xu16]. Vandermonde-type [BLdS16]. variable [BDDSC17, KY15, KY18, LY16]. variable-coefficient [BDDSC17]. variables [ACS19, BLdSV19, BP19b, Naj17]. variance [BFFN16]. variants [BS17, Far19, Lin19]. variate [Mat15a, MP17]. variation [MHS15]. variations [HW16]. varieties [BF19b, BS16d, CVV19, Gal17, Han18, Kut18]. variety [FI19b, Gal19, MT16, Oed17, Rin17]. Various [HK18b, HS19b, MR17a]. varying [BMO15, HM17b, Obe16, QWW18, Sug17, UZ19]. Vector [Gal17, BP19b, BDFR15, CFW15, CG19c, DdFR16, EM16, FS17a, GP19a, Gol17, HW15a, HHR16, LH17, Mag16, OBRA15, Pan17]. Vectorial [ÇMP18]. vectors [BJC15, BP19a, FR15, GH16a, LZ15a, RMP18, SY16a, Vec16]. Verified [Miy19]. Veronese [Gal19, Han18]. Vershik [GH15]. version [lHS16]. versions [AL19b, KMOR19, RWH17]. versus [GLMPC18]. Vertex [HXL19, AHJR18, ALOR19, CEGM16, Fer15, OL16, RMS16, TWM16, WM17a]. Vertex-connectivity [HXL19]. vertical [CJB18]. vertices [CPR18, FdC18, FJ19, QKP16]. via [BS18a, BA15, BDP^{+}18b, CM17, bCR18, CRRY15, CK15b, CS18a, CYD19, CP17, Col19, CLZ^{+}16b, DKS15, DEH16, FMR19, FS19b, GX17, HJW19, Hun16, IM19, LHC15, MAR17, MV16, MN19a, Mes17, NSCV16, OP16b, PT19, RSS17, Sab18, SSCS16, San15, UZ19, ZSWB16]. view [JP17, Riv15a]. views [BBT15]. Vladimir [BBF19]. vol [Bru16a, Zha15a]. volume [MO18]. volumes [Bil19]. Vong [Bru16a]. Voronoi [GMLdS16, GMLT19]. VSPs [Jel18].

Walk

[KKS18, QKP16, HKS19, KIS18, KSTY19]. walk-regularity [HKS19, KKS18]. walks [AGHN18, CRRY15, DMS16, HPSS19, KSTY18, LT18]. wandering [BEKS17]. Wang [Bah19]. Waring [CCGO17, Tok17]. Waring-like [CCGO17]. Wasserstein [BJL19, HL17a]. Wave [Far16, BS18a]. waveguide [RMKJ18]. wavelet [HL17c]. Weak [BMO15, BCP16, CP16a, Lju15, LD17c, NP15a, CRRY15, HHR16, LD16b, LD17b, MP15b, Nie19b, Obe16]. Weak- [CP16a, NP15a]. Weak-local [BCP16]. weakly [BmBCC17, FBH19, MC19, VK19]. Wedderburn [Mar15a, SB15]. Wedderburn-type [Mar15a]. wedge [HP15]. wedge-shaped [HP15]. weight [Abi19, GM15, LZ16c, ZHC16]. weight-regular [Abi19]. Weighted [ABvN19, BMN16, CGM17, JL16b, Mos17, BPP16, BH16a, DDF17b, DPS19, DL16, HS17, KY15, KY18, KS19, KMS16, KMMS19, LK16, LY16, Lee19a, LS19, LC19, Mac16, MMS18, PP17a, RT16, TNK16, UYY17, WLWZ19]. weights [ABK19, BS15b, SSS15, TZK17, Yus17, ZD16a, ZD16b]. well [Bar16b, Zho17]. well-conditioned [Bar16b]. well-defined [Zho17]. Weng [Bru16a]. Weyl [FKRS17]. Weyr [FMM16, OW17]. Where [Beh17, BF15, CR19a]. Which [ZLS17, BMN16, EP18a, JDS17, LwW16, PP17b]. Whitehead [BBMFM17]. whose [BS19a, DTZ16, HL19b, JZ18, JLDS18, MHL15, MWT16a, NS15c, OS17, SUY16, Ter18, TW17, WWT18]. width [CNM19, FHJT19]. Wielandt [ABR18, CJ17a, GH16a, LSZ18]. Wiener [BBdH16, BBdH18, JS18b]. Wigner [IT18, Ngu18]. Wildness [BP19b, FKPS18, FGS19]. Williamson [SZ23, IGW17]. Williford [BM19a]. windmill [Koo19]. Wirtinger [LZQ19]. Wirtinger-Northcott [LZQ19]. wise [AG17]. Wishart [ME15]. without [Eve18, GH19, SUY16, WZZ19]. witnesses [JP18]. Wold [BBEE18]. Wold-type [BBEE18]. Wong [NP15b]. Woodbury [ACM15]. work [Pan15a]. World [Bru16a, Tsa16b]. Wreath [DD17a, MX16, Xu17a]. ‘Wrong’ [AL18]. WST [BC19b]. WST-decomposition [BC19b]. Wu [CDR^{+}19, Lee15b]. Wuwen [AB15b].

Xiao [Bru16a]. Xiao-Qing [Bru16a]. Xingzhi [Zha15a].

Yakubovich [RRV15]. Yamazaki [DDF17a]. Yang [SS18]. years [CRS15]. yielding [Alf18a]. yields [DTZ16]. Young [MN15, SS19].

Zero [ALH15, BDDO16, Gao18, Lin19, TD15, AJPS16, BES16, Ben18, BOvdD17, Bot16a, Bot16b, CW18a, EJ17, FMY16, FWB^{+}19, FGG^{+}18, HLPS19, Hil17b, HL19b, KL19a, Lin16a, MM17b, MS16c, Nom15b, PM15b, Ram16, Shi16g, Shi17b, SW18a, Wan15a, WMC15]. zero-diagonal [Nom15b]. zerodivisor [Wan15a]. Zero-nonzero [Gao18, BOvdD17, FGG^{+}18, MM17b, Shi17b]. zeros [AO18, BBP19, Cas19, CMQV16, HK19a, PW18]. zeta [DM15, KMS16, KMMS19, Som17, TH16a]. Zhan [Zha15a]. Zhang [QWW18, UZ19]. zigzag [DDMV16]. Zusmanovich [LCWZ19, ZLLL18].

References

Adhikari:2016:SPP

[AA16] Bibhas Adhikari and Rafikul Alam. Structured Procrust’s problem. Linear Algebra and its Applications, 490(??): 145–161, February 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0024379515006217.

Ayupov:2017:LDA

[AA17] Shavkat Ayupov and Farhodjon Arzikulov. 2-local derivations on associative and Jordan matrix rings over commutative rings. Linear Algebra and its Applications, 522(??):28–50, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0024379517300885.

Afshari:2019:SRL

[AA19] B. Afshari and S. Akbari. Some results on the Laplacian Spread Conjecture. Linear Algebra and its Applications, 574(??):22– 29, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                                      /www.sciencedirect.com/
                                                      science/article/pii/S0024379519300977.

Aalipour:2016:DSG

[AAB^{+}16] Ghodratollah Aalipour, Aida Abiad, Zhanar Berikkyzy, Jay Cummings, Jessica De Silva, Wei Gao, Kristin Heysse, Leslie Hogben, Franklin H. J. Kenter, Jephian C.-H. Lin, and Michael Tait. On the distance spectra of graphs. Linear Algebra and its Applications, 497(??):66–87, May 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516001208.

Andelic:2015:SNC

[AAC^{+}15] M. Andelić, E. Andrade, D. M. Cardoso, C. M. da Fonseca, S. K. Simić, and D. V. Tosić. Some new considerations about double nested graphs. Linear Algebra and its Applications, 483(??):323–341, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003638.

Antezana:2017:SFO

[AAC17] Jorge Antezana, M. Laura Arias, and Gustavo Corach. On some factorizations of operators. Linear Algebra and its Applications, 515(??):226–245, February 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516304827.

Alazemi:2019:LPS

[AAdFS19] Abdullah Alazemi, Milica Andelić, Carlos M. da Fonseca, and Vladimir V. Sergeichuk. Lipschitz property for systems of linear mappings and bilinear forms. Linear Algebra and its Applications, 573(??):26–36, July 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930117X.

Adm:2018:FAC

[AAG^{+}18] Mohammad Adm, Khawla Al Muhtaseb, Ayed Abedel Ghani, Shaun Fallat, and Jürgen Garloff. Further applications of the Cauchon algorithm to rank determination and bidiagonal factorization. Linear Algebra and its Applications, 545 (??):240–255, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300491.

Aashtab:2019:MER

[AAG^{+}19] A. Aashtab, S. Akbari, E. Ghasemian, A. H. Ghodrati, M. A. Hosseinzadeh, and F. Koorepazan-Moftakhar. On the minimum energy of regular graphs. Linear Algebra and its Applications, 581(??): 51–71, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302824.

Alazemi:2017:DRG

[AAKS17] Abdullah Alazemi, Milica Andelić, Tamara Koledin, and Zoran Stanić. Distanceregular graphs with small number of distinct distance eigenvalues. Linear Algebra and its Applications, 531(??):83–97, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303336.

Abdelali:2015:MPL

[AAM15] Z. Abdelali, A. Achchi, and R. Marzouki. Maps preserving the local spectrum of skew-product of operators. Linear Algebra and its Applications, 485(??): 58–71, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004334.

Afshari:2018:ACG

[AAMM18] B. Afshari, S. Akbari, M. J. Moghaddamzadeh, and B. Mohar. The algebraic connectivity of a graph and its complement. Linear Algebra and its Applications, 555(??):157–162, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302994.

Alazemi:2016:ELC

[AAS16] Abdullah Alazemi, Milica Andelić, and Slobodan K. Simić. Eigenvalue location for chain graphs. Linear Algebra and its Applications, 505(??):194–210, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630129X.

Adam:2018:EHR

[AAS18] Maria Adam, Aikaterini Aretaki, and Ilya M. Spitkovsky. Elliptical higher rank numerical range of some Toeplitz matrices. Linear Algebra and its Applications, 549 (??):256–275, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301344.

Altinisik:2015:PCM

[AB15a] Ercan Altinisik and Serife Büyükköse. A proof of a conjecture on monotonic behavior of the smallest and the largest eigenvalues of a number theoretic matrix. Linear Algebra and its Applications, 471 (??):141–149, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008234.

Ambrozy:2015:CGW

[AB15b] Zofia Ambrozy and Iwo Biborski. On continuity of Guo Wuwen function. Linear Algebra and its Applications, 471(??): 685–692, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000208.

Alam:2016:RER

[AB16] Rafikul Alam and Namita Behera. Recovery of eigenvectors of rational matrix functions from Fiedlerlike linearizations. Linear Algebra and its Applications, 510(??):373–394, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303986.

Araujo:2017:SAG

[AB17] José O. Araujo and Tim Bratten. The spectra of arrangement graphs. Linear Algebra and its Applications, 530(??):461–469, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303324.

Azimi:2018:MPI

[AB18] A. Azimi and R. B. Bapat. Moore–Penrose inverse of the incidence matrix of a distance regular graph. Linear Algebra and its Applications, 551(??): 92–103, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301812.

Abadias:2019:GOE

[AB19] Luciano Abadias and Antonio Bonilla. Growth orders and ergodicity for absolutely Cesàro bounded operators. Linear Algebra and its Applications, 561(??): 253–267, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304762.

Ackelsberg:2016:LST

[ABC^{+}16a] Ethan Ackelsberg, Zachary Brehm, Ada Chan, Joshua Mundinger, and Christino Tamon. Laplacian state transfer in coronas. Linear Algebra and its Applications, 506(??):154–167, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630180X.

Alpay:2016:BCQ

[ABC^{+}16b] Daniel Alpay, Marek Bozejko, Fabrizio Colombo, David P. Kimsey, and Irene Sabadini. Boolean convolution in the quaternionic setting. Linear Algebra and its Applications, 506(??): 382–412, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630221X.

Anguas:2019:CEC

[ABD19] Luis Miguel Anguas, María Isabel Bueno, and Froilán M. Dopico. A comparison of eigenvalue condition numbers for matrix polynomials. Linear Algebra and its Applications, 564(??): 170–200, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305548.

Akbari:2018:SCS

[ABDN18] Saieed Akbari, Francesco Belardo, Ebrahim Dodongeh, and Mohammad Ali Nematollahi. Spectral characterizations of signed cycles. Linear Algebra and its Applications, 553(??):307– 327, September 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830243X.

Akian:2016:NAV

[ABG16a] Marianne Akian, Ravindra Bapat, and Stéphane Gaubert. Non-Archimedean valuations of eigenvalues of matrix polynomials. Linear Algebra and its Applications, 498(??):592–627, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001646.

Ambrus:2016:SSS

[ABG16b] Gergely Ambrus, Imre Bárány, and Victor Grinberg. Small subset sums. Linear Algebra and its Applications, 499(??):66–78, June 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001580.

Aron:2018:SSF

[ABGJR^{+}18] R. M. Aron, L. Bernal-González, P. Jiménez-Rodríguez, G. A. Muñoz-Fernández, and J. B. Seoane-Sepúlveda. On the size of special families of linear operators. Linear Algebra and its Applications, 544(??):186– 205, May 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300120.

Akbari:2019:LES

[ABH^{+}19] Saieed Akbari, Francesco Belardo, Farideh Heydari, Mohammad Maghasedi, and Mona Souri. On the largest eigenvalue of signed unicyclic graphs. Linear Algebra and its Applications, 581(??):145–162, November 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951930268X.

Abiad:2019:CAW

[Abi19] Aida Abiad. A characterization and an application of weight-regular partitions of graphs. Linear Algebra and its Applications, 569 (??):162–174, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300278.

Atik:2019:RMG

[ABK19] Fouzul Atik, R. B. Bapat, and M. Rajesh Kannan. Resistance matrices of graphs with matrix weights. Linear Algebra and its Applications, 571 (??):41–57, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300667.

Augat:2016:CSF

[ABM16] M. Augat, S. Balasubramanian, and Scott McCullough. Compact sets in the free topology. Linear Algebra and its Applications, 506 (??):6–9, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301604.

Abdelali:2018:LPL

[ABM18] Z. Abdelali, A. Bourhim, and M. Mabrouk. Lie product and local spectrum preservers. Linear Algebra and its Applications, 553(??):328–361, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302441.

Alves:2015:KDI

[ABO15] João F. Alves, António Bravo, and Henrique M. Oliveira. Kneading determinants of infinite order linear recurrences. Linear Algebra and its Applications, 475 (??):306–325, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001421.

Alavi:2015:TFG

[ABP15] Seyed Hassan Alavi, John Bamberg, and Cheryl E. Praeger. Triple factorisations of the general linear group and their associated geometries. Linear Algebra and its Applications, 469(??):169–203, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007526.

Arambasic:2018:ROD

[ABR18] Ljiljana Arambasić, Tomislav Berić, and Rajna Rajić. Roberts orthogonality and Davis–Wielandt shell. Linear Algebra and its Applications, 539(??):1–13, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517306122.

Arashi:2019:WDE

[ABvN19] M. Arashi, A. Bekker, and J. van Niekerk. Weighted distributions of eigenvalues. Linear Algebra and its Applications, 561(??): 24–40, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304609.

Alnajjar:2015:LTE

[AC15] Hasan Alnajjar and Brian Curtin. Leonard triples from the equitable basis of sl_{2}. Linear Algebra and its Applications, 482(??): 47–54, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003274.

Andrade:2019:CPP

[ACD19] Enide Andrade, Lorenzo Ciardo, and Geir Dahl. Combinatorial Perron parameters for trees. Linear Algebra and its Applications, 566(??):138–166, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518305998.

Aruliah:2015:BMH

[ACDT^{+}15] D. A. Aruliah, R. M. Corless, G. M. Diaz-Toca, L. Gonzalez-Vega, and A. Shakoori. The Bézout matrix for Hermite interpolants. Linear Algebra and its Applications, 474 (??):12–29, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000695.

Arauz:2015:DSP

[ACE15] C. Araúz, A. Carmona, and A. M. Encinas. Discrete Serrin’s problem. Linear Algebra and its Applications, 468 (??):107–121, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400072X.

Arias:2015:RAS

[ACM15] M. Laura Arias, Gustavo Corach, and Alejandra Maestripieri. Range additivity, shorted operator and the Sherman–Morrison– Woodbury formula. Linear Algebra and its Applications, 467(??):86–99, February 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007241.

Alyani:2017:DMH

[ACM17] Khaled Alyani, Marco Congedo, and Maher Moakher.

Diagonality measures of Hermitian positive-definite matrices with application to the approximate joint diagonalization problem. Linear Algebra and its Applications, 528(??):290–320, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303834.

Andrade:2015:FIL

[ACPR15] Enide Andrade, Domingos M. Cardoso, Germain Pastén, and Oscar Rojo. On the Faria’s inequality for the Laplacian and signless Laplacian spectra: a unified approach. Linear Algebra and its Applications, 472(??):81–96, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000555.

Andrade:2015:LSG

[ACRR15] Enide Andrade, Domingos M. Cardoso, María Robbiano, and Jonnathan Rodríguez. Laplacian spread of graphs: Lower bounds and relations with invariant parameters. Linear Algebra and its Applications, 486(??):494–503, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

(electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004954.

Alpay:2019:HFF

[ACS19] Daniel Alpay, Fabrizio Colombo, and Irene Sabadini. Herglotz functions: the Fueter variables case. Linear Algebra and its Applications, 563(??):75–97, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304932.

Abiad:2016:SBI

[ACT16] Aida Abiad, Sebastian M. Cioaba, and Michael Tait. Spectral bounds for the kindependence number of a graph. Linear Algebra and its Applications, 510(??): 160–170, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303615.

Andelic:2019:SCT

[AdF19] Milica Andelić and Carlos M. da Fonseca. Some comments on tridiagonal (p, r)-Toeplitz matrices. Linear Algebra and its Applications, 572(??):46–50, July 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519300953.

Andrade:2018:NLB

[AdFRR18] Enide Andrade, Maria Aguieiras A. de Freitas, María Robbiano, and Jonnathan Rodríguez. New lower bounds for the Randić spread. Linear Algebra and its Applications, 544 (??):254–272, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300193.

Agudelo:2016:EVT

[AdlPR16] N. Agudelo, J. A. de la Peña, and J. Rada. Extremal values of the trace norm over oriented trees. Linear Algebra and its Applications, 505(??):261–268, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301525.

Andrade:2019:NBS

[ADLR19] Enide Andrade, Geir Dahl, Laura Leal, and María Robbiano. New bounds for the signless Laplacian spread. Linear Algebra and its Applications, 566 (??):98–120, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305822.

Andriantiana:2019:AMR

[ADW19] Eric O. D. Andriantiana, Kenneth Dadedzi, and Stephan Wagner. The ancestral matrix of a rooted tree. Linear Algebra and its Applications, 575(??): 35–65, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301569.

Abo:2016:EJC

[AEKP16] Hirotachi Abo, David Eklund, Thomas Kahle, and Chris Peterson. Eigenschemes and the Jordan canonical form. Linear Algebra and its Applications, 496 (??):121–151, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000598.

Anzis:2015:ILP

[AEV15] Benjamin E. Anzis, Zachary M. Emrich, and Kaavya G. Valiveti. On the images of Lie polynomials evaluated on Lie algebras. Linear Algebra and its Applications, 469 (??):51–75, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007496.

Alaminos:2016:SPA

[AEV16] J. Alaminos, J. Extremera, and A. R. Villena. Spectral preservers and approximate spectral preservers on operator algebras. Linear Algebra and its Applications, 496(??):36–70, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000616.

Aguilar:2015:LCC

[AG15a] Cesar O. Aguilar and Bahman Gharesifard. Laplacian controllability classes for threshold graphs. Linear Algebra and its Applications, 471(??):575–586, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500049X.

Ay:2015:RSA

[AG15b] Serdar Ay and Aurelian Gheondea. Representations of ∗-semigroups associated to invariant kernels with values adjointable operators. Linear Algebra and its Applications, 486(??):361–388, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004772.

See corrigendum [AG20].

Adm:2016:TNE

[AG16] Mohammad Adm and Jürgen Garloff. Total nonnegativity of the extended Perron complement. Linear Algebra and its Applications, 508(??): 214–224, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302646.

Adm:2017:ITN

[AG17] Mohammad Adm and Jürgen Garloff. Invariance of total nonnegativity of a matrix under entry-wise perturbation and subdirect sum of totally nonnegative matrices. Linear Algebra and its Applications, 514(??): 222–233, February 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305171.

Arias:2018:PPS

[AG18] M. Laura Arias and M. Celeste Gonzalez. Products of projections and selfadjoint operators. Linear Algebra and its Applications, 555(??):70–83, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302842.

Ay:2020:CSR

[AG20] Serdar Ay and Aurelian Gheondea. Corrigendum to “Representations of ∗- semigroups associated to invariant kernels with values adjointable operators” [Linear Algebra Appl. 486 (2015) 361–388]. Linear Algebra and its Applications, 589(??):242–246, March 15, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379520300136.

See [AG15b].

Akbari:2017:IIA

[AGH17] Saieed Akbari, Amir Hossein Ghodrati, and Mohammad Ali Hosseinzadeh. Imprimitivity index of the adjacency matrix of digraphs. Linear Algebra and its Applications, 517 (??):1–10, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305833.

Arrigo:2018:EGF

[AGHN18] Francesca Arrigo, Peter Grindrod, Desmond J. Higham, and Vanni Noferini. On the exponential generating function for nonbacktracking walks. Linear Algebra and its Applications, 556(??):381–399, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303288.

Akian:2018:TCM

[AGN18] Marianne Akian, Stéphane Gaubert, and Adi Niv. Tropical compound matrix identities. Linear Algebra and its Applications, 551(??): 162–206, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301848.

Alahmadi:2017:TCC

[AGP^{+}17] Adel Alahmadi, S. P. Glasby, Cheryl E. Praeger, Patrick Solé, and Bahattin Yildiz. Twisted centralizer codes. Linear Algebra and its Applications, 524(??):235–249, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517301635.

Albahboh:2019:TCI

[AGQ19] Hussin Albahboh, Harry Gingold, and Jocelyn Quaintance. Trigonometry in complex inner product spaces.

Linear Algebra and its Applications, 575(??):216–234, August 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301661.

Andrade:2016:UBL

[AGRR16] Enide Andrade, Helena Gomes, María Robbiano, and Jonnathan Rodríguez. Upper bounds on the Laplacian spread of graphs. Linear Algebra and its Applications, 492(??):26–37, March 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515006734.

Akian:2017:LMM

[AGS17] Marianne Akian, Stéphane Gaubert, and Meisam Sharify. Log-majorization of the moduli of the eigenvalues of a matrix polynomial by tropical roots. Linear Algebra and its Applications, 528(??):394–435, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305201.

Armentia:2017:BSP

[AGV17] Gorka Armentia, Juan-Miguel Gracia, and Francisco-Enrique Velasco. On the boundaries of strict pseudospectra. Linear Algebra and its Applications, 525 (??):20–32, July 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730160X.

Armentia:2018:PSP

[AGV18] Gorka Armentia, Juan-Miguel Gracia, and Francisco-Enrique Velasco. Perforated strict pseudospectra of Demmel’s matrices. Linear Algebra and its Applications, 548(??):77–94, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301095.

Atay:2016:SLS

[AH16] Fatihcan M. Atay and Bobo Hua. On the symmetry of the Laplacian spectra of signed graphs. Linear Algebra and its Applications, 495 (??):24–37, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000306.

Arizmendi:2018:EV

[AHJR18] Octavio Arizmendi, Jorge Fernandez Hidalgo, and Oliver Juarez-Romero. Energy of a vertex. Linear Algebra and its Applications, 557(??):464–495, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518303872.

Akbari:2018:SGC

[AHMM18] Saieed Akbari, Willem H. Haemers, Hamid Reza Maimani, and Leila Parsaei Majd. Signed graphs cospectral with the path. Linear Algebra and its Applications, 553(??):104–116, September 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518302155.

anHuef:2015:STS

[aHRT15] Astrid an Huef, Iain Raeburn, and Ilija Tolich. Structure theorems for starcommuting power partial isometries. Linear Algebra and its Applications, 481(??):107–114, September 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515002657.

Albrecht:2019:FEG

[AHV19] Amie Albrecht, Phil Howlett, and Geetika Verma. The fundamental equations for the generalized resolvent of an elementary pencil in a unital Banach algebra. Linear Algebra and its Applications, 574(??): 216–251, August 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301466.

Aishima:2018:QCAa

[Ais18a]  Kensuke Aishima.  A

quadratically convergent algorithm based on matrix equations for inverse eigenvalue problems. Linear Algebra and its Applications, 542(??):310–333, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303075.

Aishima:2018:QCAb

[Ais18b]  Kensuke Aishima.  A

quadratically convergent algorithm for inverse eigenvalue problems with multiple eigenvalues. Linear Algebra and its Applications, 549(??):30–52, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301320.

Aliabadi:2019:MSA

[AJ19] Mohsen Aliabadi and Mano Vikash Janardhanan. On matchable subsets in abelian groups and their linear analogues. Linear Algebra and its Applications, 582(??):138–155, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303416.

Alahmadi:2016:QPS

[AJL16] Adel Alahmadi, S. K. Jain, and André Leroy. Quasipermutation singular matrices are products of idempotents. Linear Algebra and its Applications, 496 (??):487–495, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000884.

Abdollahi:2015:DBS

[AJO15] Alireza Abdollahi, Shahrooz Janbaz, and Mohammad Reza Oboudi. Distance between spectra of graphs. Linear Algebra and its Applications, 466(??):401–408, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006892.

Ansill:2016:FSZ

[AJPS16] Thomas Ansill, Bonnie Jacob, Jaime Penzellna, and Daniel Saavedra. Failed skew zero forcing on a graph. Linear Algebra and its Applications, 509(??): 40–63, November 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302816.

Ayupov:2016:LAF

[AK16a] Shavkat Ayupov and Karimbergen Kudaybergenov. 2local automorphisms on finite-dimensional Lie algebras. Linear Algebra and its Applications, 507(??): 121–131, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302178.

Ayupov:2016:LDF

[AK16b] Shavkat Ayupov and Karimbergen Kudaybergenov. Local derivations on finitedimensional Lie algebras. Linear Algebra and its Applications, 493(??):381–398, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007089.

Allamigeon:2017:TFP

[AK17] Xavier Allamigeon and Ricardo D. Katz. Tropicalization of facets of polytopes. Linear Algebra and its Applications, 523(??):79– 101, June 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300873.

Aldalabih:2019:HSN

[AK19] Alaa Aldalabih and Fuad Kittaneh. Hilbert–Schmidt numerical radius inequalities for operator matrices. Linear Algebra and its Applications, 581(??): 72–84, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930285X.

Alvarez:2017:ESR

[ÁKM17] Teresa Álvarez, Sonia Keskes, and Maher Mnif. On essentially semi regular linear relations. Linear Algebra and its Applications, 530(??): 518–540, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303804.

Atreas:2019:DMD

[AKPS19] N. Atreas, N. Karantzas, M. Papadakis, and T. Stavropoulos. On the design of multidimensional compactly supported Parseval framelets with directional characteristics. Linear Algebra and its Applications, 582(??): 1–36, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303155.

See erratum [KAPS20].

Ayupov:2015:LDF

[AKR15] Shavkat Ayupov, Karimbergen Kudaybergenov, and Isamiddin Rakhimov. 2local derivations on finitedimensional Lie algebras. Linear Algebra and its Applications, 474(??):1–11, June 1, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515000452.

Aggarwal:2017:CMOa

[AKS17a] Anchal Aggarwal, Yogesh Kapil, and Mandeep Singh. Contractive maps on operator ideals and norm inequalities II. Linear Algebra and its Applications, 513(??): 182–200, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630489X.

Aggarwal:2017:CMOb

[AKS17b] Anchal Aggarwal, Yogesh Kapil, and Mandeep Singh. Contractive maps on operator ideals and norm inequalities III. Linear Algebra and its Applications, 530(??): 322–343, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303129.

Altinisik:2016:CIH

[AKYD16] Ercan Altinisik, Ali Keskin, Mehmet Yildiz, and Murat Demirbüken. On a conjecture of Ilmonen, Haukkanen and Merikoski concerning the smallest eigenvalues of certain GCD related matrices. Linear Algebra and its Applications, 493 (??):1–13, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006977.

See [IHM08].

Ando:2015:PCL

[AL15] Tsuyoshi Ando and Minghua Lin. Proof of a conjectured lower bound on the chromatic number of a graph. Linear Algebra and its Applications, 485(??):480–484, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004723.

Arioli:2016:SAA

[AL16] Mario Arioli and Daniel Loghin. Spectral analysis of the anisotropic Steklov– Poincaré matrix. Linear Algebra and its Applications, 488(??):168–183, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005674.

Alpin:2017:HMR

[Al’17] Yu. A. Al’pin. The Hankel matrix rank theorem revisited. Linear Algebra and its Applications, 534(??): 97–101, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730486X.

Alpay:2018:SWS

[AL18] Daniel Alpay and Izchak Lewkowicz. ‘Wrong’ side interpolation by positive real rational functions. Linear Algebra and its Applications, 539(??):175–197, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305402.

Alpay:2019:CRF

[AL19a] Daniel Alpay and Izchak Lewkowicz. Composition of rational functions: Statespace realization and applications. Linear Algebra and its Applications, 580(??): 359–383, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302666.

Axelsson:2019:PMV

[AL19b] Owe Axelsson and Zhao-Zheng Liang. Parameter modified versions of preconditioning and iterative inner product free refinement methods for two-by-two block matrices. Linear Algebra and its Applications, 582(??):403–429, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303118.

Alfakih:2015:FLD

[Alf15] A. Y. Alfakih. On Farkas lemma and dimensional rigidity of bar frameworks. Linear Algebra and its Applications, 486(??):504–522, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500508X.

Alfakih:2018:YJY

[Alf18a] A. Y. Alfakih. On yielding and jointly yielding entries of Euclidean distance matrices. Linear Algebra and its Applications, 556(??): 144–161, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303252.

Alfaro:2018:GRA

[Alf18b] Carlos A. Alfaro. Graphs with real algebraic co-rank at most two. Linear Algebra and its Applications, 556(??):100–107, November 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518303203.

An:2015:ZJP

[ALH15] Guangyu An, Jiankui Li, and Jun He. Zero Jordan product determined algebras. Linear Algebra and its Applications, 475(??):90– 93, June 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000646.

Alizadeh:2020:CEH

[Ali20] Rahim Alizadeh. Corrigendum to “An extension of Haagerup–Harpe inequality” [Linear Algebra Appl. 577 (2019) 69–77]. Linear Algebra and its Applications, 600(??):161–164, September 1, 2020. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379520302202.

See [HAD19].

Andrade:2019:BDS

[ALMZ^{+}19] Enide Andrade, Eber Lenes, Exequiel Mallea-Zepeda, María Robbiano, and Jonnathan Rodríguez Z. Bounds for different spreads of line and total graphs. Linear Algebra and its Applications, 579(??):365–381, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302599.

Alon:2015:ASB

[Alo15] Noga Alon. Approximating sparse binary matrices in the cut-norm. Linear Algebra and its Applications, 486(??):409–418, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004929.

Arizmendi:2019:CIF

[ALOR19] Octavio Arizmendi, Beatriz Carely Luna-Olivera, and Marcelino Ramírez Ibáñez. Coulson integral formula for the vertex energy of a graph.

Linear Algebra and its Applications, 580(??):166–183, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302617.

Liu:2015:SRB

[aLwW15] Chia an Liu and Chih wen Weng. Spectral radius of bipartite graphs. Linear Algebra and its Applications, 474(??):30–43, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000701.

Agore:2015:TCA

[AM15a] A. L. Agore and G. Militaru. On a type of commutative algebras. Linear Algebra and its Applications, 485(??):222–249, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004498.

Araznibarreta:2015:JTC

[AM15b] Gerardo Araznibarreta and Manuel Mañas. A Jacobi type Christoffel–Darboux formula for multiple orthogonal polynomials of mixed type. Linear Algebra and its Applications, 468(??): 154–170, March 1, 2015.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514002560.

Aleksandrov:2016:DRS

[AM16a] Alexander Aleksandrov and Oliver Mason. Diagonal Riccati stability and applications. Linear Algebra and its Applications, 492 (??):38–51, March 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006679.

Arratia:2016:GCI

[AM16b] Argimiro Arratia and Carlos Marijuán. On graph combinatorics to improve eigenvector-based measures of centrality in directed networks. Linear Algebra and its Applications, 504(??): 325–353, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301100.

Azarija:2018:TNS

[AM18] Jernej Azarija and Tilen Marc. There is no (75, 32, 10, 16) strongly regular graph. Linear Algebra and its Applications, 557(??):62–83, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Acharya:2019:ACS

[AM19] Keshav Raj Acharya and Matt McBride. Action of complex symplectic matrices on the Siegel upper half space. Linear Algebra and its Applications, 563(??): 47–62, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830507X.

Afshin:2018:PGC

[AMC18] Hamid Reza Afshin, Mohammad Ali Mehrjoofard, and Akbar Zare Chavoshi. Pauli group: Classification and joint higher rank numerical range. Linear Algebra and its Applications, 549 (??):136–152, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301186.

Araujo:2019:CNM

[AMG19] C. Mendes Araújo and S. Mendes-Gonçalves. On a class of nonsingular matrices containing B -matrices. Linear Algebra and its Applications, 578(??):356–369, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302241.

Al-Mohy:2011:CAM

[AMH11] Awad H. Al-Mohy and Nicholas J. Higham. Computing the action of the matrix exponential, with an application to exponential integrators. SIAM Journal on Scientific Computing, 33(2):488–511, ???? 2011. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

                                        //epubs.siam.org/sisc/
                                        resource/1/sjoce3/v33/

i2/p488_s1. See [Fis17].

Andrade:2018:BMG

[AMNR18] Enide Andrade, Cristina Manzaneda, Hans Nina, and María Robbiano. Block matrices and Guo’s index for block circulant matrices with circulant blocks. Linear Algebra and its Applications, 556(??):301–322, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303410.

Agapito:2016:CGN

[AMPT16] José Agapito, Ângela Mestre, Pasquale Petrullo, and Maria M. Torres. Combinatorics of a generalized Narayana identity. Linear Algebra and its Applications, 503(??):56–82, August 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300507.

Abara:2016:EMS

[AMR^{+}16] Ma. Nerissa M. Abara, Dennis I. Merino, Viacheslav I. Rabanovich, Vladimir V. Sergeichuk, and John Patrick Sta. Maria. Each n-by-n matrix with n > 1 is a sum of 5 coninvolutory matrices. Linear Algebra and its Applications, 508(??):246–254, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302737.

Andrade:2018:RLC

[AMR18] Enide Andrade, Cristina Manzaneda, and María Robbiano. Realizable lists on a class of nonnegative matrices. Linear Algebra and its Applications, 551(??): 36–56, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301824.

Aghamollaei:2019:LPP

[AMR19] Gh. Aghamollaei, L. W. Marcoux, and H. Radjavi.

Linear preservers of polynomial numerical hulls of matrices. Linear Algebra and its Applications, 575 (??):27–34, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301363.

Aleksandrov:2017:DRS

[AMV17] Alexander Aleksandrov, Oliver Mason, and Anna Vorob’eva. Diagonal Riccati stability and the Hadamard product. Linear Algebra and its Applications, 534(??): 158–173, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304913.

Amparan:2016:CRF

[AMZ16] A. Amparan, S. Marcaida, and I. Zaballa. On coprime rational function matrices. Linear Algebra and its Applications, 507(??): 1–31, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302051.

Amparan:2017:MPS

[AMZ17] A. Amparan, S. Marcaida, and I. Zaballa. On matrix polynomials with the same finite and infinite elementary divisors. Linear Algebra and its Applications, 513 (??):1–32, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304621.

Albania:2017:SFO

[AN17] Imam Nugraha Albania and Masaru Nagisa. Some families of operator norm inequalities. Linear Algebra and its Applications, 534(??):102–121, December 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517303944.

Aron:2017:OEH

[ANAPSR17] R. M. Aron, D. Núñez-Alarcón, D. M. Pellegrino, and D. M. Serrano-Rodríguez. Optimal exponents for Hardy–Littlewood inequalities for m-linear operators. Linear Algebra and its Applications, 531(??): 399–422, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303713.

Andruchow:2015:PPS

[And15] Esteban Andruchow. Parametrizing projections with selfadjoint operators. Linear Algebra and its Applications, 466(??):307–328, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006983.

Anonymous:2015:EBa

[Ano15a]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 464(??):ii– iii, January 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006697.

Anonymous:2015:EBb

[Ano15b]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 465(??):ii– iii, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007071.

Anonymous:2015:EBc

[Ano15c]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 466(??):ii– iii, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007423.

Anonymous:2015:EBd

[Ano15d]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 467(??): ii–iii, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007848.

Anonymous:2015:EBe

[Ano15e]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 468(??):ii– iii, March 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008076.

Anonymous:2015:EBf

[Ano15f]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 469(??):ii– iii, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000051.

Anonymous:2015:EBg

[Ano15g]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 472(??):ii– iii, May 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500107X.

Anonymous:2015:EBh

[Ano15h]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 473(??):ii– iii, May 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500244X.

Anonymous:2015:EBi

[Ano15i]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 474(??):ii– iii, June 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001615.

Anonymous:2015:EBj

[Ano15j]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 475(??):ii– iii, June 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001743.

Anonymous:2015:EBk

[Ano15k]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 476(??):ii– iii, July 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001962.

Anonymous:2015:EBl

[Ano15l]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 477(??):ii– iii, July 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002281.

Anonymous:2015:EBm

[Ano15m]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 478(??):ii– iii, August 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002694.

Anonymous:2015:EBn

[Ano15n]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 479(??):ii– iii, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500302X.

Anonymous:2015:EBo

[Ano15o]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 480(??): ii–iii, September 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003183.

Anonymous:2015:EBp

[Ano15p]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 481(??): ii–iii, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003341.

Anonymous:2015:EBq

[Ano15q]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 482(??):ii– iii, October 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003882.

Anonymous:2015:EBr

[Ano15r]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 483(??):ii– iii, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004188.

Anonymous:2015:EBs

[Ano15s]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 484(??): ii–iii, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004644.

Anonymous:2015:EBt

[Ano15t]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 485(??): ii–iii, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005005.

Anonymous:2015:EBu

[Ano15u]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 486(??): ii–iii, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005480.

Anonymous:2015:EBv

[Ano15v]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 487(??): ii–iii, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005807.

Anonymous:2016:CSS

[Ano16a] Anonymous. Corrigendum to “Structured Numerical Linear and Multilinear Algebra: Analysis, Algorithms and Applications” [Linear Algebra Appl. 502 (2016) 1– 4]. Linear Algebra and its Applications, 504(??):623– 624, September 1, 2016.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301562.

See [BMVW16].

Anonymous:2016:EBa

[Ano16b]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 488(??):ii– iii, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006369.

Anonymous:2016:EBb

[Ano16c]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 489(??):ii– iii, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006515.

Anonymous:2016:EBc

[Ano16d]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 490(??):ii– iii, February 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006874.

Anonymous:2016:EBd

[Ano16e]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 491(??): ii–iii, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007363.

Anonymous:2016:EBe

[Ano16f]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 492(??):ii– iii, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007478.

Anonymous:2016:EBf

[Ano16g]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 493(??):ii– iii, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000392.

Anonymous:2016:EBg

[Ano16h]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 494(??):ii– iii, April 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000768.

Anonymous:2016:EBh

[Ano16i]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 495(??):ii– iii, April 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001051.

Anonymous:2016:EBi

[Ano16j]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 496(??):ii– iii, May 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001324.

Anonymous:2016:EBj

[Ano16k]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 497(??):ii– iii, May 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001506.

Anonymous:2016:EBk

[Ano16l]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 498(??):ii– iii, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300155.

[Ano16m]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 499(??):ii– iii, June 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300283.

Anonymous:2016:EBm

[Ano16n]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 500(??):ii– iii, July 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300568.

Anonymous:2016:EBn

[Ano16o]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 501(??):ii– iii, July 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300817.

Anonymous:2016:EBo

[Ano16p]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 502(??):ii– iii, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300994.

Anonymous:2016:EBp

[Ano16q]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 503(??):ii– iii, August 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301380.

Anonymous:2016:EBq

[Ano16r]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 504(??): ii–iii, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301653.

Anonymous:2016:EBr

[Ano16s]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 505(??): ii–iii, September 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301902.

Anonymous:2016:EBs

[Ano16t]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 506(??):ii– iii, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302853.

Anonymous:2016:EBt

[Ano16u]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 507(??):ii– iii, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303433.

Anonymous:2016:EBu

[Ano16v]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 508(??): ii–iii, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303688.

Anonymous:2016:EBv

[Ano16w]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 509(??): ii–iii, November 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304037.

Anonymous:2016:EBw

[Ano16x]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 510(??): ii–iii, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304396.

Anonymous:2016:EBx

[Ano16y]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 511(??): ii–iii, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304712.

Anonymous:2017:EBa

[Ano17a]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 512(??):ii– iii, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630502X.

Anonymous:2017:EBb

[Ano17b]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 513(??):ii– iii, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305341.

Anonymous:2017:EBc

[Ano17c]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 515(??): ii–iii, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305948.

[Ano17d]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 516(??):ii– iii, March 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306206.

Anonymous:2017:EBe

[Ano17e]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 517(??):ii– iii, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300046.

Anonymous:2017:EBf

[Ano17f]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 518(??):ii– iii, April 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300320.

Anonymous:2017:EBg

[Ano17g]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 519(??):ii– iii, April 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300745.

Anonymous:2017:EBh

[Ano17h]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 520(??):ii– iii, May 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301003.

Anonymous:2017:EBi

[Ano17i]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 521(??):ii– iii, May 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301106.

Anonymous:2017:EBj

[Ano17j]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 522(??):ii– iii, June 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301428.

Anonymous:2017:EBk

[Ano17k]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 523(??):ii– iii, June 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301672.

[Ano17l]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 524(??):ii– iii, July 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301982.

Anonymous:2017:EBm

[Ano17m]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 525(??):ii– iii, July 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302203.

Anonymous:2017:EBn

[Ano17n]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 526(??):ii– iii, August 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730246X.

Anonymous:2017:EBo

[Ano17o]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 527(??):ii– iii, August 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302756.

Anonymous:2017:EBp

[Ano17p]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 528(??): ii–iii, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303208.

Anonymous:2017:EBq

[Ano17q]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 529(??): ii–iii, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303543.

Anonymous:2017:EBr

[Ano17r]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 530(??):ii– iii, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304275.

Anonymous:2017:EBs

[Ano17s]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 531(??):ii– iii, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304573.

[Ano17t]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 532(??): ii–iii, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304779.

Anonymous:2017:EBu

[Ano17u]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 533(??): ii–iii, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305025.

Anonymous:2017:EBv

[Ano17v]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 534(??): ii–iii, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305189.

Anonymous:2017:EBw

[Ano17w]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 535(??): ii–iii, December 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305487.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 536(??):ii– iii, January 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305839.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 537(??):ii– iii, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730602X.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 538(??): iii–iv, February 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306195.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 539(??):i, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306717.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 539(??):ii, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306729.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 540(??): i, March 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306936.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 540(??): ii, March 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306948.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 541(??): i, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830003X.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 541(??): ii, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300041.

Anonymous:2018:EBj

[Ano18j]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 542 (??):i, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300296.

Anonymous:2018:EBk

[Ano18k]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 542 (??):ii, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300302.

Anonymous:2018:EBl

[Ano18l]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 543(??): i, April 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300624.

Anonymous:2018:EBm

[Ano18m]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 543(??): ii, April 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300636.

Anonymous:2018:EBn

[Ano18n]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 544(??): i, May 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830079X.

Anonymous:2018:EBo

[Ano18o]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 544(??): ii, May 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300806.

Anonymous:2018:EBp

[Ano18p]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 545(??): i, May 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300934.

Anonymous:2018:EBq

[Ano18q]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 545(??): ii, May 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300946.

Anonymous:2018:EBr

[Ano18r]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 546(??): i, June 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301125.

Anonymous:2018:EBs

[Ano18s]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 546(??): ii, June 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301137.

Anonymous:2018:EBt

[Ano18t]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 547(??): i, June 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301666.

Anonymous:2018:EBu

[Ano18u]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 547(??): ii, June 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301678.

Anonymous:2018:EBv

[Ano18v]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 548 (??):i, July 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830171X.

Anonymous:2018:EBw

[Ano18w]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 548(??): ii, July 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301721.

Anonymous:2018:EBx

[Ano18x]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 549(??): i, July 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301885.

Anonymous:2018:EBy

[Ano18y]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 549(??): ii, July 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301897.

Anonymous:2018:EBz

[Ano18z]  Anonymous.  Editorial

board. Linear Algebra and its Applications, 550(??): i, August 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302027.

Anonymous:2018:EBaa

[Ano18-27]  Anonymous.  Editorial

board. Linear Algebra and its Applications, 550(??): ii, August 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302039.

Anonymous:2018:EBab

[Ano18-28]  Anonymous.  Editorial

board. Linear Algebra and its Applications, 551(??): i, August 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302283.

Anonymous:2018:EBac

[Ano18-29]  Anonymous.  Editorial

board. Linear Algebra and its Applications, 551(??): ii, August 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302295.

Anonymous:2018:EBad

[Ano18-30]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 552(??):i, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302489.

Anonymous:2018:EBae

[Ano18-31]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 552(??):ii, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302490.

Anonymous:2018:EBaf

[Ano18-32]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 553(??):i, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302775.

Anonymous:2018:EBag

[Ano18-33]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 553(??):ii, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302787.

Anonymous:2018:EBah

[Ano18-34]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 554(??): i, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303355.

Anonymous:2018:EBai

[Ano18-35]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 554(??): ii, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303367.

Anonymous:2018:EBaj

[Ano18-36]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 555(??):i, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303628.

Anonymous:2018:EBak

[Ano18-37]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 555(??):ii, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830363X.

Anonymous:2018:EBal

[Ano18-38]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 556(??):i, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303914.

Anonymous:2018:EBam

[Ano18-39]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 556(??):ii, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303926.

Anonymous:2018:EBan

[Ano18-40]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 558(??):i, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304373.

Anonymous:2018:EBao

[Ano18-41]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 558(??):ii, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304385.

Anonymous:2018:EBap

[Ano18-42]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 559(??):i, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304543.

Anonymous:2018:EBaq

[Ano18-43]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 559(??):ii, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304555.

Anonymous:2018:NEC

[Ano18-44]  Anonymous.  Note from

the Editors-in-Chief. Linear Algebra and its Applications, 538(??):Page i, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Anonymous:2019:EBa

[Ano19a]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 560(??): i, January 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304877.

Anonymous:2019:EBb

[Ano19b]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 560(??): ii, January 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304889.

Anonymous:2019:EBc

[Ano19c]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 561(??):i, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305020.

Anonymous:2019:EBd

[Ano19d]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 561(??):ii, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305032.

Anonymous:2019:EBe

[Ano19e]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 562(??):i, February 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305202.

Anonymous:2019:EBf

[Ano19f]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 562(??):ii, February 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305214.

Anonymous:2019:EBg

[Ano19g]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 563(??):i, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305639.

Anonymous:2019:EBh

[Ano19h]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 563(??):ii, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305640.

Anonymous:2019:EBi

[Ano19i]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 564(??): i, March 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305871.

Anonymous:2019:EBj

[Ano19j]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 564(??): ii, March 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305883.

Anonymous:2019:EBk

[Ano19k]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 565(??): i, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300096.

Anonymous:2019:EBl

[Ano19l]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 565(??): ii, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300102.

Anonymous:2019:EBm

[Ano19m]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 567(??): i, April 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300540.

Anonymous:2019:EBn

[Ano19n]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 567(??): ii, April 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300552.

Anonymous:2019:EBo

[Ano19o]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 568(??): i, May 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300710.

Anonymous:2019:EBp

[Ano19p]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 568(??): ii, May 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300722.

Anonymous:2019:EBq

[Ano19q]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 569(??): i, May 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300837.

Anonymous:2019:EBr

[Ano19r]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 569(??): ii, May 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300849.

Anonymous:2019:EBs

[Ano19s]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 570(??): i, June 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301065.

Anonymous:2019:EBt

[Ano19t]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 570(??): ii, June 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301077.

Anonymous:2019:EBu

[Ano19u]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 571(??): i, June 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301260.

Anonymous:2019:EBv

[Ano19v]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 571(??): ii, June 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301272.

Anonymous:2019:EBw

[Ano19w]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 572 (??):i, July 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301417.

Anonymous:2019:EBx

[Ano19x]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 572(??): ii, July 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301429.

Anonymous:2019:EBy

[Ano19y]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 573(??): i, July 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930151X.

Anonymous:2019:EBz

[Ano19z]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 573(??): ii, July 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301521.

Anonymous:2019:EBaa

[Ano19-27]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 574(??): i, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301740.

Anonymous:2019:EBab

[Ano19-28]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 574(??): ii, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301752.

Anonymous:2019:EBac

[Ano19-29]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 575(??): i, August 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301983.

Anonymous:2019:EBad

[Ano19-30]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 575(??): ii, August 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301995.

Anonymous:2019:EBae

[Ano19-31]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 576(??):i, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302149.

Anonymous:2019:EBaf

[Ano19-32]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 576(??):ii, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302150.

Anonymous:2019:EBag

[Ano19-33]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 578(??): i, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519303039.

Anonymous:2019:EBah

[Ano19-34]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 578(??): ii, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519303040.

Anonymous:2019:EBai

[Ano19-35]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 579(??):i, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519303210.

Anonymous:2019:EBaj

[Ano19-36]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 579(??):ii, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303222.

Anonymous:2019:EBak

[Ano19-37]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 580(??):i, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303453.

Anonymous:2019:EBal

[Ano19-38]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 580(??):ii, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303465.

Anonymous:2019:EBam

[Ano19-39]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 581(??):i, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303635.

Anonymous:2019:EBan

[Ano19-40]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 581(??):ii, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303647.

Anonymous:2019:EBao

[Ano19-41]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 582(??):i, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303830.

Anonymous:2019:EBap

[Ano19-42]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 582(??):ii, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303842.

Anonymous:2019:EBaq

[Ano19-43]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 583(??):i, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519304082.

Anonymous:2019:EBar

[Ano19-44]  Anonymous.  Editorial

Board. Linear Algebra and its Applications, 583(??):ii, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519304094.

Almeida:2016:SMA

[ANP16] P. J. Almeida, D. Napp, and R. Pinto. Superregular matrices and applications to convolutional codes. Linear Algebra and its Applications, 499(??):1–25, June 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001464.

Abu-Omar:2018:SRI

[AO18] Amer Abu-Omar. Spectral radius inequalities for sums of operators with an application to the problem of bounding the zeros of polynomials. Linear Algebra and its Applications, 550 (??):28–36, August 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301435.

Abu-Omar:2015:NRI

[AOK15] Amer Abu-Omar and Fuad Kittaneh. Numerical radius inequalities for n × n operator matrices. Linear Algebra and its Applications, 468 (??):18–26, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379513006423.

Abu-Omar:2019:GNR

[AOK19] Amer Abu-Omar and Fuad Kittaneh. A generalization of the numerical radius. Linear Algebra and its Applications, 569(??): 323–334, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300436.

Araujo:2015:LBC

[AP15a] Gustavo Araújo and Daniel Pellegrino. Lower bounds for the complex polynomial Hardy–Littlewood inequality. Linear Algebra and its Applications, 474 (??):184–191, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001172.

Atik:2015:DSD

[AP15b] Fouzul Atik and Pratima Panigrahi. On the distance spectrum of distance regular graphs. Linear Algebra and its Applications, 478(??):

256–273, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002177.

Andrade:2017:AEI

[APR^{+}17] Enide Andrade, Pamela Pizarro, María Robbiano, B. San Martín, and Katherine Tapia. Applications of Estrada indices and energy to a family of compound graphs. Linear Algebra and its Applications, 532(??): 162–178, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303993.

Alfaro:2018:NMP

[APS18] Jaime H. Alfaro, Germain Pastén, and Ricardo L. Soto. On nonnegative matrices with prescribed eigenvalues and diagonal entries. Linear Algebra and its Applications, 556(??):400–420, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303513.

Aretaki:2017:ETT

[APT17] Aik. Aretaki, P. Psarrakos, and M. Tsatsomeros. The envelope of tridiagonal Toeplitz matrices and block-shift matrices. Linear Algebra and its Applications, 532(??):60–85, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730383X.

Adin:2015:MCD

[AR15a] Ron M. Adin and Yuval Roichman. Matrices, characters and descents. Linear Algebra and its Applications, 469(??):381–418, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007630.

Arambasic:2015:OPS

[AR15b] Ljiljana Arambasić and Rajna Rajić. Operators preserving the strong Birkhoff– James orthogonality on B(H). Linear Algebra and its Applications, 471(??): 394–404, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000403.

Agudelo:2016:LBN

[AR16] Natalia Agudelo and Juan Rada. Lower bounds of Nikiforov’s energy over digraphs. Linear Algebra and its Applications, 494 (??):156–164, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000100.

Aghamollaei:2018:QNR

[AR18] Gholamreza Aghamollaei and Meysam Rahjoo. On quaternionic numerical ranges with respect to nonstandard involutions. Linear Algebra and its Applications, 540 (??):11–25, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306419.

Agudelo:2018:UBT

[ARR18] Natalia Agudelo, Juan Rada, and Mauricio Rivera. Upper bound for the trace norm of the Laplacian matrix of a digraph and normally regular digraphs. Linear Algebra and its Applications, 552(??):194–209, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302167.

Adiga:2016:MAM

[ARS16] Chandrashekar Adiga, B. R. Rakshith, and Wasin So. On the mixed adjacency matrix of a mixed graph. Linear Algebra and its Applications, 495(??):223–241, April 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000483.

Andrade:2017:LBE

[ARS17] Enide Andrade, María Robbiano, and B. San Martín. A lower bound for the energy of symmetric matrices and graphs. Linear Algebra and its Applications, 513(??): 264–275, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304955.

Andreotti:2018:MLE

[ARSB18] Eleonora Andreotti, Daniel Remondini, Graziano Servizi, and Armando Bazzani. On the multiplicity of Laplacian eigenvalues and Fiedler partitions. Linear Algebra and its Applications, 544 (??):206–222, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300156.

Allman:2015:TNR

[ARSZ15] Elizabeth S. Allman, John A. Rhodes, Bernd Sturmfels, and Piotr Zwiernik. Tensors of nonnegative rank two. Linear Algebra and its Applications, 473(??):37– 53, May 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379513006812.

Alpay:2017:BLT

[AS17] Daniel Alpay and Irene Sabadini. Beurling–Lax type theorems in the complex and quaternionic setting. Linear Algebra and its Applications, 530(??): 15–46, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302860.

Ashraf:2019:EMN

[Ash19] Firouzeh Ashraf. Energy, matching number and odd cycles of graphs. Linear Algebra and its Applications, 577(??):159–167, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301879.

Ahmadizadeh:2017:CSE

[ASMN17a] Saeed Ahmadizadeh, Iman Shames, Samuel Martin, and Dragan Nesić. Corrigendum to “On eigenvalues of Laplacian matrix for a class of directed signed graphs” [Linear Algebra Appl. 523 (2017) 281– 306]. Linear Algebra and its Applications, 530(??): 541–557, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303865.

See [ASMN17b].

Ahmadizadeh:2017:ELM

[ASMN17b] Saeed Ahmadizadeh, Iman Shames, Samuel Martin, and Dragan Nesić. On eigenvalues of Laplacian matrix for a class of directed signed graphs. Linear Algebra and its Applications, 523 (??):281–306, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301301.

See corrigendum [ASMN17a].

Amooshahi:2015:CSC

[AT15] M. Amooshahi and B. Taeri. Cayley sum color and anticirculant graphs. Linear Algebra and its Applications, 466(??):409–420, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007009.

Adeli:2018:RMC

[ATM18] Iman Adeli, Maryam Taheri, and Mahmoud Mohseni Moghadam. A recursive method for constructing doubly stochastic matrices and inverse eigenvalue problem. Linear Algebra and its Applications, 537(??): 318–331, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305748.

Alvarez-Vizoso:2019:GCR

[ÁVAK^{+}19] J. Álvarez-Vizoso, Robert Arn, Michael Kirby, Chris Peterson, and Bruce Draper. Geometry of curves in R^{n} from the local singular value decomposition. Linear Algebra and its Applications, 571 (??):180–202, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300618.

Arav:2015:ISS

[AvdHS15] Marina Arav, Hein van der Holst, and John Sinkovic. On the inertia set of a signed graph with loops. Linear Algebra and its Applications, 471(??):169–183, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008258.

Arav:2016:ISS

[AvdHS16] Marina Arav, Hein van der Holst, and John Sinkovic. On the inertia set of a signed tree with loops. Linear Algebra and its Applications, 510(??):361–372, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303895.

Astuti:2015:CSH

[AW15a] Pudji Astuti and Harald K. Wimmer. Characteristic subspaces and hyperinvariant frames. Linear Algebra and its Applications, 482 (??):21–46, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003109.

Astuti:2015:HSL

[AW15b] Pudji Astuti and Harald K. Wimmer. Hyperinvariant subspaces of locally nilpotent linear transformations. Linear Algebra and its Applications, 486(??):469–474, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005157.

Altinisik:2017:NDL

[AYK17] Ercan Altinisik, Mehmet Yildiz, and Ali Keskin. Nondivisibility of LCM matrices by GCD matrices on gcdclosed sets. Linear Algebra and its Applications, 516 (??):47–68, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305547.

Aguilar:2018:SCA

[AyLPS18] Cesar O. Aguilar, Joon yeob Lee, Eric Piato, and Barbara J. Schweitzer. Spectral characterizations of anti-regular graphs. Linear Algebra and its Applications, 557(??):84–104, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303562.

Boumal:2015:LRM

[BA15] Nicolas Boumal and P.-A. Absil. Low-rank matrix completion via preconditioned optimization on the Grassmann manifold. Linear Algebra and its Applications, 475 (??):200–239, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001342.

Bahmani:2019:CCW

[Bah19] Asghar Bahmani. A counterexample to a conjecture of Wang and Hou on signed graphs. Linear Algebra and its Applications, 577(??):240–243, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301910.

Brent:2015:NBP

[BaHOS15] Richard P. Brent, Judy anne H. Osborn, and Warren D. Smith. Note on best possible bounds for determinants of matrices close to the identity matrix. Linear Algebra and its Applications, 466(??): 21–26, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006387.

Bai:2018:SCH

[Bai18] Zhong-Zhi Bai. On spectral clustering of HSS preconditioner for generalized saddle-point matrices. Linear Algebra and its Applications, 555(??):285–300, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303069.

Balaji:2015:LCR

[Bal15a] R. Balaji. Linear complementarity results for Zmatrices on Lorentz cone. Linear Algebra and its Applications, 480(??):107–114, September 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002529.

Balaji:2015:URL

[Bal15b] R. Balaji. A uniqueness result for linear complementarity problems over the Jordan spin algebra. Linear Algebra and its Applications, 479 (??):1–11, August 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002098.

Ballico:2015:LSM

[Bal15c] E. Ballico. Linear subspaces of matrices associated to a Ferrers diagram and with a prescribed lower bound for their rank. Linear Algebra and its Applications, 483(??):30–39, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500347X.

Ballico:2017:NIE

[Bal17a] E. Ballico. Non-integral elements in holomorphic algebras. Linear Algebra and its Applications, 520 (??):77–79, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300435.

Ballico:2017:NRM

[Bal17b] E. Ballico. On the numerical range of matrices over a finite field. Linear Algebra and its Applications, 512(??):162–171, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304359.

See corrigendum [?].

Ballico:2018:CNR

[Bal18] E. Ballico. Corrigendum to “On the numerical range of matrices over a finite field” [Linear Algebra Appl. 512 (2017) 162–171]. Linear Algebra and its Applications, 556(??):421–427, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303550.

See [Bal17b].

Ballico:2019:NRF

[Bal19] E. Ballico. Numerical range over finite fields: Restriction to subspaces. Linear Algebra and its Applications, 571(??):1–13, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300680.

Banasiak:2016:EFL

[Ban16] J. Banasiak. Explicit formulae for limit periodic flows on networks. Linear Algebra and its Applications, 500(??):30–42, July 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001695.

Bart:2015:RMA

[Bar15] Harm Bart. Review of Matrix Analysis, second edition, Roger A. Horn, Charles R. Johnson. Cambridge University Press, Cambridge, Melbourne etc. (2013), ISBN: 978-0-521-54823-6. Linear Algebra and its Applications, 466(??):527–529, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006922.

Barry:2016:RAG

[Bar16a] Paul Barry. Riordan arrays, generalized Narayana triangles, and series reversion. Linear Algebra and its Applications, 491(??): 343–385, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006618.

Barvinok:2016:CMD

[Bar16b] Alexander Barvinok. Concentration of the mixed discriminant of well-conditioned matrices. Linear Algebra and its Applications, 493(??): 120–133, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007223.

Barabanov:2017:NSC

[Bar17a] N. E. Barabanov. Necessary and sufficient conditions for solvability of the Riccati inequalities in the general case. Linear Algebra and its Applications, 513(??): 33–54, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303822.

Barbarino:2017:EBG

[Bar17b] Giovanni Barbarino. Equivalence between GLT sequences and measurable functions. Linear Algebra and its Applications, 529(??):397–412, September 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517302896.

Barry:2019:HRA

[Bar19] Paul Barry. On the halves of a Riordan array and their antecedents. Linear Algebra and its Applications, 582(??):114–137, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303295.

Benner:2016:SLS

[BB16a] Peter Benner and Zvonimir Bujanović. On the solution of large-scale algebraic Riccati equations by using lowdimensional invariant subspaces. Linear Algebra and its Applications, 488(??): 430–459, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500542X.

Breiding:2016:DER

[BB16b] Paul Breiding and Peter Bürgisser. Distribution of the eigenvalues of a random system of homogeneous polynomials. Linear Algebra and its Applications, 497 (??):88–107, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001221.

Beltita:2017:NOP

[BB17] Ingrid Beltita and Daniel Beltita. Nonlinear oblique projections. Linear Algebra and its Applications, 533(??):451–467, November 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951730469X.

Bazan:2018:SMI

[BB18] Fermín S. V. Bazán and Everton Boos. Schultz matrix iteration based method for stable solution of discrete ill-posed problems. Linear Algebra and its Applications, 554(??):120–145, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302623.

Bolla:2015:SPM

[BBC^{+}15] Marianna Bolla, Brian Bullins, Sorathan Chaturapruek, Shiwen Chen, and Katalin Friedl. Spectral properties of modularity matrices. Linear Algebra and its Applications, 473(??): 359–376, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007186.

Belardo:2018:SBG

[BBC18] Francesco Belardo, Maurizio Brunetti, and Adriana Ciampella. Signed bicyclic graphs minimizing the least Laplacian eigenvalue. Linear Algebra and its Applications, 557(??):201–233, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303549.

Baragana:2016:RWH

[BBdH16] Itziar Baragaña, M. Asunción Beitia, and Inmaculada de Hoyos. A refined Wiener– Hopf equivalence relation for polynomial matrices. Linear Algebra and its Applications, 506(??):342–362, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302245.

Baragana:2018:CFR

[BBdH18] Itziar Baragaña, M. Asunción Beitia, and Inmaculada de Hoyos. Canonical form for the refined Wiener–Hopf equivalence relation for nonsingular 3 × 3 polynomial matrices. Linear Algebra and its Applications, 542 (??):527–556, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304391.

Bagheri-Bardi:2018:WTD

[BBEE18] G. A. Bagheri-Bardi, A. Elyaspour, and G. H. Esslamzadeh. Wold-type decompositions in Baer ∗-rings. Linear Algebra and its Applications, 539(??):117–133, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306134.

Bagheri-Bardi:2019:RAS

[BBEE19] G. A. Bagheri-Bardi, A. Elyaspour, and G. H. Esslamzadeh. The role of algebraic structure in the invariant subspace theory. Linear Algebra and its Applications, 583(??):102– 118, December 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303696.

Bebiano:2019:PSI

[BBF19] Natalia Bebiano, Matej Bresar, and Vyacheslav Futorny. Preface to the special issue dedicated to Vladimir Sergeichuk on the occasion of his 70th birthday. Linear Algebra and its Applications, 568(??):1–9, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930062X.

Bueno:2017:SCH

[BBFF17] M. I. Bueno, J. Breen, S. Ford, and S. Furtado. On the sign characteristic of Hermitian linearizations in DL(P). Linear Algebra and its Applications, 519 (??):73–101, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306346.

Bogoya:2016:EHT

[BBGM16] J. M. Bogoya, A. Böttcher, S. M. Grudsky, and E. A. Maximenko. Eigenvectors of Hermitian Toeplitz matrices with smooth simple-loop symbols. Linear Algebra and its Applications, 493(??): 606–637, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007314.

Barrett:2016:SMI

[BBH16] Wayne Barrett, Steve Butler, and H. Tracy Hall. Symmetric 0–1 matrices with inverses having two distinct values and constant diagonal. Linear Algebra and its Applications, 491(??): 41–55, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001251.

Bucero:2016:CGC

[BBM16] Marta Abril Bucero, Chandrajit Bajaj, and Bernard Mourrain. On the construction of general cubature formula by flat extensions. Linear Algebra and its Applications, 502(??): 104–125, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500590X.

Badaoui:2019:RBC

[BBM19] Mohamad Badaoui, Alain Bretto, and Bassam Mourad. On a relationship between Cayley graphs and G-graphs with some applications. Linear Algebra and its Applications, 582(??):37–57, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303180.

Belchi:2017:HOW

[BBMFM17] Francisco Belchí, Urtzi Buijs, José M. Moreno-Fernández, and Aniceto Murillo. Higher order Whitehead products and L_{∞} structures on the homology of a DGL. Linear Algebra and its Applications, 520 (??):16–31, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300241.

Bermudez:2019:EDP

[BBMP19] T. Bermúdez, A. Bonilla, V. Müller, and A. Peris. Ergodic and dynamical properties of m-isometries. Linear Algebra and its Applications, 561(??):98–112, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304634.

Bhunia:2019:NRI

[BBP19] Pintu Bhunia, Santanu Bag, and Kallol Paul. Numerical radius inequalities and its applications in estimation of zeros of polynomials. Linear Algebra and its Applications, 573(??):166– 177, July 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301181.

Bendaoud:2016:PPS

[BBS16a] M. Bendaoud, A. Benyouness, and M. Sarih. Preservers of pseudo spectral radius of operator products. Linear Algebra and its Applications, 489(??): 186–198, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006175.

Braatvedt:2016:RBA

[BBS16b] G. Braatvedt, R. Brits, and F. Schulz. Rank in Banach algebras: a generalized Cayley–Hamilton theorem. Linear Algebra and its Applications, 507(??): 389–398, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302385.

Bendaoud:2018:PRU

[BBSS18] M. Bendaoud, A. Benyouness, M. Sarih, and S. Sekkat. Preservers of radial unitary similarity functions on products of operators. Linear Algebra and its Applications, 542 (??):484–500, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304214.

Bertolini:2015:CLP

[BBT15] Marina Bertolini, GianMario Besana, and Cristina Turrini. Critical loci for projective reconstruction from multiple views in higher dimension: a comprehensive theoretical approach. Linear Algebra and its Applications, 469(??):335–363, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007563.

Beltita:2015:CLA

[BC15a] Daniel Beltita and Benjamin Cahen. Contractions of Lie algebras with 2-dimensional generic coadjoint orbits. Linear Algebra and its Applications, 466(??):41–63, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006521.

Boussairi:2015:SSM

[BC15b] Abderrahim Boussaïri and Brahim Chergui. Skewsymmetric matrices and their principal minors. Linear Algebra and its Applications, 485(??):47–57, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004383.

Bausch:2016:CD

[BC16a] Johannes Bausch and Toby Cubitt. The complexity of divisibility. Linear Algebra and its Applications, 504(??):64–107, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300520.

Brandts:2016:GAS

[BC16b] Jan Brandts and Apo Cihangir. Geometric aspects of the symmetric inverse M-matrix problem. Linear Algebra and its Applications, 506(??):33–81, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630177X.

Borobia:2017:AMC

[BC17a] Alberto Borobia and Roberto Canogar. ACI-matrices of constant rank over arbitrary fields. Linear Algebra and its Applications, 527(??): 232–259, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730215X.

Borobia:2017:RNI

[BC17b] Alberto Borobia and Roberto Canogar. The real nonnegative inverse eigenvalue problem is NP-hard. Linear Algebra and its Applications, 522 (??):127–139, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300861.

Bhatia:2019:PPR

[BC19a] Rajendra Bhatia and Marco Congedo. Procrustes problems in Riemannian manifolds of positive definite matrices. Linear Algebra and its Applications, 563(??): 440–445, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305329.

Borobia:2019:WDP

[BC19b] Alberto Borobia and Roberto Canogar. The WSTdecomposition for partial matrices. Linear Algebra and its Applications, 564 (??):95–125, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305512.

Berman:2018:FGD

[BCC^{+}18] Abraham Berman, Dong-Mei Chen, Zhi-Bing Chen, Wen-Zhe Liang, and Xiao-Dong Zhang. A family of graphs that are determined by their normalized Laplacian spectra. Linear Algebra and its Applications, 548(??):66–76, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301034.

Butler:2016:EPR

[BCF^{+}16] Steve Butler, Minerva Catral, Shaun M. Fallat, H. Tracy Hall, Leslie Hogben, P. van den Driessche, and Michael Young. The enhanced principal rank characteristic sequence. Linear Algebra and its Applications, 498 (??):181–200, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001871.

Breen:2018:CBM

[BCF^{+}18] Jane Breen, Emanuele Crisostomi, Mahsa Faizrahnemoon, Steve Kirkland, and Robert Shorten. Clustering behaviour in Markov chains with eigenvalues close to one. Linear Algebra and its Applications, 555(??): 163–185, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302970.

Buckley:2016:QCD

[BCJ^{+}16] Shannon P. Buckley, Joseph G. Corliss, Charles R. Johnson, Cristina Araúz Lombardía, and Carlos M. Saiago. Questions, conjectures, and data about multiplicity lists for trees. Linear Algebra and its Applications, 511(??):72– 109, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303299.

Bemrose:2017:UCH

[BCKL17] Travis Bemrose, Peter G. Casazza, Victor Kaftal, and Richard G. Lynch. The unconditional constants for Hilbert space frame expansions. Linear Algebra and its Applications, 521(??):1– 18, May 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306024.

Banerjee:2017:SGH

[BCM17] Anirban Banerjee, Arnab Char, and Bibhash Mondal. Spectra of general hypergraphs. Linear Algebra and its Applications, 518 (??):14–30, April 1, 2017.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306115.

Barrachina:2015:DCF

[BCMASS15] Xavier Barrachina, J. Alberto Conejero, Marina Murillo-Arcila, and Juan B. Seoane-Sepúlveda. Distributional chaos for the Forward and Backward Control traffic model. Linear Algebra and its Applications, 479(??):202–215, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002372.

Burgos:2016:WLT

[BCP16] María J. Burgos, Juan Carlos Cabello, and Antonio M. Peralta. Weaklocal triple derivations on C^{∗}-algebras and JB^{∗}-triples. Linear Algebra and its Applications, 506(??):614–627, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302592.

Chan:2018:UBG

[bCR18] Yao ban Chan and Andrew Rechnitzer. Upper bounds on the growth rates of independent sets in two dimensions via corner transfer matrices. Linear Algebra and its Applications, 555(??): 139–156, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302921.

Bermudez:2015:CSR

[BCS15] J. M. Ancochea Bermúdez and R. Campoamor-Stursberg. Classification of solvable real rigid Lie algebras with a nilradical of dimension n ≤ 6. Linear Algebra and its Applications, 471 (??):54–75, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514008222.

Bermudez:2016:CRS

[BCS16] J. M. Ancochea Bermúdez and R. Campoamor-Stursberg. Cohomologically rigid solvable Lie algebras with a nilradical of arbitrary characteristic sequence. Linear Algebra and its Applications, 488(??):135–147, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005686.

Bru:2015:ERO

[BCU15] Rafael Bru, Rafael Cantó, and Ana M. Urbano. Eigenstructure of rank one updated matrices. Linear Algebra and its Applications, 485(??):372–391, November 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515004504.

Baaijens:2015:EDD

[BD15a] Jasmijn A. Baaijens and Jan Draisma. Euclidean distance degrees of real algebraic groups. Linear Algebra and its Applications, 467(??): 174–187, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007368.

Bracic:2015:RNR

[BD15b] Janko Bracic and Cristina Diogo. Relative numerical ranges. Linear Algebra and its Applications, 485(??):208–221, November 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515004516.

Brualdi:2019:ASM

[BD19] Richard A. Brualdi and Geir Dahl. Alternating sign matrices, related (0, 1)- matrices, and the Smith normal form. Linear Algebra and its Applications, 568 (??):10–28, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830137X.

Barragan:2016:ZSS

[BDDO16] Luis A. Barragán, Geir Dahl, Alberto Domínguez, and Arantxa Otín. Zero sum sign-central matrices and applications. Linear Algebra and its Applications, 505(??):109–125, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301331.

Bertaccini:2017:OMR

[BDDSC17] Daniele Bertaccini, Marco Donatelli, Fabio Durastante, and Stefano Serra-Capizzano. Optimizing a multigrid Runge–Kutta smoother for variable-coefficient convection-diffusion equations. Linear Algebra and its Applications, 533(??):507– 535, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304664.

Bueno:2015:LVS

[BDFR15] M. I. Bueno, F. M. Dopico, S. Furtado, and M. Rychnovsky. Large vector spaces of block-symmetric strong linearizations of matrix polynomials. Linear Algebra and its Applications, 477 (??):165–210, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002086.

Bartoszewicz:2015:MSS

[BDG15] Artur Bartoszewicz, Pratulananda Das, and Szymon Glab. On matrix summability of spliced sequences and A-density of points. Linear Algebra and its Applications, 487(??):22–42, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005091.

Barlow:2015:P

[BDIM15] Jesse L. Barlow, Zlatko Drmac, Ilse Ipsen, and Julio Moro. Preface. Linear Algebra and its Applications, 464 (??):1–2, January 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006533.

Bostan:2017:SMR

[BDK^{+}17a] A. Bostan, C. D’Andrea, T. Krick, A. Szanto, and M. Valdettaro. Subresultants in multiple roots: an extremal case. Linear Algebra and its Applications, 529(??):185–198, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302562.

Boutsidis:2017:RAA

[BDK^{+}17b] Christos Boutsidis, Petros Drineas, Prabhanjan Kambadur, Eugenia-Maria Kontopoulou, and Anastasios Zouzias. A randomized algorithm for approximating the log determinant of a symmetric positive definite matrix. Linear Algebra and its Applications, 533(??):95– 117, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304147.

Benito:2017:FNN

[BdlCL17] P. Benito, D. de-la Concepción, and J. Laliena. Free nilpotent and nilpotent quadratic Lie algebras. Linear Algebra and its Applications, 519(??): 296–326, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730023X.

Bini:2016:CEL

[BDLM16] D. A. Bini, S. Dendievel, G. Latouche, and B. Meini. Computing the exponential of large block-triangular block-Toeplitz matrices encountered in fluid queues. Linear Algebra and its Applications, 502(??):387–419, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002116.

Benitez:2018:SUP

[BDM18] Teresa Cortadellas Benítez, Carlos D’Andrea, and Eulàlia Montoro. The set of unattainable points for the Rational Hermite Interpolation Problem. Linear Algebra and its Applications, 538(??):116–142, February 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517305682.

Balbuena:2019:CIC

[BDMB19] C. Balbuena, C. Dalfó, and B. Martínez-Barona. Characterizing identifying codes from the spectrum of a graph or digraph. Linear Algebra and its Applications, 570 (??):138–147, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300655.

Bebiano:2018:ILH

[BdP18a] N. Bebiano and J. da Providência. Implications of losing Hermiticity in quantum mechanics. Linear Algebra and its Applications, 542(??): 54–65, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300629.

Bueno:2018:SAF

[BDP^{+}18b] M. I. Bueno, F. M. Dopico, J. Pérez, R. Saavedra, and B. Zykoski. A simplified approach to Fiedler-like pencils via block minimal bases pencils. Linear Algebra and its Applications, 547 (??):45–104, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300478.

Benner:2016:POS

[BDRC16] Peter Benner, Tobias Damm, Martin Redmann, and Yolanda Rocio Rodriguez Cruz. Positive operators and stable truncation. Linear Algebra and its Applications, 498 (??):74–87, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Bomze:2015:SCP

[BDS15] Immanuel M. Bomze, Peter J. C. Dickinson, and Georg Still. The structure of completely positive matrices according to their CP-rank and CP-plus-rank. Linear Algebra and its Applications, 482(??): 191–206, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003304.

Braga:2015:TDN

[BDVRT15] Rodrigo O. Braga, Renata R. Del-Vecchio, Virgínia M. Rodrigues, and Vilmar Trevisan. Trees with 4 or 5 distinct normalized Laplacian eigenvalues. Linear Algebra and its Applications, 471 (??):615–635, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000476.

Benbouziane:2015:MMC

[BE15] Hassane Benbouziane and Mustapha Ech-Cherif El Kettani. Maps on matrices compressing the local spectrum in the spectrum. Linear Algebra and its Applications, 475(??): 176–185, June 15, 2015.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001214.

Bourget:2016:GAP

[BEF16] A. Bourget, C. Espinosa, and S. Faulkinbury. A generalization of the Avram– Parter Theorem. Linear Algebra and its Applications, 492(??):98–110, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006795.

Behrends:2017:WMM

[Beh17] Ehrhard Behrends. Where is matrix multiplication locally open? Linear Algebra and its Applications, 517(??): 167–176, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306036.

Bekjan:2016:SCL

[Bek16] Turdebek N. Bekjan. A submajorization of Carlen and Lieb convexity. Linear Algebra and its Applications, 494(??):23–31, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007636.

Bhattacharjee:2017:DWS

[BEKS17] M. Bhattacharjee, J. Eschmeier, Dinesh K. Keshari, and Jaydeb Sarkar. Dilations, wandering subspaces, and inner functions. Linear Algebra and its Applications, 523(??): 263–280, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301337.

Belozyorov:2016:NSM

[Bel16] Vasiliy Ye. Belozyorov. New solution method of linear static output feedback design problem for linear control systems. Linear Algebra and its Applications, 504(??):204–227, September 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516300751.

Bendaoud:2015:PLS

[Ben15] Mohamed Bendaoud. Preservers of local spectrum of matrix Jordan triple products. Linear Algebra and its Applications, 471(??): 604–614, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000518.

Benediktovich:2016:SCH

[Ben16] Vladimir I. Benediktovich. Spectral condition for Hamiltonicity of a graph. Linear Algebra and its Applications, 494(??):70–79, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000070.

Benvenuti:2018:NFD

[Ben18] Luca Benvenuti. The NIEP for four dimensional Leslie and doubly stochastic matrices with zero trace from the coefficients of the characteristic polynomial. Linear Algebra and its Applications, 544(??):286–298, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518300235.

Bernstein:2017:CTM

[Ber17] Daniel Irving Bernstein. Completion of tree metrics and rank 2 matrices. Linear Algebra and its Applications, 533(??): 1–13, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304366.

Bergqvist:2018:CEB

[Ber18] Göran Bergqvist. Curves and envelopes that bound the spectrum of a matrix. Linear Algebra and its Applications, 557(??): 1–21, November 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303537.

Bart:2016:SIL

[BES16] Harm Bart, Torsten Ehrhardt, and Bernd Silbermann. Sums of idempotents and logarithmic residues in zero pattern matrix algebras. Linear Algebra and its Applications, 498(??):262–316, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515004000.

Bart:2019:URT

[BES19] Harm Bart, Torsten Ehrhardt, and Bernd Silbermann. Unions of rank/trace complete preorders. Linear Algebra and its Applications, 570 (??):181–224, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300503.

Brouwer:2015:DRG

[BF15] A. E. Brouwer and M. A. Fiol. Distance-regular graphs where the distanced graph has fewer distinct eigenvalues. Linear Algebra and its Applications, 480(??):115–126, September 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500261X.

Benvenuti:2017:RLR

[BF17a] Luca Benvenuti and Lorenzo Farina. Revisiting the linear recursions with nonnegative coefficients problem. Linear Algebra and its Applications, 535(??):191– 194, December 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305268.

Bottcher:2017:ASL

[BF17b] Albrecht Böttcher and Lenny Fukshansky. Addendum to “Lattices from equiangular tight frames” [Linear Algebra Appl. 510 (2016) 395– 420]. Linear Algebra and its Applications, 531(??): 592–601, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/

science/article/pii/S0024379517303877.

See [BFG^{+}16].

Baklouti:2018:JSR

[BF18a] Hamadi Baklouti and Kais Feki. On joint spectral radius of commuting operators in Hilbert spaces. Linear Algebra and its Applications, 557(??):455–463, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303963.

Bebiano:2018:SDN

[BF18b] Natália Bebiano and Susana Furtado. Structured distance to normality of tridiagonal matrices. Linear Algebra and its Applications, 552(??):239–255, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302179.

Bebiano:2019:RAS

[BF19a] Natália Bebiano and Susana Furtado. A reducing approach for symmetrically sparse banded and anti-banded matrices. Linear Algebra and its Applications, 581(??):36–50, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302873.

Bigeni:2019:SDC

[BF19b] Ange Bigeni and Evgeny Feigin. Symmetric Dellac configurations and symplectic/orthogonal flag varieties. Linear Algebra and its Applications, 573(??):54– 79, July 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301168.

Bluhm:2019:DRS

[BF19c] Andreas Bluhm and Daniel Stilck França. Dimensionality reduction of SDPs through sketching. Linear Algebra and its Applications, 563(??):461–475, February 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518305354.

Baklouti:2018:JNR

[BFA18] Hamadi Baklouti, Kais Feki, and Ould Ahmed Mahmoud Sid Ahmed. Joint numerical ranges of operators in semi-Hilbertian spaces. Linear Algebra and its Applications, 555(??):266–284, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Borges:2018:GIU

[BFD18] Alex Ramos Borges, Claudemir Fidelis, and Diogo Diniz. Graded isomorphisms on upper block triangular matrix algebras. Linear Algebra and its Applications, 543 (??):92–105, April 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730681X.

Bailey:2016:EVC

[BFFN16] R. A. Bailey, Sandra S. Ferreira, Dário Ferreira, and Célia Nunes. Estimability of variance components when all model matrices commute. Linear Algebra and its Applications, 492 (??):144–160, March 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006655.

Bottcher:2016:LET

[BFG^{+}16] Albrecht Böttcher, Lenny Fukshansky, Stephan Ramon Garcia, Hiren Maharaj, and Deanna Needell. Lattices from equiangular tight frames. Linear Algebra and its Applications, 510(??): 395–420, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303974.

See addendum [BF17b].

Bapat:2017:P

[BFH^{+}17] Ravindra B. Bapat, Shmuel Friedland, John Holbrook, Roger Horn, and Fuad Kittaneh. Preface. Linear Algebra and its Applications, 528(??):1–16, September 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951730157X.

Benzi:2019:GAN

[BFM19] Michele Benzi, Paraskevi Fika, and Marilena Mitrouli. Graphs with absorption: numerical methods for the absorption inverse and the computation of centrality measures. Linear Algebra and its Applications, 574(??):123–152, August 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519301338.

Brugger:2018:ACC

[BFMK18] Matthias Brugger, Maximilian Fiedler, Bernardo González Merino, and Anja Kirschbaum. Additive colourful Carathéodory type results with an application to radii. Linear Algebra and its Applications, 554(??):342–357, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302726.

Barrios:2015:PSC

[BFN15] Jorge Barrios, Orizon P. Ferreira, and Sándor Z. Németh. Projection onto simplicial cones by Picard’s method. Linear Algebra and its Applications, 480(??): 27–43, September 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002608.

Barrett:2017:EDG

[BFW17] Wayne Barrett, Amanda Francis, and Benjamin Webb. Equitable decompositions of graphs with symmetries. Linear Algebra and its Applications, 513(??): 409–434, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304906.

Benner:2018:SRC

[BFY18] Peter Benner, Heike Faßbender, and Chao Yang. Some remarks on the complex J-symmetric eigenproblem. Linear Algebra and its Applications, 544(??):407–442, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518300211.

Bandtlow:2015:EUB

[BG15] Oscar F. Bandtlow and Ayse Güven. Explicit upper bounds for the spectral distance of two trace class operators. Linear Algebra and its Applications, 466(??): 329–342, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006776.

Bernal-Gonzalez:2019:LCS

[BGCMASS19] L. Bernal-González, J. A. Conejero, M. Murillo-Arcila, and J. B. Seoane-Sepúlveda. [S]-linear and convex structures in function families. Linear Algebra and its Applications, 579(??):463–483, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302848.

Bernal-Gonzalez:2016:HFN

[BGCMLSSS16] L. Bernal-González, M. C. Calderón-Moreno, J. López-Salazar, and J. B. Seoane-Sepúlveda. Holomorphic functions on non-Runge domains and related problems. Linear Algebra and its Applications, 511(??):226– 236, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304207.

Bernal-Gonzalez:2015:SSF

[BGCMPB15] L. Bernal-González, M. C. Calderón-Moreno, and J. A. Prado-Bassas. The set of space-filling curves: Topological and algebraic structure. Linear Algebra and its Applications, 467(??): 57–74, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007381.

Bernal-Gonzalez:2017:IDH

[BGdCCMSS17] Luis Bernal-González, María del Carmen Calderón-Moreno, and Juan Benigno Seoane-Sepúlveda. Infinite dimensional holomorphic nonextendability and algebraic genericity. Linear Algebra and its Applications, 513(??):149–159, January 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516304815.

Bru:2016:CMS

[BGGS16] Rafael Bru, Maria T. Gassó, Isabel Giménez, and Máximo Santana. Combined matrices of sign regular matrices. Linear Algebra and its Applications, 498(??):88–98, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514008003.

Belton:2019:SKM

[BGKP19] Alexander Belton, Dominique Guillot, Apoorva Khare, and Mihai Putinar. Simultaneous kernels of matrix Hadamard powers. Linear Algebra and its Applications, 576(??):142– 157, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301459.

Batalshchikov:2019:AEL

[BGM^{+}19] A. A. Batalshchikov, S. M. Grudsky, I. S. Malisheva, S. S. Mihalkovich, E. Ramírez de Arellano, and V. A. Stukopin. Asymptotics of eigenvalues of large symmetric Toeplitz matrices with smooth simple-loop symbols. Linear Algebra and its Applications, 580(??): 292–335, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Barg:2015:FTD

[BGOY15] Alexander Barg, Alexey Glazyrin, Kasso A. Okoudjou, and Wei-Hsuan Yu. Finite two-distance tight frames. Linear Algebra and its Applications, 475(??): 163–175, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001263.

Batalshchikov:2015:AES

[BGS15] A. A. Batalshchikov, S. M. Grudsky, and V. A. Stukopin. Asymptotics of eigenvalues of symmetric Toeplitz band matrices. Linear Algebra and its Applications, 469(??): 464–486, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007691.

Bovdi:2018:RPS

[BGSS18] Victor A. Bovdi, Tatiana G. Gerasimova, Mohamed A. Salim, and Vladimir V. Sergeichuk. Reduction of a pair of skew-symmetric matrices to its canonical form under congruence. Linear Algebra and its Applications, 543(??):17–30, April 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517306833.

Berger:2019:GDS

[BGT^{+}19] Thomas Berger, Hannes Gernandt, Carsten Trunk, Henrik Winkler, and Michal Wojtylak. The gap distance to the set of singular matrix pencils. Linear Algebra and its Applications, 564 (??):28–57, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305433.

Baseilhac:2017:CTP

[BGV17] P. Baseilhac, A. M. Gainutdinov, and T. T. Vu. Cyclic tridiagonal pairs, higher order Onsager algebras and orthogonal polynomials. Linear Algebra and its Applications, 522 (??):71–110, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730085X.

Bodmann:2016:AOB

[BH16a] Bernhard G. Bodmann and John Haas. Achieving the orthoplex bound and constructing weighted complex projective 2-designs with Singer sets. Linear Algebra and its Applications, 511(??):54–71, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303949.

Butler:2016:CFG

[BH16b] Steve Butler and Kristin Heysse. A cospectral family of graphs for the normalized Laplacian found by toggling. Linear Algebra and its Applications, 507(??): 499–512, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302506.

Bolten:2016:SMA

[BHK16] Matthias Bolten, Thomas K. Huckle, and Christos D. Kravvaritis. Sparse matrix approximations for multigrid methods. Linear Algebra and its Applications, 502 (??):58–76, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006710.

Batzies:2015:GMG

[BHML15] E. Batzies, K. Hüper, L. Machado, and F. Silva Leite. Geometric mean and geodesic regression on Grassmannians. Linear Algebra and its Applications, 466(??):83–101, February 1, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379514006508.

Bose:2017:SCN

[BHP17] Snehasish Bose, Sk Monowar Hossein, and Kallol Paul. Solution of a class of nonlinear matrix equations. Linear Algebra and its Applications, 530(??):109–126, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730294X.

Bovdi:2018:SSP

[BHSS18] Victor A. Bovdi, Roger A. Horn, Mohamed A. Salim, and Vladimir V. Sergeichuk. Symplectic spaces and pairs of symmetric and nonsingular skew-symmetric matrices under congruence. Linear Algebra and its Applications, 537(??):84–99, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305608.

Bier:2015:LCS

[Bie15] Agnieszka Bier. On lattices of closed subgroups in the group of infinite triangular matrices over a field. Linear Algebra and its Applications, 485(??):132–152, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004462.

Bik:2019:NCL

[Bik19] Arthur Bik. Noetherianity up to conjugation of locally diagonal inverse limits. Linear Algebra and its Applications, 582(??): 237–290, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303337.

Bilokopytov:2019:PMA

[Bil19] Eugene Bilokopytov. Principal minor assignment, isometries of Hilbert spaces, volumes of parallelepipeds and rescaling of sesquiholomorphic functions. Linear Algebra and its Applications, 580(??):37–61, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302629.

Botelho:2015:AIP

[BJ15] Fernanda Botelho and James Jamison. Algebraic and invariance properties of the group of isometries. Linear Algebra and its Applications, 471(??):531–543, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514008210.

Baeth:2016:FBT

[BJ16] Nicholas R. Baeth and Joel Jeffries. Factorizations of block triangular matrices. Linear Algebra and its Applications, 511(??):403–420, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304116.

Barany:2015:HTT

[BJC15] Imre Bárány and Jesús Jerónimo-Castro. Helly type theorems for the sum of vectors in a normed plane. Linear Algebra and its Applications, 469(??):39– 50, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007538.

Bapat:2017:OIC

[BJKR17] Ravindra B. Bapat, Surender Kumar Jain, K. Manjunatha Prasad Karantha, and M. David Raj. Outer inverses: Characterization and applications. Linear Algebra and its Applications, 528(??):171–184, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302622.

Bhatia:2019:IWM

[BJL19] Rajendra Bhatia, Tanvi Jain, and Yongdo Lim. Inequalities for the Wasserstein mean of positive definite matrices. Linear Algebra and its Applications, 576(??):108–123, September 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518301277.

Bu:2017:BTE

[BJLD17] Changjiang Bu, Xiuquan Jin, Haifeng Li, and Chunli Deng. Brauer-type eigenvalue inclusion sets and the spectral radius of tensors. Linear Algebra and its Applications, 512(??): 234–248, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630458X.

Berg:2018:ETP

[BJS18] Sören Berg, Katharina Jochemko, and Laura Silverstein. Ehrhart tensor polynomials. Linear Algebra and its Applications, 539(??): 72–93, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306109.

Bahmani:2015:MAE

[BK15] Asghar Bahmani and Dariush Kiani. On the multiplicity of the adjacency eigenvalues of graphs. Linear Algebra and its Applications, 477(??):1–20, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001688.

Bahmani:2016:GRT

[BK16] Asghar Bahmani and Dariush Kiani. Graph reduction techniques and the multiplicity of the Laplacian eigenvalues. Linear Algebra and its Applications, 503(??): 215–232, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300933.

Bang:2017:DRG

[BK17a] Sejeong Bang and Jack Koolen. Distance-regular graphs of diameter 3 having eigenvalue −1. Linear Algebra and its Applications, 531(??):38–53, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303385.

Bhatnagar:2017:SD

[BK17b] Gaurav Bhatnagar and Christian Krattenthaler. Spiral determinants. Linear Algebra and its Applications, 529(??):374–390, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302719.

Breen:2017:MLM

[BK17c] Jane Breen and Steve Kirkland. Minimising the largest mean first passage time of a Markov chain: the influence of directed graphs. Linear Algebra and its Applications, 520(??):306–334, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517300630.

Bapat:2018:ICM

[BK18] Ravindra B. Bapat and Masoud Karimi. Integral complete multipartite graphs. Linear Algebra and its Applications, 549 (??):1–11, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301368.

Bapat:2019:CPE

[BK19] Ravindra B. Bapat and Hiroshi Kurata. On Cartesian product of Euclidean distance matrices. Linear Algebra and its Applications, 562(??):135–153, February 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304828.

Bazn:2016:SMB

[BKLP16] Fermín S. V. Bazán, Andreas Kleefeld, Koung Hee Leem, and George Pelekanos. Sampling method based projection approach for the reconstruction of 3D acoustically penetrable scatterers. Linear Algebra and its Applications, 495(??): 289–323, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007570.

Bapat:2017:CQF

[BKNS17] R. B. Bapat, D. Kalita, M. Nath, and D. Sarma. Convex and quasiconvex functions on trees and their applications. Linear Algebra and its Applications, 533(??):210–234, November 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517304226.

Bapat:2018:OIJ

[BKNS18] Ravindra B. Bapat, Manjunatha Prasad Karantha, Nupur Nandini, and Divya P. Shenoy. Outer inverses and Jacobi type identities. Linear Algebra and its Applications, 536(??): 274–294, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305591.

Burdak:2017:CI

[BKPS17] Zbigniew Burdak, Marek Kosiek, Patryk Pagacz, and Marek Slociński. On the commuting isometries. Linear Algebra and its Applications, 516(??):167–185, March 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305638.

Bahturin:2018:CIG

[BKRE18] Yuri Bahturin, Mikhail Kochetov, and Adrián Rodrigo-Escudero. Classification of involutions on gradeddivision simple real algebras. Linear Algebra and its Applications, 546(??):1– 36, June 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300557.

Burdak:2015:CPC

[BKS15] Zbigniew Burdak, Marek Kosiek, and Marek Slociński. Compatible pairs of commuting isometries. Linear Algebra and its Applications, 479(??):216–259, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002384.

See corrigendum [BKSP23].

Burdak:2023:CCP

[BKSP23] Zbigniew Burdak, Marek Kosiek, Marek Słociński, and Patryk Pagacz. Corrigendum to “Compatible pairs of commuting isometries” [Linear Algebra Appl. 479 (2015) 216–259]. Linear Algebra and its Applications, 675(??):106–117, October 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379523002367.

See [BKS15].

Bardsley:2015:DBA

[BL15] Johnathan M. Bardsley and Aaron Luttman. Dealing with boundary artifacts in MCMC-based deconvolution. Linear Algebra and its Applications, 473 (??):339–358, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006089.

Bremigan:2016:MOR

[BL16] Ralph Bremigan and John Lorch. Mutually orthogonal rectangular gerechte designs. Linear Algebra and its Applications, 497 (??):44–61, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001191.

Bai:2018:BSR

[BL18a] Shuliang Bai and Linyuan Lu. A bound on the spectral radius of hypergraphs with e edges. Linear Algebra and its Applications, 549 (??):203–218, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830140X.

Bai:2018:SRT

[BL18b] Shuliang Bai and Linyuan Lu. Spectral radius of 0,1-tensor with prescribed number of ones. Linear Algebra and its Applications, 558(??):205–235, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304026.

Bourhim:2018:MLS

[BL18c] Abdellatif Bourhim and Ji Eun Lee. Multiplicatively local spectrum-preserving maps. Linear Algebra and its Applications, 549(??):291– 308, July 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301538.

Bourin:2019:RDT

[BL19] Jean-Christophe Bourin and Eun-Young Lee. On the Russo–Dye Theorem for positive linear maps. Linear Algebra and its Applications, 571(??):92–102, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519300795.

Barone:2016:EJF

[BLdS16] M. Barone, J. B. Lima, and R. M. Campello de Souza. The eigenstructure and Jordan form of the Fourier transform over fields of characteristic 2 and a generalized Vandermonde-type formula. Linear Algebra and its Applications, 494(??):245– 262, April 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007582.

Bessades:2019:MAI

[BLdSV19] D. Bessades, G. Leal, R. B. dos Santos, and A. C. Vieira. Matrix algebras with involution and standard polynomial identities in symmetric variables. Linear Algebra and its Applications, 582(??):467–481, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303507.

Behrndt:2015:EFR

[BLPT15] Jussi Behrndt, Leslie Leben, Francisco Martínez Pería, and Carsten Trunk. The effect of finite rank perturbations on Jordan chains of linear operators. Linear Algebra and its Applications, 479(??):118–130, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002220.

Bhatia:2016:SNI

[BLY16] Rajendra Bhatia, Yongdo Lim, and Takeaki Yamazaki. Some norm inequalities for matrix means. Linear Algebra and its Applications, 501 (??):112–122, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300040.

See comment [DDF17a].

Bourhim:2015:MPL

[BM15] Abdellatif Bourhim and Mohamed Mabrouk. Maps preserving the local spectrum of Jordan product of matrices. Linear Algebra and its Applications, 484(??): 379–395, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003973.

Banerjee:2016:ELT

[BM16a] Anirban Banerjee and Ranjit Mehatari. An eigenvalue localization theorem for stochastic matrices and its application to Randić matrices. Linear Algebra and its Applications, 505(??):85– 96, September 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301227.

Breaz:2016:NCC

[BM16b] Simion Breaz and George Ciprian Modoi. Nil-clean companion matrices. Linear Algebra and its Applications, 489(??): 50–60, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500600X.

Bai:2017:LQC

[BM17a] Zhong-Zhi Bai and Cun-Qiang Miao. On local quadratic convergence of inexact simplified Jacobi– Davidson method. Linear Algebra and its Applications, 520(??):215–241, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517300460.

Banerjee:2017:NST

[BM17b] Anirban Banerjee and Ranjit Mehatari. On the normalized spectrum of threshold graphs. Linear Algebra and its Applications, 530(??): 288–304, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302951.

Botelho:2017:SNL

[BM17c] Geraldo Botelho and Ximena Mujica. Spaces of σ(p)-nuclear linear and multilinear operators on Banach spaces and their duals. Linear Algebra and its Applications, 519(??): 219–237, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300150.

Baker:2018:LED

[BM18] Charles E. Baker and Boris S. Mityagin. Localization of eigenvalues of doubly cyclic matrices. Linear Algebra and its Applications, 540 (??):160–202, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306444.

Bamberg:2019:ISP

[BM19a] John Bamberg and Klaus Metsch. On intriguing sets of the Penttila–Williford association scheme. Linear Algebra and its Applications, 582(??):327–345, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303398.

Benasseni:2019:IEA

[BM19b] Jacques Bénasséni and Alain Mom. Inequalities for the eigenvectors associated to extremal eigenvalues in rank one perturbations of symmetric matrices. Linear Algebra and its Applications, 570(??):123–137, June 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951930045X.

Bankoussou-mabiala:2017:MDW

[BmBCC17] Edward Bankoussou-mabiala, Abderrahim Boussaïri, Abdelhak Chaïchaâ, and Brahim Chergui. A matrix description of weakly bipartitive and bipartitive families. Linear Algebra and its Applications, 533(??):186– 194, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304457.

Burgos:2015:MPS

[BMGP15] M. Burgos, A. C. Márquez-García, and P. Patrício. On mappings preserving the sharp and star orders. Linear Algebra and its Applications, 483(??):268–292, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500292X.

Bermudez:2016:WSC

[BMN16] Teresa Bermúdez, Antonio Martinón, and Juan Agustín Noda. Weighted shift and composition operators on l_{p} which are (m, q)-isometries. Linear Algebra and its Applications, 505(??):152–173, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301318.

Bourles:2015:WES

[BMO15] H. Bourlès, B. Marinescu, and U. Oberst. Weak exponential stability of linear time-varying differential behaviors. Linear Algebra and its Applications, 486(??): 523–571, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005121.

Bini:2017:DDS

[BMR17] Dario A. Bini, Stefano Massei, and Leonardo Robol. On the decay of the off-diagonal singular values in cyclic reduction. Linear Algebra and its Applications, 519 (??):27–53, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306164.

Benac:2015:AOD

[BMS15] María J. Benac, Pedro G. Massey, and Demetrio Stojanoff. Aliasing and oblique dual pair designs for consistent sampling. Linear Algebra and its Applications, 487(??):112–145, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005200.

Bellalij:2016:DEK

[BMS16a] M. Bellalij, G. Meurant, and H. Sadok. The distance of an eigenvector to a Krylov subspace and the convergence of the Arnoldi method for eigenvalue problems. Linear Algebra and its Applications, 504(??): 387–405, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301173.

Brualdi:2016:HS

[BMS16b] Richard A. Brualdi, Volker Mehrmann, and Peter Semrl. Hans Schneider, 1927–2015. Linear Algebra and its Applications, 498(??):1–2, June 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516001658.

Bini:2016:SNL

[BMVW16] Dario Bini, Marilena Mitrouli, Marc Van Barel, and Joab Winkler. Structured numerical linear and multilinear algebra: Analysis, algorithms and applications. Linear Algebra and its Applications, 502(??):1–4, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300659.

See corrigendum [Ano16a].

Brown:2018:CFG

[BMZB18] Morgan V. Brown, Jackson S. Morrow, and David Zureick-Brown. Chip-firing groups of iterated cones. Linear Algebra and its Applications, 556(??):46–54, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303070.

Benoist:2017:BUM

[BN17] Tristan Benoist and Ion Nechita. On bipartite unitary matrices generating subalgebra-preserving quantum operations. Linear Algebra and its Applications, 521 (??):70–103, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300484.

Borisov:2017:SSL

[BNST17] Lev Borisov, Patrizio Neff, Suvrit Sra, and Christian Thiel. The sum of squared logarithms inequality in arbitrary dimensions. Linear Algebra and its Applications, 528(??):124–146, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302439.

Bryant:2015:AER

[BÓ15] Darryn Bryant and Padraig Ó Catháin. An asymptotic existence result on compressed sensing matrices. Linear Algebra and its Applications, 475(??): 134–150, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001159.

Bukovsek:2017:LSS

[BO17] Damjana Kokol Bukovsek and Matjaz Omladic. Linear spaces of symmetric nilpotent matrices. Linear Algebra and its Applications, 530(??):384–404, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303300.

Bobrowski:2016:HTA

[Bob16] Adam Bobrowski. On Hilletype approximation of degenerate semigroups of operators. Linear Algebra and its Applications, 511(??): 31–53, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303883.

Borobia:2017:FTI

[Bor17] Alberto Borobia. Fillmore’s theorem for integer matrices. Linear Algebra and its Applications, 531(??): 281–284, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303749.

Botha:2016:PIS

[Bot16a] J. D. Botha. Products of idempotent and square-zero matrices. Linear Algebra and its Applications, 497 (??):116–133, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001282.

Botha:2016:SZF

[Bot16b]  J. D. Botha.  Square-

zero factorization of matrices. Linear Algebra and its Applications, 488(??):

71–85, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005698.

Belardo:2019:SOT

[BOT19] Francesco Belardo, Elismar R. Oliveira, and Vilmar Trevisan. Spectral ordering of trees with small index. Linear Algebra and its Applications, 575(??): 250–272, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301648.

Bourgeois:2015:NGM

[Bou15] Gerald Bourgeois. Nonsymmetric generic matrix equations. Linear Algebra and its Applications, 476 (??):159–183, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001524.

Boumazgour:2016:UEO

[Bou16a] Mohamed Boumazgour. On the S-universal elementary operators. Linear Algebra and its Applications, 507(??):274–287, October 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516302269.

Bounds:2016:CMR

[Bou16b] Jordan Bounds. Commuting maps over the ring of strictly upper triangular matrices. Linear Algebra and its Applications, 507(??): 132–136, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302166.

Berliner:2017:SRI

[BOvdD17] Adam H. Berliner, D. D. Olesky, and P. van den Driessche. Sets of refined inertias of zero-nonzero patterns. Linear Algebra and its Applications, 516(??): 243–263, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630578X.

Belardo:2015:SCS

[BP15] Francesco Belardo and Pawel Petecki. Spectral characterizations of signed lollipop graphs. Linear Algebra and its Applications, 480(??): 144–167, September 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002633.

Balaji:2016:LCS

[BP16a] R. Balaji and K. Palpandi. On the Lipschitz continuity of the solution map in linear complementarity problems over second-order cone. Linear Algebra and its Applications, 510(??):146–159, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303561.

Barreras:2016:STP

[BP16b] A. Barreras and J. M. Peña. Similarity to totally positive matrices and accurate computations. Linear Algebra and its Applications, 491(??):317–327, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005236.

Bittens:2019:RSF

[BP19a] Sina Bittens and Gerlind Plonka. Real sparse fast DCT for vectors with short support. Linear Algebra and its Applications, 582(??): 359–390, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303404.

Bondarenko:2019:WPC

[BP19b] Vitalij M. Bondarenko and Anatoliy P. Petravchuk. Wildness of the problem of classifying nilpotent Lie algebras of vector fields in four variables. Linear Algebra and its Applications, 568 (??):165–172, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303598.

Bapat:2016:SRE

[BPP16] R. B. Bapat, S. K. Panda, and S. Pati. Strong reciprocal eigenvalue property of a class of weighted graphs. Linear Algebra and its Applications, 511(??):460–475, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304578.

Bapat:2017:SIU

[BPP17] R. B. Bapat, S. K. Panda, and S. Pati. Self-inverse unicyclic graphs and strong reciprocal eigenvalue property. Linear Algebra and its Applications, 531(??): 459–478, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303695.

Bondarenko:2017:NES

[BPR17] A. V. Bondarenko, A. Prymak, and D. Radchenko. Non-existence of (76, 30, 8, 14) strongly regular graph. Linear Algebra and its Applications, 527(??):53–72, August 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302136.

Burman:2015:MTT

[BPT15] Yurii Burman, Andrey Ploskonosov, and Anastasia Trofimova. Matrix-tree theorems and discrete path integration. Linear Algebra and its Applications, 466(??): 64–82, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006466.

Baragana:2017:GRM

[BPZ17] I. Baragaña, F. Puerta, and I. Zaballa. On the geometry of realizable Markov parameters by SIMO and MISO systems. Linear Algebra and its Applications, 518 (??):97–143, April 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306310.

Banaji:2015:SRS

[BR15] Murad Banaji and Carrie Rutherford. Some results on the structure and spectra of matrix-products. Linear Algebra and its Applications, 474(??):192–212, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515001020.

Bini:2016:CMP

[BR16a] Dario A. Bini and Leonardo Robol. On a class of matrix pencils and l-ifications equivalent to a given matrix polynomial. Linear Algebra and its Applications, 502(??):275–298, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515004310.

Bini:2016:QHR

[BR16b] Dario A. Bini and Leonardo Robol. Quasiseparable Hessenberg reduction of real diagonal plus low rank matrices and applications. Linear Algebra and its Applications, 502(??):186–213, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004942.

Bunger:2017:YME

[BR17] F. Bünger and S. M. Rump. Yet more elementary proofs that the determinant of a symplectic matrix is 1. Linear Algebra and its Applications, 515(??):87–95, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305304.

Bruhn:2018:MDC

[BR18] Henning Bruhn and Dieter Rautenbach. Maximal determinants of combinatorial matrices. Linear Algebra and its Applications, 553(??):37– 57, September 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302246.

Baragana:2019:LFS

[BR19] Itziar Baragaña and Alicia Roca. Linear feedback shift registers and the minimal realization problem. Linear Algebra and its Applications, 576(??):200– 227, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302933.

Bresar:2017:DTP

[Bre17] Matej Bresar. Derivations of tensor products of nonassociative algebras. Linear Algebra and its Applications, 530(??):244–252, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303105.

Breaz:2018:MFF

[Bre18]  Simion Breaz.  Matrices

over finite fields as sums of periodic and nilpotent elements. Linear Algebra and its Applications, 555(??): 92–97, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830301X.

Briat:2017:SPM

[Bri17] Corentin Briat. Sign properties of Metzler matrices with applications. Linear Algebra and its Applications, 515(??):53–86, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305274.

Bien:2018:ESS

[BRN18] M. H. Bien and M. Ramezan-Nassab. Engel subnormal subgroups of skew linear groups. Linear Algebra and its Applications, 558(??): 74–78, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304014.

Brualdi:2015:EC

[Bru15] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 465(??):430, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379514006545.

Brualdi:2016:IAM

[Bru16a] Richard A. Brualdi. Book review: An Introduction to Applied Matrix Analysis, vol. 20, Xiao-Qing Jin, Seak-Weng Vong, in: Series in Contemporary Applied Mathematics CAM. World Scientific Publishing Co., New Jersey (2016), ISBN: 978-981-4749-46-6. Linear Algebra and its Applications, 508(??):304–305, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630355X.

Brualdi:2016:EC

[Bru16b] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 488(??):522–523, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006229.

Brualdi:2017:EC

[Bru17] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 512(??):331–332, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304980.

Banica:2015:QAP

[BS15a] Teo Banica and Adam Skalski. The quantum algebra of partial Hadamard matrices. Linear Algebra and its Applications, 469(??): 364–380, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951400768X.

Bapat:2015:PDM

[BS15b] R. B. Bapat and Sivaramakrishnan Sivasubramanian. Product distance matrix of a tree with matrix weights. Linear Algebra and its Applications, 468 (??):145–153, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514001785.

Belardo:2015:LCS

[BS15c] Francesco Belardo and Slobodan K. Simić. On the Laplacian coefficients of signed graphs. Linear Algebra and its Applications, 475 (??):94–113, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000890.

Berndt:2015:HPI

[BS15d] Wolfgang Berndt and Suvrit Sra. Hlawka–Popoviciu inequalities on positive definite tensors. Linear Algebra and its Applications, 486(??): 317–327, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004966.

Bapat:2016:SDM

[BS16a] R. B. Bapat and Sivaramakrishnan Sivasubramanian. Squared distance matrix of a tree: Inverse and inertia. Linear Algebra and its Applications, 491(??): 328–342, February 15, 2016.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005212.

Beasley:2016:PEP

[BS16b] LeRoy B. Beasley and Seok-Zun Song. Primitive exponent preservers. Linear Algebra and its Applications, 491(??):263–275, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004747.

Bhatia:2016:PLM

[BS16c] Rajendra Bhatia and Rajesh Sharma. Positive linear maps and spreads of matrices — II. Linear Algebra and its Applications, 491(??): 30–40, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006995.

Blekherman:2016:RRR

[BS16d] Grigoriy Blekherman and Rainer Sinn. Real rank with respect to varieties. Linear Algebra and its Applications, 505(??):344–360, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Bonin:2016:IFE

[BS16e] Joseph E. Bonin and Thomas J. Savitsky. An infinite family of excluded minors for strong base-orderability. Linear Algebra and its Applications, 488(??):396–429, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515005935.

Boyle:2016:SSE

[BS16f] Mike Boyle and Scott Schmieding. Strong shift equivalence and the Generalized Spectral Conjecture for nonnegative matrices. Linear Algebra and its Applications, 498(??):231– 243, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003572.

Barik:2017:LSS

[BS17] Sasmita Barik and Gopinath Sahoo. On the Laplacian spectra of some variants of corona. Linear Algebra and its Applications, 512 (??):32–47, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304293.

Baricz:2018:HTT

[BS18a] Árpád Baricz and Frantisek Stampach. The Hurwitztype theorem for the regular Coulomb wave function via Hankel determinants. Linear Algebra and its Applications, 548(??):259– 272, July 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301228.

Brambilla:2018:ABA

[BS18b] Maria Chiara Brambilla and Giovanni Staglianò. On the algebraic boundaries among typical ranks for real binary forms. Linear Algebra and its Applications, 557(??):403–418, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303720.

Bunimovich:2018:GEI

[BS18c] Leonid Bunimovich and Longmei Shu. Generalized eigenvectors of isospectral transformations, spectral equivalence and reconstruction of original networks. Linear Algebra and its Applications, 551(??): 104–124, August 15, 2018.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830185X.

Blum-Smith:2019:RGW

[BS19a] Ben Blum-Smith. A rotation group whose subspace arrangement is not from a real reflection group. Linear Algebra and its Applications, 581(??):405–412, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303179.

Bovdi:2019:CBD

[BS19b] Victor A. Bovdi and Volodymyr P. Shchedryk. Commutative Bezout domains of stable range 1.5. Linear Algebra and its Applications, 568 (??):127–134, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302969.

Bresar:2019:CCF

[BS19c] Matej Bresar and Peter Semrl. Continuous commuting functions on matrix algebras. Linear Algebra and its Applications, 568(??):29–38, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301423.

Berman:2019:CMG

[BSMSZ19] Abraham Berman, Naomi Shaked-Monderer, Ranveer Singh, and Xiao-Dong Zhang. Complete multipartite graphs that are determined, up to switching, by their Seidel spectrum. Linear Algebra and its Applications, 564 (??):58–71, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305457.

Bhat:2018:BSG

[BSP18] Mushtaq A. Bhat, U. Samee, and S. Pirzada. Bicyclic signed graphs with minimal and second minimal energy. Linear Algebra and its Applications, 551(??): 18–35, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301782.

Belardo:2016:ESC

[BSS16] Francesco Belardo, Irene Sciriha, and Slobodan K. Simić. On eigenspaces of some compound signed graphs. Linear Algebra and its Applications, 509(??): 19–39, November 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Bovdi:2017:NRE

[BSS17] Victor A. Bovdi, Mohammed A. Salim, and Vladimir V. Sergeichuk. Neighborhood radius estimation for Arnold’s miniversal deformations of complex and p-adic matrices. Linear Algebra and its Applications, 512(??):97–112, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304256.

Belinschi:2018:ENH

[BŚS18] Serban T. Belinschi, Piotr Śniady, and Roland Speicher. Eigenvalues of non-Hermitian random matrices and Brown measure of non-normal operators: Hermitian reduction and linearization method. Linear Algebra and its Applications, 537(??):48–83, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730558X.

Bermudez:2018:IHS

[BSZ18] T. Bermúdez, A. Saddi, and H. Zaway. (A, m)-isometries on Hilbert spaces. Linear Algebra and its Applications, 540(??):95–111, March 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730633X.

Botelho:2015:HIM

[BT15] Geraldo Botelho and Ewerton R. Torres. Hyperideals of multilinear operators. Linear Algebra and its Applications, 482(??):1– 20, October 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003110.

Bai:2017:PRA

[BT17a] Zhong-Zhi Bai and Min Tao. On preconditioned and relaxed AVMM methods for quadratic programming problems with equality constraints. Linear Algebra and its Applications, 516 (??):264–285, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305766.

Baran:2017:RR

[BT17b] Fatma Yesil Baran and Naim Tuglu. q-Riordan representation. Linear Algebra and its Applications, 525 (??):105–117, July 15, 2017.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301805.

Boussaada:2018:FRE

[BTN^{+}18] Islam Boussaada, Sami Tliba, Silviu-Iulian Niculescu, Hakki Ulas Ünal, and Tomás Vyhlídal. Further remarks on the effect of multiple spectral values on the dynamics of time-delay systems. Application to the control of a mechanical system. Linear Algebra and its Applications, 542(??): 589–604, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730650X.

Berger:2016:LRK

[BTW16] Thomas Berger, Carsten Trunk, and Henrik Winkler. Linear relations and the Kronecker canonical form. Linear Algebra and its Applications, 488(??): 13–44, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005601.

Brown:2018:SRS

[BU18] Lawrence G. Brown and Mitsuru Uchiyama. Some results on strongly operator convex functions and operator monotone functions. Linear Algebra and its Applications, 553(??):238–251, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302362.

Bunger:2017:SNR

[Bün17] Florian Bünger. A short note on the ratio between sign-real and sign-complex spectral radius of a real square matrix. Linear Algebra and its Applications, 529(??):126–132, September 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517302598.

Burnik:2015:SPQ

[Bur15] Konrad Burnik. A structurepreserving QR factorization for centrosymmetric real matrices. Linear Algebra and its Applications, 484(??): 356–378, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003997.

Burgos:2018:NMA

[Bur18] J. M. Burgos. A note on Möbius algebras and applications. Linear Algebra and its Applications, 557(??):419–429, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518303707.

Butkovic:2016:TS

[But16] Peter Butkovic. On tropical supereigenvectors. Linear Algebra and its Applications, 498(??):574–591, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516001452.

Bliss:2018:MGN

[BV18a] Nathan Bliss and Jan Verschelde. The method of Gauss–Newton to compute power series solutions of polynomial homotopies. Linear Algebra and its Applications, 542(??):569–588, ???? 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517306110.

Bytsko:2018:TPM

[BV18b] Andrei Bytsko and Alexander Volkov. Totally positive matrices and dilogarithm identities. Linear Algebra and its Applications, 547 (??):148–167, June 15, 2018.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300880.

Bai:2016:NIT

[BW16] Yu Bai and Guangsheng Wei. A new inverse three spectra theorem for Jacobi matrices. Linear Algebra and its Applications, 493(??): 301–312, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007119.

Bai:2018:CRR

[BW18] Zhong-Zhi Bai and Wen-Ting Wu. On convergence rate of the randomized Kaczmarz method. Linear Algebra and its Applications, 553(??):252–269, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302404.

Bai:2019:PRE

[BW19] Zhong-Zhi Bai and Wen-Ting Wu. On partially randomized extended Kaczmarz method for solving large sparse overdetermined inconsistent linear systems. Linear Algebra and its Applications, 578(??):225–250, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302071.

Bu:2015:MMR

[BWSZ15a] Changjiang Bu, Wenzhe Wang, Lizhu Sun, and Jiang Zhou. Minimum (maximum) rank of sign pattern tensors and sign nonsingular tensors. Linear Algebra and its Applications, 483(??): 101–114, October 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003456.

Bu:2015:BTE

[BWSZ15b] Changjiang Bu, Yuanpeng Wei, Lizhu Sun, and Jiang Zhou. Brualdi-type eigenvalue inclusion sets of tensors. Linear Algebra and its Applications, 480(??):168– 175, September 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002876.

Bahturin:2019:GPI

[BY19] Yuri Bahturin and Felipe Yasumura. Graded polynomial identities as identities of universal algebras. Linear Algebra and its Applications, 562 (??):1–14, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304804.

Bahturin:2016:SGD

[BZ16a] Yuri Bahturin and Mikhail Zaicev. Simple graded division algebras over the field of real numbers. Linear Algebra and its Applications, 490(??):102–123, February 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006163.

Balan:2016:LAL

[BZ16b] Radu Balan and Dongmian Zou. On Lipschitz analysis and Lipschitz synthesis for the phase retrieval problem. Linear Algebra and its Applications, 496 (??):152–181, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000586.

Belardo:2016:SGE

[BZ16c] Francesco Belardo and Yue Zhou. Signed graphs with extremal least Laplacian eigenvalue. Linear Algebra and its Applications, 497 (??):167–180, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Choudhary:2018:SAB

[CA18] Rajendra Choudhary and Kapil Ahuja. Stability analysis of Bilinear Iterative Rational Krylov Algorithm. Linear Algebra and its Applications, 538(??): 56–88, February 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730575X.

Cheze:2019:HCC

[CA19] Guillaume Chèze and Luca Amodei. How to cut a cake with a Gram matrix. Linear Algebra and its Applications, 560(??): 114–132, January 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304695.

Calugareanu:2016:SMC

[Cal16] Grigore Calugareanu. Some matrix completions over integral domains. Linear Algebra and its Applications, 507(??):414–419, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302518.

Cariello:2017:GPE

[Car17] D. Cariello. A gap for PPT entanglement. Linear Algebra and its Applications, 529(??):89–114, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302380.

Castillo:2019:MZP

[Cas19] K. Castillo. On monotonicity of zeros of paraorthogonal polynomials on the unit circle. Linear Algebra and its Applications, 580(??): 475–490, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930299X.

Cao:2016:SGL

[CC16] Yan Cao and Liangyun Chen. On the structure of graded Leibniz triple systems. Linear Algebra and its Applications, 496(??):496– 509, May 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000859.

Casazza:2017:FSC

[CC17a] Peter G. Casazza and Xuemei Chen. Frame scalings: a condition number approach. Linear Algebra and its Applications, 523 (??):152–168, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301192.

Chan:2017:LMP

[CC17b] Jor-Ting Chan and Kong Chan. Linear mappings preserving the diameter of the numerical range. Linear Algebra and its Applications, 530(??):220–229, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303063.

Chen:2018:MP

[CC18] Zhi Chen and Lei Cao. On the maximum of the permanent of (I − A). Linear Algebra and its Applications, 555(??):412–431, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830315X.

Cao:2019:PPD

[CC19] Lei Cao and Zhi Chen. Partitions of the polytope of doubly substochastic matrices. Linear Algebra and its Applications, 563(??): 98–122, February 15, 2019.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305123.

Chan:2019:CNN

[CCCNT19] Jor-Ting Chan, Kong Chan, Choi-Nai, and Charlies Tu. The cone of nonnegative c-numerical range and its preservers. Linear Algebra and its Applications, 564 (??):225–235, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305676.

Chen:2019:CPS

[CCD19] Zhiqi Chen, Xueqing Chen, and Ming Ding. On the characteristic polynomial of ∫↕(∈, F). Linear Algebra and its Applications, 579(??): 237–243, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302538.

Cardoso:2018:SSL

[CCdFV18] Domingos M. Cardoso, Paula Carvalho, Maria Aguieiras A. de Freitas, and Cybele T. M. Vinagre. Spectra, signless Laplacian and Laplacian spectra of complementary prisms of graphs. Linear Algebra and its Applications, 544 (??):325–338, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830034X.

Catalisano:2017:WLD

[CCGO17] Maria Virginia Catalisano, Luca Chiantini, Anthony V. Geramita, and Alessandro Oneto. Waring-like decompositions of polynomials,

1. Linear Algebra and its Applications, 533(??):311– 325, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730441X.

Chan:2019:ALM

[CCGV^{+}19] Eunice Y. S. Chan, Robert M. Corless, Laureano Gonzalez-Vega, J. Rafael Sendra, and Juana Sendra. Algebraic linearizations of matrix polynomials. Linear Algebra and its Applications, 563(??): 373–399, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305160.

Caalim:2018:LAA

[CCiT18] Jonathan Caalim, Clarisson Rizzie Canlubo, and Yu ichi Tanaka. On the Lie algebra associated to S-unitary matrices. Linear Algebra and its Applications, 553(??):167–181, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302416.

Chung:2015:ORL

[CCO15] Julianne Chung, Matthias Chung, and Dianne P. O’Leary. Optimal regularized low rank inverse approximation. Linear Algebra and its Applications, 468 (??):260–269, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514004741.

Cahill:2016:GBN

[CCW16] Jameson Cahill, Xuemei Chen, and Rongrong Wang. The gap between the null space property and the restricted isometry property. Linear Algebra and its Applications, 501(??):363– 375, July 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300088.

Chang:2018:SHG

[CCW18] Zhihua Chang, Jin Cheng, and Yongjie Wang. Second homology of generalized periplectic Lie superalgebras. Linear Algebra and its Applications, 546 (??):122–153, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300569.

Cardinali:2016:RPH

[CD16a] Ilaria Cardinali and Bart De Bruyn. Regular partitions of half-spin geometries. Linear Algebra and its Applications, 511(??):203– 225, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304128.

Chen:2016:SRL

[CD16b] Xiaodan Chen and Kinkar Ch. Das. Some results on the Laplacian spread of a graph. Linear Algebra and its Applications, 505(??):245–260, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301513.

Cooper:2016:SPS

[CD16c] Joshua Cooper and Jeffrey Davis. Successful pressing sequences for a bicolored graph and binary matrices. Linear Algebra and its Applications, 490(??): 162–173, February 1, 2016.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006643.

Courtemanche:2016:AEA

[CD16d] Jordan Courtemanche and Manfred Dugas. Automorphisms of the endomorphism algebra of a free module. Linear Algebra and its Applications, 510(??): 79–91, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303548.

Chavez-Dominguez:2019:FPF

[CDFK19] J. A. Chávez-Domínguez, D. Freeman, and K. Kornelson. Frame potential for finite-dimensional Banach spaces. Linear Algebra and its Applications, 578 (??):1–26, October 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302046.

Carriegos:2016:PEM

[CDG16] Miguel V. Carriegos and Noemí DeCastro-García. Partitions of elements in a monoid and its applications to systems theory. Linear Algebra and its Applications, 491(??):161–170, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003535.

Courtemanche:2018:LAF

[CDH18] Jordan Courtemanche, Manfred Dugas, and Daniel Herden. Local automorphisms of finitary incidence algebras. Linear Algebra and its Applications, 541(??): 221–257, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730678X.

Carvalho:2019:BBQ

[CDM19] Luís Carvalho, Cristina Diogo, and Sérgio Mendes. A bridge between quaternionic and complex numerical ranges. Linear Algebra and its Applications, 581(??):496–504, November 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951930309X.

Carando:2015:IFM

[CDMP15] Daniel Carando, Verónica Dimant, Santiago Muro, and Damián Pinasco. An integral formula for multiple summing norms of operators. Linear Algebra and its Applications, 478(??):

274–293, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002074.

Cardoso:2018:DMJ

[CDR18] Domingos M. Cardoso, Roberto C. Díaz, and Oscar Rojo. Distance matrices on the H-join of graphs: a general result and applications. Linear Algebra and its Applications, 559(??): 34–53, December 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304051.

Camenga:2019:SBP

[CDR^{+}19] Kristin A. Camenga, Louis Deaett, Patrick X. Rault, Tsvetanka Sendova, Ilya M. Spitkovsky, and Rebekah B. Johnson Yates. Singularities of base polynomials and Gau–Wu numbers. Linear Algebra and its Applications, 581(??):112– 127, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302861.

Charina:2017:MMG

[CDRT17] M. Charina, M. Donatelli, L. Romani, and V. Turati. Multigrid methods:

Grid transfer operators and subdivision schemes. Linear Algebra and its Applications, 520(??):151–190, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516306140.

Cipolla:2015:AMA

[CDTZ15] S. Cipolla, C. Di Fiore, F. Tudisco, and P. Zellini. Adaptive matrix algebras in unconstrained minimization. Linear Algebra and its Applications, 471(??): 544–568, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000397.

Carmona:2016:GON

[CEGM16] A. Carmona, A. M. Encinas, S. Gago, and M. Mitjana. Green operators of networks with a new vertex. Linear Algebra and its Applications, 491(??): 419–433, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006758.

Carmona:2015:GMA

[CEM15a] A. Carmona, A. M. Encinas, and M. Mitjana. Green matrices associated with generalized linear polyominoes. Linear Algebra and its Applications, 468(??):38– 47, March 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514000238.

Carmona:2015:PDE

[CEM15b] A. Carmona, A. M. Encinas, and M. Mitjana. Perturbations of discrete elliptic operators. Linear Algebra and its Applications, 468 (??):270–285, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007216.

Carmona:2019:RDE

[CEM19] Á. Carmona, A. M. Encinas, and M. Mitjana. Resistance distances in extended or contracted networks. Linear Algebra and its Applications, 576(??): 5–34, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300405.

Chen:2018:TRT

[CF18] Lin Chen and Shmuel Friedland. The tensor rank of tensor product of two three-qubit W states is eight. Linear Algebra and its Applications, 543(??):1– 16, April 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306857.

Cvetkovic:2016:PIC

[CFH^{+}16] Ljiljana Cvetković, Andreas Frommer, Suk-Geun Hwang, Helena Smigoc, and Fuzhen Zhang. Preface to the 19th ILAS Conference Proceedings, Seoul, South Korea 2014. Linear Algebra and its Applications, 491 (??):1–3, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006837.

Cheng:2019:ICG

[CFH19] Tao Cheng, Lihua Feng, and Hualin Huang. Integral Cayley graphs over dicyclic group. Linear Algebra and its Applications, 566 (??):121–137, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300151.

Carlen:2016:SOT

[CFL16] Eric A. Carlen, Rupert L. Frank, and Elliott H. Lieb. Some operator and trace function convexity theorems. Linear Algebra and its Applications, 490(??):174–185, February 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006692.

Corach:2016:OIA

[CFM16] Gustavo Corach, Guillermina Fongi, and Alejandra Maestripieri. Optimal inverses and abstract splines. Linear Algebra and its Applications, 496 (??):182–192, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000574.

Cruz:2017:SSN

[CFNP17] J. Y. Bello Cruz, O. P. Ferreira, S. Z. Németh, and L. F. Prudente. A semismooth Newton method for projection equations and linear complementarity problems with respect to the second order cone. Linear Algebra and its Applications, 513(??):160–181, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304670.

Compta:2018:PMR

[CFP18] Albert Compta, Josep Ferrer, and Marta Peña. Perturbed marked reduced forms of invariant subspaces. Linear Algebra and its Applications, 559(??):194–226, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304427.

Chien:2015:TFC

[CFW15] Tuan-Yow Chien, Victor Flynn, and Shayne Waldron. Tight frames for cyclotomic fields and other rational vector spaces. Linear Algebra and its Applications, 476 (??):98–123, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001275.

Castro:2015:DPT

[CG15] M. Castro and F. A. Grünbaum. The Darboux process and time-and-band limiting for matrix orthogonal polynomials. Linear Algebra and its Applications, 487(??):328–341, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005273.

Cardinali:2016:MDS

[CG16] Ilaria Cardinali and Luca Giuzzi. Minimum distance of symplectic Grassmann codes. Linear Algebra and its Applications, 488(??): 124–134, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005583.

Cardinali:2019:ILH

[CG19a] Ilaria Cardinali and Luca Giuzzi. Implementing line-Hermitian Grassmann codes. Linear Algebra and its Applications, 580(??):96–120, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302721.

Castillo:2019:BFE

[CG19b] Jesús M. F. Castillo and Ricardo García. Bilinear forms and the Ext 2-problem in Banach spaces. Linear Algebra and its Applications, 566 (??):199–211, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300060.

Chmielinski:2019:CDS

[CG19c] Jacek Chmieliński and Moshe Goldberg. Continuity and discontinuity of seminorms on infinite-dimensional vector spaces. Linear Algebra and its Applications, 578(??):153–158, October 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302058.

Castro-Gonzalez:2016:MPT

[CGDM16]  Nieves  Castro-González,

Froilán M. Dopico, and Juan M. Molera. Multiplicative perturbation theory of the Moore–Penrose inverse and the least squares problem. Linear Algebra and its Applications, 503 (??):1–25, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300362.

Coutinho:2015:PST

[CGGV15] G. Coutinho, C. Godsil, K. Guo, and F. Vanhove. Perfect state transfer on distance-regular graphs and association schemes. Linear Algebra and its Applications, 478(??):108–130, August 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001883.

Connelly:2017:UPS

[CGK^{+}17] Erin Connelly, Nathaniel Grammel, Michael Kraut, Luis Serazo, and Christino Tamon. Universality in perfect state transfer. Linear Algebra and its Applications, 531(??):516–532, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303786.

Cho:2016:EPI

[CGL16] Muneo Cho, Caixing Gu, and Woo Young Lee. Elementary properties of ∞- isometries on a Hilbert space. Linear Algebra and its Applications, 511(??):378– 402, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304220.

Contino:2017:WLS

[CGM17] Maximiliano Contino, Juan Giribet, and Alejandra Maestripieri. Weighted least squares solutions of the equation AXB − C =

0. Linear Algebra and its Applications, 518(??):177– 197, April 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306279.

Castro-Gonzalez:2015:EMP

[CGMSR15] N. Castro-González, M. F. Martínez-Serrano, and J. Robles. Expressions for the Moore–Penrose inverse of block matrices involving the Schur complement. Linear Algebra and its Applications, 471(??):353–368, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000324.

Cicone:2019:SCA

[CGSC19] Antonio Cicone, Carlo Garoni, and Stefano Serra-Capizzano. Spectral and convergence analysis of the Discrete ALIF method. Linear Algebra and its Applications, 580(??):62–95, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302733.

Coutinho:2016:ELC

[CGSZ16] G. Coutinho, C. Godsil, H. Shirazi, and H. Zhan. Equiangular lines and covers of the complete graph. Linear Algebra and its Applications, 488(??):264–283, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005443.

Cheng:2019:TGF

[CH19] Chuangxun Cheng and Deguang Han. On twisted group frames. Linear Algebra and its Applications, 569 (??):285–310, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300485.

Chan:2015:NRI

[Cha15] Kong Chan. c-Numerical radius isometries on matrix algebras and triangular matrix algebras. Linear Algebra and its Applications, 466(??): 160–181, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006818.

Chamizo:2019:MPM

[Cha19] Fernando Chamizo. Morphisms and period matrices. Linear Algebra and its Applications, 582(??): 103–113, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303325.

Carriegos:2015:RTS

[CHASSSG15] M. V. Carriegos, J. A. Hermida-Alonso, A. Sáez-Schwedt, and T. Sánchez-Giralda. Rosenbrock’s theorem for systems over von Neumann regular rings. Linear Algebra and its Applications, 482(??):122–130, October 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003262.

Chouh:2015:SNR

[CHB15] Meriem Chouh, Mohamed Hanafi, and Kamel Boukhetala. Semi-nonnegative rank for real matrices and its connection to the usual rank. Linear Algebra and its Applications, 466(??):27–37, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006430.

Chen:2015:DGK

[Che15a] Hao Chen. Distance geometry for kissing spheres. Linear Algebra and its Applications, 479(??):185–201, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002396.

Cheng:2015:CTP

[Che15b] Chuangxun Cheng. A character theory for projective representations of finite groups. Linear Algebra and its Applications, 469(??):

230–242, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007629.

Chebotar:2018:SLP

[Che18a]  Mikhail Chebotar.  On

skew Laurent polynomial rings over locally nilpotent rings. Linear Algebra and its Applications, 549 (??):287–290, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301496.

Chen:2018:IRB

[Che18b] Xiaodan Chen. Improved results on Brouwer’s conjecture for sum of the Laplacian eigenvalues of a graph. Linear Algebra and its Applications, 557(??):327–338, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303756.

Chen:2018:AEA

[Che18c] Xiaodan Chen. On ABC eigenvalues and ABC energy. Linear Algebra and its Applications, 544(??):141– 157, May 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830017X.

Chen:2018:MRR

[Che18d] Yu Chen. Matrix representations of the real numbers. Linear Algebra and its Applications, 536(??): 174–185, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305360.

Chen:2019:BCS

[Che19a] Xiaodan Chen. On Brouwer’s conjecture for the sum of k largest Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 578(??):402–410, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302381.

Chen:2019:EAS

[Che19b] Xiaodan Chen. On extremality of ABC spectral radius of a tree. Linear Algebra and its Applications, 564 (??):159–169, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305585.

Chistopolskaya:2018:NGL

[Chi18] Alisa Chistopolskaya. On nilpotent generators of the Lie algebra sln. Linear Algebra and its Applications, 559(??):73–79, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304324.

Cioaba:2018:CMU

[CHJM18] Sebastian M. Cioaba, Willem H. Haemers, Travis Johnston, and Matt McGinnis. Cospectral mates for the union of some classes in the Johnson association scheme. Linear Algebra and its Applications, 539(??):219–228, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306390.

Cheon:2017:MRG

[CHK17] Gi-Sang Cheon, I-Chiau Huang, and Sooyeong Kim. Multivariate Riordan groups and their representations. Linear Algebra and its Applications, 514(??):198–207, February 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305146.

Catalano:2019:MPP

[CHK19] Louisa Catalano, Samuel Hsu, and Regan Kapalko. On maps preserving products of matrices. Linear Algebra and its Applications, 563(??):193–206, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305172.

Choi:2017:IRP

[Cho17a] Daeshik Choi. Inequalities related to partial transpose and partial trace. Linear Algebra and its Applications, 516(??):1–7, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305535.

Choi:2017:IRT

[Cho17b] Daeshik Choi. Inequalities related to trace and determinant of positive semidefinite block matrices. Linear Algebra and its Applications, 532 (??):1–7, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303890.

Choi:2019:RDM

[Cho19] Michael C. H. Choi. On resistance distance of Markov chain and its sum rules. Linear Algebra and its Applications, 571(??):14–25, June 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951930076X.

Cui:2019:GDM

[CHT19] Shu-Yu Cui, Jing-Xiang He, and Gui-Xian Tian. The generalized distance matrix. Linear Algebra and its Applications, 563(??): 1–23, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304944.

Chuysurichay:2015:SSE

[Chu15] Sompong Chuysurichay. Strong shift equivalence and positive doubly stochastic matrices. Linear Algebra and its Applications, 467(??): 15–28, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007228.

Chu:2016:FDF

[Chu16] Moody T. Chu. On the first degree Fejér–Riesz factorization and its applications to X + A^{∗}X^{−1}A = Q. Linear Algebra and its Applications, 489(??):

123–143, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005893.

Chu:2019:SEC

[Chu19] Wenchang Chu. Spectrum and eigenvectors for a class of tridiagonal matrices. Linear Algebra and its Applications, 582(??):499–516, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930357X.

Chugunov:2015:CSP

[CI15a] V. N. Chugunov and Kh. D. Ikramov. A complete solution of the permutability problem for Toeplitz and Hankel matrices. Linear Algebra and its Applications, 478(??):53–80, August 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001548.

Chugunov:2015:PTH

[CI15b] V. N. Chugunov and Kh. D. Ikramov. Permutability of Toeplitz and Hankel matrices. Linear Algebra and its Applications, 467(??): 226–242, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951400723X.

Casas:2016:TLA

[CILL16] J. M. Casas, M. A. Insua, M. Ladra, and S. Ladra. Test for Leibniz n-algebra structure. Linear Algebra and its Applications, 494 (??):138–155, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000161.

Cimpric:2015:FLM

[Cim15]  Jaka Cimpric.  Finsler’s

Lemma for matrix polynomials. Linear Algebra and its Applications, 465(??): 239–261, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951400634X.

Cvetkovic-Ilic:2015:POI

[CIMW15] Dragana S. Cvetković-Ilić, Dijana Mosić, and Yimin Wei. Partial orders on B(H). Linear Algebra and its Applications, 481(??):115– 130, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002761.

Cheng:2015:IMB

[CIY15] Tzu-Chang F. Cheng, Ching-Kang Ing, and Shu-Hui Yu. Inverse moment bounds for sample autocovariance matrices based on detrended time series and their applications. Linear Algebra and its Applications, 473 (??):180–201, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514003085.

Cheon:2015:SCA

[CJ15] Gi-Sang Cheon and Ji-Hwan Jung. Some combinatorial applications of the q-Riordan matrix. Linear Algebra and its Applications, 482(??): 241–260, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003675.

Cheon:2016:SSN

[CJ16a] Gi-Sang Cheon and Ji-Hwan Jung. The q-Sheffer sequences of a new type and associated orthogonal polynomials. Linear Algebra and its Applications, 491(??): 171–186, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004085.

Choi:2016:PDH

[CJ16b] Hayoung Choi and Farhad Jafari. Positive definite Hankel matrix completions and Hamburger moment completions. Linear Algebra and its Applications, 489(??): 217–237, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006035.

Chuong:2016:CRL

[CJ16c] T. D. Chuong and V. Jeyakumar. Characterizing robust local error bounds for linear inequality systems under data uncertainty. Linear Algebra and its Applications, 489(??):199–216, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006060.

Chuong:2016:RGE

[CJ16d] T. D. Chuong and V. Jeyakumar. Robust global error bounds for uncertain linear inequality systems with applications. Linear Algebra and its Applications, 493(??):183–205, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007090.

Carden:2017:INR

[CJ17a] Russell L. Carden and Mostafa Zahed Jahromi. The inverse q-numerical range problem and connections to the Davis–Wielandt shell and the pseudospectra of a matrix. Linear Algebra and its Applications, 531(??):479–497, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303415.

Cheon:2017:GMD

[CJ17b] Gi-Sang Cheon and Sung-Tae Jin. The group of multi-dimensional Riordan arrays. Linear Algebra and its Applications, 524 (??):263–277, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301623.

Chuong:2017:GFL

[CJ17c] T. D. Chuong and V. Jeyakumar. A generalized Farkas lemma with a numerical certificate and linear semiinfinite programs with SDP duals. Linear Algebra and its Applications, 515(??): 38–52, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305249.

Cvetkovic:2017:CGG

[CJ17d] Dragos Cvetković and Irena M. Jovanović. Constructing graphs with given spectrum and the spectral radius at most 2. Linear Algebra and its Applications, 515(??): 255–274, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305468.

Cheon:2018:HVF

[CJB18] Gi-Sang Cheon, Ji-Hwan Jung, and Paul Barry. Horizontal and vertical formulas for exponential Riordan matrices and their applications. Linear Algebra and its Applications, 541(??): 266–284, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306808.

Coons:2016:NRF

[CJK^{+}16] Jane Ivy Coons, Jack Jenkins, Douglas Knowles, Rayanne A. Luke, and Patrick X. Rault. Numerical ranges over finite fields. Linear Algebra and its Applications, 501(??):37– 47, July 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Cheon:2019:RGS

[CJKM19a] Gi-Sang Cheon, Ji-Hwan Jung, Sergey Kitaev, and Seyed Ahmad Mojallal. Riordan graphs I: Structural properties. Linear Algebra and its Applications, 579(??):89–135, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302423.

Cheon:2019:RGI

[CJKM19b] Gi-Sang Cheon, Ji-Hwan Jung, Sergey Kitaev, and Seyed Ahmad Mojallal. Riordan graphs II: Spectral properties. Linear Algebra and its Applications, 575(??):174–215, August 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519301636.

Campos:2015:RPB

[CJRMF^{+}15] J. R. Campos, P. Jiménez-Rodríguez, G. A. Muñoz-Fernández, D. Pellegrino, and J. B. Seoane-Sepúlveda. On the real polynomial Bohnenblust-Hille inequality. Linear Algebra and its Applications, 465(??):

391–400, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006375.

Cheon:2016:CER

[CJS16] Gi-Sang Cheon, Sung-Tae Jin, and Louis W. Shapiro. A combinatorial equivalence relation for formal power series. Linear Algebra and its Applications, 491(??): 123–137, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500169X.

Collao:2018:URS

[CJS18] Macarena Collao, Charles R. Johnson, and Ricardo L. Soto. Universal realizability of spectra with two positive eigenvalues. Linear Algebra and its Applications, 545 (??):226–239, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300545.

Cowen:2019:CIS

[CJW19] Carl C. Cowen, William Johnston, and Rebecca G. Wahl. Constructing invariant subspaces as kernels of commuting matrices. Linear Algebra and its Applications, 583(??):

46–62, December 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303544.

Christandl:2018:TRM

[CJZ18] Matthias Christandl, Asger Kjærulff Jensen, and Jeroen Zuiddam. Tensor rank is not multiplicative under the tensor product. Linear Algebra and its Applications, 543(??): 125–139, April 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306900.

Chapman:2015:GCA

[CK15a] Adam Chapman and Jung-Miao Kuo. On the generalized Clifford algebra of a monic polynomial. Linear Algebra and its Applications, 471(??):184–202, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500021X.

Cheon:2015:RPC

[CK15b] Gi-Sang Cheon and Hana Kim. Representing polynomials as characteristic polynomials via the Stieltjes transform. Linear Algebra and its Applications, 476 (??):184–196, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001573.

Cheon:2019:MEM

[CK19] Gi-Sang Cheon and Hana Kim. Mertens equimodular matrices of Redheffer type. Linear Algebra and its Applications, 572 (??):252–272, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301107.

Chooi:2017:CIM

[CKL17] Wai Leong Chooi, Kiam Heong Kwa, and Ming-Huat Lim. Coherence invariant maps on tensor products. Linear Algebra and its Applications, 516(??):24–46, March 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516305493.

Christou:2016:ALC

[CKM16] Dimitrios Christou, Nicos Karcanias, and Marilena Mitrouli. Approximate least common multiple of several polynomials using the ERES division algorithm. Linear Algebra and its Applications, 511(??):141–175, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303998.

Casado:2019:BIE

[CKM19] Yolanda Cabrera Casado, Müge Kanuni, and Mercedes Siles Molina. Basic ideals in evolution algebras. Linear Algebra and its Applications, 570 (??):148–180, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300266.

Cheon:2016:POT

[CKS16] Gi-Sang Cheon, Hana Kim, and Louis W. Shapiro. Profiles of ordered trees with mutation and associated Riordan matrices. Linear Algebra and its Applications, 511(??):296–317, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303950.

Chooi:2019:CMI

[CKT19] Wai Leong Chooi, Kiam Heong Kwa, and Li Yin Tan. Commuting maps on invertible triangular matrices over F_{2}. Linear Algebra and its Applications, 583(??):77–101, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303702.

Chen:2015:LLD

[CL15a] Lin Chen and Fangyan Lu. Local Lie derivations of nest algebras. Linear Algebra and its Applications, 475 (??):62–72, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000683.

Cheng:2015:SNC

[CL15b] Che-Man Cheng and Chunyu Lei. On Schatten p-norms of commutators. Linear Algebra and its Applications, 484(??):409–434, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004097.

Chorianopoulos:2015:IPG

[CL15c] Christos Chorianopoulos and Peter Lancaster. An inverse problem for gyroscopic systems. Linear Algebra and its Applications, 465(??): 188–203, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Cronin:2015:CDK

[CL15d] Anthony G. Cronin and Thomas J. Laffey. On a conjecture of Deveci and Karaduman. Linear Algebra and its Applications, 471 (??):569–574, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000439.

Chen:2016:MPH

[CL16] Chaoqun Chen and Fangyan Lu. Multiplicative preservers of higher-dimensional numerical ranges. Linear Algebra and its Applications, 490(??):186–195, February 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515006709.

Chen:2017:SPS

[CL17a] Xiaolin Chen and Huishu Lian. Solution to a problem on skew spectral radii of oriented graphs. Linear Algebra and its Applications, 524(??):61–67, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301210.

Cheng:2017:SSB

[CL17b] Che-Man Cheng and Yaru Liang. Some sharp bounds for the commutator of real matrices. Linear Algebra and its Applications, 521 (??):263–282, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300605.

Chen:2018:BD

[CL18] Ting Chen and Fangyan Lu. 2-bilocal derivations. Linear Algebra and its Applications, 555(??):222–240, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303021.

Cai:2019:PAM

[CL19] Yunfeng Cai and Ren-Cang Li. Perturbation analysis for matrix joint block diagonalization. Linear Algebra and its Applications, 581(??):163–197, November 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519302939.

Chen:2018:NUB

[CLF18] Xiaodan Chen, Jingjian Li, and Yingmei Fan. Note on an upper bound for sum of the Laplacian eigenvalues of a graph. Linear Algebra and its Applications, 541(??): 258–265, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306687.

Chen:2015:SFA

[CLHL15] Jianzhang Chen, Jianping Li, Yuanyuan Huang, and Jie Lin. Some families of asymmetric quantum codes and quantum convolutional codes from constacyclic codes. Linear Algebra and its Applications, 475 (??):186–199, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001317.

Chen:2015:LBS

[CLL15] Xiaolin Chen, Xueliang Li, and Huishu Lian. Lower bounds of the skew spectral radii and skew energy of oriented graphs. Linear Algebra and its Applications, 479(??):91–105, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002165.

Curtin:2019:MGA

[CLLS19] Eugene Curtin, Junu Lee, Andrew Lu, and Sophia Sun. A modified Grassmann algebra approach to theorems on permanents and determinants. Linear Algebra and its Applications, 581(??): 20–35, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302836.

Chen:2019:HAD

[CLM19a] Cairong Chen, Ren-Cang Li, and Changfeng Ma. Highly accurate doubling algorithm for quadratic matrix equation from quasi-birth-and-death process. Linear Algebra and its Applications, 583(??):1–45, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303581.

Chen:2019:SLE

[CLM19b] Yuanyuan Chen, Dan Li, and Jixiang Meng. On the second largest A_{α}- eigenvalues of graphs. Linear Algebra and its Applications, 580(??):343–358, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Cui:2015:PTD

[CLN15] Lu-Bin Cui, Wen Li, and Michael K. Ng. Primitive tensors and directed hypergraphs. Linear Algebra and its Applications, 471 (??):96–108, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000245.

Casas:2013:CSL

[CLOK13] J. M. Casas, M. Ladra, B. A. Omirov, and I. A. Karimjanov. Classification of solvable Leibniz algebras with naturally graded filiform nilradical. Linear Algebra and its Applications, 438 (7):2973–3000, April 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379512008300.

See corrigendum [LMO16].

Cui:2016:PUS

[CLP16] Jianlian Cui, Chi-Kwong Li, and Yiu-Tung Poon. Preservers of unitary similarity functions on Lie products of matrices. Linear Algebra and its Applications, 498 (??):160–180, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001433.

Che:2017:PST

[CLQW17] Maolin Che, Guoyin Li, Liqun Qi, and Yimin Wei. Pseudo-spectra theory of tensors and tensor polynomial eigenvalue problems. Linear Algebra and its Applications, 533(??):536–572, November 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304469.

Charlier:2016:APF

[CLR16] Émilie Charlier, Julien Leroy, and Michel Rigo. Asymptotic properties of free monoid morphisms. Linear Algebra and its Applications, 500(??):119–148, July 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001427.

Chen:2018:COH

[CLR^{+}18] Gina Chen, Vivian Liu, Ellen Robinson, Lucas J. Rusnak, and Kyle Wang. A characterization of oriented hypergraphic Laplacian and adjacency matrix coefficients. Linear Algebra and its Applications, 556(??): 323–341, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303306.

Casas:2019:MPC

[CLR19] J. M. Casas, M. Ladra, and U. A. Rozikov. Markov processes of cubic stochastic matrices: Quadratic stochastic processes. Linear Algebra and its Applications, 575(??):273–298, August 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301685.

Cui:2017:PPS

[CLS17] Jianlian Cui, Chi-Kwong Li, and Nung-Sing Sze. Products of positive semi-definite matrices. Linear Algebra and its Applications, 528(??): 17–24, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005728.

Cheng:2018:FUS

[CLS18] Qingqing Cheng, Fusheng Lv, and Wenchang Sun. Frames of uniform subframe bounds with applications to erasures. Linear Algebra and its Applications, 555(??): 186–200, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302659.

Catral:2018:MSR

[CLST18] Minerva Catral, Leila Lebtahi, Jeffrey Stuart, and Néstor Thome. Matrices A such that A^{s+1}R = RA^{∗} with R^{k} = I. Linear Algebra and its Applications, 552(??):85–104, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301964.

Chen:2015:NCC

[CLT15] Sheng Chen, Li Liang, and Yunbo Tian. The number of connected components in a graph associated with a rectangular (0, 1)- matrix. Linear Algebra and its Applications, 487(??): 74–85, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005339.

Cano:2017:PCG

[CLUP17] Angel Cano, Luis Loeza, and Alejandro Ucan-Puc. Projective cyclic groups in higher dimensions. Linear Algebra and its Applications, 531(??):169–209, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303476.

Chen:2015:TPR

[CLW15] Xi Chen, Huyile Liang, and Yi Wang. Total positivity of recursive matrices. Linear Algebra and its Applications, 471(??): 383–393, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000385.

Chen:2016:RMG

[CLZ16a] Xiaolin Chen, Xueliang Li, and Yingying Zhang. 3regular mixed graphs with optimum Hermitian energy. Linear Algebra and its Applications, 496(??):475–486, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516001130.

Cui:2016:CSR

[CLZ^{+}16b] Xiangzhao Cui, Chun Li, Jine Zhao, Li Zeng, Defei Zhang, and Jianxin Pan. Covariance structure regularization via Frobeniusnorm discrepancy. Linear Algebra and its Applications, 510(??):124–145, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303500.

Conde:2016:NIR

[CM16] Cristian Conde and Mohammad Sal Moslehian. Norm inequalities related to p-Schatten class. Linear Algebra and its Applications, 498 (??):441–449, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007053.

Chabbabi:2017:NFS

[CM17] Fadil Chabbabi and Mostafa Mbekhta. New formulas for the spectral radius via λ-Aluthge transform. Linear Algebra and its Applications, 515(??):246–254, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305481.

Castro:2018:HMS

[CM18] Francis N. Castro and Luis A. Medina. Hadamard matrices and the spectrum of quadratic symmetric polynomials over finite fields. Linear Algebra and its Applications, 549(??):153– 175, July 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830123X.

Clouatre:2019:JSC

[CM19] Raphaël Clouâtre and Diarra Mbacke. Joint similarity for commuting families of power bounded matrices. Linear Algebra and its Applications, 560(??):1– 16, January 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304440.

Cordova-Martinez:2018:GSA

[CME18] Alejandra S. Córdova-Martínez and Alberto Elduque. Gradings on semisimple algebras. Linear Algebra and its Applications, 559(??):145– 171, December 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304294.

Contino:2019:SCS

[CMM19] Maximiliano Contino, Alejandra Maestripieri, and Stefania Marcantognini. Schur complements of selfadjoint Krein space operators. Linear Algebra and its Applications, 581(??):214–246, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303003.

Cesmelioglu:2018:VBF

[ÇMP18] Ayça Çesmelioğlu, Wilfried Meidl, and Alexander Pott. Vectorial bent functions and their duals. Linear Algebra and its Applications, 548 (??):305–320, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301265.

Csajbok:2018:NFM

[CMPZ18] Bence Csajbók, Giuseppe Marino, Olga Polverino, and Corrado Zanella. A new family of MRD-codes. Linear Algebra and its Applications, 548(??):203–220, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518301022.

Chen:2016:CMC

[CMQV16] Wei Chen, Yanfang Mo, Li Qiu, and Pravin Varaiya. Constrained (0, 1)-matrix completion with a staircase of fixed zeros. Linear Algebra and its Applications, 510(??):171–185, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Casado:2016:EAA

[CMV16] Yolanda Cabrera Casado, Mercedes Siles Molina, and M. Victoria Velasco. Evolution algebras of arbitrary dimension and their decompositions. Linear Algebra and its Applications, 495 (??):122–162, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000094.

Casado:2017:CTD

[CMV17] Yolanda Cabrera Casado, Mercedes Siles Molina, and M. Victoria Velasco. Classification of three-dimensional evolution algebras. Linear Algebra and its Applications, 524(??):68–108, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300915.

Camps:2019:IFR

[CMV19] Daan Camps, Karl Meerbergen, and Raf Vandebril. An implicit filter for rational Krylov using core transformations. Linear Algebra and its Applications, 561(??): 113–140, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304622.

Chien:2015:NPC

[CN15] Mao-Ting Chien and Hiroshi Nakazato. A new Poncelet curve for the boundary generating curve of a numerical range. Linear Algebra and its Applications, 487(??): 1–21, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005145.

Chien:2017:RTF

[CN17a] Mao-Ting Chien and Hiroshi Nakazato. Reducibility of the ternary forms of unitary bordering matrices. Linear Algebra and its Applications, 527(??): 73–86, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301830.

Chien:2017:SPA

[CN17b] Mao-Ting Chien and Hiroshi Nakazato. Singular points of the algebraic curves associated to unitary bordering matrices. Linear Algebra and its Applications, 513(??): 224–239, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Chien:2018:USD

[CN18] Mao-Ting Chien and Hiroshi Nakazato. Unitary similarity of the determinantal representation of unitary bordering matrices. Linear Algebra and its Applications, 541 (??):13–35, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306560.

Caruso:2019:CAL

[CN19a] Noe Angelo Caruso and Paolo Novati. Convergence analysis of LSQR for compact operator equations. Linear Algebra and its Applications, 583(??):146–164, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303714.

Chien:2019:BGC

[CN19b] Mao-Ting Chien and Hiroshi Nakazato. Boundary generating curves of quaternionic numerical ranges. Linear Algebra and its Applications, 574(??):201–215, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301570.

Chien:2019:DWN

[CNM19] Mao-Ting Chien, Hiroshi Nakazato, and Jie Meng. The diameter and width of numerical ranges. Linear Algebra and its Applications, 582(??):76–98, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303301.

Chen:2017:BCF

[CNX17] Yuan Chen, Liujie Nie, and Yunge Xu. Belitskii’s canonical forms of linear dynamical systems. Linear Algebra and its Applications, 531(??):533–536, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303828.

Chien:2018:INR

[CNY18] Mao-Ting Chien, Hiroshi Nakazato, and Lina Yeh. Inverse numerical range and determinantal representation. Linear Algebra and its Applications, 558(??): 79–100, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Cole:2019:RNP

[Col19]  Sam Cole.  Recovering

nonuniform planted partitions via iterated projection. Linear Algebra and its Applications, 576(??):79– 107, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301198.

Costara:2016:CMP

[Cos16a] Constantin Costara. Continuous maps preserving local spectra of matrices. Linear Algebra and its Applications, 492(??):1–8, March 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006801.

Costara:2016:NMP

[Cos16b] Constantin Costara. On nonlinear maps preserving the reduced minimum modulus on differences of matrices. Linear Algebra and its Applications, 507(??): 288–299, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302336.

Collins:2018:FLM

[COS18] Benoît Collins, Hiroyuki Osaka, and Gunjan Sapra. On a family of linear maps from Mn(C) to Mn2(C). Linear Algebra and its Applications, 555(??):398–411, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302957.

Costantini:2019:LAF

[Cos19a] Mauro Costantini. Local automorphisms of finite dimensional simple Lie algebras. Linear Algebra and its Applications, 562(??): 123–134, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830483X.

Costara:2019:NDP

[Cos19b] Constantin Costara. Nonlinear determinant preserving maps on matrix algebras. Linear Algebra and its Applications, 583(??):165– 170, December 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303799.

Culos:2016:RIS

[COvdD16] G. J. Culos, D. D. Olesky, and P. van den Driessche.

Refined inertias of strongly connected orientations of the Petersen graph. Linear Algebra and its Applications, 507(??):322–343, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301756.

Coxon:2015:MMP

[Cox15]  Nicholas Coxon.  Mont-

gomery’s method of polynomial selection for the number field sieve. Linear Algebra and its Applications, 485(??):72–102, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004395.

Cabello:2016:WLS

[CP16a] Juan Carlos Cabello and Antonio M. Peralta. Weak-2-local symmetric maps on C^{⋆}-algebras. Linear Algebra and its Applications, 494 (??):32–43, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007612.

Cifuentes:2016:ETD

[CP16b] Diego Cifuentes and Pablo A. Parrilo. An efficient tree decomposition method for permanents and mixed discriminants. Linear Algebra and its Applications, 493 (??):45–81, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007144.

Cohen:2017:MPF

[CP17] Nir Cohen and Edgar Pereira. Matrix polynomials: Factorization via bisolvents. Linear Algebra and its Applications, 530 (??):57–76, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302707.

Cardoso:2018:MEG

[CPR18] Domingos M. Cardoso, Germain Pastén, and Oscar Rojo. On the multiplicity of α as an eigenvalue of A_{α}(G) of graphs with pendant vertices. Linear Algebra and its Applications, 552(??): 52–70, September 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830199X.

Casati:2017:IMF

[CPS17] Paolo Casati, Andrea Previtali, and Fernando Szechtman. Indecomposable modules of a family of solvable Lie algebras. Linear Algebra and its Applications, 531(??):423–446, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303488.

Chen:2017:HML

[CPW17] Pingge Chen, Yuejian Peng, and Suijie Wang. The Hessian matrix of Lagrange function. Linear Algebra and its Applications, 531(??): 537–546, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303750.

Castano:2016:ATI

[CPZ16a] Diego Castano, Vehbi E. Paksoy, and Fuzhen Zhang. Angles, triangle inequalities, correlation matrices and metric-preserving and subadditive functions. Linear Algebra and its Applications, 491(??):15–29, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006806.

Chang:2016:ITP

[CPZ16b] Haixia Chang, Vehbi E. Paksoy, and Fuzhen Zhang. An inequality for tensor product of positive operators and its applications. Linear Algebra and its Applications, 498 (??):99–105, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000178.

Caldeira:2015:IFP

[CQ15a] Cristina Caldeira and João Filipe Queiró. Invariant factors of products over elementary divisor domains. Linear Algebra and its Applications, 485(??):345–358, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004528.

Cluzeau:2015:NIS

[CQ15b] Thomas Cluzeau and Alban Quadrat. A new insight into Serre’s reduction problem. Linear Algebra and its Applications, 483(??): 40–100, October 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002888.

Chen:2019:BNT

[CQCZ19] Haibin Chen, Liqun Qi, Louis Caccetta, and Guanglu Zhou. Birkhoff–von Neumann theorem and decomposition for doubly stochastic tensors. Linear Algebra and its Applications, 583(??):119–133, December 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951930374X.

Cruickshank:2018:HSH

[CQS18] James Cruickshank, Rachel Quinlan, and Fernando Szechtman. Hermitian and skew Hermitian forms over local rings. Linear Algebra and its Applications, 551(??):147–161, August 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301952.

Centrone:2015:IGI

[CR15] Lucio Centrone and Viviane Ribeiro Tomaz da Silva. On Z_{2}-graded identities of UT_{2}(E) and their growth. Linear Algebra and its Applications, 471 (??):469–499, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000270.

Capaverde:2017:EMQ

[CR17a] Juliane Capaverde and Peter Rowlinson. Eigenvalue multiplicity in quartic graphs.

Linear Algebra and its Applications, 535(??):160–170, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305116.

Cardoso:2017:EPG

[CR17b] Domingos M. Cardoso and Oscar Rojo. Edge perturbation on graphs with clusters: Adjacency, Laplacian and signless Laplacian eigenvalues. Linear Algebra and its Applications, 512(??): 113–128, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630430X.

Cavenagh:2018:CML

[CR18a] Nicholas J. Cavenagh and Reshma Ramadurai. Constructing (0, 1)-matrices with large minimal defining sets. Linear Algebra and its Applications, 537(??):38–47, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305566.

Chen:2018:LSD

[CR18b] Ricky X. F. Chen and Christian M. Reidys. Linear sequential dynamical systems, incidence algebras, and Möbius functions. Linear Algebra and its Applications, 553(??):270–291, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302386.

Carando:2019:SMF

[CR19a] Daniel Carando and Jorge Tomás Rodríguez. Symmetric multilinear forms on Hilbert spaces: Where do they attain their norm? Linear Algebra and its Applications, 563(??):178–192, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305111.

Carmona:2019:ISL

[CR19b] Juan R. Carmona and Jonnathan Rodríguez. An increasing sequence of lower bounds for the Estrada index of graphs and matrices. Linear Algebra and its Applications, 580(??): 200–211, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302708.

Chatelin:2017:HMS

[CRC17] Françoise Chatelin and M. Monserrat Rincon-Camacho.

Hermitian matrices: Spectral coupling, plane geometry/trigonometry and optimisation. Linear Algebra and its Applications, 533(??):282–310, November 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517304196.

Chen:2015:COH

[CRRY15] Vinciane Chen, Angeline Rao, Lucas J. Rusnak, and Alex Yang. A characterization of oriented hypergraphic balance via signed weak walks. Linear Algebra and its Applications, 485(??):442–453, November 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515004541.

Cvetkovic:2015:GLE

[CRS15] Dragos Cvetković, Peter Rowlinson, and Slobodan Simić. Graphs with least eigenvalue −2: Ten years on. Linear Algebra and its Applications, 484(??): 504–539, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003651.

Ciesielski:2019:AME

[CRVSS19] K. C. Ciesielski, D. L. Rodríguez-Vidanes, and J. B. Seoane-Sepúlveda. Algebras of measurable extendable functions of maximal cardinality. Linear Algebra and its Applications, 565(??):258–266, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305809.

Calvetti:2015:CDO

[CRX15] D. Calvetti, L. Reichel, and H. Xu. A CS decomposition for orthogonal matrices with application to eigenvalue computation. Linear Algebra and its Applications, 476 (??):197–232, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001512.

Carmichael:2015:EIC

[CS15a] Owen Carmichael and Lyudmila Sakhanenko. Estimation of integral curves from high angular resolution diffusion imaging (HARDI) data. Linear Algebra and its Applications, 473(??): 377–403, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007976.

Choudhury:2015:TTA

[CS15b] Projesh Nath Choudhury and K. C. Sivakumar. Tucker’s theorem for almost skew-symmetric matrices and a proof of Farkas’ lemma. Linear Algebra and its Applications, 482 (??):55–69, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003286.

Choudhury:2017:EMI

[CS17] Projesh Nath Choudhury and K. C. Sivakumar. An extension of a matrix inequality of Thompson. Linear Algebra and its Applications, 535(??):151–159, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304949.

Cheon:2018:NAR

[CS18a] Gi-Sang Cheon and Minho Song. A new aspect of Riordan arrays via Krylov matrices. Linear Algebra and its Applications, 554(??): 329–341, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Cruickshank:2018:GRU

[CS18b] J. Cruickshank and F. Szechtman. Generators and relations for the unitary group of a skew Hermitian form over a local ring. Linear Algebra and its Applications, 552(??):1–28, September 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518301794.

Chatterjee:2019:IGI

[CS19a] Manami Chatterjee and K. C. Sivakumar. Inequalities for group invertible Hmatrices. Linear Algebra and its Applications, 576(??): 158–180, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301976.

Chatterjee:2019:IHM

[CS19b] Manami Chatterjee and K. C. Sivakumar. Inverse H-matrices. Linear Algebra and its Applications, 572(??):1–31, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300874.

Shi:2015:SPR

[cSfCfX15] De cai Shi, Yun feng Cai, and Shu fang Xu. Some perturbation results for a normalized Non-Orthogonal Joint Diagonalization problem. Linear Algebra and its Applications, 484(??): 457–476, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003754.

Costas-Santos:2016:MTP

[CSJ16] R. S. Costas-Santos and C. R. Johnson. Matrices totally positive relative to a tree, II. Linear Algebra and its Applications, 505(??):1– 10, September 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301203.

Chmielinski:2017:AON

[CSW17] Jacek Chmieliński, Tomasz Stypula, and Paweł Wójcik. Approximate orthogonality in normed spaces and its applications. Linear Algebra and its Applications, 531(??):305–317, October 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517303634.

Choi:2019:EFI

[CTZ19] Daeshik Choi, Tin-Yau Tam, and Pingping Zhang. Extensions of Fischer’s inequality. Linear Algebra and its Applications, 569(??):311– 322, May 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300461.

Canto:2018:MRT

[CU18] Rafael Cantó and Ana M. Urbano. On the maximum rank of totally nonnegative matrices. Linear Algebra and its Applications, 551(??): 125–146, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830168X.

Cohen:2015:AFN

[CUPW15] Stephen D. Cohen, Sartaj Ul Hasan, Daniel Panario, and Qiang Wang. An asymptotic formula for the number of irreducible transformation shift registers. Linear Algebra and its Applications, 484(??): 46–62, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003730.

Cushman:2017:UNF

[Cus17]  Richard Cushman.  The

uniform normal form of a linear mapping. Linear Algebra and its Applications, 512(??):249–255, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304311.

Chandrasekaran:2017:SLP

[CV17] Karthekeyan Chandrasekaran and Ameya Velingker. Shift lifts preserving Ramanujan property. Linear Algebra and its Applications, 529(??):199–214, September 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517302689.

Corrales:2018:ASG

[CV18] Hugo Corrales and Carlos E. Valencia. Arithmetical structures on graphs. Linear Algebra and its Applications, 536(??):120–151, January 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305426.

Caseiro:2019:PMD

[CVV19] R. Caseiro, M. A. Facas Vicente, and José Vitória. Projection method and the distance between two linear varieties. Linear Algebra and its Applications, 563(??): 446–460, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305317.

Chen:2015:LAF

[CW15a] Zhengxin Chen and Dengyin Wang. 2-local automorphisms of finite-dimensional simple Lie algebras. Linear Algebra and its Applications, 486(??):335–344, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004930.

Chretien:2015:NTI

[CW15b] Stéphane Chrétien and Tianwen Wei. Von Neumann’s trace inequality for tensors. Linear Algebra and its Applications, 482(??): 149–157, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003146.

Cordwell:2016:MPS

[CW16a] Katherine Cordwell and George Wang. Multilinear polynomials of small degree evaluated on matrices over a unital algebra. Linear Algebra and its Applications, 496 (??):262–287, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000896.

Costello:2016:NIG

[CW16b] Kevin P. Costello and Parker Williams. On the number of integral graphs. Linear Algebra and its Applications, 493(??):447–454, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007016.

Chien:2017:NEF

[CW17] Tuan-Yow Chien and Shayne Waldron. Nice error frames, canonical abstract error groups and the construction of SICs. Linear Algebra and its Applications, 516 (??):93–117, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305523.

Cao:2018:RZP

[CW18a] Lei Cao and Hugo J. Woerdeman. Real zero polynomials and A. Horn’s problem. Linear Algebra and its Applications, 552(??):147– 158, September 1, 2018.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830209X.

Chretien:2018:SSC

[CW18b] Stéphane Chrétien and Tianwen Wei. On the subdifferential of symmetric convex functions of the spectrum for symmetric and orthogonally decomposable tensors. Linear Algebra and its Applications, 542 (??):84–100, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300691.

Chen:2019:NTP

[CW19] Xi Chen and Yi Wang. Notes on the total positivity of Riordan arrays. Linear Algebra and its Applications, 569 (??):156–161, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300394.

Chen:2017:CPI

[CX17] Sheng Chen and Chao Xia. A class of posets with integral Laplacian spectrum. Linear Algebra and its Applications, 529(??):258–270, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302641.

Chen:2016:BCF

[CXLF16] Yuan Chen, Yunge Xu, Huanhuan Li, and Wenhao Fu. Belitskii’s canonical forms of upper triangular nilpotent matrices under upper triangular similarity. Linear Algebra and its Applications, 506(??): 139–153, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301823.

Chowdhury:2019:SCP

[CYD19] Agniva Chowdhury, Jiasen Yang, and Petros Drineas. Structural conditions for projection-cost preservation via randomized matrix multiplication. Linear Algebra and its Applications, 573 (??):144–165, July 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301144.

Das:2019:ASS

[DA19a] Ranjan Kumar Das and Rafikul Alam. Affine spaces of strong linearizations for rational matrices and the recovery of eigenvectors and minimal bases. Linear Algebra and its Applications, 569 (??):335–368, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300497.

Das:2019:ARE

[DA19b] Ranjan Kumar Das and Rafikul Alam. Automatic recovery of eigenvectors and minimal bases of matrix polynomials from generalized Fiedler pencils with repetition. Linear Algebra and its Applications, 569 (??):78–112, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300291.

Das:2019:RMB

[DA19c] Ranjan Kumar Das and Rafikul Alam. Recovery of minimal bases and minimal indices of rational matrices from Fiedler-like pencils. Linear Algebra and its Applications, 566(??):34– 60, April 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305846.

Delgado:2016:NRR

[DAG16] Ramón A. Delgado, Juan C. Agüero, and Graham C. Goodwin. A novel representation of rank constraints for real matrices. Linear Algebra and its Applications, 496 (??):452–462, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000999.

Dahl:2015:RPM

[Dah15a] Geir Dahl. L-rays of permutation matrices and doubly stochastic matrices. Linear Algebra and its Applications, 480(??):127–143, September 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002864.

Dahl:2015:LET

[Dah15b] Geir Dahl. The Laplacian energy of threshold graphs and majorization. Linear Algebra and its Applications, 469(??):518–530, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007927.

Dalfo:2017:ILD

[Dal17a] C. Dalfó. Iterated line digraphs are asymptotically dense. Linear Algebra and its Applications, 529(??):391– 396, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Dalfo:2017:SSC

[Dal17b] C. Dalfó. The spectra of sub-Kautz and cyclic Kautz digraphs. Linear Algebra and its Applications, 531(??): 210–219, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303464.

Dalfo:2019:SMM

[Dal19] C. Dalfó. A survey on the missing Moore graph. Linear Algebra and its Applications, 569(??):1–14, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300138.

Das:2015:PCD

[Das15] Kinkar Ch. Das. Proof of conjectures on the distance signless Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 467(??):100–115, February 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007320.

Das:2018:PCI

[Das18] Kinkar Ch. Das. Proof of conjecture involving algebraic connectivity and average degree of graphs. Linear Algebra and its Applications, 548(??):172–188, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518301083.

Dax:2017:NRK

[Dax17] Achiya Dax. The numerical rank of Krylov matrices. Linear Algebra and its Applications, 528(??):185– 205, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302944.

Das:2019:SSP

[DB19a] Sunil Das and Sriparna Bandopadhyay. On some sign patterns of algebraically positive matrices. Linear Algebra and its Applications, 562(??):91–122, February 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304816.

deBondt:2019:QPM

[dB19b] Michiel de Bondt. Quadratic polynomial maps with Jacobian rank two. Linear Algebra and its Applications, 565(??):267–286, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305792.

Dolzan:2017:DCC

[DBK17] D. Dolzan, D. Kokol Bukovsek, and B. Kuzma. On diameter of components in commuting graphs. Linear Algebra and its Applications, 522 (??):161–174, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300678.

daCruz:2017:MDS

[dCDFK17] Henrique F. da Cruz, Gregor Dolinar, Rosário Fernandes, and Bojan Kuzma. Maximal doubly stochastic matrix centralizers. Linear Algebra and its Applications, 532(??):387–396, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303932.

See correction [dCDFK18].

daCruz:2018:CSM

[dCDFK18] Henrique F. da Cruz, Gregor Dolinar, Rosário Fernandes, and Bojan Kuzma.

Corrigendum to “Maximal doubly stochastic matrix centralizers” [Linear Algebra Appl. 532 (2017) 387– 396]. Linear Algebra and its Applications, 555(??): 432–433, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830346X.

See [dCDFK17].

deCamargo:2015:CSD

[dCdSA15] André Pierro de Camargo and Felipe da Silva Alves. Can Sylvester’s determinantal identity, equivalently Muir’s law of extensible minors be generalized? Linear Algebra and its Applications, 482(??):70–83, October 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003092.

daCruz:2017:MMB

[dCFF17] Henrique F. da Cruz, Rosário Fernandes, and Susana Furtado. Minimal matrices in the Bruhat order for symmetric (0, 1)- matrices. Linear Algebra and its Applications, 530(??): 160–184, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

DeVos:2016:BTU

[DD16a] Alexis De Vos and Stijn De Baerdemacker. The Birkhoff theorem for unitary matrices of prime dimension. Linear Algebra and its Applications, 493(??): 455–468, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007156.

Djikic:2016:CPE

[DD16b] Marko S. Djikić and Dragan S. Djordjević. Coherent and precoherent elements in Rickart *-rings. Linear Algebra and its Applications, 509(??):64–81, November 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302932.

DAngeli:2017:WPM

[DD17a] Daniele D’Angeli and Alfredo Donno. Wreath product of matrices. Linear Algebra and its Applications, 513(??):276–303, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304967.

Dmytryshyn:2017:GCE

[DD17b] Andrii Dmytryshyn and Froilán M. Dopico. Generic complete eigenstructures for sets of matrix polynomials with bounded rank and degree. Linear Algebra and its Applications, 535(??):213– 230, December 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305293.

Domanov:2017:CPD

[DD17c] Ignat Domanov and Lieven De Lathauwer. Canonical polyadic decomposition of third-order tensors: Relaxed uniqueness conditions and algebraic algorithm. Linear Algebra and its Applications, 513(??):342–375, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630492X.

Dmytryshyn:2018:GSS

[DD18] Andrii Dmytryshyn and Froilán M. Dopico. Generic skew-symmetric matrix polynomials with fixed rank and fixed odd grade. Linear Algebra and its Applications, 536(??):1–18, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517305281.

DeVos:2019:BTU

[DD19] Alexis De Vos and Stijn De Baerdemacker. The Birkhoff theorem for unitary matrices of prime-power dimension. Linear Algebra and its Applications, 577 (??):27–52, October 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302095.

Duffner:2016:RNL

[DdC16] M. Antónia Duffner and Henrique F. da Cruz. Rank nonincreasing linear maps preserving the determinant of tensor product of matrices. Linear Algebra and its Applications, 510(??): 186–191, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303585.

DeBaerdemacker:2017:BTU

[DDCY17] Stijn De Baerdemacker, Alexis De Vos, Lin Chen, and Li Yu. The Birkhoff theorem for unitary matrices of arbitrary dimensions. Linear Algebra and its Applications, 514(??): 151–164, February 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305134.

Du:2015:AMS

[DdF15] Zhibin Du and Carlos M. da Fonseca. The acyclic matrices with a P -set of maximum size. Linear Algebra and its Applications, 468 (??):27–37, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379513006770.

Dinh:2017:CBL

[DDF17a] Trung Hoa Dinh, Raluca Dumitru, and Jose A. Franco. On a conjecture of Bhatia, Lim and Yamazaki. Linear Algebra and its Applications, 532(??):140–145, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304044.

See [BLY16].

Dinh:2017:MWP

[DDF17b] Trung Hoa Dinh, Raluca Dumitru, and Jose A. Franco. On the monotonicity of weighted power means for matrices. Linear Algebra and its Applications, 527(??): 128–140, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Dinh:2018:NCO

[DDF18] Trung Hoa Dinh, Raluca Dumitru, and Jose A. Franco. New characterizations of operator monotone functions. Linear Algebra and its Applications, 546(??):169– 186, June 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300594.

Dmytryshyn:2016:CPL

[DdFR16] Andrii Dmytryshyn, Carlos M. da Fonseca, and Tetiana Rybalkina. Classification of pairs of linear mappings between two vector spaces and between their quotient space and subspace. Linear Algebra and its Applications, 509(??):228– 246, November 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302786.

DeTeran:2017:EDI

[DDL17] Fernando De Terán, Froilán M. Dopico, and J. M. Landsberg. An explicit description of the irreducible components of the set of matrix pencils with bounded normal rank. Linear Algebra and its Applications, 520 (??):80–103, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300502.

DeTeran:2019:QRP

[DDMP19] Fernando De Terán, Froilán M. Dopico, D. Steven Mackey, and Vasilije Perović. Quadratic realizability of palindromic matrix polynomials. Linear Algebra and its Applications, 567(??):202–262, April 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300163.

Teran:2016:PZM

[DDMV16] Fernando De Terán, Froilán M. Dopico, D. Steven Mackey, and Paul Van Dooren. Polynomial zigzag matrices, dual minimal bases, and the realization of completely singular polynomials. Linear Algebra and its Applications, 488(??):460–504, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005303.

DellAcqua:2016:OPI

[DDSC^{+}16] P. Dell’Acqua, M. Donatelli, S. Serra-Capizzano, D. Sesana, and C. Tablino-Possio. optimal preconditioning for image deblurring with anti-reflective boundary conditions. Linear Algebra and its Applications, 502(??):159–185, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005078.

Das:2015:BEP

[DdSJdFDV15] Kinkar Ch. Das, Celso M. da Silva Junior, Maria Aguieiras A. de Freitas, and Renata R. Del-Vecchio. Bounds on the entries of the principal eigenvector of the distance signless Laplacian matrix. Linear Algebra and its Applications, 483(??): 200–220, October 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003560.

Tern:2016:CSI

[DDV16] Fernando De Terán, Froilán M. Dopico, and Paul Van Dooren. Constructing strong ℓ-ifications from dual minimal bases. Linear Algebra and its Applications, 495 (??):344–372, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000495.

DeBruyn:2016:PED

[De 16a] Bart De Bruyn. Projective embeddings of dual polar spaces of mixed type. Linear Algebra and its Applications, 501(??):263–289, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516300416.

DeParis:2016:ALT

[De 16b] Alessandro De Paris. The asymptotic leading term for maximum rank of ternary forms of a given degree. Linear Algebra and its Applications, 500(??):15–29, July 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001713.

Dubbs:2015:IRM

[DE15] Alexander Dubbs and Alan Edelman. Infinite random matrix theory, tridiagonal bordered Toeplitz matrices, and the moment problem. Linear Algebra and its Applications, 467(??):188–201, February 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007307.

Dewilde:2016:FMQ

[DEH16] P. Dewilde, Y. Eidelman, and I. Haimovici. LU factorization for matrices in quasiseparable form via orthogonal transformations. Linear Algebra and its Applications, 502(??):5–40, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516000197.

Dere:2019:SUL

[Der19] Jonas Deré. The structure of underlying Lie algebras. Linear Algebra and its Applications, 581(??):471– 495, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930326X.

DeVille:2019:OGS

[DeV19]  Lee DeVille.  Optimiz-

ing Gershgorin for symmetric matrices. Linear Algebra and its Applications, 577(??):360–383, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301922.

Deng:2015:PLM

[DF15] Guodong Deng and Yun Fan. Permutation-like matrix groups with a maximal cycle of power of odd prime length. Linear Algebra and its Applications, 480(??): 1–10, September 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002591.

Dalfo:2016:CDL

[DF16] C. Dalfó and M. A. Fiol. Cospectral digraphs from locally line digraphs. Linear Algebra and its Applications, 500(??):52–62, July 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001737.

Deng:2018:PLM

[DF18] Guodong Deng and Yun Fan. Permutation-like matrix groups with a maximal cycle of length power of two. Linear Algebra and its Applications, 553(??):292– 306, September 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301800.

daFonseca:2017:TCS

[dFFRS17] Carlos M. da Fonseca, Vyacheslav Futorny, Tetiana Rybalkina, and Vladimir V. Sergeichuk. Topological classification of systems of bilinear and sesquilinear forms. Linear Algebra and its Applications, 515(??):1– 5, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305286.

Dmytryshyn:2015:CCC

[DFK^{+}15] Andrii Dmytryshyn, Vyacheslav Futorny, Bo Kågström, Lena Klimenko, and Vladimir V. Sergeichuk. Change of the congruence canonical form of 2-by-2 and 3-by-3 matrices under perturbations and bundles of matrices under congruence. Linear Algebra and its Applications, 469(??):305–334, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007289.

Dmytryshyn:2017:GRS

[DFKS17] Andrii Dmytryshyn, Vyacheslav Futorny, Tetiana Klymchuk, and Vladimir V. Sergeichuk. Generalization of Roth’s solvability criteria to systems of matrix equations. Linear Algebra and its Applications, 527(??): 294–302, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302367.

Djikic:2017:MOR

[DFM17] Marko S. Djikić, Guillermina Fongi, and Alejandra Maestripieri. The minus order and range additivity. Linear Algebra and its Applications, 531(??): 234–256, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303452.

deFreitas:2016:MQI

[dFNP16] Maria Aguieiras A. de Freitas, Vladimir Nikiforov, and Laura Patuzzi. Maxima of the Q-index: Graphs with no K_{s,t}. Linear Algebra and its Applications, 496 (??):381–391, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000902.

daFonseca:2016:TCS

[dFRS16] Carlos M. da Fonseca, Tetiana Rybalkina, and Vladimir V. Sergeichuk. Topological classification of sesquilinear forms: Reduction to the nonsingular case. Linear Algebra and its Applications, 504(??):581–589, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Das:2018:LEL

[DG18] Kinkar Ch. Das and Ivan Gutman. On Laplacian energy, Laplacian-energy-like invariant and Kirchhoff index of graphs. Linear Algebra and its Applications, 554(??):170–184, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302702.

DeCastro-Garca:2016:CNR

[DGCC16] Noemí DeCastro-García, Miguel V. Carriegos, and Ángel Luis Muñoz Castañeda. A characterization of von Neumann rings in terms of linear systems. Linear Algebra and its Applications, 494 (??):236–244, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000227.

DeCastro-Garcia:2018:CLS

[DGGP18] Noemí DeCastro-García and M. I. García-Planas. Concatenated linear systems over rings and their application to construction of concatenated families of convolutional codes. Linear Algebra and its Applications, 542(??):624–647, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306791.

Dantas:2017:BPB

[DGK^{+}17] Sheldon Dantas, Domingo García, Sun Kwang Kim, Han Ju Lee, and Manuel Maestre. On the Bishop– Phelps–Bollobás theorem for multilinear mappings. Linear Algebra and its Applications, 532(??):406–431, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304123.

Das:2018:DBE

[DGM^{+}18] Kinkar Ch. Das, Ivan Gutman, Igor Milovanović, Emina Milovanović, and Boris Furtula. Degree-based energies of graphs. Linear Algebra and its Applications, 554(??):185–204, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302672.

Dunkl:2015:RNS

[DGP^{+}15] Charles F. Dunkl, Piotr Gawron, Lukasz Pawela, Zbigniew Puchala, and Karol Zyczkowski. Real numerical shadow and generalized B-splines. Linear Algebra and its Applications, 479 (??):12–51, August 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002050.

Dahl:2018:MMC

[DGS18] Geir Dahl, Alexander Guterman, and Pavel Shteyner. Majorization for matrix classes. Linear Algebra and its Applications, 555(??): 201–221, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302878.

Daepp:2017:MTB

[DGSV17] Ulrich Daepp, Pamela Gorkin, Andrew Shaffer, and Karl Voss. Möbius transformations and Blaschke products: the geometric connection. Linear Algebra and its Applications, 516 (??):186–211, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305584.

DaSilva:2015:OHT

[DH15] Curt Da Silva and Felix J. Herrmann. Optimization on the Hierarchical Tucker manifold — applications to tensor completion. Linear Algebra and its Applications, 481(??):131– 173, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002530.

Drensky:2019:NIT

[DH19] Vesselin Drensky and Elitza Hristova. Noncommutative invariant theory of symplectic and orthogonal groups. Linear Algebra and its Applications, 581(??):198–213, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302940.

Das:2019:SBT

[DHHZI19] Kinkar Ch. Das, Mohammad Ali Hosseinzadeh, Samaneh Hossein-Zadeh, and Ali Iranmanesh. Some bounds for total communicability of graphs. Linear Algebra and its Applications, 569 (??):266–284, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300473.

Dairyko:2017:NNR

[DHL^{+}17] Michael Dairyko, Leslie Hogben, Jephian C.-H. Lin, Joshua Lockhart, David Roberson, Simone Severini, and Michael Young. Note on von Neumann and Rényi entropies of a graph. Linear Algebra and its Applications, 521(??):240–253, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517300666.

Dendievel:2019:TDE

[DHLT19] Sarah Dendievel, Sophie Hautphenne, Guy Latouche, and Peter G. Taylor. The time-dependent expected reward and deviation matrix of a finite QBD process. Linear Algebra and its Applications, 570(??):61–92, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300515.

Ducey:2018:CGK

[DHS18] Joshua E. Ducey, Ian Hill, and Peter Sin. The critical group of the Kneser graph on 2-subsets of an nelement set. Linear Algebra and its Applications, 546 (??):154–168, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300661.

Dutkay:2016:SPS

[DHW16] Dorin Ervin Dutkay, John Haussermann, and Eric Weber. Spectral properties of small Hadamard matrices. Linear Algebra and its Applications, 506(??): 363–381, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302233.

Du:2018:TGI

[DHW18] Kui Du, Yunqing Huang, and Yiwei Wang. On two generalized inverse eigenvalue problems for Hessenberg-upper triangular pencils and their application to the study of GMRES convergence. Linear Algebra and its Applications, 553(??):16–36, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302258.

DeTeran:2016:USG

[DI16] Fernando De Terán and Bruno Iannazzo. Uniqueness of solution of a generalized ∗-Sylvester matrix equation. Linear Algebra and its Applications, 493(??): 323–335, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Dickinson:2019:NCC

[Dic19] Peter J. C. Dickinson. A new certificate for copositivity. Linear Algebra and its Applications, 569(??):15– 37, May 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305962.

DeTeran:2018:CSU

[DIPR18a] Fernando De Terán, Bruno Iannazzo, Federico Poloni, and Leonardo Robol. Corrigendum to “Solvability and uniqueness criteria for generalized Sylvester-type equations”. Linear Algebra and its Applications, 542(??): 522–526, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305979.

DeTeran:2018:SUC

[DIPR18b] Fernando De Terán, Bruno Iannazzo, Federico Poloni, and Leonardo Robol. Solvability and uniqueness criteria for generalized Sylvestertype equations. Linear Algebra and its Applications, 542(??):501–521, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304202.

Datt:2016:OTA

[DIR16] M. Sumanth Datt, Uma N. Iyer, and G. Koteswara Rao. Operators on triangular algebras. Linear Algebra and its Applications, 501 (??):254–262, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300490.

Dykema:2016:IKL

[DK16] Kenneth J. Dykema and Igor Klep. Instances of the Kaplansky–Lvov multilinear conjecture for polynomials of degree three. Linear Algebra and its Applications, 508(??):272–288, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303329.

Draga:2018:WMB

[DK18] Szymon Draga and Tomasz Kania. When is multiplication in a Banach algebra open? Linear Algebra and its Applications, 538(??): 149–165, February 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Dolinar:2016:RSM

[DKNS16] G. Dolinar, B. Kuzma, G. Nagy, and P. Szokol. Restricted skew-morphisms on matrix algebras. Linear Algebra and its Applications, 490(??):1–17, February 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515005984.

Dufosse:2018:FNB

[DKPU18] Fanny Dufossé, Kamer Kaya, Ioannis Panagiotas, and Bora Uçar. Further notes on Birkhoff–von Neumann decomposition of doubly stochastic matrices. Linear Algebra and its Applications, 554(??):68–78, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302568.

Dehghani:2015:DMI

[DKS15] Mahdi Dehghani, Mohsen Kian, and Yuki Seo. Developed matrix inequalities via positive multilinear mappings. Linear Algebra and its Applications, 484(??): 63–85, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500381X.

Dehghani:2017:MPM

[DKS17] Mahdi Dehghani, Mohsen Kian, and Yuki Seo. Matrix power means and the information monotonicity. Linear Algebra and its Applications, 521(??):57–69, May 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951730054X.

DAndrea:2019:CFU

[DKSV19] Carlos D’Andrea, Teresa Krick, Agnes Szanto, and Marcelo Valdettaro. Closed formula for univariate subresultants in multiple roots. Linear Algebra and its Applications, 565(??):123–155, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305731.

Drury:2015:GNM

[DL15] Stephen Drury and Huiqiu Lin. On graphs with near minimal distance energy. Linear Algebra and its Applications, 477(??):76– 101, July 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Dong:2016:LMG

[DL16] Jiu-Gang Dong and Lin Lin. Laplacian matrices of general complex weighted directed graphs. Linear Algebra and its Applications, 510 (??):1–9, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630338X.

Deng:2018:BTE

[DLB18] Chunli Deng, Haifeng Li, and Changjiang Bu. Brauertype eigenvalue inclusion sets of stochastic/irreducible tensors and positive definiteness of tensors. Linear Algebra and its Applications, 556(??):55–69, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303161.

delaCruz:2015:ESM

[dlC15] Ralph John de la Cruz. Each symplectic matrix is a product of four symplectic involutions. Linear Algebra and its Applications, 466(??): 382–400, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006855.

delaCruz:2017:ECS

[dlCdR17] Ralph John de la Cruz and Kennett dela Rosa. Each 2n-by-2n complex symplectic matrix is a product of n + 1 commutators of Jsymmetries. Linear Algebra and its Applications, 517 (??):53–62, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305857.

delaCruz:2016:DMS

[dlCF16] Ralph John de la Cruz and Heike Faßbender. On the diagonalizability of a matrix by a symplectic equivalence, similarity or congruence transformation. Linear Algebra and its Applications, 496 (??):288–306, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000331.

delaCruz:2018:PSN

[dlCG18] Ralph John de la Cruz and Daryl Q. Granario. Products of symplectic normal matrices. Linear Algebra and its Applications, 543 (??):162–172, April 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

delaCruz:2015:OMS

[dlCMP15] Ralph John de la Cruz, Dennis I. Merino, and Agnes T. Paras. S orthogonal matrices and S symmetries. Linear Algebra and its Applications, 474(??):213–229, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515001305.

delaCruz:2017:ECM

[dlCMP17a] Ralph John de la Cruz, Dennis I. Merino, and Agnes T. Paras. Every 2n-by-2n complex matrix is a sum of three symplectic matrices. Linear Algebra and its Applications, 517(??): 199–206, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630605X.

delaCruz:2017:SPD

[dlCMP17b] Ralph John de la Cruz, Dennis I. Merino, and Agnes T. Paras. Skew φ polar decompositions. Linear Algebra and its Applications, 531(??):129–140, October 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517303294.

Deng:2016:MBG

[DLG16] Bo Deng, Xueliang Li, and Ivan Gutman. More on borderenergetic graphs. Linear Algebra and its Applications, 497(??):199–208, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001415.

Donatelli:2015:SFT

[DLM15] Marco Donatelli, Alessandra Luati, and Andrea Martinelli. Spectral filtering for trend estimation. Linear Algebra and its Applications, 473(??): 217–235, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514004042.

delaPena:2018:TCP

[dlP18] José A. de la Peña. On the trace of the Coxeter polynomial of an algebra. Linear Algebra and its Applications, 538(??):103–115, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305943.

delaPena:2018:BRC

[dlPJG18] José Antonio de la Peña and Jesús Arturo Jiménez-González. Bounds related to Coxeter spectral measures of graphs. Linear Algebra and its Applications, 545 (??):32–54, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300442.

Ding:2018:PTP

[DLQ18] Weiyang Ding, Ziyan Luo, and Liqun Qi. P -tensors, P_{0}-tensors, and their applications. Linear Algebra and its Applications, 555(??): 336–354, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303124.

Drton:2015:PSI

[DLW15] Mathias Drton, Lek-Heng Lim, and Wei Biao Wu. Preface to the special issue on statistics. Linear Algebra and its Applications, 473(??):1–2, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002402.

Durfee:2015:DGZ

[DM15] Christina Durfee and Kimball Martin. Distinguishing graphs with zeta functions and generalized spectra. Linear Algebra and its Applications, 481(??):54– 82, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002797.

Das:2016:RBS

[DM16a] Kinkar Ch. Das and Seyed Ahmad Mojallal. Relation between signless Laplacian energy, energy of graph and its line graph. Linear Algebra and its Applications, 493 (??):91–107, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007168.

Deaett:2016:MRP

[DM16b] Louis Deaett and Seth A. Meyer. The minimum rank problem for circulants. Linear Algebra and its Applications, 491(??):386–418, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500662X.

Dratman:2018:IBI

[DM18] Ezequiel Dratman and Guillermo Matera. Improved bounds for the inverses of diagonally dominant tridiagonal matrices. Linear Algebra and its Applications, 551 (??):57–82, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301307.

Das:2016:ELG

[DMG16] Kinkar Ch. Das, Seyed Ahmad Mojallal, and Ivan Gutman. On energy of line graphs. Linear Algebra and its Applications, 499 (??):79–89, June 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001610.

Das:2017:RBD

[DMG17] Kinkar Ch. Das, Seyed Ahmad Mojallal, and Ivan Gutman. Relations between degrees, conjugate degrees and graph energies. Linear Algebra and its Applications, 515(??): 24–37, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305250.

Goulart:2019:MRM

[dMGC19] J. H. de M. Goulart and P. Comon. On the minimal ranks of matrix pencils and the existence of a best approximate blockterm tensor decomposition. Linear Algebra and its Applications, 561(??):161–186, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304725.

Dopico:2019:SLR

[DMQ19] Froilán M. Dopico, Silvia Marcaida, and María C. Quintana. Strong linearizations of rational matrices with polynomial part expressed in an orthogonal basis. Linear Algebra and its Applications, 570(??):1– 45, June 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300588.

Dellacherie:2016:PRW

[DMS16] Claude Dellacherie, Servet Martinez, and Jaime San Martin. Potentials of random walks on trees. Linear Algebra and its Applications, 501(??):123–161, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516300118.

Das:2019:SLE

[DMS19] Kinkar Chandra Das, Seyed Ahmad Mojallal, and Shaowei Sun. On the sum of the k largest eigenvalues of graphs and maximal energy of bipartite graphs. Linear Algebra and its Applications, 569 (??):175–194, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300400.

Das:2016:DLE

[DMT16] Kinkar Ch. Das, Seyed Ahmad Mojallal, and Vilmar Trevisan. Distribution of Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 508(??): 48–61, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302567.

Dmytryshyn:2016:MDP

[Dmy16] Andrii Dmytryshyn. Miniversal deformations of pairs of skew-symmetric matrices under congruence. Linear Algebra and its Applications, 506(??):506–534, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302324.

Dmytryshyn:2017:SPS

[Dmy17] Andrii Dmytryshyn. Structure preserving stratification of skew-symmetric matrix polynomials. Linear Algebra and its Applications, 532(??):266–286, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304081.

Dmytryshyn:2019:MDP

[Dmy19] Andrii Dmytryshyn. Miniversal deformations of pairs of symmetric matrices under congruence. Linear Algebra and its Applications, 568 (??):84–105, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830274X.

Dubey:2016:HZM

[DN16] Dipti Dubey and S. K. Neogy. On hidden Z-matrices and the linear complementarity problem. Linear Algebra and its Applications, 496 (??):81–100, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000872.

Dubey:2018:TDI

[DN18] Dipti Dubey and S. K. Neogy. Total dual integrality and integral solutions of the linear complementarity problem. Linear Algebra and its Applications, 557(??):359–374, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303768.

Dolzan:2016:MSN

[DO16] David Dolzan and Polona Oblak. Maximal simultaneously nilpotent sets of matrices over antinegative semirings. Linear Algebra and its Applications, 510(??): 222–229, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303512.

Dodig:2016:CQR

[Dod16]  Marija Dodig.  Comple-

tion of quasi-regular matrix pencils. Linear Algebra and its Applications, 501 (??):198–241, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300064.

Dodig:2017:MCP

[Dod17] Marija Dodig. Minimal completion problem for quasiregular matrix pencils. Linear Algebra and its Applications, 525(??):84–104, July 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517301611.

Dogan:2018:DIM

[Dog18] Hamide Dogan. Differing instructional modalities and cognitive structures: Linear algebra. Linear Algebra and its Applications, 542 (??):464–483, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304172.

Dzhumadildaev:2016:CEA

[DOR16] A. Dzhumadil’daev, B. A. Omirov, and U. A. Rozikov. Constrained evolution algebras and dynamical systems of a bisexual population. Linear Algebra and its Applications, 496 (??):351–380, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000938.

DeSimone:2015:FLH

[DP15a] Anna De Simone and Pavel Pták. On the Farkas lemma and the Horn–Tarski measure-extension theorem. Linear Algebra and its Applications, 481(??):243–248, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Dilip:2015:YAC

[DP15b] A. Sanand Amita Dilip and Harish K. Pillai. Yet another characterization of solutions of the Algebraic Riccati Equation. Linear Algebra and its Applications, 481(??):1–35, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515002773.

Dilip:2016:CSN

[DP16] A. Sanand Amita Dilip and Harish K. Pillai. Characterization of solutions of non-symmetric algebraic Riccati equations. Linear Algebra and its Applications, 507(??):356–372, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630249X.

Dey:2018:CCD

[DP18] Papri Dey and Harish K. Pillai. A complete characterization of determinantal quadratic polynomials. Linear Algebra and its Applications, 543(??):106–124, April 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306845.

Dilip:2017:DAR

[DPJ17] A. Sanand Amita Dilip, Harish K. Pillai, and Raphaël M. Jungers. On discrete algebraic Riccati equations: a rank characterization of solutions. Linear Algebra and its Applications, 527(??): 184–215, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302148.

Diaz:2019:NRM

[DPR19] Roberto C. Díaz, Germain Pastén, and Oscar Rojo. New results on the D_{α}- matrix of connected graphs. Linear Algebra and its Applications, 577(??):168–185, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301880.

Domelevo:2019:FMW

[DPS19] Komla Domelevo, Stefanie Petermichl, and Kristina Ana Skreb. Failure of the matrix weighted bilinear Carleson embedding theorem. Linear Algebra and its Applications, 582(??):452–466, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303519.

Dopico:2019:BMB

[DPV19] Froilán M. Dopico, Javier Pérez, and Paul Van Dooren. Block minimal bases lifications of matrix polynomials. Linear Algebra and its Applications, 562(??): 163–204, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304841.

Deadman:2016:TTM

[DR16] Edvin Deadman and Samuel D. Relton. Taylor’s theorem for matrix functions with applications to condition number estimation. Linear Algebra and its Applications, 504(??):354–371, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300957.

Douglas:2017:SRT

[DR17] Andrew Douglas and Joe Repka. The subalgebras of the rank two symplectic Lie algebra. Linear Algebra and its Applications, 527(??): 303–348, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302112.

Diaz:2018:SUB

[DR18] Roberto C. Díaz and Oscar Rojo. Sharp upper bounds on the distance energies of a graph. Linear Algebra and its Applications, 545 (??):55–75, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300466.

Duttweiler:2019:SCP

[DR19] Luke Duttweiler and Nathan Reff. Spectra of cycle and path families of oriented hypergraphs. Linear Algebra and its Applications, 578(??):251–271, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302101.

Drazin:2016:LRG

[Dra16a] Michael P. Drazin. Left and right generalized inverses. Linear Algebra and its Applications, 510(??): 64–78, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303378.

Drazin:2016:RJJ

[Dra16b] Michael P. Drazin. On a result of J. J. Sylvester. Linear Algebra and its Applications, 505(??):361–366, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301574.

Dragicevic:2018:NNE

[Dra18] Davor Dragicević. A note on the nonuniform exponential stability and dichotomy for nonautonomous difference equations. Linear Algebra and its Applications, 552(??):105–126, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830212X.

Drnovsek:2019:SBS

[Drn19] Roman Drnovsek. Sequences of bounds for the spectral radius of a positive operator. Linear Algebra and its Applications, 574(??):40– 45, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301314.

Duffner:2015:III

[DS15] M. Graça Duffner and Fernando C. Silva. On the interlacing inequalities for invariant factors. Linear Algebra and its Applications, 486(??):443–448, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004565.

Das:2017:DBN

[DS17a] Kinkar Ch. Das and Shaowei Sun. Distance between the normalized Laplacian spectra of two graphs. Linear Algebra and its Applications, 530(??):305–321, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303130.

Duffner:2017:EUM

[DS17b] M. Graça Duffner and Fernando C. Silva. On the existence of unimodular matrices with a prescribed submatrix. Linear Algebra and its Applications, 515(??): 321–330, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305419.

deSa:2018:NPN

[dS18a] Eduardo Marques de Sá. Noncrossing partitions, noncrossing graphs, and qpermanental equations. Linear Algebra and its Applications, 541(??):36–53, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306523.

Diao:2018:MCC

[DS18b] Huai-An Diao and Yang Sun. Mixed and componentwise condition numbers for a linear function of the solution of the total least squares problem. Linear Algebra and its Applications, 544(??):1–29, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300144.

deSa:2019:LPP

[dS19a] Eduardo Marques de Sá. Linear preservers for the qpermanent, cycle q-permanent expansions, and positive crossings in digraphs. Linear Algebra and its Applications, 561(??):228–252, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304683.

Didenko:2019:ITP

[DS19b] Victor D. Didenko and Bernd Silbermann. The invertibility of Toeplitz plus Hankel operators with subordinated operators of even index. Linear Algebra and its Applications, 578(??): 425–445, October 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930237X.

Dodig:2019:FIS

[DS19c] Marija Dodig and Marko Stosić. Feedback invariants of series connected systems. Linear Algebra and its Applications, 577(??):244– 269, September 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301892.

Dou:2017:GED

[DSCD17] Yan-Ni Dou, Wei-Juan Shi, Miao-Miao Cui, and Hong-Ke Du. General explicit descriptions for intertwining operators and direct rotations of two orthogonal projections. Linear Algebra and its Applications, 531(??): 575–591, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304007.

Silva:2015:GFM

[dSN15] Celso Marques da Silva, Jr. and Vladimir Nikiforov. Graph functions maximized on a path. Linear Algebra and its Applications, 485(??):21–32, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004127.

Pazzis:2015:RCH

[dSP15] Clément de Seguins Pazzis. Range-compatible homomorphisms on matrix spaces. Linear Algebra and its Applications, 484(??):237–289, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003778.

Pazzis:2016:ASS

[dSP16a] Clément de Seguins Pazzis. Affine spaces of symmetric or alternating matrices with bounded rank. Linear Algebra and its Applications, 504(??):503–558, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301252.

Pazzis:2016:FTD

[dSP16b] Clément de Seguins Pazzis. The Flanders theorem over division rings. Linear Algebra and its Applications, 493(??):313–322, March 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515006965.

Pazzis:2016:LSB

[dSP16c] Clément de Seguins Pazzis. Large spaces of bounded rank matrices revisited. Linear Algebra and its Applications, 504(??):124–189, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630074X.

Pazzis:2016:LSS

[dSP16d] Clément de Seguins Pazzis. Large spaces of symmetric or alternating matrices with bounded rank. Linear Algebra and its Applications, 508(??):146–189, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302671.

Pazzis:2016:RCH

[dSP16e] Clément de Seguins Pazzis. Range-compatible homomorphisms on spaces of symmetric or alternating matrices. Linear Algebra and its Applications, 503(??):

135–163, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300702.

Pazzis:2019:LPS

[dSP19a] Clément de Seguins Pazzis. On the linear preservers of Schur matrix functionals. Linear Algebra and its Applications, 567(??):63– 117, April 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300035.

Pazzis:2019:SGPa

[dSP19b] Clément de Seguins Pazzis. The structured Gerstenhaber problem (I). Linear Algebra and its Applications, 567(??):263–298, April 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303884.

Pazzis:2019:SGPb

[dSP19c] Clément de Seguins Pazzis. The structured Gerstenhaber problem (II). Linear Algebra and its Applications, 569(??):113–145, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519300217.

daSilva:2017:NSC

[dSRST17] J. A. Dias da Silva, Fátima Rodrigues, Pedro C. Silva, and Maria M. Torres. A note on symmetry classes of tensors and Specht modules. Linear Algebra and its Applications, 531(??): 118–128, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730335X.

Dodig:2015:MQF

[DSX15] Marija Dodig, Marko Stosić, and João Xavier. On minimizing a quadratic function on Stiefel manifold. Linear Algebra and its Applications, 475(??):251–264, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515001354.

Duarte:2015:EIG

[DT15] Pedro Duarte and Maria Joana Torres. Eigenvectors of isospectral graph transformations. Linear Algebra and its Applications, 474 (??):110–123, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

DiFiore:2016:LTT

[DTZ16] Carmine Di Fiore, Francesco Tudisco, and Paolo Zellini. Lower triangular Toeplitz– Ramanujan systems whose solution yields the Bernoulli numbers. Linear Algebra and its Applications, 496 (??):510–526, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000987.

Dopico:2016:CMS

[DU16a] Froilán Dopico and Frank Uhlig. Computing matrix symmetrizers, part 2: New methods using eigendata and linear means; a comparison. Linear Algebra and its Applications, 504(??): 590–622, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003845.

Dufosse:2016:NBN

[DU16b] Fanny Dufossé and Bora Uçar. Notes on Birkhoff– von Neumann decomposition of doubly stochastic matrices. Linear Algebra and its Applications, 497 (??):108–115, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001257.

Dukes:2012:RDD

[Duk12] Peter Dukes. Rational decomposition of dense hypergraphs and some related eigenvalue estimates. Linear Algebra and its Applications, 436(9):3736–3746, May 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379512000626.

See corrigendum [Duk15].

Dukes:2015:CSD

[Duk15] Peter J. Dukes. Corrigendum to “Rational decomposition of dense hypergraphs and some related eigenvalue estimates” [Linear Algebra Appl. 436 (9) (2012) 3736– 3746]. Linear Algebra and its Applications, 467(??): 267–269, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007587.

See [Duk12].

DeCarli:2018:SRG

[DV18] Laura De Carli and Pierluigi Vellucci. Stability results for Gabor frames and the p-order hold models. Linear Algebra and its Applications, 536(??):

186–200, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730544X.

Devriendt:2019:TSB

[DV19a] Karel Devriendt and Piet Van Mieghem. Tighter spectral bounds for the cut size, based on Laplacian eigenvectors. Linear Algebra and its Applications, 572 (??):68–91, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301016.

Dopico:2019:RPM

[DV19b] Froilán M. Dopico and Paul Van Dooren. Robustness and perturbations of minimal bases II: the case with given row degrees. Linear Algebra and its Applications, 576(??):268–300, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304968.

Duvenhage:2018:ASP

[DvSW18] Rocco Duvenhage, Wernd van Staden, and Jan Wuzyk. Analyticity and spectral properties of noncommutative Ricci flow in a matrix geometry. Linear Algebra and its Applications, 539(??): 160–174, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306328.

Deng:2015:SRP

[DY15] Chunyuan Deng and Anqi Yu. Some relations of projection and star order in Hilbert space. Linear Algebra and its Applications, 474 (??):158–168, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001238.

Dyachenko:2017:HMD

[Dya17] Alexander Dyachenko. Hurwitz matrices of doubly infinite series. Linear Algebra and its Applications, 530(??):266–287, October 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517303002.

Dymarsky:2016:CSB

[Dym16] Anatoly Dymarsky. Convexity of a small ball under quadratic map. Linear Algebra and its Applications, 488(??):109–123, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Elphick:2017:NGO

[EA17] Clive Elphick and Mustapha Aouchiche. Nordhaus– Gaddum and other bounds for the sum of squares of the positive eigenvalues of a graph. Linear Algebra and its Applications, 530(??): 150–159, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303014.

Estrada:2017:CSG

[EB17] Ernesto Estrada and Michele Benzi. Core-satellite graphs: Clustering, assortativity and spectral properties. Linear Algebra and its Applications, 517(??):30–52, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305869.

Egan:2019:MBC

[EC19] Ronan Egan and Padraig Ó Catháin. Morphisms of Butson classes. Linear Algebra and its Applications, 577(??):78–93, September 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519301788.

Eshkaftaki:2015:CMD

[EE15] Ali Bayati Eshkaftaki and Noha Eftekhari. Convex majorization on discrete ℓ^{p} spaces. Linear Algebra and its Applications, 474 (??):124–140, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007204.

Eftekhari:2016:ILO

[EE16] Noha Eftekhari and Ali Bayati Eshkaftaki. Isotonic linear operators on the space of all convergent real sequences. Linear Algebra and its Applications, 506(??): 535–550, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302221.

Eggers:2017:CCT

[EEV17] John Eggers, Ron Evans, and Mark Van Veen. Classification of certain types of maximal matrix subalgebras. Linear Algebra and its Applications, 532(??): 397–405, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304184.

Evans:2017:ROD

[EF17] Tyler J. Evans and Alice Fialowski. Restricted onedimensional central extensions of restricted simple Lie algebras. Linear Algebra and its Applications, 513(??): 96–102, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304542.

Evans:2019:ROD

[EF19] Tyler J. Evans and Alice Fialowski. Restricted onedimensional central extensions of the restricted filiform Lie algebras m_{0}λ(p). Linear Algebra and its Applications, 565(??):244–257, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305603.

Feng:2015:CAN

[eFLYL15] Jun e Feng, James Lam, Guowu Yang, and Zhao Li. On a conjecture about the norm of Lyapunov mappings. Linear Algebra and its Applications, 465(??): 88–103, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006041.

Elsheikh:2015:RAE

[EG15] Mustafa Elsheikh and Mark Giesbrecht. Relating p-adic eigenvalues and the local Smith normal form. Linear Algebra and its Applications, 481(??):330–349, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002906.

Elsworth:2019:CBB

[EG19] Steven Elsworth and Stefan Güttel. Conversions between barycentric, RK-FUN, and Newton representations of rational interpolants. Linear Algebra and its Applications, 576(??): 246–257, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304774.

Eger:2016:LPC

[Ege16] Steffen Eger. On limits of powers of certain absolutely row-stochastic matrices. Linear Algebra and its Applications, 508(??): 1–13, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630252X.

Eggermont:2015:FPC

[Egg15] Rob H. Eggermont. Finiteness properties of congruence classes of infinite-by-infinite matrices. Linear Algebra and its Applications, 484(??):290–303, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003985.

Escribano:2015:IIM

[EGT15] C. Escribano, R. Gonzalo, and E. Torrano. On the inversion of infinite moment matrices. Linear Algebra and its Applications, 475 (??):292–305, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500138X.

Elsner:2015:MSR

[EH15] L. Elsner and K. P. Hadeler. Maximizing the spectral radius of a matrix product. Linear Algebra and its Applications, 469(??): 153–168, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007551.

Elyseeva:2018:DOT

[EH18] Julia Elyseeva and Roman Simon Hilscher. Discrete oscillation theorems for symplectic eigenvalue problems with general boundary conditions depending nonlinearly on spectral parameter. Linear Algebra and its Applications, 558(??): 108–145, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303860.

Estrada:2017:PLO

[EHHL17] Ernesto Estrada, Ehsan Hameed, Naomichi Hatano, and Matthias Langer. Path Laplacian operators and superdiffusive processes on graphs. I. One-dimensional case. Linear Algebra and its Applications, 523(??): 307–334, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301283.

Eggermont:2016:ABM

[EHK16] Rob H. Eggermont, Emil Horobet, and Kaie Kubjas. Algebraic boundary of matrices of nonnegative rank at most three. Linear Algebra and its Applications, 508(??):62–80, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Estrada:2018:PLO

[EHLP18] Ernesto Estrada, Ehsan Hameed, Matthias Langer, and Aleksandra Puchalska. Path Laplacian operators and superdiffusive processes on graphs. II. Twodimensional lattice. Linear Algebra and its Applications, 555(??):373–397, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303100.

Ehrmann:2016:GSM

[EHN16] Daniel Ehrmann, Zach Higgins, and Viorel Nitica. The geometric structure of max-plus hemispaces. Linear Algebra and its Applications, 489(??):144–166, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006023.

Elhodaibi:2017:MPO

[EJ17] M. Elhodaibi and A. Jaatit. On maps preserving operators of local spectral radius zero. Linear Algebra and its Applications, 512(??): 191–201, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630461X.

Encinas:2018:EIT

[EJ18] A. M. Encinas and M. J. Jiménez. Explicit inverse of a tridiagonal (p, r)-Toeplitz matrix. Linear Algebra and its Applications, 542 (??):402–421, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303737.

Encinas:2019:TSC

[EJ19] A. M. Encinas and M. J. Jiménez. Triangular sequences, combinatorial recurrences and linear difference equations. Linear Algebra and its Applications, 576(??):301–323, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304956.

Eastman:2014:CMP

[EKSV14] B. Eastman, I.-J. Kim, B. L. Shader, and K. N. Vander Meulen. Companion matrix patterns. Linear Algebra and its Applications, 463(??): 255–272, December 15, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/

science/article/pii/S0024379514005953.

See corrigendum [EKSV18].

Eastman:2018:CSC

[EKSV18] B. Eastman, I.-J. Kim, B. L. Shader, and K. N. Vander Meulen. Corrigendum to “Companion matrix patterns” [Linear Algebra Appl. 463 (2014) 255–272]. Linear Algebra and its Applications, 538(??):225–227, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305967.

See [EKSV14].

Elduque:2016:NEA

[EL16] Alberto Elduque and Alicia Labra. On nilpotent evolution algebras. Linear Algebra and its Applications, 505(??):11–31, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301240.

Elden:2015:CFD

[Eld15] Lars Eldén. Computing Frechet derivatives in partial least squares regression. Linear Algebra and its Applications, 473(??): 316–338, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006028.

Echeverria:2018:BDD

[ELN18] Carlos Echeverría, Jörg Liesen, and Reinhard Nabben. Block diagonal dominance of matrices revisited: Bounds for the norms of inverses and eigenvalue inclusion sets. Linear Algebra and its Applications, 553(??):365–383, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302192.

Eshkaftaki:2015:SPO

[EM15] A. Bayati Eshkaftaki and S. M. Manjegani. Some properties of operators preserving convex majorization on discrete ℓ^{p} spaces. Linear Algebra and its Applications, 484(??):130–140, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003705.

Ellia:2016:SMC

[EM16] Ph. Ellia and P. Menegatti. Spaces of matrices of constant rank and uniform vector bundles. Linear Algebra and its Applications, 507(??):474–485, October 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516302361.

Erway:2017:SLS

[EM17] Jennifer B. Erway and Roummel F. Marcia. On solving large-scale limitedmemory quasi-Newton equations. Linear Algebra and its Applications, 515(??): 196–225, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305195.

Emami:2018:RES

[Ema18] Hadi Emami. Ridge estimation in semiparametric linear measurement error models. Linear Algebra and its Applications, 552(??):127– 146, September 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302106.

Essaleh:2018:LMC

[EP18a] Ahlem Ben Ali Essaleh and Antonio M. Peralta. Linear maps on C^{∗}-algebras which are derivations or triple derivations at a point. Linear Algebra and its Applications, 538(??):1–21, February 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517305785.

Essaleh:2018:PAT

[EP18b] Ahlem Ben Ali Essaleh and Antonio M. Peralta. Preservers of λ-Aluthge transforms. Linear Algebra and its Applications, 554(??): 86–119, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830260X.

Eremita:2016:FIU

[Ere16] Daniel Eremita. Functional identities in upper triangular matrix rings. Linear Algebra and its Applications, 493(??):580–605, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007594.

Ernst:2018:SSM

[Ern18] Thomas Ernst. On several q-special matrices, including the q-Bernoulli and q-Euler matrices. Linear Algebra and its Applications, 542(??):422–440, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303919.

Ellard:2016:CSC

[ES16] Richard Ellard and Helena Smigoc. Connecting sufficient conditions for the Symmetric Nonnegative Inverse Eigenvalue Problem. Linear Algebra and its Applications, 498(??):521– 552, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000215.

Estrada:2016:WLG

[Est16] Ernesto Estrada. When local and global clustering of networks diverge. Linear Algebra and its Applications, 488(??):249–263, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005753.

Elsasser:2016:IGH

[ET16] Robert Elsässer and Horst Trinker. On the isomorphism of graphs having some eigenvalues of moderate multiplicity. Linear Algebra and its Applications, 488(??): 377–395, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005388.

Et-Taoui:2018:IFE

[ET18] Boumediene Et-Taoui. Infinite family of equi-isoclinic planes in Euclidean odd dimensional spaces and of complex symmetric conference matrices of odd orders. Linear Algebra and its Applications, 556(??): 373–380, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830332X.

Et-Taoui:2018:SCG

[ETF18] Boumediene Et-Taoui and Augustin Fruchard. On switching classes of graphs. Linear Algebra and its Applications, 549(??):246–255, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518301472.

Ellers:2015:HTC

[EV15] Erich W. Ellers and Oliver Villa. Half turns in characteristic 2. Linear Algebra and its Applications, 483(??):221–226, October 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515003547.

Evert:2018:MCS

[Eve18] Eric Evert. Matrix convex sets without absolute extreme points. Linear Algebra and its Applications, 537(??):287–301, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305670.

Evans:2019:PTP

[EW19] Ron Evans and Nolan Wallach. Pfaffians of Toeplitz payoff matrices. Linear Algebra and its Applications, 577(??):114–120, September 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519301843.

Fagundes:2019:IMP

[Fag19] Pedro S. Fagundes. The images of multilinear polynomials on strictly upper triangular matrices. Linear Algebra and its Applications, 563(??):287–301, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305378.

Farashahi:2016:WPT

[Far16] Arash Ghaani Farashahi. Wave packet transforms over finite cyclic groups. Linear Algebra and its Applications, 489(??):75–92, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005960.

Farid:2019:TCM

[Far19] Farid O. Farid. On three classes of matrices with variants of the diagonal dominance property. Linear Algebra and its Applications, 579(??):382–418, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302502.

Fasi:2019:OPS

[Fas19]  Massimiliano Fasi.  Op-

timality of the Paterson– Stockmeyer method for evaluating matrix polynomials and rational matrix functions. Linear Algebra and its Applications, 574(??): 182–200, August 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301454.

Faybusovich:2018:PDP

[Fay18] Leonid Faybusovich. Primaldual potential reduction algorithm for symmetric programming problems with nonlinear objective functions. Linear Algebra and its Applications, 536(??): 228–249, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305414.

Fan:2019:ENS

[FBH19] Yi-Zheng Fan, Yan-Hong Bao, and Tao Huang. Eigenvariety of nonnegative symmetric weakly irreducible tensors associated with spectral radius and its application to hypergraphs. Linear Algebra and its Applications, 564(??):72–94, March 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518305500.

Feng:2016:SRE

[FCL^{+}16] Lihua Feng, Jianxiang Cao, Weijun Liu, Shifeng Ding, and Henry Liu. The spectral radius of edge chromatic critical graphs. Linear Algebra and its Applications, 492 (??):78–88, March 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006825.

Fabila-Carrasco:2018:SGD

[FCLP18] John Stewart Fabila-Carrasco, Fernando Lledó, and Olaf Post. Spectral gaps and discrete magnetic Laplacians. Linear Algebra and its Applications, 547(??): 183–216, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300673.

Fernandes:2015:CCN

[FdC15] Rosário Fernandes and Henrique F. da Cruz. A canonical construction for nonnegative integral matrices with given line sums. Linear Algebra and its Applications, 484(??):304–321, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003961.

Fernandes:2018:NPV

[FdC18] Rosário Fernandes and Henrique F. da Cruz. The number of P -vertices in a matrix with maximum nullity. Linear Algebra and its Applications, 547(??): 168–182, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300855.

Fernandes:2018:MDL

[FdFdSDV18] Rosário Fernandes, Maria Aguieiras A. de Freitas, Celso M. da Silva, and Renata R. Del-Vecchio. Multiplicities of distance Laplacian eigenvalues and forbidden subgraphs. Linear Algebra and its Applications, 541 (??):81–93, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306596.

Fernandes:2015:CDV

[Fer15] Rosário Fernandes. Computing the degree of a vertex in the skeleton of acyclic Birkhoff polytopes. Linear Algebra and its Applications, 475(??):119–133, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515000877.

Fongi:2016:PIP

[FG16] Guillermina Fongi and M. Celeste Gonzalez. Partial isometries and pseudoinverses in semi-Hilbertian spaces. Linear Algebra and its Applications, 495(??): 324–343, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951600029X.

Francisco:2017:FPM

[FG17] Juliano B. Francisco and Douglas S. Gonçalves. A fixed-point method for approximate projection onto the positive semidefinite cone. Linear Algebra and its Applications, 523(??):59– 78, June 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300903.

Fang:2018:MRS

[FGG^{+}18] Wei Fang, Wei Gao, Yubin Gao, Fei Gong, Guangming Jing, Zhongshan Li, Yanling Shao, and Lihua Zhang. Minimum ranks of sign patterns and zerononzero patterns and pointhyperplane configurations. Linear Algebra and its Applications, 558(??):44–62, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304002.

Fiedler:2015:RDA

[FGH^{+}15] Miroslav Fiedler, Wei Gao, Frank J. Hall, Guangming Jing, Zhongshan Li, and Mikhail Stroev. Ranks of dense alternating sign matrices and their sign patterns. Linear Algebra and its Applications, 471(??): 109–121, April 15, 2015.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000257.

Fang:2016:GCI

[FGS^{+}16] Wei Fang, Yubin Gao, Yanling Shao, Wei Gao, Guangming Jing, and Zhongshan Li. The generalized competition indices of primitive minimally strong digraphs. Linear Algebra and its Applications, 493(??): 206–226, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007107.

Futorny:2019:WT

[FGS19] Vyacheslav Futorny, Joshua A. Grochow, and Vladimir V. Sergeichuk. Wildness for tensors. Linear Algebra and its Applications, 566 (??):212–244, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305937.

Futorny:2019:CKF

[FH19] Vyacheslav Futorny and Jonas T. Hartwig. De Concini–Kac filtration and Gelfand–Tsetlin generators for quantum glN. Linear Algebra and its Applications, 568(??):173–188, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518303847.

Francetic:2015:MNR

[FHH15] Nevena Francetić, Sarada Herke, and Daniel Horsley. More nonexistence results for symmetric pair coverings. Linear Algebra and its Applications, 487(??): 43–73, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005194.

Furer:2019:ELG

[FHJT19] Martin Fürer, Carlos Hoppen, David P. Jacobs, and Vilmar Trevisan. Eigenvalue location in graphs of small clique-width. Linear Algebra and its Applications, 560 (??):56–85, January 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304506.

Foata:2017:EPM

[FHS17a] Dominique Foata, Guo-Niu Han, and Volker Strehl. The Entringer–Poupard matrix sequence. Linear Algebra and its Applications, 512 (??):71–96, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304153.

Futorny:2017:SCS

[FHS17b] Vyacheslav Futorny, Roger A. Horn, and Vladimir V. Sergeichuk. Specht’s criterion for systems of linear mappings. Linear Algebra and its Applications, 519 (??):278–295, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300198.

Forrester:2016:RES

[FI16] Peter J. Forrester and Jesper R. Ipsen. Real eigenvalue statistics for products of asymmetric real Gaussian matrices. Linear Algebra and its Applications, 510(??): 259–290, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303524.

Fasi:2019:CPS

[FI19a] Massimiliano Fasi and Bruno Iannazzo. Computing primary solutions of equations involving primary matrix functions. Linear Algebra and its Applications, 560 (??):17–42, January 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304439.

Fonseca:2019:DPV

[FI19b] Claudia Cavalcante Fonseca and Kostiantyn Iusenko. On dimension of poset variety. Linear Algebra and its Applications, 568(??):155– 164, May 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303033.

Fiedler:2015:MEG

[Fie15a] Miroslav Fiedler. Majorization in Euclidean geometry and beyond. Linear Algebra and its Applications, 466(??):233–240, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006843.

Fiedler:2015:ONA

[Fie15b] Miroslav Fiedler. Old and new about positive definite matrices. Linear Algebra and its Applications, 484(??): 496–503, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004279.

Fikioris:2018:SPK

[Fik18] George Fikioris. Spectral properties of Kac–Murdock– Szegő matrices with a complex parameter. Linear Algebra and its Applications, 553(??):182–210, September 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518302350.

Filipovski:2018:NEA

[Fil18] Slobodan Filipovski. On the non-existence of antipodal cages of even girth. Linear Algebra and its Applications, 546(??):261–273, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518300715.

Fiol:2016:QPG

[Fio16]  M. A. Fiol.  Quotient-

polynomial graphs. Linear Algebra and its Applications, 488(??):363–376, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005911.

Fischer:2017:AAM

[Fis17] Thomas M. Fischer. On the algorithm by Al-Mohy and Higham for computing the action of the matrix exponential: a posteriori roundoff error estimation. Linear Algebra and its Applications, 531(??): 141–168, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303427.

See [AMH11].

Farnik:2017:QPM

[FJ17] M. Farnik and Z. Jelonek. On quadratic polynomial mappings of the plane. Linear Algebra and its Applications, 529(??):441–456, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303087.

Filipovski:2019:NEF

[FJ19] Slobodan Filipovski and Robert Jajcay. On the non-existence of families of (d, k, δ)-digraphs containing only selfrepeat vertices. Linear Algebra and its Applications, 563(??):302–312, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305305.

Fickus:2018:ETF

[FJKM18] Matthew Fickus, John Jasper, Emily J. King, and Dustin G. Mixon. Equiangular tight frames that contain regular simplices. Linear Algebra and its Applications, 555(??):98–138, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830288X.

Farnik:2018:QPM

[FJM18] M. Farnik, Z. Jelonek, and P. Migus. On quadratic polynomial mappings from the plane into the n dimensional space. Linear Algebra and its Applications, 554(??): 249–274, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302581.

Fickus:2015:GTC

[FJMP15] Matthew Fickus, John Jasper, Dustin G. Mixon, and Jesse Peterson. Grouptheoretic constructions of erasure-robust frames. Linear Algebra and its Applications, 479(??):131–154, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002190.

Foulis:2015:PES

[FJP15a] David J. Foulis, Anna Jencová, and Sylvia Pulmannová. A projection and an effect in a synaptic algebra. Linear Algebra and its Applications, 485(??):417– 441, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500453X.

Foulis:2015:TPS

[FJP15b] David J. Foulis, Anna Jencová, and Sylvia Pulmannová. Two projections in a synaptic algebra. Linear Algebra and its Applications, 478(??):162–187, August 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001858.

Fallat:2017:TPS

[FJS17] Shaun Fallat, Charles R. Johnson, and Alan D. Sokal. Total positivity of sums, Hadamard products and Hadamard powers: Results and counterexamples. Linear Algebra and its Applications, 520 (??):242–259, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/

science/article/pii/S0024379517300411.

See corrigendum [FJS21].

Fallat:2021:CTP

[FJS21] Shaun Fallat, Charles R. Johnson, and Alan D. Sokal. Corrigendum to “Total positivity of sums, Hadamard products and Hadamard powers: Results and counterexamples” [Linear Algebra Appl. 520 (2017) 242– 259]. Linear Algebra and its Applications, 613(??): 393–396, March 15, 2021. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379520305863.

See [FJS17].

Fernandes:2017:CEG

[FJT17] Rafael Fernandes, Joaquim Judice, and Vilmar Trevisan. Complementary eigenvalues of graphs. Linear Algebra and its Applications, 527(??):216–231, August 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302094.

Forster:2016:PFT

[FK16] K.-H. Förster and P. Kallus. Perron–Frobenius theorems for the numerical range of semi-monic matrix polynomials. Linear Algebra and its Applications, 498(??):471– 489, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007533.

Filipiak:2018:AKP

[FK18] Katarzyna Filipiak and Daniel Klein. Approximation with a Kronecker product structure with one component as compound symmetry or autoregression. Linear Algebra and its Applications, 559(??): 11–33, December 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304208.

Fox:2015:CIA

[FKD15] Christopher J. Fox, Andreas Käufl, and Mathias Drton. On the causal interpretation of acyclic mixed graphs under multivariate normality. Linear Algebra and its Applications, 473 (??):93–113, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514000998.

Fritzsche:2017:SAEa

[FKM17a] Bernd Fritzsche, Bernd Kirstein, and Conrad Mädler. On a simultaneous approach to the even and odd truncated matricial Stieltjes moment problem I: an α-Schur– Stieltjes-type algorithm for sequences of complex matrices. Linear Algebra and its Applications, 521(??): 142–216, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300575.

Fritzsche:2017:SAEb

[FKM17b] Bernd Fritzsche, Bernd Kirstein, and Conrad Mädler. On a simultaneous approach to the even and odd truncated matricial Stieltjes moment problem II: an α-Schur–Stieltjes-type algorithm for sequences of holomorphic matrix-valued functions. Linear Algebra and its Applications, 520(??):335– 398, May 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300599.

Fritzsche:2018:TMS

[FKMS18] Bernd Fritzsche, Bernd Kirstein, Conrad Mädler, and Torsten Schröder. On the truncated matricial Stieltjes moment problem M[[α, ∞); (s_{j})_{j=0}^{m}, ≤]. Linear Algebra and its Applications, 544(??):30–114, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518300107.

Futorny:2018:WPC

[FKPS18] Vyacheslav Futorny, Tetiana Klymchuk, Anatolii P. Petravchuk, and Vladimir V. Sergeichuk. Wildness of the problems of classifying two-dimensional spaces of commuting linear operators and certain Lie algebras. Linear Algebra and its Applications, 536(??): 201–209, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305438.

Fritzsche:2017:SPE

[FKRS17] Bernd Fritzsche, Bernd Kirstein, Inna Roitberg, and Alexander Sakhnovich. Stability of the procedure of explicit recovery of skewselfadjoint Dirac systems from rational Weyl matrix functions. Linear Algebra and its Applications, 533(??):428–450, November 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517304640.

Futorny:2016:RSC

[FKS16] Vyacheslav Futorny, Tetiana Klymchuk, and Vladimir V. Sergeichuk. Roth’s solvability criteria for the matrix equations AX − X̄B = C and X − AX̄B = C over the skew field of quaternions with an involutive automorphism q ↦ q̂. Linear Algebra and its Applications, 510(??):246–258, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303597.

Forbes:2015:LEE

[FL15] Peter G. M. Forbes and Steffen Lauritzen. Linear estimating equations for exponential families with application to Gaussian linear concentration models. Linear Algebra and its Applications, 473(??):261–283, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379514005485.

Friedland:2016:EPQ

[FL16] Shmuel Friedland and Raphael Loewy. On the extreme points of quantum channels. Linear Algebra and its Applications, 498(??):553– 573, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000914.

Flinth:2016:GSC

[Fli16] Axel Flinth. A geometrical stability condition for compressed sensing. Linear Algebra and its Applications, 504(??):406–432, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301161.

Farenick:2018:PPI

[FLVV18] D. Farenick, B. Lemmens, M. Van Barel, and R. Vandebril. Preface to the proceedings of the 20th ILAS meeting, Leuven, Belgium, 2016. Linear Algebra and its Applications, 542(??): 1–3, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305396.

Fika:2016:EBF

[FM16] Paraskevi Fika and Marilena Mitrouli. Estimation of the bilinear form y^{∗}f(A)x for Hermitian matrices. Linear Algebra and its Applications, 502(??): 140–158, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500511X.

Fenzi:2017:RSO

[FM17] Luca Fenzi and Wim Michiels. Robust stability optimization for linear delay systems in a probabilistic framework. Linear Algebra and its Applications, 526(??):1– 26, August 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301829.

Ferrari:2019:DSH

[FM19a] Margherita Maria Ferrari and Emanuele Munarini. Decomposition of some Hankel matrices generated by the generalized rencontres polynomials. Linear Algebra and its Applications, 567 (??):180–201, April 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300011.

Fikioris:2019:DBE

[FM19b] George Fikioris and Themistoklis K. Mavrogordatos. Double, borderline, and extraordinary eigenvalues of Kac–Murdock–Szegő matrices with a complex parameter. Linear Algebra and its Applications, 575(??): 314–333, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301697.

Ferrer:2016:CDH

[FMM16] Josep Ferrer, David Mingueza, and M. Eulàlia Montoro. Counting d-dimensional hyperinvariant subspaces of a Weyr matrix. Linear Algebra and its Applications, 493 (??):14–27, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007004.

Fallat:2018:QPR

[FMR18] Shaun M. Fallat and Xavier Martínez-Rivera. The quasi principal rank characteristic sequence. Linear Algebra and its Applications, 548(??):42–56, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300995.

Feliu:2019:CIC

[FMR19] E. Feliu, S. Müller, and G. Regensburger. Characterizing injectivity of classes of maps via classes of matrices. Linear Algebra and its Applications, 580(??): 236–261, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Fallat:2016:CPS

[FMY16] Shaun Fallat, Karen Meagher, and Boting Yang. On the complexity of the positive semidefinite zero forcing number. Linear Algebra and its Applications, 491(??): 101–122, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500155X.

Funk-Neubauer:2017:BT

[FN17] Darren Funk-Neubauer. Bidiagonal triples. Linear Algebra and its Applications, 521 (??):104–134, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300551.

Ferreira:2019:SQC

[FNX19] O. P. Ferreira, S. Z. Németh, and L. Xiao. On the spherical quasi-convexity of quadratic functions. Linear Algebra and its Applications, 562(??):205–222, February 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304919.

Fallat:2016:EPR

[FOvdD16] Shaun M. Fallat, Dale D. Olesky, and Pauline van den Driessche. The enhanced principal rank characteristic sequence for skew-symmetric matrices. Linear Algebra and its Applications, 498 (??):366–377, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004796.

Fernandez-Polo:2017:TPS

[FPGPV17] Francisco J. Fernández-Polo, Jorge J. Garcés, Antonio M. Peralta, and Ignacio Villanueva. Tingley’s problem for spaces of trace class operators. Linear Algebra and its Applications, 529(??):294– 323, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302616.

Fuhr:2015:BVU

[FR15] Hartmut Führ and Ziemowit Rzeszotnik. On biunimodular vectors for unitary matrices. Linear Algebra and its Applications, 484(??): 86–129, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003729.

Fassbender:2016:CFS

[FR16] H. Faßbender and M. Rozlozník. On the conditioning of factors in the SR decomposition. Linear Algebra and its Applications, 505(??):224–244, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301495.

Fuhr:2018:NFU

[FR18] Hartmut Führ and Ziemowit Rzeszotnik. A note on factoring unitary matrices. Linear Algebra and its Applications, 547(??):32– 44, June 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300843.

Frederick:2018:SEP

[Fre18] Christina Frederick. An L^{2}- stability estimate for periodic nonuniform sampling in higher dimensions. Linear Algebra and its Applications, 555(??):361–372, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302982.

Fassbender:2017:VSL

[FS17a] Heike Faßbender and Philip Saltenberger. On vector spaces of linearizations for matrix polynomials in orthogonal bases. Linear Algebra and its Applications, 525(??):59–83, July 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301799.

Fawzi:2017:LCT

[FS17b] Hamza Fawzi and James Saunderson. Lieb’s concavity theorem, matrix geometric means, and semidefinite optimization. Linear Algebra and its Applications, 513(??):240–263, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304852.

Fassbender:2018:BKA

[FS18a] Heike Faßbender and Philip Saltenberger. Block Kronecker ansatz spaces for matrix polynomials. Linear Algebra and its Applications, 542(??):118–148, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301817.

Fujii:2018:ROE

[FS18b] Jun Ichi Fujii and Yuki Seo. The relative operator entropy and the Karcher mean. Linear Algebra and its Applications, 542(??): 4–34, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305444.

Foucart:2019:IHT

[FS19a] Simon Foucart and Srinivas Subramanian. Iterative hard thresholding for lowrank recovery from rank-one projections. Linear Algebra and its Applications, 572 (??):117–134, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301028.

Fujimoto:2019:SIO

[FS19b] Masayuki Fujimoto and Yuki Seo. The Schwarz inequality via operator-valued inner product and the geometric operator mean. Linear Algebra and its Applications, 561(??):141–160, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830466X.

Francis:2017:EAE

[FSSW17] Amanda Francis, Dallas Smith, Derek Sorensen, and Benjamin Webb. Extensions and applications of equitable decompositions for graphs with symmetries. Linear Algebra and its Applications, 532(??):432–462, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304093.

Francis:2019:GED

[FSW19] Amanda Francis, Dallas Smith, and Benjamin Webb. General equitable decompositions for graphs with symmetries. Linear Algebra and its Applications, 577(??):287–316, September 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519301934.

Fasino:2016:GMM

[FT16] Dario Fasino and Francesco Tudisco. Generalized modularity matrices. Linear Algebra and its Applications, 502(??):327–345, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003663.

Fasino:2018:MBS

[FT18] Dario Fasino and Francesco Tudisco. A modularity based spectral method for simultaneous community and anti-community detection. Linear Algebra and its Applications, 542 (??):605–623, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306638.

Fan:2016:LES

[FuIKT16] Yi-Zheng Fan, Murad ul Islam Khan, and Ying-Ying Tan. The largest Heigenvalue and spectral radius of Laplacian tensor of non-odd-bipartite generalized power hypergraphs. Linear Algebra and its Applications, 504(??):487–502, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300921.

Fan:2017:BPN

[FW17] Yi-Zheng Fan and Long Wang. Bounds for the positive and negative inertia index of a graph. Linear Algebra and its Applications, 522(??):15–27, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300812.

Fan:2019:ELS

[FWB^{+}19] Yi-Zheng Fan, Yi Wang, Yan-Hong Bao, Jiang-Chao Wan, Min Li, and Zhu Zhu. Eigenvectors of Laplacian or signless Laplacian of hypergraphs associated with zero eigenvalue. Linear Algebra and its Applications, 579(??):244–261, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302514.

Fu:2019:RPI

[FX19] Chunhong Fu and Qingxiang Xu. A remark on the partial isometry associated to the generalized polar decomposition of a matrix. Linear Algebra and its Applications, 574(??):30– 39, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301326.

Feng:2017:SCS

[FZL^{+}17] Lihua Feng, Pengli Zhang, Henry Liu, Weijun Liu, Minmin Liu, and Yuqin Hu. Spectral conditions for some graphical properties. Linear Algebra and its Applications, 524(??):182– 198, July 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301581.

Fan:2016:QLQ

[FZwCW16] H.-Y. Fan, L. Zhang, E. K. w. Chu, and Y. Wei. Qless QR decomposition in inner product spaces. Linear Algebra and its Applications, 491(??):292–316, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005133.

Gaetz:2016:CGG

[Gae16] Christian Gaetz. Critical groups of group representations. Linear Algebra and its Applications, 508(??): 91–99, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302634.

Galazka:2017:VBG

[Gal17] Maciej Galazka. Vector bundles give equations of cactus varieties. Linear Algebra and its Applications, 521 (??):254–262, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305845.

Galuppi:2019:RVV

[Gal19] Francesco Galuppi. The Rough Veronese variety. Linear Algebra and its Applications, 583(??):282–299, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303775.

Gao:2018:ZNP

[Gao18] Wei Gao. Zero-nonzero patterns of order n ≥ 4 do not require H_{n}^{∗}. Linear Algebra and its Applications, 554 (??):1–14, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302544.

Ganie:2016:SLE

[GAP16] Hilal A. Ganie, Ahmad M. Alghamdi, and S. Pirzada. On the sum of the Laplacian eigenvalues of a graph and Brouwer’s conjecture. Linear Algebra and its Applications, 501(??):376–389, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516300453.

See corrigendum [GP18].

Garoni:2017:TFA

[Gar17] Carlo Garoni. Topological foundations of an asymptotic approximation theory for sequences of matrices with increasing size. Linear Algebra and its Applications, 513(??):324–341, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304943.

Gasiorek:2019:CTC

[Gas19] Marcin Gasiorek. A Coxeter type classification of one-peak principal posets. Linear Algebra and its Applications, 582(??):197–217, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303350.

Griewank:2015:SPL

[GBRS15] Andreas Griewank, Jens-Uwe Bernt, Manuel Radons, and Tom Streubel. Solving piecewise linear systems in abs-normal form. Linear Algebra and its Applications, 471(??):500–530, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008209.

Gokulraj:2019:SLG

[GC19] S. Gokulraj and A. Chandrashekaran. On symmetric linear games. Linear Algebra and its Applications, 562(??):44–54, February 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304786.

Guan:2018:CAS

[GCC18] Yu Guan, Moody T. Chu, and Delin Chu. Convergence analysis of an SVDbased algorithm for the best rank-1 tensor approximation. Linear Algebra and its Applications, 555(??): 53–69, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302908.

Ganie:2018:SLE

[GCP18] Hilal A. Ganie, Bilal A. Chat, and S. Pirzada. Signless Laplacian energy of a graph and energy of a line graph. Linear Algebra and its Applications, 544 (??):306–324, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300351.

Guo:2015:PAE

[GCQX15] Pei-Chang Guo, Yun-Feng Cai, Jiang Qian, and Shu-Fang Xu. Perturbation analysis of the extinction probability of a Markovian binary tree. Linear Algebra and its Applications, 475 (??):11–27, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000889.

Gribling:2017:MHC

[GdLL17] Sander Gribling, David de Laat, and Monique Laurent. Matrices with high completely positive semidefinite rank. Linear Algebra and its Applications, 513(??):122–148, January 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516304888.

Geher:2015:CAC

[Geh15] György Pál Gehér. A contribution to the Aleksandrov conservative distance problem in two dimensions. Linear Algebra and its Applications, 481(??):280–287, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002943.

Gnang:2017:SHD

[GF17] Edinah K. Gnang and Yuval Filmus. On the spectra of hypermatrix direct sum and Kronecker products constructions. Linear Algebra and its Applications, 519 (??):238–277, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300162.

Gaal:2018:IGS

[GG18] Marcell Gaál and Robert M. Guralnick. On isometry groups of self-adjoint traceless and skew-symmetric matrices. Linear Algebra and its Applications, 541(??): 211–220, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306778.

Gao:2015:TAG

[GGH15] Xiaojuan Gao, Suogang Gao, and Bo Hou. The Terwilliger algebras of Grassmann graphs. Linear Algebra and its Applications, 471 (??):427–448, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000415.

Garza:2015:MCS

[GGMPC15] Lino G. Garza, Luis E. Garza, Francisco Marcellán, and Natalia C. Pinzón-Cortés. A matrix characterization for the D_{ν}-semiclassical and D_{ν}- coherent orthogonal polynomials. Linear Algebra and its Applications, 487(??):242– 259, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005297.

Gupta:2015:CSV

[GH15] Chander K. Gupta and Waldemar Holubowski. Commutator subgroup of Vershik– Kerov group II. Linear Algebra and its Applications, 471 (??):85–95, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000294.

Gautier:2016:TNM

[GH16a] Antoine Gautier and Matthias Hein. Tensor norm and maximal singular vectors of nonnegative tensors — a Perron–Frobenius theorem, a Collatz–Wielandt characterization and a generalized power method. Linear Algebra and its Applications, 505(??):313–343, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301239.

Gerard:2016:HOL

[GH16b] David Gerard and Peter Hoff. A higher-order LQ decomposition for separable covariance models. Linear Algebra and its Applications, 505(??):57–84, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630132X.

Georgieva:2017:BUA

[GH17] I. Georgieva and C. Hofreither. On best uniform approximation by low-rank matrices. Linear Algebra and its Applications, 518 (??):159–176, April 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306334.

Gao:2019:SRG

[GH19] Jun Gao and Xinmin Hou. The spectral radius of graphs without long cycles. Linear Algebra and its Applications, 566(??):17–33, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305949.

Gumus:2017:NII

[GHK17] Ibrahim Halil Gumus, Omar Hirzallah, and Fuad Kittaneh. Norm inequalities involving accretive-dissipative 2 × 2 block matrices. Linear Algebra and its Applications, 528(??):76–93, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301148.

Ghorbani:2016:PCP

[Gho16] Ebrahim Ghorbani. Proof of a conjecture on ‘plateaux’ phenomenon of graph Laplacian eigenvalues. Linear Algebra and its Applications, 506(??):274–278, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302063.

Ghorbani:2019:EFI

[Gho19] Ebrahim Ghorbani. Eigenvaluefree interval for threshold graphs. Linear Algebra and its Applications, 583(??): 300–305, December 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303751.

Guttel:2019:PPI

[GHPS19] Stefan Güttel, Leslie Hogben, Miklós Pálfia, and Tatjana Stykel. Preface to the Proceedings of ILAS 2017: Connections, Twenty-first Conference of the International Linear Algebra Society, Iowa State University, 24–28 July 2017. Linear Algebra and its Applications, 576(??):1–4, September 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518305445.

Gilbert:2016:DSG

[GHS16] Anna C. Gilbert, Jeremy G. Hoskins, and John C. Schotland. Diffuse scattering on graphs. Linear Algebra and its Applications, 496(??):1–35, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000148.

Guiver:2016:NEP

[GHT16] Chris Guiver, Dave Hodgson, and Stuart Townley. A note on the eigenvectors of perturbed matrices with applications to linear positive systems. Linear Algebra and its Applications, 509(??):143–167, November 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302725.

Gil:2016:BST

[Gil16] Michael Gil’. Bounds for the spectrum of a two parameter matrix eigenvalue problem. Linear Algebra and its Applications, 498 (??):201–218, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002979.

Giambruno:2016:SPM

[GIM16] Antonio Giambruno, Antonio Ioppolo, and Fabrizio Martino. Standard polynomials and matrices with superinvolutions. Linear Algebra and its Applications, 504(??):272–291, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630115X.

Girard:2019:DDT

[Gir19] Mark W. Girard. On directional derivatives of trace functionals of the form A → Tr(P f(A)). Linear Algebra and its Applications, 569 (??):62–77, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930028X.

Gowda:2018:CSS

[GJ18] M. Seetharama Gowda and Juyoung Jeong. On the connectedness of spectral sets and irreducibility of spectral cones in Euclidean Jordan algebras. Linear Algebra and its Applications, 559(??): 181–193, December 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304336.

Guterman:2018:LIP

[GJK18] Alexander Guterman, Marianne Johnson, and Mark Kambites. Linear isomorphisms preserving Green’s relations for matrices over anti-negative semifields. Linear Algebra and its Applications, 545(??):1–14, May 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518300375.

Gu:2016:SCR

[GJLS16] Jiao Gu, Jürgen Jost, Shiping Liu, and Peter F. Stadler. Spectral classes of regular, random, and empirical graphs. Linear Algebra and its Applications, 489(??):30–49, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500587X.

Gavrilyuk:2015:TPQ

[GK15] Alexander L. Gavrilyuk and Jack H. Koolen. The Terwilliger polynomial of a Qpolynomial distance-regular graph and its application to pseudo-partition graphs. Linear Algebra and its Applications, 466(??):117–140, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006454.

Goel:2016:CLA

[GKR16] Charu Goel, Salma Kuhlmann, and Bruce Reznick. On the Choi–Lam analogue of Hilbert’s 1888 theorem for symmetric forms. Linear Algebra and its Applications, 496(??):114–120, May 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516000276.

Greenbaum:2018:CRD

[GKS18] Anne Greenbaum, Faranges Kyanfar, and Abbas Salemi. On the convergence rate of DGMRES. Linear Algebra and its Applications, 552(??):219–238, September 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518302210.

Gillis:2019:ANS

[GKS19] Nicolas Gillis, Michael Karow, and Punit Sharma. Approximating the nearest stable discrete-time system. Linear Algebra and its Applications, 573(??):37–53, July 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301156.

Guo:2015:SATb

[GL15a] Jun Guo and Fenggao Li. Several Anzahl theorems of alternate matrices over Galois rings. Linear Algebra and its Applications, 474 (??):169–183, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001226.

Guo:2015:SATa

[GL15b] Jun Guo and Fenggao Li. Several Anzahl theorems of matrices over Galois rings and their applications. Linear Algebra and its Applications, 465(??):296–311, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006338.

Guo:2016:ARE

[GL16] Chun-Hua Guo and Di Lu. On algebraic Riccati equations associated with regular singular M-matrices. Linear Algebra and its Applications, 493(??):108–119, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006989.

Gao:2017:PSP

[GL17] Wei Gao and Zhongshan Li. Path sign patterns of order n ≥= 5 do not require H_{n}. Linear Algebra and its Applications, 532(??): 99–126, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304032.

Gau:2019:ENRa

[GL19] Hwa-Long Gau and Yueh-Hua Lu. Extremality of numerical radii of tensor products of matrices. Linear Algebra and its Applications, 565(??):82–98, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305718.

Garcia:2019:JCF

[GLAdS19] Esther García, Miguel Gómez Lozano, Rubén Muñoz Alcázar, and Guillermo Vera de Salas. A Jordan canonical form for nilpotent elements in an arbitrary ring. Linear Algebra and its Applications, 581(??):324–335, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302976.

Garcia-Lopez:2018:SPM

[GLMPC18] J. García-López, C. Marijuán, and L. M. Pozo-Coronado. Structural properties of minimal strong digraphs versus trees. Linear Algebra and its Applications, 540(??):203–220, March 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306559.

Guterman:2018:RPC

[GLMS18] Alexander Guterman, Thomas Laffey, Olga Markova, and Helena Smigoc. A resolution of Paz’s conjecture in the presence of a nonderogatory matrix. Linear Algebra and its Applications, 543 (??):234–250, April 15, 2018.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300089.

Gumus:2018:PSD

[GLRT18] Mehmet Gumus, Jianzhen Liu, Samir Raouafi, and Tin-Yau Tam. Positive semidefinite 2 × 2 block matrices and norm inequalities. Linear Algebra and its Applications, 551(??): 83–91, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301770.

Gong:2017:HMT

[GLW17] Chao Gong, Benjian Lv, and Kaishun Wang. The Hilton–Milner theorem for the distance-regular graphs of bilinear forms. Linear Algebra and its Applications, 515(??):130–144, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305420.

Gao:2016:CSS

[GLZ16a] Wei Gao, Zhongshan Li, and Lihua Zhang. Characterization of star sign patterns that require H_{n}. Linear Algebra and its Applications, 499(??):43–65, June 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001622.

Gao:2016:SPR

[GLZ16b] Wei Gao, Zhongshan Li, and Lihua Zhang. Sign patterns that require H_{n} exist for each n ≥ 4. Linear Algebra and its Applications, 489(??):15–23, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005881.

Gao:2015:EDL

[GM15] Chenghua Gao and Ruyun Ma. Eigenvalues of discrete linear second-order periodic and antiperiodic eigenvalue problems with sign-changing weight. Linear Algebra and its Applications, 467(??): 40–56, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007265.

Gao:2016:EDS

[GM16a] Chenghua Gao and Ruyun Ma. Eigenvalues of discrete Sturm–Liouville problems with eigenparameter dependent boundary conditions. Linear Algebra and its Applications, 503(??): 100–119, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300660.

Ginzberg:2016:QSM

[GM16b] Paul Ginzberg and Christiana Mavroyiakoumou. The QRD and SVD of matrices over a real algebra. Linear Algebra and its Applications, 504(??):27–47, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300386.

Grover:2017:PBM

[GM17] Priyanka Grover and Pradip Mishra. Perturbation bounds for Mostow’s decomposition and the bipolar decomposition. Linear Algebra and its Applications, 529(??):12–29, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302070.

Glassett:2019:SAP

[GM19] Jillian L. Glassett and Judith J. McDonald. Spectrally arbitrary patterns over rings with unity. Linear Algebra and its Applications, 576(??):228–245, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830418X.

Goberna:2016:VIM

[GMLdS16] M. A. Goberna, J. E. Martínez-Legaz, and V. N. Vera de Serio. The Voronoi inverse mapping. Linear Algebra and its Applications, 504(??):248–271, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300891.

Goberna:2019:FVC

[GMLT19] M. A. Goberna, J. E. Martínez-Legaz, and M. I. Todorov. On farthest Voronoi cells. Linear Algebra and its Applications, 583(??):306–322, December 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519303787.

Guterman:2016:LQC

[GMM16] A. E. Guterman, O. V. Markova, and V. Mehrmann. Lengths of quasi-commutative pairs of matrices. Linear Algebra and its Applications, 498(??):450–470, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515007028.

Guterman:2019:LRP

[GMM19] A. E. Guterman, O. V. Markova, and V. Mehrmann. Length realizability for pairs of quasi-commuting matrices. Linear Algebra and its Applications, 568(??):135– 154, May 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303045.

Greaves:2019:IFH

[GMO19] Gary Greaves, Bojan Mohar, and Suil O. Interlacing families and the Hermitian spectral norm of digraphs. Linear Algebra and its Applications, 564(??): 201–208, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305597.

Granario:2015:IPD

[GMP15] Daryl Q. Granario, Dennis I. Merino, and Agnes T. Paras. The ϕ_{S} polar decomposition when the cosquare of S is normal. Linear Algebra and its Applications, 467(??): 75–85, February 15, 2015.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007332.

Giribet:2016:ONP

[GMP16a] Juan I. Giribet, Alejandra Maestripieri, and Francisco Martínez Pería. Optimal normal projections in Krein spaces. Linear Algebra and its Applications, 490(??):77–101, February 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006102.

Granario:2016:PDW

[GMP16b] Daryl Q. Granario, Dennis I. Merino, and Agnes T. Paras. The ϕ_{S} polar decomposition when the cosquare of S is normal. Linear Algebra and its Applications, 495 (??):51–66, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000288.

Granario:2016:STO

[GMP16c] Daryl Q. Granario, Dennis I. Merino, and Agnes T. Paras. The sum of two ϕ_{S} orthogonal matrices when S^{T} S is normal and −1 ∉ σ(S^{−T} S). Linear Algebra and its Applications, 495 (??):67–89, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000264.

Guglielmi:2018:BNL

[GMW18] Nicola Guglielmi, Oliver Mason, and Fabian Wirth. Barabanov norms, Lipschitz continuity and monotonicity for the max algebraic joint spectral radius. Linear Algebra and its Applications, 550 (??):37–58, August 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301526.

Greenbaum:2018:NIC

[GO18] Anne Greenbaum and Michael L. Overton. Numerical investigation of Crouzeix’s conjecture. Linear Algebra and its Applications, 542 (??):225–245, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302847.

Goldberg:2016:RMA

[Gol16] Moshe Goldberg. The radius in matrix algebras — examples and remarks. Linear Algebra and its Applications, 492(??):136–143, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006199.

Goldberg:2017:CSF

[Gol17] Moshe Goldberg. Continuity of seminorms on finitedimensional vector spaces. Linear Algebra and its Applications, 515(??):175–179, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305298.

Gonzalez:2015:DQA

[Gon15] Jesús Arturo Jiménez González. Differential quiver algebras and tree modules. Linear Algebra and its Applications, 485(??):267–288, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004486.

Gora:2017:SMC

[Gór17] Michal Góra. Some methods of construction of a common Lyapunov solution to a finite set of complex systems. Linear Algebra and its Applications, 530 (??):77–93, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Gowda:2016:CML

[Gow16] M. Seetharama Gowda. Completely mixed linear games on a self-dual cone. Linear Algebra and its Applications, 498(??):219–230, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515003444.

Gowda:2017:PDS

[Gow17] M. Seetharama Gowda. Positive and doubly stochastic maps, and majorization in Euclidean Jordan algebras. Linear Algebra and its Applications, 528(??): 40–61, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001269.

Ganie:2018:CSS

[GP18] Hilal A. Ganie and S. Pirzada. Corrigendum to “On the sum of the Laplacian eigenvalues of a graph and Brouwer’s conjecture” [Linear Algebra Appl. 501 (2016) 376–389]. Linear Algebra and its Applications, 538(??):228–230, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306092.

See [GAP16].

Gernandt:2019:SRT

[GP19a] H. Gernandt and J. P. Pade. Schur reduction of trees and extremal entries of the Fiedler vector. Linear Algebra and its Applications, 570 (??):93–122, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300631.

Goel:2019:CGI

[GP19b] Gopal Goel and David Perkinson. Critical groups of iterated cones. Linear Algebra and its Applications, 567 (??):138–142, April 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300254.

Grishkov:2018:LCC

[GPI18] A. Grishkov and J. M. Pérez-Izquierdo. Lie’s correspondence for commutative automorphic formal loops. Linear Algebra and its Applications, 544(??):460– 501, May 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300429.

Greif:2015:PIC

[GPW15] Chen Greif, Panayiotis J. Psarrakos, and Hugo J. Woerdeman. Preface to the 18th ILAS Conference Proceedings, Providence, 2013. Linear Algebra and its Applications, 468(??):1– 2, March 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008039.

Gowda:2015:GTV

[GR15] M. Seetharama Gowda and G. Ravindran. On the game-theoretic value of a linear transformation relative to a self-dual cone. Linear Algebra and its Applications, 469(??):440–463, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007678.

Gorla:2016:ESC

[GR16] Elisa Gorla and Alberto Ravagnani. Equidistant subspace codes. Linear Algebra and its Applications, 490(??):48–65, February 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515006448.

Grcar:2018:MLR

[Grc18] Joseph F. Grcar. Monotonic linear recurrences. Linear Algebra and its Applications, 537(??):12–21, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305621.

Greaves:2018:ELS

[Gre18] Gary R. W. Greaves. Equiangular line systems and switching classes containing regular graphs. Linear Algebra and its Applications, 536 (??):31–51, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305323.

Gutman:2015:UBR

[GRS15] Ivan Gutman, María Robbiano, and Bernardo San Martín. Upper bound on Randić energy of some graphs. Linear Algebra and its Applications, 478(??): 241–255, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002153.

Guo:2017:MLS

[GRS17] Ji-Ming Guo, Ji-Yun Ren, and Jin-Song Shi. Maximizing the least signless Laplacian eigenvalue of unicyclic graphs. Linear Algebra and its Applications, 519 (??):136–145, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300113.

Goodall:2016:MIC

[GRV16] Andrew Goodall, Guus Regts, and Lluís Vena. Matroid invariants and counting graph homomorphisms. Linear Algebra and its Applications, 494(??):263–273, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516000252.

Gu:2015:SRH

[GS15] Caixing Gu and Mark Stankus. Some results on higher order isometries and symmetries: Products and sums with a nilpotent operator. Linear Algebra and its Applications, 469(??): 500–509, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006417.

Guterman:2016:RFT

[GS16] Alexander Guterman and Yaroslav Shitov. Rank functions of tropical matrices. Linear Algebra and its Applications, 498(??):326– 348, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004048.

Garcia:2017:MSP

[GS17] Stephan Ramon Garcia and David Sherman. Matrices similar to partial isometries. Linear Algebra and its Applications, 526(??):35– 41, August 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301787.

Gillis:2018:SAA

[GS18] Nicolas Gillis and Punit Sharma. A semi-analytical approach for the positive semidefinite Procrustes problem. Linear Algebra and its Applications, 540 (??):112–137, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306511.

Gao:2019:MAE

[GS19a] Yubin Gao and Yanling Shao. The minimum ABC energy of trees. Linear Algebra and its Applications, 577(??):186–203, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301909.

Gillis:2019:LRM

[GS19b] Nicolas Gillis and Yaroslav Shitov. Low-rank matrix approximation in the infinity norm. Linear Algebra and its Applications, 581(??):367– 382, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302988.

Ghosh:2016:OBL

[GSP16] Puja Ghosh, Debmalya Sain, and Kallol Paul. Orthogonality of bounded linear operators. Linear Algebra and its Applications, 500 (??):43–51, July 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001683.

Gajula:2015:TDT

[GSS15] Sailaja Gajula, Ivan Soprunov, and Jenya Soprunova. Tropical determinant on transportation polytopes. Linear Algebra and its Applications, 475 (??):28–44, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000658.

Garnett:2016:CFG

[GSSvdD16] C. Garnett, B. L. Shader, C. L. Shader, and P. van den Driessche. Characterization of a family of generalized companion matrices. Linear Algebra and its Applications, 498(??):360– 365, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004450.

Garcia:2016:SBN

[GSW16] Stephan Ramon Garcia, David Sherman, and Gary Weiss. On the similarity of AB and BA for normal and other matrices. Linear Algebra and its Applications, 508(??): 14–21, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302610.

Gasiorek:2015:SCD

[GSZ15] Marcin Gasiorek, Daniel Simson, and Katarzyna Zajac. Structure and a Coxeter-Dynkin type classification of corank two non-negative posets. Linear Algebra and its Applications, 469(??): 76–113, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007277.

Gasiorek:2016:GCN

[GSZ16] Marcin Gasiorek, Daniel Simson, and Katarzyna Zajac. A Gram classification of non-negative coranktwo loop-free edge-bipartite graphs. Linear Algebra and its Applications, 500 (??):88–118, July 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951600166X.

Gongopadhyay:2016:DCH

[GT16] Krishnendu Gongopadhyay and Cigole Thomas. Decomposition of complex hyperbolic isometries by involutions. Linear Algebra and its Applications, 500 (??):63–76, July 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001701.

Gonzalez-Torres:2017:GSC

[GT17] Raúl E. González-Torres. A geometric study of cores of idempotent stochastic matrices. Linear Algebra and its Applications, 527(??): 87–127, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302124.

Guterman:2018:VPM

[GT18] A. E. Guterman and K. A. Taranin. On the values of the permanent of (0, 1)- matrices. Linear Algebra and its Applications, 552(??): 256–276, September 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302209.

Guo:2018:EEC

[GTL^{+}18] Ji-Ming Guo, Pan-Pan Tong, Jianxi Li, Wai Chee Shiu, and Zhi-Wen Wang. The effect on eigenvalues of connected graphs by adding edges. Linear Algebra and its Applications, 548(??):57– 65, July 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300739.

See correction [GWL^{+}18].

Gu:2016:SCE

[Gu16] Xiaofeng Gu. Spectral conditions for edge connectivity and packing spanning trees in multigraphs. Linear Algebra and its Applications, 493 (??):82–90, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Guan:2018:BEM

[Gua18] Yonghui Guan. Brill’s equations as a GL(V)-module. Linear Algebra and its Applications, 548(??):273–292, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518301010.

Guo:2019:ECR

[Guo19] Chun-Hua Guo. Explicit convergence regions of Newton’s method and Chebyshev’s method for the matrix p-th root. Linear Algebra and its Applications, 583(??):63–76, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930360X.

Gau:2016:ENR

[GW16a] Hwa-Long Gau and Pei Yuan Wu. Extremality of numerical radii of matrix products. Linear Algebra and its Applications, 501(??):17– 36, July 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300076.

Gau:2016:NRH

[GW16b] Hwa-Long Gau and Pei Yuan Wu. Numerical radius of Hadamard product of matrices. Linear Algebra and its Applications, 504(??): 292–308, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301124.

Gau:2017:ENR

[GW17a] Hwa-Long Gau and Kuo-Zhong Wang. Extremality of numerical radii of matrix commutators and Jordan products. Linear Algebra and its Applications, 533(??):258–281, November 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304445.

Guo:2017:OSE

[GW17b] Lifeng Guo and Ligong Wang. Optimum skew energy of a tournament. Linear Algebra and its Applications, 530(??):405–413, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303166.

Gau:2018:ETN

[GW18] Hwa-Long Gau and Pei Yuan Wu. Equality of three numerical radius inequalities. Linear Algebra and its Applications, 554(??): 51–67, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302611.

Grossmann:2019:PME

[GW19a] Ben W. Grossmann and Hugo J. Woerdeman. On the preservers of maximally entangled states. Linear Algebra and its Applications, 583(??):171–194, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303805.

Guillot:2019:TNG

[GW19b] Dominique Guillot and Jiaru Wu. Total nonnegativity of GCD matrices and kernels. Linear Algebra and its Applications, 578(??): 446–461, October 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302253.

Guo:2018:CSE

[GWL^{+}18] Ji-Ming Guo, Zhi-Wen Wang, Jianxi Li, Wai Chee Shiu, and Pan-Pan Tong. Corrigendum to “The effect on eigenvalues of connected graphs by adding edges” [Linear Algebra Appl. 548 (2018) 57–65]. Linear Algebra and its Applications, 555(??):434–436, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303525.

See [GTL^{+}18].

Gau:2016:NRR

[GWW16] Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Numerical ranges of row stochastic matrices. Linear Algebra and its Applications, 506(??):478–505, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302270.

Gau:2019:ENRb

[GWW19a] Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Equality of numerical ranges of matrix powers. Linear Algebra and its Applications, 577(??):95–110, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302228.

Gau:2019:ENR

[GWW19b] Hwa-Long Gau, Kuo-Zhong Wang, and Pei Yuan Wu. Equality of numerical ranges of matrix powers. Linear Algebra and its Applications, 578(??):95–110, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302228.

Gumus:2016:CDL

[GX16] Mehmet Gumus and Jianhong Xu. On common diagonal Lyapunov solutions. Linear Algebra and its Applications, 507(??): 32–50, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302075.

Gumus:2017:NCS

[GX17] Mehmet Gumus and Jianhong Xu. A new characterization of simultaneous Lyapunov diagonal stability via Hadamard products. Linear Algebra and its Applications, 531(??): 220–233, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730349X.

Gumus:2019:CSL

[GX19] Mehmet Gumus and Jianhong Xu. On common αscalar Lyapunov solutions. Linear Algebra and its Applications, 563(??):123–141, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305147.

Ge:2017:DTI

[GXYZ17] Jianquan Ge, Song Xu, Hangyu You, and Yi Zhou. DDVV-type inequality for Hermitian matrices. Linear Algebra and its Applications, 529(??):133–147, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302653.

Guo:2018:SRU

[GZ18] Haiyan Guo and Bo Zhou. On the spectral radius of uniform hypertrees. Linear Algebra and its Applications, 558(??):236–249, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303719.

Giuzzi:2019:IMC

[GZ19] Luca Giuzzi and Ferdinando Zullo. Identifiers for MRDcodes. Linear Algebra and its Applications, 575(??): 66–86, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301375.

Golnari:2019:MFT

[GZB19] Golshan Golnari, Zhi-Li Zhang, and Daniel Boley. Markov fundamental tensor and its applications to network analysis. Linear Algebra and its Applications, 564 (??):126–158, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305470.

Hegazi:2015:NEA

[HA15a] A. S. Hegazi and Hani Abdelwahab. Nilpotent evolution algebras over arbitrary fields. Linear Algebra and its Applications, 486(??): 345–360, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004838.

Heydari:2015:NRF

[HA15b] M. T. Heydari and A. Abdollahi. The numerical range of finite order elliptic automorphism composition operators. Linear Algebra and its Applications, 483(??):

128–138, October 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003420.

Hegazi:2016:CFD

[HA16] A. S. Hegazi and Hani Abdelwahab. Classification of five-dimensional nilpotent Jordan algebras. Linear Algebra and its Applications, 494(??):165–218, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516000173.

Hadjidimos:2017:EOR

[Had17] Apostolos Hadjidimos. On equivalence of optimal relaxed block iterative methods for the singular nonsymmetric saddle point problem. Linear Algebra and its Applications, 522(??):175– 202, June 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300642.

Hamedani:2019:EHH

[HAD19] Ali Memar Hamedani, Rahim Alizadeh, and Silvestru Sever Dragomir. An extension of Haagerup–Harpe inequality. Linear Algebra and its Applications, 577(??):69– 77, September 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930179X.

See corrigendum [Ali20].

Hegazi:2016:CDN

[HAM16] A. S. Hegazi, Hani Abdelwahab, and A. J. Calderon Martin. The classification of N-dimensional non-Lie Malcev algebras with (N − 4)-dimensional annihilator. Linear Algebra and its Applications, 505(??):32– 56, September 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301288.

Han:2018:STS

[Han18] Kangjin Han. On singularities of third secant varieties of Veronese embeddings. Linear Algebra and its Applications, 544(??):391– 406, May 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830020X.

Hanaki:2019:DCA

[Han19a] Akihide Hanaki. Double centralizers of association schemes. Linear Algebra and its Applications, 564 (??):1–14, March 1, 2019.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830541X.

Hansen:2019:CMO

[Han19b] Frank Hansen. Convex multivariate operator means. Linear Algebra and its Applications, 564(??):209–224, March 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830555X.

He:2016:TSC

[HAWM16] Zhuo-Heng He, Oscar Mauricio Agudelo, Qing-Wen Wang, and Bart De Moor. Two-sided coupled generalized Sylvester matrix equations solving using a simultaneous decomposition for fifteen matrices. Linear Algebra and its Applications, 496 (??):549–593, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001142.

Hayashi:2019:NIP

[Hay19] Tomohiro Hayashi. On a norm inequality for a positive block-matrix. Linear Algebra and its Applications, 566(??):86–97, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Hou:2015:IDMa

[HC15] Yaoping Hou and Jing Chen. Inverse of the distance matrix of a cactoid digraph. Linear Algebra and its Applications, 475 (??):1–10, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000713.

Huang:2019:UED

[HC19] Chien-Hao Huang and Jein-Shan Chen. On unitary elements defined on Lorentz cone and their applications. Linear Algebra and its Applications, 565(??):1– 24, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305688.

Han:2017:MAI

[HCS17] Jianzhi Han, Qiufan Chen, and Yucai Su. Modules over the algebra Vir(a, b). Linear Algebra and its Applications, 515(??):11–23, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305183.

He:2015:MCR

[He15] Tian-Xiao He. Matrix characterizations of Riordan arrays. Linear Algebra and its Applications, 465(??): 15–42, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951400593X.

He:2016:ARM

[He16a] Tian-Xiao He. Applications of Riordan matrix functions to Bernoulli and Euler polynomials. Linear Algebra and its Applications, 507(??): 208–228, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302312.

He:2016:SOD

[He16b] Tian-Xiao He. Shift operators defined in the Riordan group and their applications. Linear Algebra and its Applications, 496 (??):331–350, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951600094X.

He:2018:SCD

[He18] Tian-Xiao He. Sequence characterizations of double Riordan arrays and their compressions. Linear Algebra and its Applications, 549 (??):176–202, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301393.

Herbrich:2016:LGT

[Her16] Peter Herbrich. Line graphs and the transplantation method. Linear Algebra and its Applications, 507(??): 158–168, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301835.

Heysse:2017:CDC

[Hey17] Kristin Heysse. A construction of distance cospectral graphs. Linear Algebra and its Applications, 535(??): 195–212, December 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730527X.

Hardy:2018:LMP

[HF18] Yorick Hardy and Ajda Fosner. Linear maps preserving Kronecker quotients. Linear Algebra and its Applications, 556(??):200–209, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303471.

Hou:2015:IDMb

[HFS15] Yaoping Hou, Aixiang Fang, and Yajing Sun. Inverse of the distance matrix of a cycle-clique graph. Linear Algebra and its Applications, 485(??):33–46, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500436X.

Huang:2017:RGF

[HH17] Xueyi Huang and Qiongxiang Huang. On regular graphs with four distinct eigenvalues. Linear Algebra and its Applications, 512(??):219–233, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304608.

Hawkins:2015:FIO

[HHJM15] Kyle Hawkins, Ursula Hebert-Johnson, and Ben Mathes. The Fibonacci identities of orthogonality. Linear Algebra and its Applications, 475 (??):80–89, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000531.

Huang:2015:CGE

[HHL15] Xueyi Huang, Qiongxiang Huang, and Lu Lu. Construction of graphs with exactly k main eigenvalues. Linear Algebra and its Applications, 486(??):204–218, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004784.

He:2019:BMN

[HHL19] Shengjie He, Rong-Xia Hao, and Hong-Jian Lai. Bounds for the matching number and cyclomatic number of a signed graph in terms of rank. Linear Algebra and its Applications, 572 (??):273–291, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301132.

Herden:2016:EWB

[HHR16] Daniel Herden, Michal Hrbek, and Pavel Ruzicka. On the existence of weak bases for vector spaces. Linear Algebra and its Applications, 501(??):98– 111, July 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001592.

Huttenhain:2016:BDC

[HI16] Jesko Hüttenhain and Christian Ikenmeyer. Binary determinantal complexity. Linear Algebra and its Applications, 504(??):559–573, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301264.

Hiai:2016:CCM

[Hia16a] Fumio Hiai. Concavity of certain matrix trace and norm functions. II. Linear Algebra and its Applications, 496(??):193–220, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516000628.

Hiai:2016:GAL

[Hia16b]  Fumio Hiai.  A gener-

alization of Araki’s logmajorization. Linear Algebra and its Applications, 501(??):1–16, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630009X.

Hiai:2019:LMR

[Hia19] Fumio Hiai. Log-majorization related to Rényi divergences. Linear Algebra and its Applications, 563(??):255–276, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305275.

Hilberdink:2015:OPC

[Hil15] Titus Hilberdink. An optimization problem concerning multiplicative functions. Linear Algebra and its Applications, 485(??):289–304, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500405X.

Hilberdink:2017:MME

[Hil17a] Titus Hilberdink. Matrices with multiplicative entries are tensor products. Linear Algebra and its Applications, 532(??):179–197, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304019.

Hildebrand:2017:CMC

[Hil17b] Roland Hildebrand. Copositive matrices with circulant zero support set. Linear Algebra and its Applications, 514(??):1–46, February 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516305110.

Hilberdink:2018:QKP

[Hil18] Titus Hilberdink. Quasi Kronecker products and a determinant formula. Linear Algebra and its Applications, 536(??):87–102, January 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305359.

Hirai:2018:CDD

[Hir18] Hiroshi Hirai. Computing DM-decomposition of a partitioned matrix with rank-1 blocks. Linear Algebra and its Applications, 547 (??):105–123, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300697.

Higgins:2016:ISP

[HJ16] Vijay Higgins and Charles Johnson. Inverse spectral problems for collections of leading principal submatrices of tridiagonal matrices. Linear Algebra and its Applications, 489(??): 104–122, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005996.

Han:2019:PRO

[HJ19] Deguang Han and Ted Juste. Phase-retrievable operatorvalued frames and representations of quantum channels. Linear Algebra and its Applications, 579(??): 148–168, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302265.

Herr:2019:CBR

[HJW19] John E. Herr, Palle E. T. Jorgensen, and Eric S. Weber. A characterization of boundary representations of positive matrices in the Hardy space via the Abel product. Linear Algebra and its Applications, 576(??): 51–66, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300909.

Hwang:2017:LTM

[HK17] Jinmi Hwang and Sejong Kim. Lie–Trotter means of positive definite operators. Linear Algebra and its Applications, 531(??): 268–280, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303725.

Hadjiantoni:2018:RTS

[HK18a] Stella Hadjiantoni and Erricos John Kontoghiorghes. A recursive three-stage least squares method for largescale systems of simultaneous equations. Linear Algebra and its Applications, 536(??):210–227, January 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304950.

Hwang:2018:VNM

[HK18b] Jin Ok Hwang and Donggyun Kim. Various 3×3 nonnegative matrices with prescribed eigenvalues and diagonal entries. Linear Algebra and its Applications, 554 (??):15–48, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302337.

Heo:2019:CIZ

[HK19a] Jaeseong Heo and Young-Hoon Kiem. On characterizing integral zeros of Krawtchouk polynomials by quantum entanglement. Linear Algebra and its Applications, 567(??):167–179, April 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Hu:2019:CMG

[HK19b] Yuxiang Hu and Steve Kirkland. Complete multipartite graphs and Braess edges. Linear Algebra and its Applications, 579(??):284–301, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302526.

Huhtanen:2019:FOQ

[HK19c] Marko Huhtanen and Vesa Kotila. Field of optimal quotients and Hermitianity. Linear Algebra and its Applications, 563(??): 527–547, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305391.

Hayat:2016:NGE

[HKLQ16] Sakander Hayat, Jack H. Koolen, Fenjin Liu, and Zhi Qiao. A note on graphs with exactly two main eigenvalues. Linear Algebra and its Applications, 511(??):318– 327, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304189.

Harmouch:2018:SLR

[HKM18] J. Harmouch, H. Khalil, and B. Mourrain. Structured low rank decomposition of multivariate Hankel matrices. Linear Algebra and its Applications, 542 (??):162–185, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302409.

Hnetynkova:2017:NRR

[HKP17] Iveta Hnetynková, Marie Kubínová, and Martin Plesinger. Noise representation in residuals of LSQR, LSMR, and CRAIG regularization. Linear Algebra and its Applications, 533(??):357–379, November 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517304512.

Horton:2019:SCW

[HKS19] Eric Horton, Kyle Kloster, and Blair D. Sullivan. Subgraph centrality and walkregularity. Linear Algebra and its Applications, 570 (??):225–244, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300606.

Hiai:2017:GMF

[HL17a] Fumio Hiai and Yongdo Lim. Geometric mean flows and the Cartan barycenter on the Wasserstein space over positive definite matrices. Linear Algebra and its Applications, 533(??):118– 131, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304378.

Hiai:2017:EII

[HL17b] Fumio Hiai and Minghua Lin. On an eigenvalue inequality involving the Hadamard product. Linear Algebra and its Applications, 515(??):313–320, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305432.

Hur:2017:NCN

[HL17c] Youngmi Hur and Zachary Lubberts. New constructions of nonseparable tight wavelet frames. Linear Algebra and its Applications, 534(??):13–35, December 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304688.

Hiai:2018:LMS

[HL18] Fumio Hiai and Yongdo Lim. Log-majorizations for the (symplectic) eigenvalues of the Cartan barycenter. Linear Algebra and its Applications, 553(??):129– 144, September 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302234.

Huang:2019:SLE

[HL19a] Xueyi Huang and Huiqiu Lin. Signless Laplacian eigenvalue problems of Nordhaus–Gaddum type. Linear Algebra and its Applications, 581(??):336–353, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303131.

Huang:2019:MZT

[HL19b] Zejun Huang and Zhenhua Lyu. 0–1 matrices with zero trace whose squares are 0–1 matrices. Linear Algebra and its Applications, 565(??): 156–176, March 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305494.

Hladik:2015:ASA

[Hla15] Milan Hladík. AE solutions and AE solvability to general interval linear systems. Linear Algebra and its Applications, 465(??): 221–238, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006156.

Hou:2015:LTH

[HLG15] Bo Hou, Yanhua Liu, and Suogang Gao. The Leonard triples having classical type. Linear Algebra and its Applications, 467(??):202–225, February 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007356.

Hu:2017:SDR

[HLLZ17] Dan Hu, Xueliang Li, Xiaogang Liu, and Shenggui Zhang. The spectral distribution of random mixed graphs. Linear Algebra and its Applications, 519(??): 343–365, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300277.

Holohan:2017:EPL

[HLM17] Naoise Holohan, Douglas J. Leith, and Oliver Mason.

Extreme points of the local differential privacy polytope. Linear Algebra and its Applications, 534(??): 78–96, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304871.

Hambric:2019:MNN

[HLPS19] Christopher Logan Hambric, Chi-Kwong Li, Diane Christine Pelejo, and Junping Shi. Minimum number of nonzero-entries in a 7 × 7 stable matrix. Linear Algebra and its Applications, 572 (??):135–152, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300965.

Huang:2016:TDP

[HLS^{+}16] Huajun Huang, Chih-Neng Liu, Patrícia Szokol, Ming-Cheng Tsai, and Jun Zhang. Trace and determinant preserving maps of matrices. Linear Algebra and its Applications, 507(??):373–388, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302403.

He:2017:TSG

[HLSP17] Changxiang He, Lin Lei, Haiying Shan, and Anni Peng. Two subgraph grafting theorems on the energy of bipartite graphs. Linear Algebra and its Applications, 515(??):96–110, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305262.

Huang:2015:MEIa

[HLW15] Fei Huang, Xueliang Li, and Shujing Wang. On maximum Estrada indices of bipartite graphs with some given parameters. Linear Algebra and its Applications, 465(??):283–295, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006302.

Hou:2017:EIM

[HLZ17] Xin Hou, Shangzhi Li, and Qingqing Zheng. Expressing infinite matrices over rings as products of involutions. Linear Algebra and its Applications, 532(??): 257–265, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304111.

See corrigendum [Hou20].

Hockman:2016:OPM

[HM16] Meira Hockman and Carminda Mennen. Orientation preserving Möbius transformations in R_{∞}^{4} and quaternionic determinants. Linear Algebra and its Applications, 505(??):367–386, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301586.

Hassani:2017:CME

[HM17a] Sara Hassani and Hossein Mohebi. Characterizations of minimal elements of topical functions on semimodules with applications. Linear Algebra and its Applications, 520 (??):104–124, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300423.

Hu:2017:EST

[HM17b] Guang-Da Hu and Taketomo Mitsui. Exponential stability of time-varying linear discrete systems. Linear Algebra and its Applications, 528(??):384–393, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305237.

Holbrook:2015:BPH

[HMNP15] John Holbrook, Nishan Mudalige, Mike Newman, and Rajesh Pereira. Bounds on polygons of higher rank numerical ranges. Linear Algebra and its Applications, 474 (??):230–242, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000166.

Hoover:2018:RCP

[HMPT18] Sarah L. Hoover, Daniel A. McCormick, Pietro Paparella, and Amber R. Thrall. On the realizability of the critical points of a realizable list. Linear Algebra and its Applications, 555(??): 301–313, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303082.

Hanaki:2017:MRT

[HMS17] Akihide Hanaki, Yasuaki Miyazaki, and Osamu Shimabukuro. Modular representation theory of BIB designs. Linear Algebra and its Applications, 514(??):174–197, February 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305158.

Hon:2019:NSD

[HMSC19] Sean Hon, Mohammad Ayman Mursaleen, and Stefano Serra-Capizzano. A note on the spectral distribution of symmetrized Toeplitz sequences. Linear Algebra and its Applications, 579(??): 32–50, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302368.

Hoppen:2019:SNO

[HMT19] Carlos Hoppen, Juan Monsalve, and Vilmar Trevisan. Spectral norm of oriented graphs. Linear Algebra and its Applications, 574(??): 167–181, August 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301223.

Huang:2018:PTD

[HN18] Peng-Ruei Huang and Hiroshi Nakazato. Product of two diagonal entries of a 3-by-3 normal matrix. Linear Algebra and its Applications, 544 (??):115–140, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300090.

Huang:2016:RMD

[HNR16] Guangxin Huang, Silvia Noschese, and Lothar Reichel. Regularization matrices determined by matrix nearness problems. Linear Algebra and its Applications, 502(??):41–57, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007211.

Helton:2015:FSR

[HNS15] J. William Helton, Jiawang Nie, and Jeremy S. Semko. Free semidefinite representation of matrix power functions. Linear Algebra and its Applications, 465(??): 347–362, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006351.

Holbrook:2015:STG

[HO15] J. Holbrook and K. C. O’Meara. Some thoughts on Gerstenhaber’s theorem. Linear Algebra and its Applications, 466(??):267–295, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006788.

Hoa:2015:COM

[Hoa15]  Dinh Trung Hoa.  On

characterization of operator monotone functions. Linear Algebra and its Applications, 487(??):260–267, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005376.

Hoa:2017:SIM

[Hoa17] Dinh Trung Hoa. Some inequalities for the matrix Heron mean. Linear Algebra and its Applications, 528(??):321–330, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303962.

Holubowski:2016:NSL

[Hol16] Waldemar Holubowski. New simple Lie algebra of uncountable dimension. Linear Algebra and its Applications, 492(??):9–12, March 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500676X.

Holbrook:2018:DPD

[Hol18] Andrew Holbrook. Differentiating the pseudo determinant. Linear Algebra and its Applications, 548 (??):293–304, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301289.

Hon:2018:OPS

[Hon18] Sean Hon. Optimal preconditioners for systems defined by functions of Toeplitz matrices. Linear Algebra and its Applications, 548 (??):148–171, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301058.

Hook:2015:MPS

[Hoo15]  James Hook.  Max-plus

singular values. Linear Algebra and its Applications, 486(??):419–442, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004863.

Hook:2019:MPL

[Hoo19] James Hook. Max-plus linear inverse problems: 2norm regression and system identification of maxplus linear dynamical systems with Gaussian noise. Linear Algebra and its Applications, 579(??):1–31, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302332.

Horvath:2017:IMG

[Hor17] Ákos G. Horváth. Isometries of Minkowski geometries. Linear Algebra and its Applications, 512(??): 172–190, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304554.

Hou:2015:DCK

[Hou15] Chengjun Hou. Derivations of a class of Kadison–Singer algebras. Linear Algebra and its Applications, 466(??): 241–253, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006946.

Hou:2019:DIM

[Hou19] Xin Hou. Decomposition of infinite matrices into products of commutators of involutions. Linear Algebra and its Applications, 563(??): 231–239, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830524X.

Hou:2020:CSE

[Hou20] Xin Hou. Corrigendum to “Expressing infinite matrices over rings as products of involutions” [Linear Algebra Appl. 532 (2017) 257– 265]. Linear Algebra and its Applications, 587(??): 387–391, February 15, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519304896.

See [HLZ17].

Hnetynkova:2015:CWS

[HP15] Iveta Hnetynková and Martin Plesinger. Complex wedge-shaped matrices: a generalization of Jacobi matrices. Linear Algebra and its Applications, 487(??):203– 219, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005327.

Higuchi:2019:EEO

[HPSS19] Yusuke Higuchi, Renato Portugal, Iwao Sato, and Etsuo Segawa. Eigenbasis of the evolution operator of 2tessellable quantum walks. Linear Algebra and its Applications, 583(??):257–281, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303593.

Hnetynkova:2018:TFC

[HPZ18] Iveta Hnetynková, Martin Plesinger, and Jana Záková. TLS formulation and core reduction for problems with structured right-hand sides. Linear Algebra and its Applications, 555(??):241–265, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303008.

Hnetynkova:2019:TFC

[HPZ19] Iveta Hnetynková, Martin Plesinger, and Jana Záková. On TLS formulation and core reduction for data fitting with generalized models. Linear Algebra and its Applications, 577(??):1– 20, September 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301703.

Hu:2015:LLS

[HQX15] Shenglong Hu, Liqun Qi, and Jinshan Xie. The largest Laplacian and signless Laplacian H-eigenvalues of a uniform hypergraph. Linear Algebra and its Applications, 469(??):1–27, March 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951400754X.

Hladik:2016:RSU

[HR16] Milan Hladík and Jirí Rohn. Radii of solvability and unsolvability of linear systems. Linear Algebra and its Applications, 503(??): 120–134, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300374.

Heuberger:2017:CIM

[HR17] Clemens Heuberger and Roswitha Rissner. Computing J-ideals of a matrix over a principal ideal domain. Linear Algebra and its Applications, 527 (??):12–31, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302082.

Habibzadeh:2018:OKF

[HRM18] S. Habibzadeh, J. Rooin, and M. S. Moslehian. Operator Ky Fan type inequalities. Linear Algebra and its Applications, 556(??): 220–237, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303276.

Hrubes:2016:FAM

[Hru16] Pavel Hrubes. On families of anticommuting matrices. Linear Algebra and its Applications, 493(??): 494–507, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007296.

He:2016:RSA

[HS16] Tian-Xiao He and Louis W. Shapiro. Row sums and alternating sums of Riordan arrays. Linear Algebra and its Applications, 507(??): 77–95, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302105.

He:2017:FCM

[HS17] Tian-Xiao He and Louis W. Shapiro. Fuss–Catalan matrices, their weighted sums, and stabilizer subgroups of the Riordan group. Linear Algebra and its Applications, 532(??):25–42, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303889.

Hall:2018:DM

[HS18] Frank J. Hall and Mikhail Stroev. On discriminant matrices. Linear Algebra and its Applications, 554(??): 316–328, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302696.

He:2019:STG

[HS19a] Changxiang He and Haiying Shan. The signature of two generalizations of line graphs. Linear Algebra and its Applications, 575(??): 159–173, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301582.

Hladik:2019:RBV

[HS19b] Milan Hladík and Iwona Skalna. Relations between various methods for solving linear interval and parametric equations. Linear Algebra and its Applications, 574 (??):1–21, August 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930120X.

Huang:2016:MSR

[HSS16a] Peng Huang, Wai Chee Shiu, and Pak Kiu Sun.

Maximizing the spectral radius of k-connected graphs with given diameter. Linear Algebra and its Applications, 488(??):350–362, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005959.

Huang:2016:LRP

[HSS16b] Zejun Huang, Shiyu Shi, and Nung-Sing Sze. Linear rank preservers of tensor products of rank one matrices. Linear Algebra and its Applications, 508(??): 255–271, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630324X.

Huang:2017:BFG

[HSTW17] Li-Ping Huang, Huadong Su, Gaohua Tang, and Jia-Bin Wang. Bilinear forms graphs over residue class rings. Linear Algebra and its Applications, 523(??):13– 32, June 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300617.

Helmberg:2017:STD

[HT17] Christoph Helmberg and Vilmar Trevisan. Spectral threshold dominance, Brouwer’s conjecture and maximality of Laplacian energy. Linear Algebra and its Applications, 512(??): 18–31, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304281.

Haukkanen:2018:IPD

[HT18a] Pentti Haukkanen and László Tóth. Inertia, positive definiteness and ℓ_{p} norm of GCD and LCM matrices and their unitary analogs. Linear Algebra and its Applications, 558(??):1–24, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304038.

Hnetynkova:2018:CHR

[HT18b] Iveta Hnetynková and Petr Tichý. Characterization of half-radial matrices. Linear Algebra and its Applications, 559(??):227–243, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304452.

Hu:2015:RNN

[Hu15] Shenglong Hu. Relations of the nuclear norm of a tensor and its matrix flattenings. Linear Algebra and its Applications, 478(??): 188–199, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002189.

Huang:2015:RSP

[Hua15] Rong Huang. Rank structure properties of rectangular matrices admitting bidiagonal-type factorizations. Linear Algebra and its Applications, 465(??): 1–14, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006016.

Huang:2019:GLT

[Hua19] De Huang. A generalized Lieb’s theorem and its applications to spectrum estimates for a sum of random matrices. Linear Algebra and its Applications, 579(??): 419–448, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302654.

Huhtanen:2015:PMS

[Huh15]  Marko Huhtanen.  The

product of matrix subspaces. Linear Algebra and its Applications, 471(??):

150–168, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500018X.

Hunter:2016:CKP

[Hun16] Jeffrey J. Hunter. The computation of key properties of Markov chains via perturbations. Linear Algebra and its Applications, 511(??):176– 202, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303937.

Hunter:2018:CMF

[Hun18a] Jeffrey J. Hunter. The computation of the mean first passage times for Markov chains. Linear Algebra and its Applications, 549(??): 100–122, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301204.

Hunter:2018:KFM

[Hun18b] Jeffrey J. Hunter. Kemeny’s function for Markov chains and Markov renewal processes. Linear Algebra and its Applications, 559(??): 54–72, December 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830421X.

Hurlimann:2015:MSK

[Hür15]  Werner Hürlimann.  On

Mardia skewness and kurtosis of Soules basis matrices in ROM simulation. Linear Algebra and its Applications, 473(??):284–296, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379514005679.

Hutchinson:2017:CMS

[Hut17] George Hutchinson. On complex matrix scalings of extremal permanent. Linear Algebra and its Applications, 522(??):111–126, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517300848.

He:2018:LSC

[HvD18] Changxiang He and Edwin R. van Dam. Laplacian spectral characterization of roses. Linear Algebra and its Applications, 536 (??):19–30, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304883.

Hastings:2016:ISA

[HvdD16] Alan Hastings and P. van den Driessche. Inequalities on the spectral abscissa for matrices arising in a stage-structured population model. Linear Algebra and its Applications, 494(??):90– 104, April 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000033.

Hemasinha:2015:OVS

[HW15a] Rohan Hemasinha and James R. Weaver. Ordered vector spaces. Linear Algebra and its Applications, 468(??): 171–183, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400319X.

Hillar:2015:HTL

[HW15b] Christopher J. Hillar and Andre Wibisono. A Hadamardtype lower bound for symmetric diagonally dominant positive matrices. Linear Algebra and its Applications, 472(??):135–141, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951500066X.

Huang:2015:MEIb

[HW15c] Fei Huang and Shujing Wang. On maximum Estrada indices of k-trees. Linear Algebra and its Applications, 487(??):316–327, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005595.

Hoffman:2016:GVI

[HW16] Alan J. Hoffman and Chai Wah Wu. Gersgorin variations IV: a left eigenvector approach. Linear Algebra and its Applications, 498 (??):136–144, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001147.

Huang:2018:SNI

[HWXC18] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. Some new inequalities for the minimum Heigenvalue of nonsingular Mtensors. Linear Algebra and its Applications, 558(??): 146–173, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830404X.

Huang:2015:MPI

[HWZ15] Liji Huang, Qing-Wen Wang, and Yang Zhang. The Moore–Penrose inverses of matrices over quaternion polynomial rings. Linear Algebra and its Applications, 475(??):45–61, June 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515000737.

Hong:2019:VCE

[HXL19] Zhen-Mu Hong, Zheng-Jiang Xia, and Hong-Jian Lai. Vertex-connectivity and eigenvalues of graphs. Linear Algebra and its Applications, 579(??):72–88, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302393.

Hu:2018:CPM

[HY18] Zhiguang Hu and Rongwei Yang. On the characteristic polynomials of multiparameter pencils. Linear Algebra and its Applications, 558(??):250–263, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304257.

He:2015:EPM

[HYH15] Kan He, Qing Yuan, and Jinchuan Hou. Entropypreserving maps on quantum states. Linear Algebra and its Applications, 467(??): 243–253, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951400737X.

He:2015:ESN

[HYY15] Zilong He, Pingzhi Yuan, and Lihua You. On the exponent set of nonnegative primitive tensors. Linear Algebra and its Applications, 465(??):376–390, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006442.

Huang:2015:NNE

[HZ15] Zejun Huang and Xingzhi Zhan. Nonsymmetric normal entry patterns with the maximum number of distinct indeterminates. Linear Algebra and its Applications, 485(??):359–371, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004577.

Huo:2017:SOG

[HZ17] Lijun Huo and Gengsheng Zhang. Subconstituents of the orthogonal graph of type (m, m − 1, 0) of odd characteristic. Linear Algebra and its Applications, 524(??):1–21, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301349.

Hu:2019:DCP

[HZ19] Zhiguang Hu and Philip B. Zhang. Determinants and characteristic polynomials of Lie algebras. Linear Algebra and its Applications, 563(??):426–439, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830538X.

Hu:2015:MMM

[HZC15] Yong-Jian Hu, Xu-Zhou Zhan, and Gong-Ning Chen. On maximum mass measures in truncated matricial Hamburger moment problems and related interpolation problems. Linear Algebra and its Applications, 466(??):429–456, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006831.

Idel:2017:PBW

[IGW17] Martin Idel, Sebastián Soto Gaona, and Michael M. Wolf. Perturbation bounds for Williamson’s symplectic normal form. Linear Algebra and its Applications, 525 (??):45–58, July 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301751.

See note [SZ23].

Ilmonen:2008:EMJ

[IHM08] Pauliina Ilmonen, Pentti Haukkanen, and Jorma K. Merikoski. On eigenvalues of meet and join matrices associated with incidence functions. Linear Algebra and its Applications, 429(4): 859–874, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See comment [AKYD16].

Ilmonen:2018:GEP

[IK18] Pauliina Ilmonen and Vesa Kaarnioja. Generalized eigenvalue problems for meet and join matrices on semilattices. Linear Algebra and its Applications, 536(??): 250–273, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Izhakian:2016:MOQ

[IKR16] Zur Izhakian, Manfred Knebusch, and Louis Rowen. Minimal orderings and quadratic forms on a free module over a supertropical semiring. Linear Algebra and its Applications, 507(??): 420–461, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630218X.

Iarrobino:2018:NMH

[IKVZ18] Anthony Iarrobino, Leila Khatami, Bart Van Steirteghem, and Rui Zhao. Nilpotent matrices having a given Jordan type as maximum commuting nilpotent orbit. Linear Algebra and its Applications, 546(??):210– 260, June 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300685.

Israel:2016:AGM

[IKW16] Arie Israel, Felix Krahmer, and Rachel Ward. An arithmetic-geometric mean inequality for products of three matrices. Linear Algebra and its Applications, 488 (??):1–12, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005285.

Ito:2017:BCN

[IL17a] Masaru Ito and Bruno F. Lourenço. A bound on the Carathéodory number. Linear Algebra and its Applications, 532(??):347–363, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730407X.

Ito:2017:CDH

[IL17b] Masaru Ito and Bruno F. Lourenço. The p-cones in dimension n ≥ 3 are not homogeneous when p ≠

2. Linear Algebra and its Applications, 533(??):326– 335, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304494.

Ihringer:2019:NSR

[IM19] Ferdinand Ihringer and Akihiro Munemasa. New strongly regular graphs from finite geometries via switching. Linear Algebra and its Applications, 580(??): 464–474, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Ioppolo:2018:ESS

[Iop18] Antonio Ioppolo. The exponent for superalgebras with superinvolution. Linear Algebra and its Applications, 555(??):1–20, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830291X.

Ismail:2019:TRG

[IR19] M. E. H. Ismail and A. Sri Ranga. R_{II} type recurrence, generalized eigenvalue problem and orthogonal polynomials on the unit circle. Linear Algebra and its Applications, 562(??): 63–90, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304798.

Tsukada:2016:APC

[iT16] Shin ichi Tsukada. Asymptotic properties of a correlation matrix under a twostep monotone incomplete sample. Linear Algebra and its Applications, 488(??): 86–101, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005625.

Ilisevic:2018:SWE

[IT18] Dijana Ilisević and Aleksej Turnsek. On superstability of the Wigner equation. Linear Algebra and its Applications, 542(??): 391–401, ???? 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303646.

Tsukada:2019:UEC

[iT19] Shin ichi Tsukada. Unbiased estimator for a covariance matrix in a three-step monotone incomplete sample. Linear Algebra and its Applications, 579(??): 169–205, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302320.

Ito:2017:SET

[Ito17] Keiji Ito. The skew energy of tournaments. Linear Algebra and its Applications, 518 (??):144–158, April 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306292.

Idel:2015:SNF

[IW15] Martin Idel and Michael M. Wolf. Sinkhorn normal form for unitary matrices. Linear Algebra and its Applications, 471(??):76– 84, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000221.

Jain:2017:HPS

[Jai17] Tanvi Jain. Hadamard powers of some positive matrices. Linear Algebra and its Applications, 528(??):147– 158, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302476.

Javanshiri:2018:MNS

[JC18] Hossein Javanshiri and Mehdi Choubin. Multipliers for von Neumann–Schatten Bessel sequences in separable Banach spaces. Linear Algebra and its Applications, 545(??):108–138, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518300430.

Jung:2019:CSG

[JCW19] Dohee Jung, Changbum Chun, and Xiaofeng Wang.

Construction of stable and globally convergent schemes for the matrix sign function. Linear Algebra and its Applications, 580(??): 14–36, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930271X.

Johnson:2017:NDE

[JDS17] Charles R. Johnson, António Leal Duarte, and Carlos M. Saiago. The number of distinct eigenvalues for which an index decreases multiplicity. Linear Algebra and its Applications, 516(??):8– 23, March 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305596.

Jelisiejew:2018:VCF

[Jel18] Joachim Jelisiejew. VSPs of cubic fourfolds and the Gorenstein locus of the Hilbert scheme of 14 points on A6. Linear Algebra and its Applications, 557(??):265–286, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518303744.

Jensen:2017:GHH

[Jen17] Tommy R. Jensen. Grassmann homomorphism and Hajós-type theorems. Linear Algebra and its Applications, 522(??):140–152, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517300897.

Jeong:2016:SCE

[JG16] Juyoung Jeong and M. Seetharama Gowda. Spectral cones in Euclidean Jordan algebras. Linear Algebra and its Applications, 509(??):286– 305, November 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303317.

Jeong:2017:SSF

[JG17] Juyoung Jeong and M. Seetharama Gowda. Spectral sets and functions on Euclidean Jordan algebras. Linear Algebra and its Applications, 518 (??):31–56, April 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306097.

Jiang:2015:DEL

[Jia15] Tiefeng Jiang. Distributions of eigenvalues of large Euclidean matrices generated from ℓ_{p} balls and spheres. Linear Algebra and its Applications, 473 (??):14–36, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379513006198.

Jing:2015:UOE

[Jin15]  Naihuan Jing.  Unitary

and orthogonal equivalence of sets of matrices. Linear Algebra and its Applications, 481(??):235–242, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500289X.

Johnson:2015:CTP

[JK15] Marianne Johnson and Mark Kambites. Convexity of tropical polytopes. Linear Algebra and its Applications, 485(??):531–544, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004607.

Jung:2015:DMS

[JKKL15] Wooseok Jung, Jon-Lark Kim, Yeonho Kim, and Kisun Lee. The dimension of magic squares over fields of characteristics two and three. Linear Algebra and its Applications, 472 (??):118–134, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000609.

Junghanns:2016:CQM

[JKP16] P. Junghanns, R. Kaiser, and D. Potts. Collocationquadrature methods and fast summation for Cauchy singular integral equations with fixed singularities. Linear Algebra and its Applications, 491(??):187–238, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004061.

Johnston:2017:PQS

[JKP^{+}17] Nathaniel Johnston, Steve Kirkland, Sarah Plosker, Rebecca Storey, and Xiaohong Zhang. Perfect quantum state transfer using Hadamard diagonalizable graphs. Linear Algebra and its Applications, 531(??): 375–398, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303373.

Jovovic:2015:NBG

[JKS15] Ivana Jovović, Tamara Koledin, and Zoran Stanić. Non-bipartite graphs of fixed order and size that minimize the least eigenvalue. Linear Algebra and its Applications, 477(??):148–164, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515001913.

Ji:2016:SRT

[JL16a] Chunyu Ji and Mei Lu. On the spectral radius of trees with given independence number. Linear Algebra and its Applications, 488(??): 102–108, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005571.

Joswig:2016:WDT

[JL16b] Michael Joswig and Georg Loho. Weighted digraphs and tropical cones. Linear Algebra and its Applications, 501(??):304–343, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001397.

Johnson:2017:SDM

[JLD17] Charles R. Johnson and António Leal-Duarte. On the solvability of derived matrix problems, including completions and duals. Linear Algebra and its Applications, 514(??): 165–173, February 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303639.

Johnson:2018:MNE

[JLDS18] Charles R. Johnson, António Leal-Duarte, and Carlos M. Saiago. The minimum number of eigenvalues of multiplicity one in a diagonalizable matrix, over a field, whose graph is a tree. Linear Algebra and its Applications, 559(??):1–10, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304221.

Jocic:2018:NIC

[JLM18] Danko R. Jocić, Milan Lazarević, and Stefan Milosević. Norm inequalities for a class of elementary operators generated by analytic functions with non-negative Taylor coefficients in ideals of compact operators related to p-modified unitarily invariant norms. Linear Algebra and its Applications, 540 (??):60–83, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306432.

Jivulescu:2015:PRS

[JLNR15] Maria Anastasia Jivulescu, Nicolae Lupa, Ion Nechita, and David Reeb. Positive reduction from spectra. Linear Algebra and its Applications, 469(??): 276–304, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007666.

Jiang:2017:UPA

[JLT17] Shuxia Jiang, Yuanyuan Liu, and Yingchun Tang. A unified perturbation analysis framework for countable Markov chains. Linear Algebra and its Applications, 529(??):413–440, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302902.

Jocic:2016:ROC

[JM16] Danko R. Jocić and Stefan Milosević. Refinements of operator Cauchy–Schwarz and Minkowski inequalities for p-modified norms and related norm inequalities. Linear Algebra and its Applications, 488(??): 284–301, January 1, 2016.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005741.

Jovanovic:2016:SRL

[JMM16] Z. Jovanović, E. I. Milovanović, and I. Z. Milovanović. Some remarks on Laplacian eigenvalues of connected graphs. Linear Algebra and its Applications, 503 (??):48–55, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300696.

Johnson:2017:RCS

[JMP17] C. R. Johnson, C. Marijuán, and M. Pisonero. Ruling out certain 5-spectra for the symmetric nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 512(??): 129–135, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304219.

Jaume:2018:SBN

[JMPS18] Daniel A. Jaume, Gonzalo Molina, Adrián Pastine, and Martín D. Safe. A {−1, 0, 1}- and sparsest basis for the null space of a forest in optimal time. Linear Algebra and its Applications, 549 (??):53–66, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301290.

Julio:2019:URS

[JMPS19] Ana I. Julio, Carlos Marijuán, Miriam Pisonero, and Ricardo L. Soto. On universal realizability of spectra. Linear Algebra and its Applications, 563(??): 353–372, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305366.

Jagels:2016:CRI

[JMRV16] Carl Jagels, Thomas Mach, Lothar Reichel, and Raf Vandebril. Convergence rates for inverse-free rational approximation of matrix functions. Linear Algebra and its Applications, 510(??):291–310, December 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516303810.

Joe:2016:HNF

[Joe16] Stephen Joe. The Hermite normal form for certain rank-1 circulant and skew-circulant lattice rules. Linear Algebra and its Applications, 496(??):438–451, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516001002.

Jorgensen:2015:DSR

[Jør15] Leif K. Jørgensen. Directed strongly regular graphs with rank 5. Linear Algebra and its Applications, 477 (??):102–111, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001834.

Johnson:2016:PSR

[JP16] Charles R. Johnson and Pietro Paparella. Perron spectratopes and the real nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 493(??):281–300, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007077.

Johnson:2017:MVK

[JP17] Charles R. Johnson and Pietro Paparella. A matricial view of the Karpelevic Theorem. Linear Algebra and its Applications, 520(??):1–15, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300253.

Johnston:2018:IEP

[JP18] Nathaniel Johnston and Everett Patterson. The inverse eigenvalue problem for entanglement witnesses. Linear Algebra and its Applications, 550(??):1–27, August 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951830154X.

Julio:2019:CNR

[JRS19] Ana I. Julio, Oscar Rojo, and Ricardo L. Soto. Centrosymmetric nonnegative realization of spectra. Linear Algebra and its Applications, 581(??):260–284, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302897.

Julio:2015:PBN

[JS15] Ana I. Julio and Ricardo L. Soto. Persymmetric and bisymmetric nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 469(??):130–152, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Jeyaraman:2016:CPS

[JS16] I. Jeyaraman and K. C. Sivakumar. Complementarity properties of singular Mmatrices. Linear Algebra and its Applications, 510(??): 42–63, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303305.

Jiang:2018:UAO

[JS18a] Tianpei Jiang and Hristo Sendov. A unified approach to operator monotone functions. Linear Algebra and its Applications, 541(??): 185–210, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306626.

Johnson:2018:GPW

[JS18b] Charles R. Johnson and Carlos M. Saiago. Geometric Parter–Wiener, etc. theory. Linear Algebra and its Applications, 537(??): 332–347, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305694.

Jaume:2019:CND

[JS19] Daniel A. Jaume and Rodrigo Sota. On the corenilpotent decomposition of trees. Linear Algebra and its Applications, 563(??): 207–214, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304920.

Johnson:2016:NMS

[JSST16] Charles R. Johnson, Ilya M. Spitkovsky, Mirjana Stevanovic, and Morrison Turnansky. Normal matrices subordinate to a tree and flat portions of the field of values. Linear Algebra and its Applications, 507(??): 198–207, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302257.

Johnson:2019:CME

[JST19] Charles R. Johnson, Carlos M. Saiago, and Kenji Toyonaga. The change in multiplicity of an eigenvalue due to adding or removing edges. Linear Algebra and its Applications, 560 (??):86–99, January 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304580.

Johnson:2017:NMS

[JT17] Charles R. Johnson and Morrison Turnansky. Normal matrices subordinate to a graph. Linear Algebra and its Applications, 531(??): 54–63, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302963.

Jacobsen:2018:GIG

[JT18a] Karly A. Jacobsen and Joseph H. Tien. A generalized inverse for graphs with absorption. Linear Algebra and its Applications, 537(??):118–147, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305633.

Jankauskas:2018:CRM

[JT18b] J. Jankauskas and J. M. Thuswaldner. Characterization of rational matrices that admit finite digit representations. Linear Algebra and its Applications, 557(??):350– 358, November 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303793.

Jacobs:2015:EET

[JTT15] David P. Jacobs, Vilmar Trevisan, and Fernando Tura. Eigenvalues and energy in threshold graphs. Linear Algebra and its Applications, 465(??):412–425, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006405.

Johnson:2016:TPT

[JZ16] Charles R. Johnson and Yulin Zhang. TP perturbation of TN matrices and totally positive directions. Linear Algebra and its Applications, 490(??): 66–76, February 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006205.

Johnson:2018:MLS

[JZ18] Charles R. Johnson and Yulin Zhang. Multiplicity lists for symmetric matrices whose graphs have few missing edges. Linear Algebra and its Applications, 540 (??):221–233, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306602.

Jiang:2019:ERK

[JZC19] Xiaoyu Jiang, Yanpeng Zheng, and Xiaoting Chen. Extending a refinement of Koteljanskii’s inequality. Linear Algebra and its Applications, 574(??): 252–261, August 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301594.

Jin:2018:EPT

[JZZ18] Ya-Lei Jin, Jie Zhang, and Xiao-Dong Zhang. Equitable partition theorem of tensors and spectrum of generalized power hypergraphs. Linear Algebra and its Applications, 555(??): 21–38, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302866.

Kalita:2016:EFE

[Kal16] Debajit Kalita. Extremizing first eigenvalue of 3colored digraphs made with given blocks. Linear Algebra and its Applications, 503 (??):83–99, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300684.

Kaleibary:2018:SMR

[Kal18] V. Kaleibary. On some matrix reverse Cauchy–Schwarz and Hölder inequalities. Linear Algebra and its Applications, 552(??):210–218, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302180.

Kandi:2016:SMS

[Kan16] Marko Kandić. Sets of matrices with singleton spectra generated by positive matrices. Linear Algebra and its Applications, 496 (??):463–474, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001129.

Karantzas:2020:ESD

[KAPS20] N. Karantzas, N. Atreas, M. Papadakis, and T. Stavropoulos. Erratum to “On the design of multi-dimensional compactly supported Parseval framelets with directional characteristics [Linear Algebra Appl. 582 (2019) 1–36]. Linear Algebra and its Applications, 584(??): 483–484, January 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/

science/article/pii/S0024379519304197.

See [AKPS19].

Karlsson:2016:BCH

[Kar16] Bengt R. Karlsson. BCCB complex Hadamard matrices of order 9, and MUBs. Linear Algebra and its Applications, 504(??):309–324, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301112.

Kurata:2015:MPI

[KB15] Hiroshi Kurata and Ravindra B. Bapat. Moore– Penrose inverse of a Euclidean distance matrix. Linear Algebra and its Applications, 472(??):106–117, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515000610.

Ke:2016:BDI

[KC16] Yuanyuan Ke and Jianlong Chen. The Bott– Duffin (e, f)-inverses and their applications. Linear Algebra and its Applications, 489(??):61–74, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006047.

Kenter:2016:SSB

[Ken16] Franklin H. J. Kenter. Sharp spectral bounds of several graph parameters using eigenvector norms. Linear Algebra and its Applications, 504(??):48–63, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300350.

Kovac:2017:GCJ

[KH17] Erna Begović Kovac and Vjeran Hari. On the global convergence of the Jacobi method for symmetric matrices of order 4 under parallel strategies. Linear Algebra and its Applications, 524 (??):199–234, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301532.

Koolen:2016:CSH

[KHI16a] Jack H. Koolen, Sakander Hayat, and Quaid Iqbal. Corrigendum to “Hypercubes are determined by their distance spectra” [Linear Algebra Appl. 505 (2016) 97–108]. Linear Algebra and its Applications, 506(??):628–629, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516302609.

See [KHI16b].

Koolen:2016:HDT

[KHI16b] Jack H. Koolen, Sakander Hayat, and Quaid Iqbal. Hypercubes are determined by their distance spectra. Linear Algebra and its Applications, 505(??):97–108, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301471.

See corrigendum [KHI16a].

Kim:2015:EEH

[Kim15a]  Dong-Soo Kim.  Ellip-

soids and elliptic hyperboloids in the Euclidean space E^{n+1}. Linear Algebra and its Applications, 471 (??):28–45, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008179.

Kim:2015:CIB

[Kim15b] Hwa Kyung Kim. Characterization of irreducible Boolean matrices with the largest generalized competition index. Linear Algebra and its Applications, 466(??):218–232, February 1, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951400651X.

Kim:2018:QOB

[Kim18] Sun Kwang Kim. Quantity of operators with Bhatia– Semrl property. Linear Algebra and its Applications, 537(??):22–37, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517305554.

Kirkland:2015:SAP

[Kir15] Steve Kirkland. Sensitivity analysis of perfect state transfer in quantum spin networks. Linear Algebra and its Applications, 472 (??):1–30, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000427.

Kisi:2015:CTL

[Kis15] Emre Kisi. Corrigendum to “Tripotency of a linear combination of two involutory matrices and a tripotent matrix that mutually commute” [Linear Algebra Appl. 437 (9) (2012) 2091–2109]. Linear Algebra and its Applications, 477 (??):211–212, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Konno:2018:SAU

[KIS18] Norio Konno, Yusuke Ide, and Iwao Sato. The spectral analysis of the unitary matrix of a 2-tessellable staggered quantum walk on a graph. Linear Algebra and its Applications, 545 (??):207–225, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300363.

Joji:2016:SRA

[kJ16]  Du ko Joji.  Some re-

marks about acyclic and tridiagonal Birkhoff polytopes. Linear Algebra and its Applications, 495(??): 108–121, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000501.

Kocharovsky:2017:PCD

[KK17] Vitaly V. Kocharovsky and Vladimir V. Kocharovsky. On the permanents of circulant and degenerate Schur matrices. Linear Algebra and its Applications, 519 (??):366–381, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300538.

Kim:2018:DMM

[KK18] Ik-Pyo Kim and Arnold R. Kräuter. Decompositions of a matrix by means of its dual matrices with applications. Linear Algebra and its Applications, 537(??): 100–117, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305657.

Kernfeld:2015:TTP

[KKA15] Eric Kernfeld, Misha Kilmer, and Shuchin Aeron. Tensortensor products with invertible linear transforms. Linear Algebra and its Applications, 485(??):545–570, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004358.

Kho:2017:PIM

[KKL^{+}17] Dong Yeong Kho, Woong Kook, JaeHoon Lee, Jinhyeong Lee, and Kang-Ju Lee. Path intersection matrices and applications to networks. Linear Algebra and its Applications, 524 (??):278–292, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Keiper:2017:CSF

[KKLP17] Sandra Keiper, Gitta Kutyniok, Dae Gwan Lee, and Götz E. Pfander. Compressed sensing for finitevalued signals. Linear Algebra and its Applications, 532(??):570–613, November 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517304160.

Karimjanov:2015:SLA

[KKO15] I. A. Karimjanov, A. Kh. Khudoyberdiyev, and B. A. Omirov. Solvable Leibniz algebras with triangular nilradicals. Linear Algebra and its Applications, 466(??): 530–546, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007150.

Kazakov:2019:RHR

[KKP19] Matthew Kazakov, David W. Kribs, and Rajesh Pereira. Real higher rank numerical ranges and ellipsoids. Linear Algebra and its Applications, 577(??):204–213, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301867.

Kloster:2018:WEW

[KKS18] Kyle Kloster, Daniel Král’, and Blair D. Sullivan. Walk entropy and walk-regularity. Linear Algebra and its Applications, 546(??):115–121, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518300703.

Kim:2016:STG

[KKY16] Hyun Kwang Kim, Jack H. Koolen, and Jae Young Yang. A structure theory for graphs with fixed smallest eigenvalue. Linear Algebra and its Applications, 504(??):1–13, September 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516300672.

Kim:2015:PML

[KL15] Sejong Kim and Hosoo Lee. The power mean and the least squares mean of probability measures on the space of positive definite matrices. Linear Algebra and its Applications, 465(??): 325–346, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Kittaneh:2017:TIP

[KL17] Fuad Kittaneh and Minghua Lin. Trace inequalities for positive semidefinite block matrices. Linear Algebra and its Applications, 524 (??):153–158, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301568.

Kauffman:2018:CBF

[KL18a] Louis H. Kauffman and Pedro Lopes. Colorings beyond Fox: The other linear Alexander quandles. Linear Algebra and its Applications, 548(??):221–258, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517306535.

Kuo:2018:CUC

[KL18b] Yueh-Cheng Kuo and Tsung-Lin Lee. Computing the unique CANDECOMP/ PARAFAC decomposition of unbalanced tensors by homotopy method. Linear Algebra and its Applications, 556(??):238–264, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303227.

Kenter:2019:EPS

[KL19a] Franklin H. J. Kenter and Jephian C.-H. Lin. On the error of a priori sampling: zero forcing sets and propagation time. Linear Algebra and its Applications, 576(??):124–141, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301411.

Kim:2019:BJO

[KL19b] Sun Kwang Kim and Han Ju Lee. The Birkhoff-James orthogonality of operators on infinite dimensional Banach spaces. Linear Algebra and its Applications, 582(??): 440–451, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303520.

Kline:2019:SMR

[Kli19] Jeffery Kline. A sparser matrix representation of the Mertens function. Linear Algebra and its Applications, 581(??):354–366, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Kim:2016:FPM

[KLL16] Sejong Kim, Hosoo Lee, and Yongdo Lim. A fixed point mean approximation to the Cartan barycenter of positive definite matrices. Linear Algebra and its Applications, 496(??):420– 437, May 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000963.

Kum:2017:RAS

[KLL17] Sangho Kum, Hosoo Lee, and Yongdo Lim. The resolvent average on symmetric cones of JB-algebras. Linear Algebra and its Applications, 520(??):260–273, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517300514.

Kim:2018:RIG

[KLL18] Sejong Kim, Hosoo Lee, and Yongdo Lim. Repetition invariant geometric means. Linear Algebra and its Applications, 544 (??):443–459, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300387.

Kim:2019:GMB

[KLL19] Sejong Kim, Hosoo Lee, and Yongdo Lim. Geometric mean block matrices. Linear Algebra and its Applications, 575(??): 299–313, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301600.

Ke:2016:NRT

[KLN16] Rihuan Ke, Wen Li, and Michael K. Ng. Numerical ranges of tensors. Linear Algebra and its Applications, 508(??):100–132, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302658.

Kaygorodov:2015:CAD

[KLP15] Ivan Kaygorodov, Artem Lopatin, and Yury Popov. Conservative algebras of 2dimensional algebras. Linear Algebra and its Applications, 486(??):255–274, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004760.

Kaygorodov:2018:SIM

[KLP18] Ivan Kaygorodov, Artem Lopatin, and Yury Popov. Separating invariants for 2 × 2 matrices. Linear Algebra and its Applications, 559(??):114–124, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303835.

Kuo:2017:AAS

[KLS17] Yueh-Cheng Kuo, Wen-Wei Lin, and Shih-Feng Shieh. The asymptotic analysis of the structure-preserving doubling algorithms. Linear Algebra and its Applications, 531(??):318–355, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303683.

Kalauch:2019:SIM

[KLS19a] A. Kalauch, S. Lavanya, and K. C. Sivakumar. Singular irreducible M-matrices revisited. Linear Algebra and its Applications, 565 (??):47–64, March 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305536.

Koberda:2019:ELK

[KLS19b] Thomas Koberda, Feng Luo, and Hongbin Sun. An effective Lie–Kolchin Theorem for quasi-unipotent matrices. Linear Algebra and its Applications, 581(??):304– 323, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303106.

Ku:2015:CGS

[KLW15] Cheng Yeaw Ku, Terry Lau, and Kok Bin Wong. Cayley graph on symmetric group generated by elements fixing k points. Linear Algebra and its Applications, 471 (??):405–426, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500035X.

Ku:2016:SEP

[KLW16] Cheng Yeaw Ku, Terry Lau, and Kok Bin Wong. The smallest eigenvalues of the 1-point fixing graph. Linear Algebra and its Applications, 493(??):433–446, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007259.

Ku:2018:SEC

[KLW18] Cheng Yeaw Ku, Terry Lau, and Kok Bin Wong. The spectrum of eigenvalues for certain subgraphs of the kpoint fixing graph. Linear Algebra and its Applications, 543(??):72–91, April 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517306882.

Koolen:2018:FHG

[KLY18] Jack H. Koolen, Yan-Ran Li, and Qianqian Yang. On fat Hoffman graphs with smallest eigenvalue at least −3, part II. Linear Algebra and its Applications, 550(??): 121–143, August 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730647X.

Karow:2015:MPS

[KM15a] Michael Karow and Emre Mengi. Matrix polynomials with specified eigenvalues. Linear Algebra and its Applications, 466(??): 457–482, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400679X.

Kiani:2015:UCG

[KM15b] Dariush Kiani and Mohsen Mollahajiaghaei. On the unitary Cayley graphs of matrix algebras. Linear Algebra and its Applications, 466(??): 421–428, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006934.

Klotz:2016:SFP

[KM16] Lutz Klotz and Conrad Mädler. Some functions preserving positive semidefiniteness of 2 × 2 block matrices. Linear Algebra and its Applications, 507(??): 68–76, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630204X.

Kubkowski:2018:PGB

[KM18] Mariusz Kubkowski and Jan Mielniczuk. Projections of a general binary model on a logistic regression. Linear Algebra and its Applications, 536(??):152–173, January 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305372.

Kim:2019:CNS

[KM19] Se-Jin Kim and Arthur Mehta. Chromatic numbers, Sabidussi’s Theorem and Hedetniemi’s conjecture for non-commutative graphs. Linear Algebra and its Applications, 582(??): 291–309, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303362.

Kwon:2017:JGC

[KMM17] Y. S. Kwon, A. D. Mednykh, and I. A. Mednykh. On Jacobian group and complexity of the generalized Petersen graph GP(n, k) through Chebyshev polynomials. Linear Algebra and its Applications, 529(??):355– 373, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302690.

Konno:2019:NWI

[KMMS19] Norio Konno, Hideo Mitsuhashi, Hideaki Morita, and Iwao Sato. A new weighted Ihara zeta function for a graph. Linear Algebra and its Applications, 571 (??):154–179, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300904.

Kuzma:2019:AVR

[KMOR19] Bojan Kuzma, Mitja Mastnak, Matjaz Omladic, and Heydar Radjavi. On approximate versions of reducibility results for matrix groups and semigroups. Linear Algebra and its Applications, 577(??):41–52, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301715.

Konno:2016:QSW

[KMS16] Norio Konno, Hideo Mitsuhashi, and Iwao Sato. The quaternionic second weighted zeta function of a graph and the Study determinant. Linear Algebra and its Applications, 510(??): 92–109, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303342.

Kittaneh:2017:UIN

[KMS17a] Fuad Kittaneh, Mohammad Sal Moslehian, and Mohammad Sababheh. Unitarily invariant norm inequalities for elementary operators involving G_{1} operators. Linear Algebra and its Applications, 513(??):

84–95, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304839.

Kulkarni:2017:CPA

[KMS17b] Rajesh S. Kulkarni, Yusuf Mustopa, and Ian Shipman. The characteristic polynomial of an algebra and representations. Linear Algebra and its Applications, 530 (??):47–56, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302884.

Kittaneh:2015:CDN

[KMY15] Fuad Kittaneh, Mohammad Sal Moslehian, and Takeaki Yamazaki. Cartesian decomposition and numerical radius inequalities. Linear Algebra and its Applications, 471(??):46–53, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008192.

Kehl:2016:ASC

[KN16] René Kehl and Reinhard Nabben. Avoiding singular coarse grid systems. Linear Algebra and its Applications, 507(??):137–152, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301975.

Kumar:2018:KTT

[KNR18] V. B. Kiran Kumar, M. N. N. Namboodiri, and Rahul Rajan. A Korovkin-type theory for non-self-adjoint Toeplitz operators. Linear Algebra and its Applications, 543 (??):140–161, April 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306821.

Kang:2015:SRP

[KNY15] L. Kang, V. Nikiforov, and X. Yuan. The p-spectral radius of k-partite and kchromatic uniform hypergraphs. Linear Algebra and its Applications, 478(??): 81–107, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001706.

Kuran:2016:GAM

[KÖ16] Özge Kuran and M. Revan Özkale. Gilmour’s approach to mixed and stochastic restricted ridge predictions in linear mixed models. Linear Algebra and its Applications, 508(??):22–47, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302579.

Kobayashi:2017:DGS

[Kob17] Masato Kobayashi. A directed graph structure of alternating sign matrices. Linear Algebra and its Applications, 519(??):164–190, April 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306309.

Kolokolnikov:2015:MAC

[Kol15]  T. Kolokolnikov.  Maxi-

mizing algebraic connectivity for certain families of graphs. Linear Algebra and its Applications, 471(??): 122–140, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000142.

Kooij:2019:GWG

[Koo19] Robert Kooij. On generalized windmill graphs. Linear Algebra and its Applications, 565(??):25–46, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305482.

Kuzma:2015:EOP

[KOST15] Bojan Kuzma, Matjaz Omladic, Klemen Sivic, and Josef Teichmann. Exotic one-parameter semigroups of endomorphisms of a symmetric cone. Linear Algebra and its Applications, 477 (??):42–75, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001500.

Kozyakin:2016:HAF

[Koz16] Victor Kozyakin. Hourglass alternative and the finiteness conjecture for the spectral characteristics of sets of non-negative matrices. Linear Algebra and its Applications, 489(??): 167–185, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006126.

Karder:2017:MSP

[KP17] Mahdi Karder and Tatjana Petek. Maps on states preserving generalized entropy of convex combinations. Linear Algebra and its Applications, 532(??): 86–98, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730366X.

Kircheis:2019:DIN

[KP19] Melanie Kircheis and Daniel Potts. Direct inversion of the nonequispaced fast Fourier transform. Linear Algebra and its Applications, 575(??):106–140, August 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301351.

Kokabifar:2018:DMP

[KPL18] E. Kokabifar, P. J. Psarrakos, and G. B. Loghmani. On the distance from a matrix polynomial to matrix polynomials with some prescribed eigenvalues. Linear Algebra and its Applications, 544(??):158– 185, May 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300168.

Kunis:2016:MGP

[KPRvdO16] Stefan Kunis, Thomas Peter, Tim Römer, and Ulrich von der Ohe. A multivariate generalization of Prony’s method. Linear Algebra and its Applications, 490(??): 31–47, February 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006187.

Kirkland:2016:APM

[KQZ16] Steve Kirkland, Pu Qiao, and Xingzhi Zhan. Algebraically positive matrices. Linear Algebra and its Applications, 504(??): 14–26, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300726.

Kahl:2017:DCG

[KR17] K. Kahl and H. Rittich. The deflated conjugate gradient method: Convergence, perturbation and accuracy. Linear Algebra and its Applications, 515(??):111–129, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305122.

Krivulin:2015:EPT

[Kri15] Nikolai Krivulin. Extremal properties of tropical eigenvalues and solutions to tropical optimization problems. Linear Algebra and its Applications, 468(??): 211–232, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514004133.

Koziol:2017:BUE

[KRZ^{+}17] Arkadiusz Koziol, Anuradha Roy, Roman Zmyślony, Ricardo Leiva, and Miguel Fonseca. Best unbiased estimates for parameters of three-level multivariate data with doubly exchangeable covariance structure. Linear Algebra and its Applications, 535(??):87–104, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304718.

Koziol:2018:FCE

[KRZ^{+}18] Arkadiusz Koziol, Anuradha Roy, Roman Zmyślony, Ricardo Leiva, and Miguel Fonseca. Free-coordinate estimation for doubly multivariate data. Linear Algebra and its Applications, 547 (??):217–239, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300867.

Karami:2015:PNH

[KS15a] Saeed Karami and Abbas Salemi. Polynomial numerical hulls of the direct sum of a normal matrix and a Jordan block. Linear Algebra and its Applications, 465(??):143–160, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514005965.

Kitson:2015:MLC

[KS15b] Derek Kitson and Bernd Schulze. Maxwell–Laman counts for bar-joint frameworks in normed spaces. Linear Algebra and its Applications, 481(??):313–329, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002967.

Kandic:2017:DAG

[KS17] Marko Kandić and Klemen Sivic. On the dimension of the algebra generated by two positive semicommuting matrices. Linear Algebra and its Applications, 512(??):136–161, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304268.

Klee:2019:LAT

[KS19] Steven Klee and Matthew T. Stamps. Linear algebraic techniques for weighted spanning tree enumeration. Linear Algebra and its Applications, 582(??):391–402, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303490.

Kannan:2015:SPS

[KSMB15] M. Rajesh Kannan, Naomi Shaked-Monderer, and Abraham Berman. Some properties of strong H-tensors and general H-tensors. Linear Algebra and its Applications, 476(??):42–55, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500141X.

Kong:2015:SRE

[KSTX15] Lingchen Kong, Jie Sun, Jiyuan Tao, and Naihua Xiu. Sparse recovery on Euclidean Jordan algebras. Linear Algebra and its Applications, 465(??):65–87, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400603X.

Kubota:2018:PGW

[KSTY18] Sho Kubota, Etsuo Segawa, Tetsuji Taniguchi, and Yusuke Yoshie. Periodicity of Grover walks on generalized Bethe trees. Linear Algebra and its Applications, 554(??):371–391, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302635.

Kubota:2019:QWI

[KSTY19] Sho Kubota, Etsuo Segawa, Tetsuji Taniguchi, and Yusuke Yoshie. A quantum walk induced by Hoffman graphs and its periodicity. Linear Algebra and its Applications, 579(??): 217–236, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302356.

Kim:2017:SAP

[KSVW17] In-Jae Kim, Bryan L. Shader, Kevin N. Vander Meulen, and Matthew West. Spectrally arbitrary pattern extensions. Linear Algebra and its Applications, 517(??):120–128, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305894.

Kurata:2015:CMC

[KT15] Hiroshi Kurata and Pablo Tarazaga. The cell matrix closest to a given Euclidean distance matrix. Linear Algebra and its Applications, 485(??):194–207, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004449.

Kim:2017:PEI

[KT17] Ik-Pyo Kim and Michael J. Tsatsomeros. Pascal eigenspaces and invariant sequences of the first or second kind. Linear Algebra and its Applications, 535(??):171–190, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305141.

Khare:2019:CST

[KT19] Apoorva Khare and Akaki Tikaradze. A Carlitz–von Staudt type theorem for finite rings. Linear Algebra and its Applications, 568 (??):106–126, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302830.

Kressner:2016:LRA

[KU16] Daniel Kressner and André Uschmajew. On low-rank approximability of solutions to high-dimensional operator equations and eigenvalue problems. Linear Algebra and its Applications, 493(??):556–572, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007302.

Kulkarni:2015:NST

[Kul15] S. H. Kulkarni. The null space theorem. Linear Algebra and its Applications, 472 (??):97–105, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000592.

Kumar:2019:NCS

[Kum19] P. Santhosh Kumar. A note on convexity of sections of quaternionic numerical range. Linear Algebra and its Applications, 572 (??):92–116, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300990.

Kuratko:2019:FSP

[Kur19] Jan Kurátko. Factorization of saddle-point matrices in dynamical systems optimization-reusing pivots. Linear Algebra and its Applications, 566(??):61– 85, April 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305974.

Kushel:2015:MIS

[Kus15] O. Y. Kushel. On matrices with Q^{2}-scalings. Linear Algebra and its Applications, 465(??):139–142, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006132.

Kushel:2016:PSP

[Kus16] Olga Y. Kushel. On the positive stability of P^{2}- matrices. Linear Algebra and its Applications, 503(??): 190–214, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300763.

Kutschan:2018:TCT

[Kut18] Benjamin Kutschan. Tangent cones to tensor train varieties. Linear Algebra and its Applications, 544 (??):370–390, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300181.

Khosravi:2019:ABE

[KVP19] Amin A. Khosravi, Hamid Reza Ebrahimi Vishki, and Antonio M. Peralta. Aron– Berner extensions of triple maps with application to the bidual of Jordan Banach triple systems. Linear Algebra and its Applications, 580(??):436–463, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302903.

Klep:2017:NPR

[KVV17] Igor Klep, Victor Vinnikov, and Jurij Volcic. Nulland Positivstellensätze for rationally resolvable ideals. Linear Algebra and its Applications, 527(??):260–293, August 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302343.

Kwon:2016:TEA

[Kwo16] Hyukmin Kwon. Terminating Euclidean algorithm for a non-Noetherian Bézout domain. Linear Algebra and its Applications, 506 (??):10–32, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301793.

Kim:2015:FPV

[KY15] Jaewoong Kim and Jasang Yoon. Flat phenomena of 2-variable weighted shifts. Linear Algebra and its Applications, 486(??):234–254, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004759.

Koshlukov:2017:GGJ

[KY17] Plamen Koshlukov and Felipe Yukihide Yasumura. Group gradings on the Jordan algebra of upper triangular matrices. Linear Algebra and its Applications, 534 (??):1–12, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304706.

Kim:2018:NCS

[KY18] Jaewoong Kim and Jasang Yoon. A new characterization of subnormality for a class of 2-variable weighted shifts with 1-atomic core. Linear Algebra and its Applications, 538(??):22–42, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305797.

Krelifa:2017:RAO

[KZ17] Ali Krelifa and Ebtissem Zerouki. Riordan arrays and d-orthogonality. Linear Algebra and its Applications, 515(??):331–353, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305778.

LaGuardia:2016:OCC

[La 16] Giuliano G. La Guardia. On optimal constacyclic codes. Linear Algebra and its Applications, 496(??):594– 610, May 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001154.

Guardia:2016:CCC

[LA16] Giuliano G. La Guardia and Marcelo M. S. Alves. On cyclotomic cosets and code constructions. Linear Algebra and its Applications, 488(??):302–319, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005613.

Lange:2015:RBS

[Lan15] M. Lange. Residual bounds for some or all singular values. Linear Algebra and its Applications, 464(??): 28–37, January 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Lau:2018:LIJ

[Lau18] Pan-Shun Lau. Linear images of joint unitary orbits of Hermitian matrices. Linear Algebra and its Applications, 542(??):35–53, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306103.

Lax:2016:APS

[Lax16] David C. Lax. Accessible proof of standard monomial basis for coordinatization of Schubert sets of flags. Linear Algebra and its Applications, 494(??):105– 137, April 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000057.

Liang:2019:IRW

[LC19] Kaihao Liang and Michael J. Clay. Iterative re-weighted least squares algorithm for l_{p}-minimization with tight frame and 0 < p ≤ 1. Linear Algebra and its Applications, 581(??):413–434, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303167.

Liu:2019:SCO

[LCF19] Shuangqing Liu, Yanxun Chang, and Tao Feng. Several classes of optimal Ferrers diagram rank-metric codes. Linear Algebra and its Applications, 581(??):128– 144, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302927.

Li:2015:NEI

[LCL15] Chaoqian Li, Zhen Chen, and Yaotang Li. A new eigenvalue inclusion set for tensors and its applications. Linear Algebra and its Applications, 481(??):36– 53, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002645.

Liu:2018:DPW

[LCLW18] Jung-Hui Liu, Chun-Yen Chou, Ching-Jou Liao, and Ngai-Ching Wong. Disjointness preservers of AW^{∗}- algebras. Linear Algebra and its Applications, 552(??): 71–84, September 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301988.

Li:2019:MMS

[LCNZ19] Yuan Li, Xiaomei Cai, Jiajia Niu, and Jiaxin Zhang. The minimal and maximal symmetries for J - contractive projections. Linear Algebra and its Applications, 563(??):313–330, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305330.

Li:2019:INB

[LCWZ19] Chaoqian Li, Ljiljana Cvetković, Yimin Wei, and Jianxing Zhao. An infinity norm bound for the inverse of Dashnic–Zusmanovich type matrices with applications. Linear Algebra and its Applications, 565(??):99–122, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305767.

Lin:2016:CEG

[LD16a] Huiqiu Lin and Kinkar Ch. Das. Characterization of extremal graphs from distance signless Laplacian eigenvalues. Linear Algebra and its Applications, 500(??):77– 87, July 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300039.

Ljubenovic:2016:LPW

[LD16b] Martin Ljubenović and Dragan S. Djordjević. Linear preservers of weak majorization on ℓ^{p}(I)^{+}, when p ∈ (1, ∞). Linear Algebra and its Applications, 497 (??):181–198, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001439.

Lin:2017:DSC

[LD17a] Huiqiu Lin and Stephen Drury. The distance spectrum of complements of trees. Linear Algebra and its Applications, 530(??): 185–201, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730304X.

Ljubenovic:2017:LPW

[LD17b] Martin Ljubenović and Dragan S. Djordjević. Linear preservers of weak majorization on l^{1}(I)^{+}, when I is an infinite set. Linear Algebra and its Applications, 517(??):177–198, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Ljubenovic:2017:WSF

[LD17c] Martin Ljubenović and Dragan S. Djordjević. Weak supermajorization and families as doubly superstochastic operators on l^{p}(I). Linear Algebra and its Applications, 532(??):312–346, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730410X.

Li:2018:CIE

[LDL18] Jicheng Li, Liqiang Dong, and Guo Li. A class of inverse eigenvalue problems for real symmetric banded matrices with odd bandwidth. Linear Algebra and its Applications, 541(??): 131–162, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306663.

Linde:2015:MCG

[LdlP15] J. Linde and M. J. de la Puente. Matrices commuting with a given normal tropical matrix. Linear Algebra and its Applications, 482(??): 101–121, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002852.

Lin:2018:FTL

[LDZ18] Hongying Lin, Zhibin Du, and Bo Zhou. On the first two largest distance Laplacian eigenvalues of unicyclic graphs. Linear Algebra and its Applications, 546 (??):289–307, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300740.

Lee:2015:DBP

[Lee15a] Eun-Young Lee. The offdiagonal block of a PPT matrix. Linear Algebra and its Applications, 486(??): 449–453, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500484X.

Lee:2015:GWN

[Lee15b] Hsin-Yi Lee. Gau–Wu numbers of nonnegative matrices. Linear Algebra and its Applications, 469(??): 594–608, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007939.

Lee:2015:UPN

[Lee15c] Hsin-Yi Lee. Uniqueness problem on numerical ranges of 3-by-3 companion matrices. Linear Algebra and its Applications, 482(??): 84–100, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003134.

Lee:2016:PSD

[Lee16] Hwangrae Lee. Power sum decompositions of elementary symmetric polynomials. Linear Algebra and its Applications, 492(??):89– 97, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006813.

Lee:2019:HTO

[Lee19a] Jongrak Lee. Hyponormality of Toeplitz operators on the weighted Bergman space with matrix-valued circulant symbols. Linear Algebra and its Applications, 576(??): 35–50, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300727.

Leenheer:2019:SDC

[Lee19b] Patrick De Leenheer. Stability of diffusively coupled linear systems with an invariant cone. Linear Algebra and its Applications, 580(??): 396–416, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302769.

Lka:2016:SIC

[Lék16] Zoltán Léka. Some inequalities for central moments of matrices. Linear Algebra and its Applications, 496 (??):246–261, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000860.

Lev:2015:DNM

[Lev15]  Vsevolod F. Lev.  Dis-

crete norms of a matrix and the converse to the expander mixing lemma. Linear Algebra and its Applications, 483(??):158–181, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500350X.

Levick:2018:FQC

[Lev18]  Jeremy Levick.  Factor-

izations of quantum channels. Linear Algebra and its Applications, 553(??):145– 166, September 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302325.

Fan:2016:CSR

[lFwW16] Feng lei Fan and Chih wen Weng. A characterization of strongly regular graphs in terms of the largest signless Laplacian eigenvalues. Linear Algebra and its Applications, 506 (??):1–5, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301598.

Lu:2018:MPC

[LG18] Di Lu and Chun-Hua Guo. Monotonicity and positivity of coefficients of power series expansions associated with Newton and Halley methods for the matrix p th root. Linear Algebra and its Applications, 556(??): 131–143, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303240.

Liu:2018:BBL

[LGZ18] Xuewen Liu, Xiangqian Guo, and Kaiming Zhao. Biderivations of the block Lie algebras. Linear Algebra and its Applications, 538(??):43–55, February 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517305906.

Li:2017:FVM

[LH17] Zhongyan Li and Deguang Han. Frame vector multipliers for finite group representations. Linear Algebra and its Applications, 519 (??):191–207, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300149.

Liu:2015:EPE

[LHC15] Fuxiang Liu, Jianhua Hu, and Guanglei Chu. Estimation of parameters in the extended growth curve model via outer product least squares for covariance. Linear Algebra and its Applications, 473(??): 236–260, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514005308.

Liu:2015:LTR

[LHG15] Huan Liu, Bo Hou, and Suogang Gao. Leonard triples, the Racah algebra, and some distance-regular graphs of Racah type. Linear Algebra and its Applications, 484(??):435–456, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004036.

Lu:2017:GDL

[LHH17] Lu Lu, Qiongxiang Huang, and Xueyi Huang. On graphs with distance Laplacian spectral radius of multiplicity n − 3. Linear Algebra and its Applications, 530(??):485–499, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303440.

Lu:2018:DST

[LHL18a] Lu Lu, Qiongxiang Huang, and Zhenzhen Lou. On the distance spectra of threshold graphs. Linear Algebra and its Applications, 553(??):223–237, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302453.

Luo:2018:RBS

[LHL18b] Wenjun Luo, Jing Huang, and Shuchao Li. On the relationship between the skew-rank of an oriented graph and the rank of its underlying graph. Linear Algebra and its Applications, 554(??):205–223, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830257X.

Huang:2016:OVH

[lHS16] Wen ling Huang and Peter Semrl. The optimal version of Hua’s fundamental theorem of geometry of square matrices — the low dimensional case. Linear Algebra and its Applications, 498 (??):21–57, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514005473.

Lin:2018:NSR

[LHX18] Huiqiu Lin, Xing Huang, and Jie Xue. A note on the A_{α}-spectral radius of graphs. Linear Algebra and its Applications, 557(??):430–437, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303811.

Li:2015:NSC

[Li15a] Haohao Li. Necessary and sufficient conditions for unified optimality of interval linear program in the general form. Linear Algebra and its Applications, 484(??): 154–174, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003821.

Li:2015:NPF

[Li15b] Jiaojiao Li. L^{2} norm preserving flow in matrix geometry. Linear Algebra and its Applications, 487(??):220– 231, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005248.

Li:2019:ASI

[Li19] Haohao Li. AE solutions to interval linear systems over max-plus algebra. Linear Algebra and its Applications, 578(??):297–313, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302289.

Liao:2017:BPR

[Lia17]  Ping Liao.  Bounds for

the Perron root of nonnegative matrices and spectral radius of iteration matrices. Linear Algebra and its Applications, 530(??): 253–265, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303099.

Lieb:2018:UPN

[Lie18] Julia Lieb. Uniform probability and natural density of mutually left coprime polynomial matrices over finite fields. Linear Algebra and its Applications, 539(??): 134–159, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306341.

Lin:2015:PCI

[Lin15a] Huiqiu Lin. Proof of a conjecture involving the second largest D-eigenvalue and the number of triangles. Linear Algebra and its Applications, 472(??):48–53, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000634.

Lin:2015:OSN

[Lin15b] Minghua Lin. Orthogonal sets of normal or conjugatenormal matrices. Linear Algebra and its Applications, 483(??):227–235, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Lin:2015:SAM

[Lin15c] Minghua Lin. Some applications of a majorization inequality due to Bapat and Sunder. Linear Algebra and its Applications, 469(??): 510–517, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007915.

Lin:2016:UNZ

[Lin16a] Jephian C.-H. Lin. Using a new zero forcing process to guarantee the Strong Arnold Property. Linear Algebra and its Applications, 507(??):229–250, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302348.

Lin:2016:HMI

[Lin16b] Minghua Lin. The Hua matrix and inequalities related to contractive matrices. Linear Algebra and its Applications, 511(??): 22–30, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303925.

Lin:2016:RTR

[Lin16c] Minghua Lin. Remarks on two recent results of Audenaert. Linear Algebra and its Applications, 489(??): 24–29, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005972.

Lin:2017:NPC

[Lin17a] Minghua Lin. New properties for certain positive semidefinite matrices. Linear Algebra and its Applications, 520(??):32–43, May 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517300265.

Lin:2017:DSB

[Lin17b] Minghua Lin. On Drury’s solution of Bhatia & Kittaneh’s question. Linear Algebra and its Applications, 528(??):33–39, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001270.

Lin:2019:ZFN

[Lin19] Jephian C.-H. Lin. Zero forcing number, Grundy domination number, and their variants. Linear Algebra and its Applications, 563(??): 240–254, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305263.

Liu:2015:RNC

[Liu15] Yue Liu. Ray nonsingularity of cycle tree matrices. Linear Algebra and its Applications, 465(??): 275–282, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006363.

Liu:2016:GBM

[Liu16a]  Juntong Liu.  General-

izations of the Brunn– Minkowski inequality. Linear Algebra and its Applications, 508(??):206–213, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302683.

Liu:2016:ASP

[Liu16b] Wen Liu. The attenuated space poset A_{q}(N, M). Linear Algebra and its Applications, 506(??):244–273, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301768.

Liu:2017:BPP

[Liu17] Shunyi Liu. On the bivariate permanent polynomials of graphs. Linear Algebra and its Applications, 529(??):148–163, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302586.

Liu:2018:RNC

[Liu18] Yue Liu. The ray nonsingularity of certain uniformly random ray patterns. Linear Algebra and its Applications, 537(??):302–317, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305724.

Laali:2016:SGC

[LJK16] A. R. Fiuj Laali, H. Haj Seyyed Javadi, and Dariush Kiani. Spectra of generalized corona of graphs. Linear Algebra and its Applications, 493(??):411–425, March 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515007065.

Li:2016:TEL

[LJL16] Chaoqian Li, Aiquan Jiao, and Yaotang Li. An Stype eigenvalue localization set for tensors. Linear Algebra and its Applications, 493(??):469–483, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007326.

Ljubenovic:2015:WMD

[Lju15] Martin Ljubenović. Weak majorization and doubly substochastic operators on ℓ^{p}(I). Linear Algebra and its Applications, 486(??): 295–316, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004875.

Lee:2016:HPW

[LK16] Hosoo Lee and Sejong Kim. The Hadamard product for the weighted Karcher means. Linear Algebra and its Applications, 501(??): 290–303, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300398.

Liu:2016:PEU

[LKY16] Lele Liu, Liying Kang, and Xiying Yuan. On the principal eigenvectors of uniform hypergraphs. Linear Algebra and its Applications, 511(??):430–446, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304141.

Li:2015:DTQ

[LL15a] Chaoqian Li and Yaotang Li. Double B-tensors and quasi-double B-tensors. Linear Algebra and its Applications, 466(??):343–356, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400696X.

Li:2015:SCT

[LL15b] Yuan Li and Yu-E Li. Some characterizations of the trace norm triangle equality. Linear Algebra and its Applications, 484(??):396–408, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004115.

Liu:2015:HAM

[LL15c] Jianxi Liu and Xueliang Li. Hermitian-adjacency matrices and Hermitian energies of mixed graphs. Linear Algebra and its Applications, 466(??):182–207, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006971.

Li:2016:RBB

[LL16a] Chaoqian Li and Yaotang Li. Relationships between Brauer-type eigenvalue inclusion sets and a Brualditype eigenvalue inclusion set for tensors. Linear Algebra and its Applications, 496 (??):71–80, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000926.

Liesen:2016:FRA

[LL16b] Jörg Liesen and Robert Luce. Fast recovery and approximation of hidden Cauchy structure. Linear Algebra and its Applications, 493(??):261–280, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007120.

Liu:2016:IT

[LL16c] Weihui Liu and Wen Li. On the inverse of a tensor. Linear Algebra and its Applications, 495(??): 199–205, April 15, 2016.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000136.

Lau:2017:ADP

[LL17] Jinting Lau and Ming Huat Lim. Additive decomposability preservers and related results on tensor products of matrices. Linear Algebra and its Applications, 519 (??):54–72, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630636X.

Liu:2018:BIN

[LL18a] Huiqing Liu and Mei Lu. Bounds on the independence number and signless Laplacian index of graphs. Linear Algebra and its Applications, 539(??):44–59, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306158.

Liu:2018:CPG

[LL18b] Xiaogang Liu and Shunyi Liu. On the A_{α}-characteristic polynomial of a graph. Linear Algebra and its Applications, 546(??):274–288, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518300752.

Leites:2019:ICM

[LL19] Dimitry Leites and Oleksandr Lozhechnyk. Inverses of Cartan matrices of Lie algebras and Lie superalgebras. Linear Algebra and its Applications, 583(??):195– 256, December 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303738.

Liu:2017:JSE

[LLH17] Xiaoqian Liu, Liangyuan Liu, and Jianhua Hu. James–Stein estimation problem for a multivariate normal random matrix and an improved estimator. Linear Algebra and its Applications, 532(??):231–256, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303968.

Lin:2015:IED

[LLL15] Huiqiu Lin, Ruifang Liu, and Xiwen Lu. The inertia and energy of the distance matrix of a connected graph. Linear Algebra and its Applications, 467(??): 29–39, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007253.

Li:2017:SSP

[LLL17] Tiexiang Li, Ren-Cang Li, and Wen-Wei Lin. A symmetric structure-preserving Γ QR algorithm for linear response eigenvalue problems. Linear Algebra and its Applications, 520(??):191– 214, May 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300186.

Lau:2019:GNR

[LLPS19] Pan-Shun Lau, Chi-Kwong Li, Yiu-Tung Poon, and Nung-Sing Sze. The generalized numerical range of a set of matrices. Linear Algebra and its Applications, 563(??):24–46, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304646.

Liu:2017:CSN

[LLS17a] Shuting Liu, Huiqiu Lin, and Jinlong Shu. Corrigendum to “A note on (signless) Laplacian spectral ordering with maximum degrees of graphs” [Linear Algebra Appl. 521 (2017) 135– 141]. Linear Algebra and its Applications, 529(??):457– 459, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302550.

See [LLS17b].

Liu:2017:NSL

[LLS17b] Shuting Liu, Huiqiu Lin, and Jinlong Shu. A note on (signless) Laplacian spectral ordering with maximum degrees of graphs. Linear Algebra and its Applications, 521 (??):135–141, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300137.

See corrigendum [LLS17a].

Liao:2015:MIA

[LLSX15] Jun Liao, Heguo Liu, Minfeng Shao, and Xingzhong Xu. A matrix identity and its applications. Linear Algebra and its Applications, 471 (??):346–352, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000233.

Liu:2019:STP

[LLT19] Ruifang Liu, Hong-Jian Lai, and Yingzhi Tian. Spanning tree packing number and eigenvalues of graphs with given girth. Linear Algebra and its Applications, 578(??):411–424, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302319.

Li:2015:EBI

[LLV15] Wen Li, Dongdong Liu, and Seak-Weng Vong. Zeigenpair bounds for an irreducible nonnegative tensor. Linear Algebra and its Applications, 483(??): 182–199, October 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003523.

Liu:2017:SLI

[LLZ17] Huiqing Liu, Mei Lu, and Shunzhe Zhang. On the signless Laplacian index and radius of graphs. Linear Algebra and its Applications, 519 (??):327–342, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300472.

Laface:2016:LSB

[LM16a] Antonio Laface and Joaquín Moraga. Linear systems on the blow-up of (P^{1})^{n}. Linear Algebra and its Applications, 492(??):52–67, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006722.

Li:2016:GLD

[LM16b] Dan Li and Jixiang Meng. The graphs with the least distance eigenvalue at least −(1 + \sqrt{17})/2. Linear Algebra and its Applications, 493(??):358–380, March 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515006436.

Lu:2016:CHS

[LM16c] Linyuan Lu and Shoudong Man. Connected hypergraphs with small spectral radius. Linear Algebra and its Applications, 509(??):206–227, November 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516302750.

Liu:2017:SPR

[LM17] Lily Li Liu and Dan Ma. Some polynomials related to Dowling lattices and x-Stieltjes moment sequences. Linear Algebra and its Applications, 533(??):195–209, November 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304433.

Li:2019:GSG

[LM19] Jiaojiao Li and Li Ma. Global solutions and ground states of a nonlinear Schrödinger equation in matrix geometry. Linear Algebra and its Applications, 573(??):1– 11, July 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301193.

Luzon:2016:SIL

[LMM^{+}16] Ana Luzón, Donatella Merlini, Manuel A. Morón, L. Felipe Prieto-Martinez, and Renzo Sprugnoli. Some inverse limit approaches to the Riordan group. Linear Algebra and its Applications, 491(??):239–262, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004425.

Ladra:2016:CSC

[LMO16] M. Ladra, K. K. Masutova, and B. A. Omirov. Corrigendum to “Classification of solvable Leibniz algebras with naturally graded filiform nilradical”

[Linear Algebra Appl. 438

(7) (2013) 2973–3000]. Linear Algebra and its Applications, 507(??):513–517, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302749.

See [CLOK13].

Lopez:2017:CRH

[LMP^{+}17] Antonio Vera López, Luis Martínez, Antonio Vera Pérez, Beatriz Vera Pérez, and Olga Basova. Combinatorics related to Higman’s conjecture I: Parallelogramic digraphs and dispositions. Linear Algebra and its Applications, 530(??): 414–444, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303154.

Luzon:2017:FCA

[LMPM17] Ana Luzón, Manuel A. Morón, and L. Felipe Prieto-Martínez. A formula to construct all involutions in Riordan matrix groups. Linear Algebra and its Applications, 533(??):397–417, November 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304536.

Luzon:2016:DPR

[LMR16] Ana Luzón, Manuel A. Morón, and José L. Ramírez. Double parameter recurrences for polynomials in bi-infinite Riordan matrices and some derived identities. Linear Algebra and its Applications, 511(??):237– 258, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304190.

Lang:2015:BIF

[LMS15] Norman Lang, Hermann Mena, and Jens Saak. On the benefits of the LDL^{T} factorization for large-scale differential matrix equation solvers. Linear Algebra and its Applications, 480(??): 44–71, September 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002219.

Laffey:2016:EAI

[LMS16] Thomas Laffey, Olga Markova, and Helena Smigoc. The effect of assuming the identity as a generator on the length of the matrix algebra. Linear Algebra and its Applications, 498(??):378– 393, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Lazzarin:2019:NTG

[LMT19] João Lazzarin, Oscar F. Márquez, and Fernando C. Tura. No threshold graphs are cospectral. Linear Algebra and its Applications, 560(??):133–145, January 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518304749.

Luo:2019:HTP

[LMX19] Wei Luo, Mohammad Sal Moslehian, and Qingxiang Xu. Halmos’ two projections theorem for Hilbert C^{∗}-module operators and the Friedrichs angle of two closed submodules. Linear Algebra and its Applications, 577(??):134–158, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301831.

Lenes:2018:LGM

[LMZRZ18] Eber Lenes, Exequiel Mallea-Zepeda, María Robbiano, and Jonnathan Rodríguez Z. On line graphs with maximum energy. Linear Algebra and its Applications, 545 (??):15–31, May 15, 2018.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300399.

Li:2017:SAM

[LN17] Binlong Li and Bo Ning. Spectral analogues of Moon– Moser’s theorem on Hamilton paths in bipartite graphs. Linear Algebra and its Applications, 515(??): 180–195, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630550X.

Looseova:2018:FSD

[LN18] Iveta Looseová and Petr Necesal. The Fucík spectrum of the discrete Dirichlet operator. Linear Algebra and its Applications, 553(??):58– 103, September 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302118.

Lau:2016:CSS

[LNT16a] Pan-Shun Lau, Tuen-Wai Ng, and Nam-Kiu Tsing. Convexity and starshapedness of real linear images of special orthogonal orbits. Linear Algebra and its Applications, 507(??): 51–67, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302099.

Lau:2016:SSG

[LNT16b] Pan-Shun Lau, Tuen-Wai Ng, and Nam-Kiu Tsing. The star-shapedness of a generalized numerical range. Linear Algebra and its Applications, 506(??):308–315, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302191.

Ludwig:2016:DPM

[LNT16c] E. Ludwig, R. Nabben, and J. M. Tang. Deflation and projection methods applied to symmetric positive semi-definite systems. Linear Algebra and its Applications, 489(??): 253–273, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005947.

Leung:2017:PCP

[LNW17] Chi-Wai Leung, Chi-Keung Ng, and Ngai-Ching Wong. The positive contractive part of a noncommutative L^{p}- space is a complete Jordan invariant. Linear Algebra and its Applications, 519 (??):102–110, April 15, 2017.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306322.

Lomadze:2017:RHO

[Lom17] Vakhtang Lomadze. On the reduction of high order linear PDEs to first order. Linear Algebra and its Applications, 530(??):1– 14, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302938.

Loperfido:2015:SVD

[Lop15] Nicola Loperfido. Singular value decomposition of the third multivariate moment. Linear Algebra and its Applications, 473(??): 202–216, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514003632.

Loperfido:2017:NKM

[Lop17] Nicola Loperfido. A new kurtosis matrix, with statistical applications. Linear Algebra and its Applications, 512 (??):1–17, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304323.

Levene:2019:NGC

[LOS19] Rupert H. Levene, Polona Oblak, and Helena Smigoc. A Nordhaus–Gaddum conjecture for the minimum number of distinct eigenvalues of a graph. Linear Algebra and its Applications, 564 (??):236–263, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305561.

Lototsky:2015:SSB

[Lot15] S. V. Lototsky. Simple spectral bounds for sums of certain Kronecker products. Linear Algebra and its Applications, 469(??): 114–129, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007617.

Lin:2018:SPR

[LOvdD18] Jephian C.-H. Lin, D. D. Olesky, and P. van den Driessche. Sign patterns requiring a unique inertia. Linear Algebra and its Applications, 546(??):67– 85, June 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300454.

Lee:2015:SIS

[LP15] Sang-Gu Lee and Jin-Woo Park. Sign idempotent sign pattern matrices that allow idempotence. Linear Algebra and its Applications, 487(??):232–241, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005352.

Lentzos:2018:DRI

[LP18] Konstantinos Lentzos and Lillian F. Pasley. Determinantal representations of invariant hyperbolic plane curves. Linear Algebra and its Applications, 556(??): 108–130, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303173.

Li:2015:SGI

[LPP15] Jin Li, Rajesh Pereira, and Sarah Plosker. Some geometric interpretations of quantum fidelity. Linear Algebra and its Applications, 487(??):158–171, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005261.

Lee:2019:ICS

[LPPZ19] Dae Gwan Lee, Götz E. Pfander, Volker Pohl, and Weiqi Zhou. Identification of channels with single and multiple inputs and outputs under linear constraints. Linear Algebra and its Applications, 581(??):435–470, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930312X.

Li:2016:PTP

[LPW16] Chi-Kwong Li, Diane Christine Pelejo, and Kuo-Zhong Wang. Product of two positive contractions. Linear Algebra and its Applications, 501(??):409–423, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516300027.

Li:2015:MTI

[LQL15] Chaoqian Li, Liqun Qi, and Yaotang Li. MB-tensors and MB_{0}-tensors. Linear Algebra and its Applications, 484(??):141–153, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003833.

Lan:2015:GOD

[LS15a] Jingfen Lan and Lingsheng Shi. Graphs of order n and diameter 2(n − 1)/3 minimizing the spectral radius. Linear Algebra and its Applications, 486(??): 219–233, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004802.

Lavrauw:2015:CFI

[LS15b] Michel Lavrauw and John Sheekey. Canonical forms of 2 × 2 × 3 tensors over the real field, algebraically closed fields, and finite fields. Linear Algebra and its Applications, 476(??):133– 147, July 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001494.

Lv:2015:CRF

[LS15c] Fusheng Lv and Wenchang Sun. Construction of robust frames in erasure recovery. Linear Algebra and its Applications, 479(??): 155–170, August 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002554.

Larson:2017:NBU

[LS17a] David Larson and Sam Scholze. Nilpotent bridging for unions of two bases. Linear Algebra and its Applications, 529(??):164–184, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302355.

Loewy:2017:NCS

[LS17b] Raphael Loewy and Oren Spector. A necessary condition for the spectrum of nonnegative symmetric 5×5 matrices. Linear Algebra and its Applications, 528(??): 206–272, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303330.

Lemos:2018:SLM

[LS18a] Rute Lemos and Graça Soares. Some log-majorizations and an extension of a determinantal inequality. Linear Algebra and its Applications, 547(??):19–31, June 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518300764.

Liu:2018:RAS

[LS18b] Yan Liu and Yuming Shi. Regular approximations of spectra of singular discrete linear Hamiltonian systems with one singular endpoint. Linear Algebra and its Applications, 541(??):94–130, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306572.

Li:2019:RBP

[LS19] Shuchao Li and Wanting Sun. On the relation between the positive inertia index and negative inertia index of weighted graphs. Linear Algebra and its Applications, 563(??):411–425, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305408.

Lungten:2016:SBF

[LSM16] S. Lungten, W. H. A. Schilders, and J. M. L. Maubach. Sparse block factorization of saddle point matrices. Linear Algebra and its Applications, 502(??): 214–242, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004851.

Lungten:2018:TIF

[LSM18] Sangye Lungten, Wil H. A. Schilders, and Joseph M. L. Maubach. Threshold incomplete factorization constraint preconditioners for saddlepoint matrices. Linear Algebra and its Applications, 545 (??):76–107, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830048X.

Liu:2015:SSC

[LSX15] Ruifang Liu, Wai Chee Shiu, and Jie Xue. Sufficient spectral conditions on Hamiltonian and traceable graphs. Linear Algebra and its Applications, 467(??):254–266, February 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007514.

Lins:2018:NNR

[LSZ18] Brian Lins, Ilya M. Spitkovsky, and Siyu Zhong. The normalized numerical range and the Davis–Wielandt shell. Linear Algebra and its Applications, 546(??):187– 209, June 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300417.

Levene:2012:CBN

[LT12] Rupert H. Levene and Richard M. Timoney. Completely bounded norms of right module maps. Linear Algebra and its Applications, 436(5):1406–1424, March 1, 2012. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379511006203.

See corrigendum [LT16a].

Levene:2016:CSC

[LT16a] Rupert H. Levene and Richard M. Timoney. Corrigendum to “Completely bounded norms of right module maps” [Linear Algebra Appl. 436 (5) (2012) 1406–1424]. Linear Algebra and its Applications, 505(??):387–389, September 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516301781.

See [LT12].

Li:2016:IMM

[LT16b] Cailan Li and Man Cheung Tsui. On the images of multilinear maps of matrices over finite-dimensional division algebras. Linear Algebra and its Applications, 493(??):399–410, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500717X.

Luo:2018:GWS

[LT18] Xin Luo and Tatsuya Tate. Up and down Grover walks on simplicial complexes. Linear Algebra and its Applications, 545(??):174–206, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518300508.

Lin:2019:TPI

[LT19a] Bo Lin and Ngoc Mai Tran. Two-player incentive compatible outcome functions are affine maximizers. Linear Algebra and its Applications, 578(??):133–152, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519301855.

Liu:2019:RPM

[LT19b] Yue Liu and Lidan Tang. Ray pattern matrices requiring Perron–Frobenius properties. Linear Algebra and its Applications, 571(??): 103–109, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300886.

Li:2015:SPO

[LTWW15] Chi-Kwong Li, Ming-Cheng Tsai, Kuo-Zhong Wang, and Ngai-Ching Wong. The spectrum of the product of operators, and the product of their numerical ranges. Linear Algebra and its Applications, 469(??):487–499, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007599.

Li:2019:ADT

[LUC19] Jianze Li, Konstantin Usevich, and Pierre Comon. On approximate diagonalization of third order symmetric tensors by orthogonal transformations. Linear Algebra and its Applications, 576(??): 324–351, September 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301004.

Liu:2015:FRM

[LW15] Jinwang Liu and Mingsheng Wang. Further remarks on multivariate polynomial matrix factorizations. Linear Algebra and its Applications, 465(??):204–213, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006120.

Li:2018:SCM

[LWC18] Yuan Li, Shuaijie Wang, and Xiaomei Cai. Some characterizations of minimal compact normal operators. Linear Algebra and its Applications, 549 (??):233–245, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301447.

Liu:2019:SRC

[LWCL19] Jianping Liu, Xianzhang Wu, Jinsong Chen, and Bolian Liu. The A_{α} spectral radius characterization of some digraphs. Linear Algebra and its Applications, 563(??):63–74, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304993.

Lin:2016:DDL

[LWCS16] Huiqiu Lin, Baoyindureng Wu, Yingying Chen, and Jinlong Shu. On the distance and distance Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 492(??):128–135, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006771.

Lin:2016:BLS

[LwW16] Fan-Hsuan Lin and Chih wen Weng. A bound on the Laplacian spread which is tight for strongly regular graphs. Linear Algebra and its Applications, 494 (??):11–22, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007284.

Lu:2017:CUG

[LWX17] Yong Lu, Ligong Wang, and Peng Xiao. Complex unit gain bicyclic graphs with rank 2, 3 or 4. Linear Algebra and its Applications, 523 (??):169–186, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301325.

Li:2016:NRP

[LWY16] Hanyu Li, Yimin Wei, and Yanfei Yang. New rigorous perturbation bounds for the Cholesky-like factorization of skew-symmetric matrix. Linear Algebra and its Applications, 491(??): 83–100, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001391.

Lu:2018:RSG

[LWZ18] Yong Lu, Ligong Wang, and Qiannan Zhou. The rank of a signed graph in terms of the rank of its underlying graph. Linear Algebra and its Applications, 538(??): 166–186, February 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730592X.

Li:2017:FTC

[LX17] Haohao Li and Mengxue Xia. Farkas-type conditions of general interval linear systems for AE solvability. Linear Algebra and its Applications, 514(??): 208–221, February 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305225.

Lin:2018:SG

[LXS18] Huiqiu Lin, Jie Xue, and Jinlong Shu. On the A_{α}- spectra of graphs. Linear Algebra and its Applications, 556(??):210–219, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303215.

Lee:2016:RPV

[LY16] Sang Hoon Lee and Jasang Yoon. Recursiveness and propagation for 2-variable weighted shifts. Linear Algebra and its Applications, 504(??):228–247, September 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951630091X.

Liu:2017:PCA

[LY17] Cheng-Kai Liu and Jheng-Jie Yang. Power commuting additive maps on invertible or singular matrices. Linear Algebra and its Applications, 530(??): 127–149, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302872.

Lynch:2015:CPC

[Lyn15] Alison Gordon Lynch. Cauchy pairs and Cauchy matrices. Linear Algebra and its Applications, 471(??): 320–345, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000154.

Li:2015:SPV

[LZ15a] Chi-Kwong Li and Shixiao Zhang. Stationary probability vectors of higher-order Markov chains. Linear Algebra and its Applications, 473 (??):114–125, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514001876.

Li:2015:CGM

[LZ15b] Jianping Li and Jianbin Zhang. Characterizing graphs with maximal Laplacian Estrada index. Linear Algebra and its Applications, 465(??):312–324, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006144.

Lin:2015:DLS

[LZ15c] Hongying Lin and Bo Zhou. On the distance Laplacian spectral radius of graphs. Linear Algebra and its Applications, 475(??):265–275, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515001408.

Liu:2015:QUC

[LZ15d] Xiaogang Liu and Sanming Zhou. Quadratic unitary Cayley graphs of finite commutative rings. Linear Algebra and its Applications, 479 (??):73–90, August 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500213X.

Lu:2015:FDS

[LZ15e] Rencai Lü and Kaiming Zhao. Finite-dimensional simple modules over generalized Heisenberg algebras. Linear Algebra and its Applications, 475(??): 276–291, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001299.

Lin:2016:DSR

[LZ16a] Hongying Lin and Bo Zhou. Distance spectral radius of uniform hypergraphs. Linear Algebra and its Applications, 506(??):564–578, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302282.

Lin:2016:EGT

[LZ16b] Hongying Lin and Bo Zhou. The effect of graft transformations on distance signless Laplacian spectral radius. Linear Algebra and its Applications, 504(??): 433–461, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301197.

Liu:2016:CIB

[LZ16c] Genqiang Liu and Kaiming Zhao. Classification of irreducible bounded weight modules over the derivation Lie algebras of quantum tori. Linear Algebra and its Applications, 495 (??):11–23, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000112.

Lin:2017:SRU

[LZ17a] Hongying Lin and Bo Zhou. Spectral radius of uniform hypergraphs. Linear Algebra and its Applications, 527 (??):32–52, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302288.

Lin:2017:URT

[LZ17b] Minghua Lin and Pingping Zhang. Unifying a result of Thompson and a result of Fiedler and Markham on block positive definite matrices. Linear Algebra and its Applications, 533(??):380– 385, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304524.

Liu:2017:LLD

[LZ17c] Dan Liu and Jianhua Zhang. Local Lie derivations of factor von Neumann algebras. Linear Algebra and its Applications, 519(??): 208–218, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300174.

Li:2018:PES

[LZB18] Haifeng Li, Jiang Zhou, and Changjiang Bu. Principal eigenvectors and spectral radii of uniform hypergraphs. Linear Algebra and its Applications, 544 (??):273–285, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300247.

Li:2016:SNR

[LZH16] Yuan Li, Mei Zheng, and Shasha Hu. Some new results for Hua-type operator matrices. Linear Algebra and its Applications, 506(??): 212–225, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Lin:2016:UBH

[LZM16] Hongying Lin, Bo Zhou, and Biao Mo. Upper bounds for H- and Z-spectral radii of uniform hypergraphs. Linear Algebra and its Applications, 510(??):205–221, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303366.

Leng:2019:DWN

[LZQ19] Tuo Leng, Qingyu Zhao, and Xiaolin Qin. On discrete Wirtinger-Northcott problems. Linear Algebra and its Applications, 575(??): 141–158, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301612.

Liu:2017:SCB

[LZSD17] Muhuo Liu, Yanli Zhu, Haiying Shan, and Kinkar Ch. Das. The spectral characterization of butterfly-like graphs. Linear Algebra and its Applications, 513(??): 55–68, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304633.

Liu:2018:CSS

[LZSD18] Muhuo Liu, Yanli Zhu, Haiying Shan, and Kinkar Ch. Das. Corrigendum to “The spectral characterization of butterfly-like graphs” [Linear Algebra Appl. 513 (2017) 55–68]. Linear Algebra and its Applications, 539(??):274–276, February 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517306274.

Lin:2017:DSL

[LZW17] Hongying Lin, Bo Zhou, and Yanna Wang. Distance (signless) Laplacian spectral radius of uniform hypergraphs. Linear Algebra and its Applications, 529(??):271–293, September 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517302574.

Ma:2015:BES

[Ma15] Wei Ma. A backward error for the symmetric generalized inverse eigenvalue problem. Linear Algebra and its Applications, 464 (??):90–99, January 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379513007246.

Ma:2016:DSI

[Ma16] Jianmin Ma. The distance signatures of the incidence graphs of affine resolvable designs. Linear Algebra and its Applications, 493 (??):37–44, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007132.

MacCaig:2016:SIM

[Mac16] M. MacCaig. On scaling to an integer matrix and graphs with integer weighted cycles. Linear Algebra and its Applications, 498(??):490– 520, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000203.

Mackenzie:2018:AOM

[Mac18] Michael Mackenzie. Another orthogonal matrix, revisited. Linear Algebra and its Applications, 555(??): 355–360, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303136.

Magnusson:2016:IPE

[Mag16] Gunnar Tór Magnússon. The inner product on exterior powers of a complex vector space. Linear Algebra and its Applications, 504(??):372–386, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301136.

Manuilov:2019:AUL

[Man19]  V. Manuilov.  Approxi-

mately uniformly locally finite graphs. Linear Algebra and its Applications, 569 (??):146–155, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300382.

Martin:2015:SWT

[Mar15a] Antonio J. Calderón Martín. A second Wedderburn-type theorem for some classes of linearly structured matrices. Linear Algebra and its Applications, 481(??):249– 262, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002980.

Marx:2015:DTS

[Mar15b] Benoît Marx. A descriptor Takagi–Sugeno approach to nonlinear model reduction. Linear Algebra and its Applications, 479(??): 52–72, August 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002062.

Maroulas:2016:FHM

[Mar16] John Maroulas. Factorization of Hessenberg matrices. Linear Algebra and its Applications, 506(??): 226–243, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301987.

Manzaneda:2017:RLS

[MAR17] Cristina Manzaneda, Enide Andrade, and María Robbiano. Realizable lists via the spectra of structured matrices. Linear Algebra and its Applications, 534(??): 51–72, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304731.

Martin:2018:DLS

[Mar18] Antonio J. Calderón Martín. Decompositions of linear spaces induced by bilinear maps. Linear Algebra and its Applications, 542 (??):209–224, ???? 2018.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302665.

Masuyama:2017:CTB

[Mas17] Hiroyuki Masuyama. Continuoustime block-monotone Markov chains and their blockaugmented truncations. Linear Algebra and its Applications, 514(??):105–150, February 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304876.

Mathai:2015:FDO

[Mat15a] A. M. Mathai. Fractional differential operators in the complex matrix-variate case. Linear Algebra and its Applications, 478(??):200–217, August 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002141.

Matsuki:2015:CPR

[Mat15b] Norichika Matsuki. Counting problems and ranks of matrices. Linear Algebra and its Applications, 465(??): 104–106, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Mattila:2015:ECM

[Mat15c] Mika Mattila. On the eigenvalues of combined meet and join matrices. Linear Algebra and its Applications, 466 (??):1–20, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400648X.

Mayer:2015:TSD

[May15] Günter Mayer. Three short descriptions of the symmetric and of the skewsymmetric solution set. Linear Algebra and its Applications, 475(??):73–79, June 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515000725.

Maxwell:2015:PKH

[MB15] Andrew S. Maxwell and Stephen Brierley. On properties of Karlsson Hadamards and sets of mutually unbiased bases in dimension six. Linear Algebra and its Applications, 466(??): 296–306, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006867.

Mukherjee:2018:SEP

[MB18] Sajal Kumar Mukherjee and Sudip Bera. A simple elementary proof of “The Unimodularity Theorem” of Oliver Knill. Linear Algebra and its Applications, 547 (??):124–127, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300831.

Mendez:2018:BPF

[MBB^{+}18] Robert P. Mendez, Bernhard G. Bodmann, Zachery J. Baker, Micah G. Bullock, and Jacob E. McLaney. Binary Parseval frames from group orbits. Linear Algebra and its Applications, 556(??): 265–300, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303422.

Martinhao:2019:NSC

[MC19] A. N. Martinhão and E. L. Monte Carmelo. New short coverings of F_{q}^{3} from pairwise weakly linearly independent sets. Linear Algebra and its Applications, 567 (??):1–13, April 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305950.

McKee:2018:CTL

[McK18] James McKee. A contraction theorem for the largest eigenvalue of a multigraph. Linear Algebra and its Applications, 557(??):339–349, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303665.

Mroz:2016:PBL

[MdlP16] Andrzej Mróz and José Antonio de la Peña. Periodicity in bilinear lattices and the Coxeter formalism. Linear Algebra and its Applications, 493(??): 227–260, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006953.

Movassagh:2015:CNI

[ME15] Ramis Movassagh and Alan Edelman. Condition numbers of indefinite rank 2 ghost Wishart matrices. Linear Algebra and its Applications, 483(??):342–351, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003468.

Melman:2015:NMP

[Mel15] A. Melman. Nonscalar matrix polynomial representation of some scalar polynomials. Linear Algebra and its Applications, 474 (??):141–157, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500124X.

Melman:2016:BEM

[Mel16a] A. Melman. Bounds for eigenvalues of matrix polynomials with applications to scalar polynomials. Linear Algebra and its Applications, 504(??):190–203, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300908.

Melman:2016:CLP

[Mel16b] A. Melman. Cauchy-like and Pellet-like results for polynomials. Linear Algebra and its Applications, 505(??):174– 193, September 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301483.

Melman:2019:ELU

[Mel19]  A. Melman.  Eigenvalue

localization under partial spectral information. Linear Algebra and its Applications, 573(??):12–25, July 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301119.

Mousavi:2017:OEH

[MEMM17] Z. Mousavi, R. Eskandari, M. S. Moslehian, and F. Mirzapour. Operator equations AX + Y B = C and AXA^{∗} + BY B^{∗} = C in Hilbert C^{∗}-modules. Linear Algebra and its Applications, 517(??):85–98, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305808.

Meshulam:2017:MRM

[Mes17] Roy Meshulam. Maximal rank in matrix spaces via graph matchings. Linear Algebra and its Applications, 529(??):1–11, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517302410.

Moslehian:2016:SOP

[MF16] Mohammad Sal Moslehian and Xiaohui Fu. Squaring operator Pólya–Szegő and Diaz–Metcalf type inequalities. Linear Algebra and its Applications, 491(??): 73–82, February 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001160.

Marti-Farre:2015:CDC

[MFdM15] Jaume Martí-Farré and Anna de Mier. Completion and decomposition of a clutter into representable matroids. Linear Algebra and its Applications, 472 (??):31–47, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500052X.

Ma:2018:SPG

[MG18a] Xiaobin Ma and Xianya Geng. Signature of power graphs. Linear Algebra and its Applications, 545 (??):139–147, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300521.

Mira:2018:TH

[MG18b] José Antonio Cuenca Mira and Francisca Miguel García. On a Theorem by Hopf. Linear Algebra and its Applications, 552(??):42–51, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301940.

Moghaddam:2017:IMF

[MGM17] Mahsa R. Moghaddam, Kazem Ghanbari, and Angelo B. Mingarelli. Isospectral matrix flow maintaining staircase structure and total positivity of an initial matrix. Linear Algebra and its Applications, 517(??): 134–147, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305882.

Matejas:2015:HRA

[MH15] Josip Matejas and Vjeran Hari. On high relative accuracy of the Kogbetliantz method. Linear Algebra and its Applications, 464(??): 100–129, January 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514000913.

Massart:2018:MGM

[MHA18] Estelle M. Massart, Julien M. Hendrickx, and P.-A. Absil. Matrix geometric means based on shuffled inductive sequences. Linear Algebra and its Applications, 542 (??):334–359, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303361.

Ma:2015:SCU

[MHL15] Xiaoling Ma, Qiongxiang Huang, and Fenjin Liu. Spectral characterization of unicyclic graphs whose second largest eigenvalue does not exceed 1. Linear Algebra and its Applications, 471 (??):587–603, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951400799X.

Muller-Hermes:2018:OSR

[MHN18] Alexander Müller-Hermes and Ion Nechita. Operator Schmidt ranks of bipartite unitary matrices. Linear Algebra and its Applications, 557(??):174–187, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303446.

Muller-Hermes:2015:SVB

[MHS15] Alexander Müller-Hermes and Oleg Szehr. Spectral variation bounds in hyperbolic geometry. Linear Algebra and its Applications, 482(??):131–148, October 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003298.

Martin:2016:AAM

[MI16] Antonio J. Calderón Martín and Francisco J. Navarro Izquierdo. Arbitrary algebras with a multiplicative basis. Linear Algebra and its Applications, 498 (??):106–116, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000506.

Martin:2017:CSM

[MI17] Antonio J. Calderón Martín and Francisco J. Navarro Izquierdo. A class of simple matrix algebras. Linear Algebra and its Applications, 528(??):159–170, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302543.

Michael:2016:SPN

[Mic16] T. S. Michael. A short proof of the nonexistence of a 2-resolvable balanced incomplete block design with parameters (v, b, r, k, λ) = (10, 15, 6, 4, 2). Linear Algebra and its Applications, 498 (??):399–403, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006114.

Miguel:2016:DCG

[Mig16] C. Miguel. On the diameter of the commuting graph of the matrix ring over a centrally finite division ring. Linear Algebra and its Applications, 509(??):276– 285, November 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303287.

Minh:2017:IDL

[Min17] Hà Quang Minh. Infinitedimensional Log-Determinant divergences between positive definite trace class operators. Linear Algebra and its Applications, 528(??):331– 383, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304177.

Mirafzal:2018:NCI

[Mir18] S. Morteza Mirafzal. A new class of integral graphs constructed from the hypercube. Linear Algebra and its Applications, 558(??): 186–194, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Miyajima:2015:FEM

[Miy15] Shinya Miyajima. Fast enclosure for a matrix inverse square root. Linear Algebra and its Applications, 467(??):116–135, February 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007319.

Miyajima:2019:VCM

[Miy19] Shinya Miyajima. Verified computation for the matrix principal logarithm. Linear Algebra and its Applications, 569(??):38–61, May 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519300229.

Mizutani:2016:RAP

[Miz16] Tomohiko Mizutani. Robustness analysis of preconditioned successive projection algorithm for general form of separable NMF problem. Linear Algebra and its Applications, 497 (??):1–22, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951600118X.

Monfared:2018:SIS

[MK18] Keivan Hassani Monfared and Ehssan Khanmohammadi. A structured inverse spectrum problem for infinite graphs. Linear Algebra and its Applications, 539(??):28–43, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306146.

Mao:2015:NMC

[MLW15] Lihuan Mao, Fenjin Liu, and Wei Wang. A new method for constructing graphs determined by their generalized spectrum. Linear Algebra and its Applications, 477 (??):112–127, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001901.

Manivel:2015:SMC

[MM15a] Laurent Manivel and Mateusz Michalek. Secants of minuscule and cominuscule minimal orbits. Linear Algebra and its Applications, 481(??):288–312, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002785.

Monfared:2015:CRS

[MM15b] Keivan Hassani Monfared and Sudipta Mallik. Construction of real skewsymmetric matrices from interlaced spectral data, and graph. Linear Algebra and its Applications, 471(??): 241–263, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000269.

Marco:2016:BDR

[MM16a] Ana Marco and José-Javier Martínez. Bidiagonal decomposition of rectangular totally positive Said– Ball–Vandermonde matrices: Error analysis, perturbation theory and applications. Linear Algebra and its Applications, 495 (??):90–107, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000525.

Monfared:2016:SCM

[MM16b] Keivan Hassani Monfared and Sudipta Mallik. Spectral characterization of matchings in graphs. Linear Algebra and its Applications, 496 (??):407–419, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000951.

MacLean:2017:BDR

[MM17a] Mark S. MacLean and Stefko Miklavic. On bipartite distance-regular graphs with exactly two irreducible T-modules with endpoint two. Linear Algebra and its Applications, 515(??): 275–297, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630547X.

McDonald:2017:SAZ

[MM17b] Judith J. McDonald and Timothy C. Melvin. Spectrally arbitrary zero-nonzero patterns and field extensions. Linear Algebra and its Applications, 519(??): 146–155, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306395.

Michalek:2017:ERM

[MM17c] Mateusz Michalek and Chris Miller. Examples of kregular maps and interpolation spaces. Linear Algebra and its Applications, 530(??):94–108, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

Makuracki:2019:RSI

[MM19a] Bartosz Makuracki and Andrzej Mróz. Root systems and inflations of nonnegative quasi-Cartan matrices. Linear Algebra and its Applications, 580(??): 128–165, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302587.

Monfared:2019:AMT

[MM19b] Keivan Hassani Monfared and Sudipta Mallik. An analog of Matrix Tree Theorem for signless Laplacians. Linear Algebra and its Applications, 560(??): 43–55, January 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304518.

Milovanovic:2019:NLS

[MMM19] E. I. Milovanović, M. M. Matejić, and I. Z. Milovanović. On the normalized Laplacian spectral radius, Laplacian incidence energy and Kemeny’s constant. Linear Algebra and its Applications, 582(??): 181–196, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303386.

MacLean:2016:TAB

[MMP16] Mark S. MacLean, Štefko Miklavič, and Safet Penjić. On the Terwilliger algebra of bipartite distance-regular graphs with Δ_{2} = 0 and c_{2} = 1. Linear Algebra and its Applications, 496 (??):307–330, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000550.

Marco:2017:ABD

[MMP17] Ana Marco, José-Javier Martínez, and Juan Manuel Peña. Accurate bidiagonal decomposition of totally positive Cauchy–Vandermonde matrices and applications. Linear Algebra and its Applications, 517(??):63–84, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305821.

Mingueza:2016:CSL

[MMR16] David Mingueza, M. Eulàlia Montoro, and Alicia Roca. The characteristic subspace lattice of a linear transformation. Linear Algebra and its Applications, 506(??): 329–341, October 1, 2016.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302208.

Mingueza:2017:CCL

[MMR17] David Mingueza, M. Eulàlia Montoro, and Alicia Roca. Computing the cardinality of the lattice of characteristic subspaces. Linear Algebra and its Applications, 514(??):82–104, February 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630516X.

Mingueza:2018:LCS

[MMR18] David Mingueza, M. Eulàlia Montoro, and Alicia Roca. The lattice of characteristic subspaces of an endomorphism with Jordan– Chevalley decomposition. Linear Algebra and its Applications, 558(??):63–73, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303781.

Majdak:2016:OII

[MMS16] Witold Majdak, Mostafa Mbekhta, and Laurian Suciu. Operators intertwining with isometries and Brownian parts of 2-isometries.

Linear Algebra and its Applications, 509(??):168–190, November 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302762.

Mitsuhashi:2018:WKI

[MMS18] Hideo Mitsuhashi, Hideaki Morita, and Iwao Sato. The weighted Kirchhoff index of a graph. Linear Algebra and its Applications, 547 (??):1–18, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830051X.

Marco:2019:ABD

[MMV19] Ana Marco, José-Javier Martínez, and Raquel Viaña. Accurate bidiagonal decomposition of totally positive h-Bernstein–Vandermonde matrices and applications. Linear Algebra and its Applications, 579(??):320–335, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302551.

Mu:2017:RPM

[MMW17] Lili Mu, Jianxi Mao, and Yi Wang. Row polynomial matrices of Riordan arrays. Linear Algebra and its Applications, 522(??):1– 14, June 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300824.

Manjegani:2015:MFI

[MN15] S. M. Manjegani and A. Norouzi. Matrix form of the inverse Young inequalities. Linear Algebra and its Applications, 486(??):484–493, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004905.

Medina:2019:RLS

[MN19a] Luis Medina and Hans Nina. Realizable lists via the spectra of block matrices. Linear Algebra and its Applications, 581(??): 1–19, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302885.

Merlini:2019:FJC

[MN19b] Donatella Merlini and Massimo Nocentini. Functions and Jordan canonical forms of Riordan matrices. Linear Algebra and its Applications, 565(??): 177–207, March 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305743.

Merlevede:2019:ULE

[MNT19] Florence Merlevède, Jamal Najim, and Peng Tian. Unbounded largest eigenvalue of large sample covariance matrices: Asymptotics, fluctuations and applications. Linear Algebra and its Applications, 577(??):317–359, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302034.

Mehrmann:2016:SCH

[MNTX16] Volker Mehrmann, Vanni Noferini, Françoise Tisseur, and Hongguo Xu. On the sign characteristics of Hermitian matrix polynomials. Linear Algebra and its Applications, 511(??):328–364, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303913.

Ma:2016:IJP

[MO16] Alexander Ma and Jamie Oliva. On the images of Jordan polynomials evaluated over symmetric matrices. Linear Algebra and its Applications, 492(??):13– 25, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006783.

Mikhalev:2018:RMV

[MO18] A. Mikhalev and I. V. Oseledets. Rectangular maximum-volume submatrices and their applications. Linear Algebra and its Applications, 538(??):187–211, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305931.

Mohar:2016:HAS

[Moh16]  Bojan Mohar.  Hermi-

tian adjacency spectrum and switching equivalence of mixed graphs. Linear Algebra and its Applications, 489(??):324–340, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006138.

Molitierno:2018:TUB

[Mol18a] Jason J. Molitierno. A tight upper bound on the spectral radius of bottleneck matrices for graphs. Linear Algebra and its Applications, 551 (??):1–17, August 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301630.

Molnar:2018:BIB

[Mol18b] Lajos Molnár. Bures isometries between density spaces of C^{∗}-algebras. Linear Algebra and its Applications, 557(??):22–33, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518303264.

Molnar:2019:CCM

[Mol19] Lajos Molnár. Characterizations of certain means of positive operators. Linear Algebra and its Applications, 567 (??):143–166, April 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930014X.

Monfared:2017:ENS

[Mon17] Keivan Hassani Monfared. Existence of a not necessarily symmetric matrix with given distinct eigenvalues and graph. Linear Algebra and its Applications, 527 (??):1–11, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730229X.

Montaner:2019:GFA

[Mon19] Josep Àlvarez Montaner. Generating functions associated to Frobenius algebras. Linear Algebra and its Applications, 582(??): 310–326, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303428.

Mastnak:2016:NIS

[MOR16a] Mitja Mastnak, Matjaz Omladic, and Heydar Radjavi. Near-invariant subspaces for matrix groups are nearly invariant. Linear Algebra and its Applications, 505(??):269–281, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301549.

Morales:2016:LAS

[Mor16b] John Vincent S. Morales. On Lee association schemes over Z_{4} and their Terwilliger algebra. Linear Algebra and its Applications, 510(??): 311–328, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303858.

Morris:2017:GPL

[Mor17] Ian D. Morris. Generic properties of the lower spectral radius for some low-rank pairs of matrices. Linear Algebra and its Applications, 524(??):35–60, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301222.

Mosic:2017:WPO

[Mos17] Dijana Mosić. Weighted preorders in a Banach algebra. Linear Algebra and its Applications, 533(??):161– 185, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304470.

Muller:2015:SMA

[MP15a] Vladimir Müller and Aljosa Peperko. On the spectrum in max algebra. Linear Algebra and its Applications, 485(??):250–266, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004139.

Myskova:2015:WRI

[MP15b] Helena Mysková and Ján Plavka. On the weak robustness of interval fuzzy matrices. Linear Algebra and its Applications, 474 (??):243–259, June 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001329.

Mackey:2016:LMP

[MP16a] D. Steven Mackey and Vasilije Perović. Linearizations of matrix polynomials in Bernstein bases. Linear Algebra and its Applications, 501 (??):162–197, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300052.

McDonald:2016:JCC

[MP16b] Judith J. McDonald and Pietro Paparella. Jordan chains of h-cyclic matrices. Linear Algebra and its Applications, 498(??):145– 159, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001366.

McDonald:2016:MRI

[MP16c] Judith J. McDonald and Pietro Paparella. Matrix roots of imprimitive irreducible nonnegative matrices. Linear Algebra and its Applications, 498(??):244– 261, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003584.

Mathai:2017:ARA

[MP17] A. M. Mathai and T. Princy. Analogues of reliability analysis for matrix-variate cases. Linear Algebra and its Applications, 532(??):287–311, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303853.

Meerbergen:2018:MFB

[MP18] Karl Meerbergen and Javier Pérez. Mixed forwardbackward stability of the two-level orthogonal Arnoldi method for quadratic problems. Linear Algebra and its Applications, 553(??):1– 15, September 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302143.

Myskova:2019:REI

[MP19] H. Mysková and J. Plavka. Reachability of eigenspaces for interval matrices in maxmin algebra. Linear Algebra and its Applications, 578(??):314–333, October 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302277.

Marijuan:2017:MSC

[MPS17] C. Marijuán, M. Pisonero, and Ricardo L. Soto. A map of sufficient conditions for the symmetric nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 530(??): 344–365, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303117.

Molnr:2016:MPD

[MPV16] Lajos Molnár, József Pitrik, and Dániel Virosztek. Maps on positive definite matrices preserving Bregman and Jensen divergences. Linear Algebra and its Applications, 495(??):174–189, April 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000124.

Ma:2018:GSD

[MQTW18] Xiaobin Ma, Linming Qi, Fenglei Tian, and Dein Wong. Graphs with some distance Laplacian eigenvalue of multiplicity n −

3. Linear Algebra and its Applications, 557(??):307– 326, November 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303690.

Marciniak:2017:MPM

[MR17a] Marcin Marciniak and Adam Rutkowski. Merging of positive maps: a construction of various classes of positive maps on matrix algebras. Linear Algebra and its Applications, 529(??):215– 257, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730263X.

Massei:2017:DBN

[MR17b] Stefano Massei and Leonardo Robol. Decay bounds for the numerical quasiseparable preservation in matrix functions. Linear Algebra and its Applications, 516 (??):212–242, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305791.

Mezo:2019:CAD

[MR19] István Mező and José L. Ramírez. A combinatorial approach to derangement matrix of type B. Linear Algebra and its Applications, 582(??):156–180, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303374.

Mroz:2016:CEG

[Mró16]  Andrzej Mróz.  Coxeter

energy of graphs. Linear Algebra and its Applications, 506(??):279–307, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302129.

Massey:2015:MLI

[MRS15] Pedro G. Massey, Mariano A. Ruiz, and Demetrio Stojanoff. Multiplicative Lidskii’s inequalities and optimal perturbations of frames. Linear Algebra and its Applications, 469(??): 539–568, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007940.

Massey:2018:LLT

[MRS18] Pedro G. Massey, Noelia B. Rios, and Demetrio Stojanoff. Local Lidskii’s theorems for unitarily invariant norms. Linear Algebra and its Applications, 557(??): 34–61, November 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303501.

Micheli:2015:LSS

[MRV15] G. Micheli, J. Rosenthal, and P. Vettori. Linear spanning sets for matrix spaces. Linear Algebra and its Applications, 483(??):309–322, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003614.

Marcellan:2015:OCM

[MS15a] Francisco Marcellán and Nikta Shayanfar. OPUC, CMV matrices and perturbations of measures supported on the unit circle. Linear Algebra and its Applications, 485(??):305–344, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004401.

Molnar:2015:TPD

[MS15b] Lajos Molnár and Patrícia Szokol. Transformations on positive definite matrices preserving generalized distance measures. Linear Algebra and its Applications, 466(??):141–159, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006429.

Mallik:2016:INM

[MS16a] Sudipta Mallik and Bryan L. Shader. Integrally normalizable matrices with respect to a given set. Linear Algebra and its Applications, 498 (??):317–325, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004012.

Mashreghi:2016:NMP

[MS16b] Javad Mashreghi and Anush Stepanyan. Nonlinear maps preserving the reduced minimum modulus of operators. Linear Algebra and its Applications, 493(??): 426–432, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007247.

Monfared:2016:NZE

[MS16c] Keivan Hassani Monfared and Bryan L. Shader. The nowhere-zero eigenbasis problem for a graph. Linear Algebra and its Applications, 505(??):296–312, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301276.

Mei:2017:IFD

[MS17a] Wei Mei and Yisheng Song. Infinite and finite dimensional generalized Hilbert tensors. Linear Algebra and its Applications, 532(??): 8–24, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303920.

Momihara:2017:UBS

[MS17b] Koji Momihara and Sho Suda. Upper bounds on the size of transitive subtournaments in digraphs. Linear Algebra and its Applications, 530(??):230–243, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302677.

Martino:2019:FGG

[MS19] Ivan Martino and Rahul Singh. Finite groups generated in low real codimension. Linear Algebra and its Applications, 570 (??):245–281, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300175.

Meleiro:2017:CLM

[MSSZ17] Juan Meleiro, Vladimir V. Sergeichuk, Thiago Solovera, and André Zaidan. Classification of linear mappings between indefinite inner product spaces. Linear Algebra and its Applications, 531(??):356–374, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303658.

Munemasa:2015:SEL

[MST15] Akihiro Munemasa, Yoshio Sano, and Tetsuji Taniguchi. On the smallest eigenvalues of the line graphs of some trees. Linear Algebra and its Applications, 466(??): 501–511, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007162.

Morton:2015:GCC

[MT15] Jason Morton and Jacob Turner. Generalized counting constraint satisfaction problems with determinantal circuits. Linear Algebra and its Applications, 466(??):357–381, February 1, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379514006478.

Morrison:2016:TCV

[MT16] Ralph Morrison and Ngoc M. Tran. The tropical commuting variety. Linear Algebra and its Applications, 507(??):300–321, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302142.

Mastronardi:2016:RRD

[MV16] Nicola Mastronardi and Paul Van Dooren. Rank-revealing decomposition of symmetric indefinite matrices via block anti-triangular factorization. Linear Algebra and its Applications, 502(??): 126–139, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005649.

Macias-Virgos:2019:RSV

[MVPST19] E. Macías-Virgós, M. J. Pereira-Sáez, and Daniel Tanré. Relative singular value decomposition and applications to LS-category. Linear Algebra and its Applications, 582(??):58–75, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Mao:2015:CGD

[MW15] Lihuan Mao and Wei Wang. On the construction of graphs determined by their generalized characteristic polynomials. Linear Algebra and its Applications, 485(??):454–466, November 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515004474.

Ma:2016:CGW

[MWT16a] Xiaobin Ma, Dein Wong, and Fenglei Tian. Characterization of graphs whose signature equals the number of odd cycles. Linear Algebra and its Applications, 511(??):259–273, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304165.

Ma:2016:SRO

[MWT16b] Xiaobin Ma, Dein Wong, and Fenglei Tian. Skewrank of an oriented graph in terms of matching number. Linear Algebra and its Applications, 495(??): 242–255, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000513.

Muzychuk:2016:TAW

[MX16] Mikhail Muzychuk and Bangteng Xu. Terwilliger algebras of wreath products of association schemes. Linear Algebra and its Applications, 493(??):146–163, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007041.

Ma:2016:PFN

[MY16] Jun Ma and Yeong-Nan Yeh. Parking functions on nonsingular M-matrices. Linear Algebra and its Applications, 489(??):1–14, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515005923.

Munshi:2019:SAE

[MY19] Sachin Munshi and Rongwei Yang. Self-adjoint elements in the pseudo-unitary group U(p, p). Linear Algebra and its Applications, 560(??): 100–113, January 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304750.

Myskova:2016:IMP

[Mys16]  Helena Mysková.  Inter-

val max-plus matrix equations. Linear Algebra and its Applications, 492(??): 111–127, March 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006606.

Nadoshan:2015:MLP

[NA15] Mohammad Ali Hadian Nadoshan and Ali Armandnejad. B-majorization and its linear preservers. Linear Algebra and its Applications, 478(??):218–227, August 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002128.

Nikiforov:2017:MTN

[NA17] V. Nikiforov and N. Agudelo. On the minimum trace norm/energy of (0, 1)-matrices. Linear Algebra and its Applications, 526(??):42–59, August 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301866.

Nagy:2015:ISS

[Nag15] Gergö Nagy. Isometries of the spaces of self-adjoint traceless operators. Linear Algebra and its Applications, 484(??):1–12, November 1, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515003791.

Nagar:2019:EML

[Nag19] Mukesh Kumar Nagar. Eigenvalue monotonicity of q-Laplacians of trees along a poset. Linear Algebra and its Applications, 571 (??):110–131, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300801.

Najafi:2017:MOM

[Naj17] Hamed Najafi. More on operator monotone and operator convex functions of several variables. Linear Algebra and its Applications, 532(??):127–139, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730397X.

Neyshaburi:2018:EPI

[NAS18] Fahimeh Arabyani Neyshaburi, Ali Akbar Arefijamaal, and Ghadir Sadeghi. Extreme points and identification of optimal alternate dual frames. Linear Algebra and its Applications, 549 (??):123–135, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301319.

Nathanson:2019:MSE

[Nat19] Melvyn B. Nathanson. Matrix scaling and explicit doubly stochastic limits. Linear Algebra and its Applications, 578(??):111–132, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S002437951930206X.

Nevanlinna:2018:NHF

[Nev18] Olavi Nevanlinna. A nonholomorphic functional calculus and the complex conjugate of a matrix. Linear Algebra and its Applications, 537(??):191–220, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305736.

Nguyen:2018:CDW

[Ngu18] Hoi H. Nguyen. Concentration of distances in Wigner matrices. Linear Algebra and its Applications, 536 (??):52–86, January 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304974.

Niezgoda:2015:MRJ

[Nie15a] Marek Niezgoda. Majorization and refined Jensen– Mercer type inequalities for self-adjoint operators. Linear Algebra and its Applications, 467(??):1–14, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379514007198.

Niezgoda:2015:SLI

[Nie15b]  Marek Niezgoda.  Shan-

non like inequalities for fconnections of positive linear maps and positive operators. Linear Algebra and its Applications, 481(??):186– 201, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002955.

Niezgoda:2018:ECP

[Nie18] Marek Niezgoda. Extended commutation principles for normal decomposition systems. Linear Algebra and its Applications, 539(??): 251–273, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Niezgoda:2019:ADR

[Nie19a] Marek Niezgoda. An analog of Drury’s result in normal decomposition systems. Linear Algebra and its Applications, 567(??): 118–137, April 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300205.

Niezgoda:2019:ESO

[Nie19b] Marek Niezgoda. An extension of Schur–Ostrowski’s condition, weak Eaton triples and generalized AI functions. Linear Algebra and its Applications, 580(??):212–235, November 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519302770.

Niezgoda:2019:SVB

[Nie19c] Marek Niezgoda. On singular values of block circulant matrices. Linear Algebra and its Applications, 563(??): 400–410, February 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305421.

Nikiforov:2015:EGE

[Nik15] Vladimir Nikiforov. Extrema of graph eigenvalues. Linear Algebra and its Applications, 482(??): 158–190, October 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003250.

Nikiforov:2016:BGE

[Nik16a] V. Nikiforov. Beyond graph energy: Norms of graphs and matrices. Linear Algebra and its Applications, 506(??): 82–138, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301616.

Nikiforov:2016:MCS

[Nik16b] V. Nikiforov. Max k-cut and the smallest eigenvalue. Linear Algebra and its Applications, 504(??):462–467, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301185.

Nikiforov:2016:RER

[Nik16c] V. Nikiforov. Remarks on the energy of regular graphs. Linear Algebra and its Applications, 508(??):133–145, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302695.

Nikiforov:2017:CMS

[Nik17a] V. Nikiforov. Combinatorial methods for the spectral p-norm of hypermatrices. Linear Algebra and its Applications, 529(??):324– 354, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302604.

Nikiforov:2017:HHS

[Nik17b] V. Nikiforov. Hypergraphs and hypermatrices with symmetric spectrum. Linear Algebra and its Applications, 519(??):1–18, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306371.

Nikiforov:2017:SRG

[Nik17c] V. Nikiforov. The spectral radius of graphs with no K_{2,t} minor. Linear Algebra and its Applications, 531(??): 510–515, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303774.

Nasr-Isfahani:2016:FDI

[NINS16] Rasoul Nasr-Isfahani, Mehdi Nemati, and Somayeh Shahmoradi. Finite dimensional invariant subspaces for algebras of linear operators and amenable Banach algebras. Linear Algebra and its Applications, 510(??): 329–345, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303809.

Niv:2015:PIM

[Niv15]  Adi Niv.  On pseudo-

inverses of matrices and their characteristic polynomials in supertropical algebra. Linear Algebra and its Applications, 471(??): 264–290, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000300.

Nobari:2018:GMT

[NK18] Elham Nobari and Bijan Ahmadi Kakavandi. A geometric mean for Toeplitz and Toeplitz-block block-Toeplitz matrices. Linear Algebra and its Applications, 548(??):189–202, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

Ning:2018:MSR

[NLW18] Wenjie Ning, Mei Lu, and Kun Wang. Maximizing the spectral radius of graphs with fixed minimum degree and edge connectivity. Linear Algebra and its Applications, 540(??): 138–148, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306493.

Ning:2018:SLS

[NLWJ18] Wenjie Ning, Mei Lu, Kun Wang, and Daqing Jiang. The signless Laplacian spectral radius of k-connected irregular graphs. Linear Algebra and its Applications, 553(??):117–128, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302349.

Nader:2019:NRI

[NMBA19] Rafic Nader, Bassam Mourad, Alain Bretto, and Hassan Abbas. A note on the real inverse spectral problem for doubly stochastic matrices. Linear Algebra and its Applications, 569 (??):206–240, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300412.

Nmeth:2016:IRI

[NN16] A. B. Németh and S. Z. Németh. Isotonic regression and isotonic projection. Linear Algebra and its Applications, 494(??):80– 89, April 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000082.

Nakatsukasa:2019:ILL

[NN19a] Yuji Nakatsukasa and Vanni Noferini. Inertia laws and localization of real eigenvalues for generalized indefinite eigenvalue problems. Linear Algebra and its Applications, 578(??):272–296, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302198.

Narayanan:2019:LSO

[NN19b] H. Narayanan and Hariharan Narayanan. On the linear static output feedback problem: the annihilating polynomial approach. Linear Algebra and its Applications, 579(??):336–364, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302575.

Nobari:2016:MGM

[Nob16] Elham Nobari. A monotone geometric mean for a class of Toeplitz matrices. Linear Algebra and its Applications, 511(??): 1–18, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303846.

Nomura:2015:LPHa

[Nom15a] Kazumasa Nomura. Leonard pairs having specified endentries. Linear Algebra and its Applications, 465(??): 43–64, January 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006090.

Nomura:2015:LPHb

[Nom15b] Kazumasa Nomura. Leonard pairs having zero-diagonal TD–TD form. Linear Algebra and its Applications, 478 (??):1–52, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001846.

Nomura:2015:LTT

[Nom15c] Kazumasa Nomura. Linear transformations that are tridiagonal with respect to the three decompositions for an LR triple. Linear Algebra and its Applications, 486(??):173–203, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004814.

Nomura:2016:PCE

[Nom16] Kazumasa Nomura. An LR pair that can be extended to an LR triple. Linear Algebra and its Applications, 493(??):336–357, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007235.

Nordstrom:2018:NCM

[Nor18]  Kenneth Nordström.  A

note on the convexity of the Moore–Penrose inverse. Linear Algebra and its Applications, 538(??):143–148, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305955.

Niazi:2015:WLD

[NP15a] Mohsen Niazi and Antonio M. Peralta. Weak-2-local ⋆-derivations on B(H) are linear ⋆-derivations. Linear Algebra and its Applications, 487(??):276–300, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005431.

Noferini:2015:DMP

[NP15b] Vanni Noferini and Federico Poloni. Duality of matrix pencils, Wong chains and linearizations. Linear Algebra and its Applications, 471 (??):730–767, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000440.

Nogueira:2015:SCS

[NP15c] Tony Nogueira and Daniel Pellegrino. On the size of certain subsets of invariant Banach sequence spaces. Linear Algebra and its Applications, 487(??):172–183, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005224.

Nikiforov:2017:ST

[NPRS17] Vladimir Nikiforov, Germain Pastén, Oscar Rojo, and Ricardo L. Soto. On the A_{α}-spectra of trees. Linear Algebra and its Applications, 520(??):286–305, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517300587.

Nikolayevsky:2016:DPO

[NR16a] Y. Nikolayevsky and Z. Rakić. The duality principle for Osserman algebraic curvature tensors. Linear Algebra and its Applications, 504(??): 574–580, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300775.

Noschese:2016:SMN

[NR16b] Silvia Noschese and Lothar Reichel. Some matrix nearness problems suggested by Tikhonov regularization. Linear Algebra and its Applications, 502(??):366–386, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002232.

Nikiforov:2017:NPS

[NR17] Vladimir Nikiforov and Oscar Rojo. A note on the positive semidefiniteness of A_{α}(G). Linear Algebra and its Applications, 519(??): 156–163, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300125.

Nikiforov:2018:IGP

[NR18] Vladimir Nikiforov and Oscar Rojo. On the αindex of graphs with pendent paths. Linear Algebra and its Applications, 550 (??):87–104, August 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301460.

Nelson:2015:MET

[NS15a] Curtis Nelson and Bryan Shader. A (0, 1)-matrix existence theorem and equivalent tiling problems with dimers and monomers. Linear Algebra and its Applications, 485(??):503–526, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004589.

Nelson:2015:APS

[NS15b] Curtis Nelson and Bryan Shader. All pairs suffice for a P -set. Linear Algebra and its Applications, 475 (??):114–118, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001196.

Nelson:2015:MSM

[NS15c] Curtis Nelson and Bryan Shader. Maximal P -sets of matrices whose graph is a tree. Linear Algebra and its Applications, 485(??):485– 502, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004590.

Nagar:2017:HIH

[NS17] Mukesh Kumar Nagar and Sivaramakrishnan Sivasubramanian. Hook immanantal and Hadamard inequalities for q-Laplacians of trees. Linear Algebra and its Applications, 523(??): 131–151, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300927.

Nastase:2018:CCM

[NS18] Esmeralda L. Nastase and Papa A. Sissokho. The complete characterization of the minimum size supertail in a subspace partition. Linear Algebra and its Applications, 559(??):172– 180, December 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304348.

Nagar:2019:LIP

[NS19] Mukesh Kumar Nagar and Sivaramakrishnan Sivasubramanian. Laplacian immanantal polynomials and the GTS poset on trees. Linear Algebra and its Applications, 561(??):1–23, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518304610.

Noutsos:2016:ESE

[NSCV16] D. Noutsos, S. Serra-Capizzano, and P. Vassalos. Essential spectral equivalence via multiple step preconditioning and applications to ill conditioned Toeplitz matrices. Linear Algebra and its Applications, 491(??):276–291, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004899.

Nomura:2017:UDT

[NT17] Kazumasa Nomura and Paul Terwilliger. The universal DAHA of type C_{1}^{∨}, C_{1} and Leonard pairs of q-Racah type. Linear Algebra and its Applications, 533(??): 14–83, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304342.

Nagaoka:2018:RPA

[NT18] Takahiro Nagaoka and Akiyoshi Tsuchiya. Reflexive polytopes arising from edge polytopes. Linear Algebra and its Applications, 557(??):438– 454, November 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303859.

Ngoan:2019:SLG

[NT19] Ngô Thı. Ngoan and Nguyêñ Quôć Thǎńg. On some localglobal principles for linear algebraic groups over infinite algebraic extensions of global fields. Linear Algebra and its Applications, 568(??):39–83, May 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302660.

Nunes:2017:NEC

[Nun17] Antonio Gomes Nunes. A new estimate for the constants of an inequality due to Hardy and Littlewood. Linear Algebra and its Applications, 526(??):27–34, August 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301854.

Nagisa:2015:OMS

[NW15] Masaru Nagisa and Shuhei Wada. Operator monotonicity of some functions. Linear Algebra and its Applications, 486(??):389–408, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004887.

Nicoara:2019:FRC

[NW19] Remus Nicoara and Chase Worley. A finiteness result for circulant core complex Hadamard matrices. Linear Algebra and its Applications, 571(??):143–153, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519300783.

Nikiforov:2015:MQI

[NY15] Vladimir Nikiforov and Xiying Yuan. Maxima of the Q-index: Forbidden even cycles. Linear Algebra and its Applications, 471(??): 636–653, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000488.

Needell:2015:RBK

[NZZ15] Deanna Needell, Ran Zhao, and Anastasios Zouzias. Randomized block Kaczmarz method with projection for solving least squares. Linear Algebra and its Applications, 484(??):322–343, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003808.

O:2016:ECR

[O16] Suil O. Edge-connectivity in regular multigraphs from eigenvalues. Linear Algebra and its Applications, 491(??):4–14, February 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379514006004.

Oberst:2016:TIW

[Obe16] Ulrich Oberst. Two invariants for weak exponential stability of linear timevarying differential behaviors. Linear Algebra and its Applications, 504(??):468– 486, September 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300945.

Oboudi:2016:TLE

[Obo16] Mohammad Reza Oboudi. On the third largest eigenvalue of graphs. Linear Algebra and its Applications, 503(??):164–179, August 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300489.

Oboudi:2019:DSR

[Obo19a] Mohammad Reza Oboudi. Distance spectral radius of complete multipartite graphs and majorization. Linear Algebra and its Applications, 583(??):134–145, December 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303684.

Oboudi:2019:NLB

[Obo19b] Mohammad Reza Oboudi. A new lower bound for the energy of graphs. Linear Algebra and its Applications, 580(??):384–395, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302782.

Oboudi:2019:RBS

[Obo19c] Mohammad Reza Oboudi. A relation between the signless Laplacian spectral radius of complete multipartite graphs and majorization. Linear Algebra and its Applications, 565(??): 225–238, March 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305755.

Ould-Baba:2015:CFC

[OBRA15] Hanany Ould-Baba, Vincent Robin, and Jérôme Antoni. Concise formulae for the cumulant matrices of a random vector. Linear Algebra and its Applications, 485(??):392– 416, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004413.

Oeding:2017:QV

[Oed17] Luke Oeding. The quadrifocal variety. Linear Algebra and its Applications, 512(??):306–330, January 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516304335.

Oi:2019:ARI

[Oi19] Shiho Oi. Algebraic reflexivity of isometry groups of algebras of Lipschitz maps. Linear Algebra and its Applications, 566(??):167– 182, April 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300059.

Oikhberg:2019:SSP

[Oik19] T. Oikhberg. On the stability of some preservers. Linear Algebra and its Applications, 563(??):494–526, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305299.

Ozkale:2018:PCR

[ÖK18] M. Revan Özkale and Özge Kuran. Principal components regression and r-k class predictions in linear mixed models. Linear Algebra and its Applications, 543 (??):173–204, April 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300077.

Oh:2016:EIV

[OL16] Seungsang Oh and Sangyop Lee. Enumerating independent vertex sets in grid graphs. Linear Algebra and its Applications, 510(??): 192–204, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303627.

Olkin:2015:LAB

[Olk15] Ingram Olkin. A linear algebra biography. Linear Algebra and its Applications, 473(??):3–13, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000361.

Ong:2018:DTH

[Ong18] Darren C. Ong. On a description of the Toda hierarchy using cocycle maps. Linear Algebra and its Applications, 549(??):12–29, July 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

Ortiz:2015:LTT

[OP15] Carlos M. Ortiz and Vern I. Paulsen. Lovász theta type norms and operator systems. Linear Algebra and its Applications, 477(??):128– 147, July 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500186X.

Oquendo:2016:BNI

[OP16a] Higidio Portillo Oquendo and Patricia Sánez Pacheco. Bounds for the 1-norm of the inverses of some triangular matrices. Linear Algebra and its Applications, 495 (??):163–173, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000318.

Ortiz:2016:QGH

[OP16b] Carlos M. Ortiz and Vern I. Paulsen. Quantum graph homomorphisms via operator systems. Linear Algebra and its Applications, 497 (??):23–43, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951600121X.

Orera:2019:T

[OP19a] H. Orera and J. M. Peña. BπR-tensors. Linear Algebra and its Applications, 581(??):247–259, November 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519302964.

Orera:2019:AIN

[OP19b] H. Orera and J. M. Peña. Accurate inverses of Nekrasov Z-matrices. Linear Algebra and its Applications, 574 (??):46–59, August 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301211.

Opozda:2016:SCS

[Opo16] Barbara Opozda. A sectional curvature for statistical structures. Linear Algebra and its Applications, 497 (??):134–161, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001233.

OMahony:2018:EMG

[OQ18] Olga O’Mahony and Rachel Quinlan. Edge-minimal graphs of exponent 2. Linear Algebra and its Applications, 542(??):66–83, ???? 2018.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300708.

Ouyang:2017:FFU

[OQY17] Chen Ouyang, Liqun Qi, and Xiying Yuan. The first few unicyclic and bicyclic hypergraphs with largest spectral radii. Linear Algebra and its Applications, 527(??): 141–162, August 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302331.

Orel:2016:APIa

[Ore16a] Marko Orel. Adjacency preservers on invertible Hermitian matrices I. Linear Algebra and its Applications, 499 (??):99–128, June 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007034.

Orel:2016:APIb

[Ore16b] Marko Orel. Adjacency preservers on invertible Hermitian matrices II. Linear Algebra and its Applications, 499(??):129–146, June 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379514007022.

Oblak:2017:MMM

[OS17] Polona Oblak and Helena Smigoc. The maximum of the minimal multiplicity of eigenvalues of symmetric matrices whose pattern is constrained by a graph. Linear Algebra and its Applications, 512(??): 48–70, January 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630413X.

Ostrander:2015:LMA

[Ost15] Aaron Ostrander. Laplacian matrices and Alexandrov topologies of digraphs. Linear Algebra and its Applications, 481(??):174–185, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002840.

Otake:2015:BAO

[Ota15]  Shuichi Otake.  A Be-

zoutian approach to orthogonal decompositions of trace forms or integral trace forms of some classical polynomials. Linear Algebra and its Applications, 471(??): 291–319, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Ouyang:2017:PIQ

[Ouy17] Yingkai Ouyang. Permutationinvariant qudit codes from polynomials. Linear Algebra and its Applications, 532(??):43–59, November 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517303956.

Olesky:2017:BSA

[OvdDV17] D. D. Olesky, P. van den Driessche, and K. N. Vander Meulen. Bordering for spectrally arbitrary sign patterns. Linear Algebra and its Applications, 534(??): 36–50, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730472X.

ORourke:2018:RPL

[OVW18] Sean O’Rourke, Van Vu, and Ke Wang. Random perturbation of low rank matrices: Improving classical bounds. Linear Algebra and its Applications, 540 (??):26–59, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306420.

OMeara:2017:WSM

[OW17] K. C. O’Meara and J. Watanabe. Weyr structures of matrices and relevance to commutative finite-dimensional algebras. Linear Algebra and its Applications, 532(??): 364–386, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303841.

Odegard:2015:QID

[OZ15a] Issac Odegard and Ryan J. Zerr. The quadratic irrationals and Ducci matrix sequences. Linear Algebra and its Applications, 484(??): 344–355, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004024.

Oswald:2015:CAK

[OZ15b] Peter Oswald and Weiqi Zhou. Convergence analysis for Kaczmarz-type methods in a Hilbert space framework. Linear Algebra and its Applications, 478(??): 131–161, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001925.

Osinsky:2018:PSA

[OZ18] A. I. Osinsky and N. L. Zamarashkin. Pseudo-skeleton approximations with better accuracy estimates. Linear Algebra and its Applications, 537(??):221–249, January 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305669.

Palacios:2016:SMI

[Pal16] José Luis Palacios. Some more interplay of the three Kirchhoffian indices. Linear Algebra and its Applications, 511(??):421–429, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304244.

Pan:2015:TMS

[Pan15a] Victor Y. Pan. Transformations of matrix structures work again. Linear Algebra and its Applications, 465(??):107–138, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514005898.

Pang:2015:AAP

[Pan15b] C. H. Jeffrey Pang. Accelerating the alternating projection algorithm for the case of affine subspaces using supporting hyperplanes. Linear Algebra and its Applications, 469(??):419–439, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007708.

Pankov:2016:EGG

[Pan16a] Mark Pankov. On embeddings of Grassmann graphs in polar Grassmann graphs. Linear Algebra and its Applications, 488(??):184–198, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005340.

Pankov:2016:OAH

[Pan16b] Mark Pankov. Orthogonal apartments in Hilbert Grassmannians. Linear Algebra and its Applications, 506(??):168–182, October 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516301859.

Pankov:2017:APT

[Pan17] Mark Pankov. Apartments preserving transformations of Grassmannians of infinitedimensional vector spaces.

Linear Algebra and its Applications, 531(??):498–509, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303798.

Pan:2019:BRK

[Pan19a] Zhidong Pan. Bounded reflexivity of kernels of elementary operators. Linear Algebra and its Applications, 574(??):153–166, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301235.

Pankov:2019:CPT

[Pan19b] Mark Pankov. Commutativity preserving transformations on conjugacy classes of finite rank self-adjoint operators. Linear Algebra and its Applications, 582(??): 430–439, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303568.

Pascoe:2018:NLT

[Pas18] J. E. Pascoe. The noncommutative Löwner theorem for matrix monotone functions over operator systems. Linear Algebra and its Applications, 541(??):54– 59, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730664X.

Pascoe:2019:EMC

[Pas19] J. E. Pascoe. An elementary method to compute the algebra generated by some given matrices and its dimension. Linear Algebra and its Applications, 571 (??):132–142, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300898.

Pellegrini:2016:MRS

[PB16] M. A. Pellegrini and M. C. Tamburini Bellani. More on regular subgroups of the affine group. Linear Algebra and its Applications, 505(??):126–151, September 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516301306.

Pate:2015:QHK

[PC15] Kevin Pate and Charles Ching-An Cheng. Quadratic homogeneous Keller maps of rank two. Linear Algebra and its Applications, 476(??):16–27, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001202.

Pavlovic:2015:ACO

[PCI15] V. Pavlović and D. S. Cvetković-Ilić. Applications of completions of operator matrices to reverse order law for 1-inverses of operators on Hilbert spaces. Linear Algebra and its Applications, 484(??):219–236, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003687.

Palchak:2015:STD

[PCL15] Elisabeth M. Palchak, Alexandru Cibotarica, and James V. Lambers. Solution of timedependent PDE through rapid estimation of block Gaussian quadrature nodes. Linear Algebra and its Applications, 468(??):233–259, March 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514004583.

Poloni:2017:CFP

[PD17] Federico Poloni and Gianna M. Del Corso. Counting Fiedler pencils with repetitions. Linear Algebra and its Applications, 532(??):

463–499, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304068.

Pearson:2015:SHT

[Pea15] Kelly J. Pearson. Spectral hypergraph theory of the adjacency hypermatrix and matroids. Linear Algebra and its Applications, 465(??):176–187, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006107.

Pellegrini:2017:ICF

[Pel17] Marco Antonio Pellegrini. Isomorphism classes of four dimensional nilpotent associative algebras over a field. Linear Algebra and its Applications, 533(??):132– 160, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304354.

Peperko:2017:BJG

[Pep17] Aljosa Peperko. Bounds on the joint and generalized spectral radius of the Hadamard geometric mean of bounded sets of positive kernel operators. Linear Algebra and its Applications, 533(??):418–427, November 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304408.

Perera:2018:SFG

[Per18] Sirani M. Perera. Signal flow graph approach to efficient and forward stable DST algorithms. Linear Algebra and its Applications, 542(??):360–390, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303506.

Petek:2015:MPK

[Pet15] Tatjana Petek. Maps preserving a Ky–Fan norm of Jordan product. Linear Algebra and its Applications, 487(??):86–111, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005170.

Perdigao:2018:MRP

[PF18] Cecília Perdigão and Amélia Fonseca. Minimum rank and path cover number for generalized and double generalized cycle star graphs. Linear Algebra and its Applications, 554(??):146–169, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302556.

Pirzada:2015:LEG

[PG15] S. Pirzada and Hilal A. Ganie. On the Laplacian eigenvalues of a graph and Laplacian energy. Linear Algebra and its Applications, 486(??):454–468, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005108.

Plonka:2016:PID

[PHW16] Gerlind Plonka, Sebastian Hoffmann, and Joachim Weickert. Pseudo-inverses of difference matrices and their application to sparse signal approximation. Linear Algebra and its Applications, 503 (??):26–47, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300428.

Pietrzycki:2018:RCM

[Pie18] Paweł Pietrzycki. Reduced commutativity of moduli of operators. Linear Algebra and its Applications, 557(??):375–402, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951830380X.

Pivovarchik:2019:MEB

[Piv19] Vyacheslav Pivovarchik. On multiplicities of eigenvalues of a boundary value problem on a snowflake graph. Linear Algebra and its Applications, 571(??):78– 91, June 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300679.

Plestenjak:2017:MDR

[Ple17] Bor Plestenjak. Minimal determinantal representations of bivariate polynomials. Linear Algebra and its Applications, 532(??): 550–569, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304238.

Pastravanu:2015:SCS

[PM15a] Octavian Pastravanu and Mihaela-Hanako Matcovschi. Sufficient conditions for Schur and Hurwitz diagonal stability of complex interval matrices. Linear Algebra and its Applications, 467(??): 149–173, February 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007290.

Pelaez:2015:FOS

[PM15b] María José Peláez and Julio Moro. First order structured perturbation theory for multiple zero eigenvalues of skew-adjoint matrices. Linear Algebra and its Applications, 464(??):3– 27, January 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379513003674.

Perovic:2018:LMP

[PM18] Vasilije Perović and D. Steven Mackey. Linearizations of matrix polynomials in Newton bases. Linear Algebra and its Applications, 556(??):1–45, November 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518303148.

Phat:2015:RFT

[PMB15] V. N. Phat, N. H. Muoi, and M. V. Bulatov. Robust finite-time stability of linear differential-algebraic delay equations. Linear Algebra and its Applications, 487(??):146–157, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500525X.

Paul:2019:SBJ

[PMW19] Kallol Paul, Arpita Mal, and Pawel Wójcik. Symmetry of Birkhoff–James orthogonality of operators defined between infinite dimensional Banach spaces. Linear Algebra and its Applications, 563(??):142–153, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305081.

Polyanskii:2019:AES

[Pol19] A. Polyanskii. On almostequidistant sets. Linear Algebra and its Applications, 563(??):220–230, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305287.

Poon:2015:PME

[Poo15] Edward Poon. Preservers of maximally entangled states. Linear Algebra and its Applications, 468(??):122–144, March 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514001402.

Popa:2016:HGA

[Pop16a] Dumitru Popa. The Heisenberg group associated to a Hilbert space. Linear Algebra and its Applications, 508(??):234–245, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302774.

Popa:2016:NMO

[Pop16b] Dumitru Popa. Nuclear multilinear operators with respect to a Banach ideal of multilinear functionals. Linear Algebra and its Applications, 489(??):93–103, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006011.

Pereira:2015:ECM

[PP15] Rajesh Pereira and Sarah Plosker. Extending a characterization of majorization to infinite dimensions. Linear Algebra and its Applications, 468(??):80–86, March 1, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379514000445.

Panda:2017:IWG

[PP17a] S. K. Panda and S. Pati. Inverses of weighted graphs. Linear Algebra and its Applications, 532(??):222–230, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304056.

Panda:2017:SGW

[PP17b] S. K. Panda and S. Pati. On some graphs which satisfy reciprocal eigenvalue properties. Linear Algebra and its Applications, 530(??): 445–460, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302422.

Plesinger:2018:EEC

[PP18] Martin Plesinger and Ivana Pultarová. On the extreme eigenvalues of certain matrices of non-standard inner products of Hermite polynomials. Linear Algebra and its Applications, 546 (??):50–66, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300582.

Pilipovic:2019:IMT

[PPZ19] S. Pilipović, B. Prangoski, and M. Zigić. Invertibility of matrix type operators of infinite order with exponential off-diagonal decay. Linear Algebra and its Applications, 582(??): 346–358, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303532.

Pan:2015:RMN

[PQY15] Victor Y. Pan, Guoliang Qian, and Xiaodong Yan. Random multipliers numerically stabilize Gaussian and block Gaussian elimination: Proofs and an extension to low-rank approximation. Linear Algebra and its Applications, 481(??):202–234, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002621.

Pavlova:2018:IPE

[PR18a] N. G. Pavlova and A. O. Remizov. On isomorphisms of pseudo-Euclidean spaces with signature (p, n − p) for p = 2, 3. Linear Algebra and its Applications, 541 (??):60–80, March 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306651.

Pluta:2018:PNA

[PR18b] Robert Pluta and Bernard Russo. Projections on von Neumann algebras as limits of elementary operators. Linear Algebra and its Applications, 557(??):188– 200, November 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303586.

Perez:2019:STR

[PR19] Claudia Pérez and Daniel Rivera. Serre type relations for complex semisimple Lie algebras associated to positive definite quasi-Cartan matrices. Linear Algebra and its Applications, 567 (??):14–44, April 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300047.

Puche:2018:CCD

[PRS18] Marc Puche, Timo Reis, and Felix L. Schwenninger. Constant-coefficient differentialalgebraic operators and the Kronecker form. Linear Algebra and its Applications, 552(??):29–41, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301836.

Parnas:2019:BRU

[PRS19] Michal Parnas, Dana Ron, and Adi Shraibman. The Boolean rank of the uniform intersection matrix and a family of its submatrices. Linear Algebra and its Applications, 574(??):67– 83, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930134X.

Petrovic:2015:NCB

[PS15] Miroslav Petrović and Slobodan K. Simić. A note on connected bipartite graphs of fixed order and size with maximal index. Linear Algebra and its Applications, 483(??):21–29, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003122.

Plavka:2016:SIE

[PS16] Ján Plavka and Sergei Sergeev. X-simple image eigencones of tropical matrices. Linear Algebra and its Applications, 507(??): 169–190, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302300.

Pavlikova:2017:CII

[PS17] Sona Pavlíková and Daniel Sevcovic. On a construction of integrally invertible graphs and their spectral properties. Linear Algebra and its Applications, 532(??):512–533, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304159.

Parnas:2018:APB

[PS18a] Michal Parnas and Adi Shraibman. The augmentation property of binary matrices for the binary and Boolean rank. Linear Algebra and its Applications, 556(??):70–99, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303197.

Plavka:2018:REI

[PS18b] Ján Plavka and Sergei Sergeev. Reachability of eigenspaces for interval circulant matrices in maxalgebra. Linear Algebra and its Applications, 550 (??):59–86, August 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301514.

Park:2019:IGE

[PS19a] Jongyook Park and Yoshio Sano. On Q-integral graphs with edge-degrees at most six. Linear Algebra and its Applications, 577(??):384– 411, September 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301673.

Passbani:2019:SRG

[PS19b] Hossein Passbani and Abbas Salemi. On spectral radius of graphs with pendant paths. Linear Algebra and its Applications, 571 (??):26–40, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300643.

Passer:2019:CCT

[PS19c] Benjamin Passer and Orr Moshe Shalit. Compressions of compact tuples. Linear Algebra and its Applications, 564 (??):264–283, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305573.

Pena:2019:SUM

[PS19d] Juan Manuel Peña and Tomas Sauer. SVD update methods for large matrices and applications. Linear Algebra and its Applications, 561(??):41–62, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304476.

Pozza:2019:ADF

[PS19e] S. Pozza and Z. Strakos. Algebraic description of the finite Stieltjes moment problem. Linear Algebra and its Applications, 561(??): 207–227, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304671.

Paul:2016:BJO

[PSG16] K. Paul, D. Sain, and P. Ghosh. Birkhoff–James orthogonality and smoothness of bounded linear operators. Linear Algebra and its Applications, 506(??): 551–563, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302415.

Paul:2018:ABJ

[PSM18] Kallol Paul, Debmalya Sain, and Arpita Mal. Approximate Birkhoff–James orthogonality in the space of bounded linear operators. Linear Algebra and its Applications, 537(??): 348–357, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305773.

Pan:2015:ENR

[PSZ15] Victor Y. Pan, John Svadlenka, and Liang Zhao. Estimating the norms of random circulant and Toeplitz matrices and their inverses. Linear Algebra and its Applications, 468(??):197–210, March 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514003966.

Paparella:2019:RSS

[PT19] Pietro Paparella and Amber R. Thrall. Realizing Suleimanova spectra via permutative matrices, II. Linear Algebra and its Applications, 566(??):183– 198, April 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300023.

Pudlak:2016:LTC

[Pud16] Pavel Pudlák. Linear tree codes and the problem of explicit constructions. Linear Algebra and its Applications, 490(??):124–144, February 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500645X.

Protasov:2017:MSC

[PV17] V. Yu. Protasov and A. S. Voynov. Matrix semigroups with constant spectral radius. Linear Algebra and its Applications, 513(??): 376–408, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304864.

Perotti:2018:MMP

[PW18] Luca Perotti and Michal Wojtylak. Matrix methods for Padé approximation: Numerical calculation of poles, zeros and residues. Linear Algebra and its Applications, 548(??):95–122, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951830106X.

Price:2019:STM

[PW19] Geoffrey Price and Myles Wortham. On sequences of Toeplitz matrices over finite fields. Linear Algebra and its Applications, 561(??): 63–80, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304464.

See corrigendum [PW20].

Price:2020:CSS

[PW20] Geoffrey Price and Myles Wortham. Corrigendum to: “On sequences of Toeplitz matrices over finite fields” [Linear Algebra Appl. 561 (2019) 63–80]. Linear Algebra and its Applications, 590 (??):330–332, April 1, 2020. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379520300124.

See [PW19].

Pan:2017:NSR

[PZ17a] Victor Y. Pan and Liang Zhao. New studies of randomized augmentation and additive preprocessing. Linear Algebra and its Applications, 512(??):256–305, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304347.

Pan:2017:NSG

[PZ17b] Victor Y. Pan and Liang Zhao. Numerically safe Gaussian elimination with no pivoting. Linear Algebra and its Applications, 527(??):349–383, August 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951730232X.

Qi:2015:CCA

[QH15] Xiaofei Qi and Jinchuan Hou. Characterization of k-commuting additive maps on rings. Linear Algebra and its Applications, 468 (??):48–62, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514000202.

Qiao:2016:WRG

[QKP16] Zhi Qiao, Jack H. Koolen, and Jongyook Park. 2walk-regular graphs with a small number of vertices compared to the valency. Linear Algebra and its Applications, 510(??): 10–24, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303275.

Qi:2015:TDS

[QWC15] Liqun Qi, Qun Wang, and Yannan Chen. Three dimensional strongly symmetric circulant tensors. Linear Algebra and its Applications, 482(??):207–220, October 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003432.

Qiao:2018:TFT

[QWW18] Sanzheng Qiao, Xue-Zhong Wang, and Yimin Wei. Two finite-time convergent Zhang neural network models for time-varying complex matrix Drazin inverse. Linear Algebra and its Applications, 542(??):101–117, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301763.

Rodrguez-Andrade:2016:GAR

[RAAGA16] M. A. Rodríguez-Andrade, G. Aragón-González, and J. L. Aragón. The generation of all rational orthogonal matrices in R^{p,q}. Linear Algebra and its Applications, 496(??):101–113, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516000604.

Radjabalipour:2017:SJC

[Rad17] M. Radjabalipour. A symmetrization of the Jordan canonical form. Linear Algebra and its Applications, 528(??):94–112, September 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301999.

Rainbolt:2019:REB

[Rai19] Julianne Rainbolt. Regular elements in bases of Hecke algebras. Linear Algebra and its Applications, 575(??): 87–105, August 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301557.

Rakic:2017:NRM

[Rak17] Dragan S. Rakić. A note on Rao and Mitra’s constrained inverse and Drazin’s (b, c) inverse. Linear Algebra and its Applications, 523 (??):102–108, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730126X.

Ram:2016:CZK

[Ram16] Samrith Ram. Counting zero kernel pairs over a finite field. Linear Algebra and its Applications, 495 (??):1–10, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951600032X.

Ram:2017:NLT

[Ram17] Samrith Ram. The number of linear transformations defined on a subspace with given invariant factors. Linear Algebra and its Applications, 532(??): 146–161, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303981.

Rao:2017:SPS

[Rao17] T. S. S. R. K. Rao. Smooth points in spaces of operators. Linear Algebra and its Applications, 517(??): 129–133, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305900.

Raouafi:2018:GKM

[Rao18] Samir Raouafi. A generalization of the Kreiss Matrix Theorem. Linear Algebra and its Applications, 549 (??):86–99, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301216.

Reyes-Bustos:2016:CTG

[RB16] Cid Reyes-Bustos. Cayleytype graphs for groupsubgroup pairs. Linear Algebra and its Applications, 488(??):320–349, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515005765.

Rump:2019:ASE

[RB19] Siegfried M. Rump and Prashant Batra. Addendum to “Estimates of the determinant of a perturbed identity matrix [Linear Algebra Appl. 558 (2018) 101– 107]. Linear Algebra and its Applications, 565(??): 309–312, March 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305524.

See [Rum18].

Roozbeh:2016:CBE

[RBKA16] Mahdi Roozbeh, Saman Babaie-Kafaki, and Mohammad Arashi. A class of biased estimators based on QR decomposition. Linear Algebra and its Applications, 508(??):190–205, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302713.

Rakic:2015:POR

[RD15] Dragan S. Rakić and Dragan S. Djordjević. Partial orders in rings based on generalized inverses — unified theory. Linear Algebra and its Applications, 471 (??):203–223, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000336.

Rodrigo-Escudero:2016:CDG

[RE16] Adrián Rodrigo-Escudero. Classification of division gradings on finite-dimensional simple real algebras. Linear Algebra and its Applications, 493(??):164–182, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006990.

Rebs:2019:BCM

[Reb19] Christian Rebs. On the best constants in Markov inequalities for the Laplace operator on polynomials with the Hermite and Legendre norm. Linear Algebra and its Applications, 572 (??):32–45, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300989.

Reff:2016:OGG

[Ref16]  Nathan Reff.  Oriented

gain graphs, line graphs and eigenvalues. Linear Algebra and its Applications, 506(??):316–328, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302154.

Reese:2018:MFF

[RFPS18] Tyler M. Reese, Joseph D. Fehribach, Randy C. Paffenroth, and Brigitte Servatius. Matrices over finite fields and their Kirchhoff graphs. Linear Algebra and its Applications, 547 (??):128–147, June 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300879.

Reese:2019:UKG

[RFPS19] Tyler M. Reese, Joseph D. Fehribach, Randy C. Paffenroth, and Brigitte Servatius. Uniform Kirchhoff graphs. Linear Algebra and its Applications, 566(??):1–16, April 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305810.

Rusu:2016:PRC

[RGPH16] Cristian Rusu, Nuria González-Prelcic, and Robert W.

Heath, Jr. Properties of real and complex ETFs and their application to the design of low coherence frames. Linear Algebra and its Applications, 508(??): 81–90, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302555.

Roman:2015:GEG

[RH15] Jorge Carlos Román and James P. Hobert. Geometric ergodicity of Gibbs samplers for Bayesian general linear mixed models with proper priors. Linear Algebra and its Applications, 473 (??):54–77, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379513008215.

Ricard:2019:EIA

[Ric19] Éric Ricard. An extension of inequalities by Ando. Linear Algebra and its Applications, 580(??):336–342, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302812.

Ringel:2017:EVM

[Rin17] Claus Michael Ringel. The eigenvector variety of a matrix pencil. Linear Algebra and its Applications, 531(??):447–458, October 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517302926.

Rissner:2016:NIM

[Ris16] Roswitha Rissner. Null ideals of matrices over residue class rings of principal ideal domains. Linear Algebra and its Applications, 494 (??):44–69, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000069.

Rivero:2015:DRM

[Riv15a] Abdon Eddy Choque Rivero. On Dyukarev’s resolvent matrix for a truncated Stieltjes matrix moment problem under the view of orthogonal matrix polynomials. Linear Algebra and its Applications, 474(??):44– 109, June 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000567.

Rivero:2015:MHT

[Riv15b] Abdon Eddy Choque Rivero. On matrix Hurwitz type polynomials and their interrelations to Stieltjes positive definite sequences and orthogonal matrix polynomials. Linear Algebra and its Applications, 476(??):56– 84, July 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001457.

Rizzolo:2019:DSM

[Riz19] Douglas Rizzolo. Determinants of Seidel matrices and a conjecture of Ghorbani. Linear Algebra and its Applications, 579(??): 51–54, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302344.

Rocha:2018:RSR

[RJK18] Israel Rocha, Jeannette Janssen, and Nauzer Kalyaniwalla. Recovering the structure of random linear graphs. Linear Algebra and its Applications, 557(??):234–264, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518303574.

Rising:2015:EAS

[RKT15] Justin Rising, Alex Kulesza, and Ben Taskar. An efficient algorithm for the symmetric principal minor assignment problem. Linear Algebra and its Applications, 473 (??):126–144, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400233X.

Rosenmann:2019:PCM

[RLP19] Amnon Rosenmann, Franz Lehner, and Aljosa Peperko. Polynomial convolutions in max-plus algebra. Linear Algebra and its Applications, 578(??):370–401, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302290.

Ringh:2018:SBP

[RMKJ18] Emil Ringh, Giampaolo Mele, Johan Karlsson, and Elias Jarlebring. Sylvesterbased preconditioning for the waveguide eigenvalue problem. Linear Algebra and its Applications, 542 (??):441–463, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303907.

Rosa:2018:SSH

[RMP18] Kennett L. Dela Rosa, Dennis I. Merino, and Agnes T.

Paras. The subspaces spanned by Householder vectors associated with an orthogonal or a symplectic matrix. Linear Algebra and its Applications, 546 (??):37–49, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300570.

Robbiano:2016:EGB

[RMS16] María Robbiano, Katherine Tapia Morales, and Bernardo San Martín. Extremal graphs with bounded vertex bipartiteness number. Linear Algebra and its Applications, 493(??):28– 36, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500703X.

Rosado:2017:PET

[RNC17] James Rosado, Hieu D. Nguyen, and Lei Cao. Partitions of equiangular tight frames. Linear Algebra and its Applications, 526(??): 95–120, August 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301842.

Robbiano:2019:GIS

[Rob19]  María Robbiano.  Guo’s

index for some classes of matrices. Linear Algebra and its Applications, 564 (??):15–27, March 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305469.

Rodriguez:2019:NNB

[Rod19] Jonnathan Rodríguez. A note on new bounds for the Estrada Index. Linear Algebra and its Applications, 580(??):121–127, November 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302745.

Rohn:2018:DSM

[Roh18] Jiri Rohn. Diagonally singularizable matrices. Linear Algebra and its Applications, 555(??):84–91, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302945.

Roos:2019:GHT

[Roo19] Bero Roos. Generalization of a Hadamard type inequality for permanents. Linear Algebra and its Applications, 571 (??):58–77, June 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300771.

Rowlinson:2015:SCE

[Row15]  Peter Rowlinson.  Star

complements and edgeconnectivity in finite graphs. Linear Algebra and its Applications, 476(??):124–132, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515001470.

Rowlinson:2016:EMT

[Row16a] Peter Rowlinson. Eigenvalue multiplicity in triangle-free graphs. Linear Algebra and its Applications, 493(??): 484–493, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007260.

Rowlinson:2016:GJT

[Row16b]  Peter Rowlinson.  On

graphs with just three distinct eigenvalues. Linear Algebra and its Applications, 507(??):462–473, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302488.

Rowlinson:2018:ESC

[Row18] Peter Rowlinson. An extension of the star complement technique for regular graphs. Linear Algebra and its Applications, 557(??):496–507, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303975.

Ribeiro:2018:CPD

[RR18a] Ademir Alves Ribeiro and Peter Richtárik. The complexity of primal-dual fixed point methods for ridge regression. Linear Algebra and its Applications, 556(??): 342–372, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303434.

Richter:2018:LRC

[RR18b] Sebastian Richter and Israel Rocha. Layout of random circulant graphs. Linear Algebra and its Applications, 559(??):95–113, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304300.

Rakosnik:2016:MF

[RRS16] Jiří Rákosník, Miroslav Rozložník, and Zdenek Strakoš.

Miroslav Fiedler, 1926– 2015. Linear Algebra and its Applications, 497(??): 162–166, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001178.

Reichel:2016:NBQ

[RRT16] Lothar Reichel, Giuseppe Rodriguez, and Tunan Tang. New block quadrature rules for the approximation of matrix functions. Linear Algebra and its Applications, 502(??):299–326, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004073.

Reis:2015:KYP

[RRV15] Timo Reis, Olaf Rendel, and Matthias Voigt. The Kalman–Yakubovich–Popov inequality for differential– algebraic systems. Linear Algebra and its Applications, 485(??):153–193, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003742.

Reff:2017:CBH

[RS17a] Nathan Reff and Howard Skogman. A connection between Hadamard matrices, oriented hypergraphs and signed graphs. Linear Algebra and its Applications, 529(??):115–125, September 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302379.

Ren:2017:CMN

[RS17b] Ling-Zhi Ren and Hai-Ying Shan. Characterizations of matrices with nonzero signed row compound. Linear Algebra and its Applications, 513(??):210–223, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304669.

Rohn:2018:IMR

[RS18] Jiri Rohn and Sergey P. Shary. Interval matrices: Regularity generates singularity. Linear Algebra and its Applications, 540(??): 149–159, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306481.

Rauhut:2017:LRT

[RSS17] Holger Rauhut, Reinhold Schneider, and Zeljka Stojanac. Low rank tensor recovery via iterative hard thresholding. Linear Algebra and its Applications, 523 (??):220–262, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301295.

Rossi:2016:ILG

[RT16] Michele Rossi and Lea Terracini. z-linear Gale duality and poly weighted spaces (PWS). Linear Algebra and its Applications, 495 (??):256–288, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000549.

Rubei:2019:RRI

[Rub19]  Elena Rubei.  On rank

range of interval matrices. Linear Algebra and its Applications, 561(??): 81–97, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304592.

Rukhin:2015:RLR

[Ruk15] Andrew L. Rukhin. Restricted likelihood representation for multivariate heterogeneous random effects models. Linear Algebra and its Applications, 473 (??):78–92, May 15, 2015.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514000263.

Rump:2018:EDP

[Rum18] Siegfried M. Rump. Estimates of the determinant of a perturbed identity matrix. Linear Algebra and its Applications, 558(??): 101–107, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303823.

See addendum [RB19].

Rump:2019:BDG

[Rum19] Siegfried M. Rump. Bounds for the determinant by Gershgorin circles. Linear Algebra and its Applications, 563(??):215–219, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305068.

Robol:2018:EEA

[RV18] Leonardo Robol and Raf Vandebril. Efficient Ehrlich– Aberth iteration for finding intersections of interpolating polynomials and rational functions. Linear Algebra and its Applications, 542(??):282–309, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302987.

Ran:2016:ALI

[RW16] Yu-Hong Ran and Jun-Gang Wang. On ADI-like iteration method for fractional diffusion equations. Linear Algebra and its Applications, 493(??):544–555, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007624.

Rzezuchowski:2017:CAS

[RW17] T. Rzezuchowski and J. Wasowski. Characterization of AE solution sets of parametric linear systems based on the techniques of convex sets. Linear Algebra and its Applications, 533(??):468– 490, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304652.

Rozanski:2017:MSV

[RWH17] Michal Rózański, Roman Witula, and Edyta Hetmaniok. More subtle versions of the Hadamard inequality. Linear Algebra and its Applications, 532(??): 500–511, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304135.

Radjavi:2015:TKR

[RY15] Heydar Radjavi and Bamdad R. Yahaghi. A theorem of Kaplansky revisited. Linear Algebra and its Applications, 487(??):268– 275, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500539X.

Repovs:2017:PGL

[RZ17] Dusan D. Repovs and Mikhail V. Zaicev. Pauli gradings on Lie superalgebras and graded codimension growth. Linear Algebra and its Applications, 520 (??):134–150, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300526.

Shakeri:2017:NEP

[SA17] F. Shakeri and R. Alizadeh. Nonnegative and eventually positive matrices. Linear Algebra and its Applications, 519(??):19– 26, April 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306358.

Sababheh:2015:IIU

[Sab15] M. Sababheh. Interpolated inequalities for unitarily invariant norms. Linear Algebra and its Applications, 475 (??):240–250, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001330.

Sababheh:2016:CMM

[Sab16] M. Sababheh. Convexity and matrix means. Linear Algebra and its Applications, 506(??):588–602, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302440.

Sababheh:2018:NRI

[Sab18]  M. Sababheh.  Numeri-

cal radius inequalities via convexity. Linear Algebra and its Applications, 549 (??):67–78, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301356.

Saburov:2019:EMN

[Sab19] Mansoor Saburov. Ergodicity of p-majorizing nonlinear Markov operators on the finite dimensional space. Linear Algebra and its Applications, 578(??):53–74, October 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302204.

Simic:2015:NSL

[SAdFZ15] Slobodan K. Simić, Milica Andelić, Carlos M. da Fonseca, and Dejan Zivković. Notes on the second largest eigenvalue of a graph. Linear Algebra and its Applications, 465(??):262–274, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951400617X.

Salce:2016:CPR

[Sal16] Luigi Salce. Completely positive rooted matrices. Linear Algebra and its Applications, 505(??):211–223, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301501.

Sambou:2017:SCE

[Sam17] Diomba Sambou. A simple criterion for the existence of nonreal eigenvalues for a class of 2D and 3D Pauli operators. Linear Algebra and its Applications, 529(??):51– 88, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302276.

Sander:2015:ICR

[San15] J. W. Sander. Integral circulant Ramanujan graphs via multiplicativity and ultrafriable integers. Linear Algebra and its Applications, 477 (??):21–41, July 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001561.

Sander:2018:KIC

[San18] J. W. Sander. On the kernel of integral circulant graphs. Linear Algebra and its Applications, 549 (??):79–85, July 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301332.

Sarin:2019:GTN

[Sar19] Parth Sarin. On the geometry of tensor network states of 2 × N grids. Linear Algebra and its Applications, 575(??):235–249, August 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930148X.

Sastre:2018:EEM

[Sas18] J. Sastre. Efficient evaluation of matrix polynomials. Linear Algebra and its Applications, 539(??): 229–250, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306389.

Sawyer:2016:CID

[Saw16] P. Sawyer. Computing the Iwasawa decomposition of the classical Lie groups of noncompact type using the QR decomposition. Linear Algebra and its Applications, 493(??):573–579, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515007272.

Schulz:2015:CCD

[SB15] F. Schulz and R. Brits. Commutators, commutativity and dimension in the socle of a Banach algebra: a generalized Wedderburn– Artin and Shoda’s theorem. Linear Algebra and its Applications, 484(??): 175–198, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500395X.

Solak:2018:SPC

[SB18] Süleyman Solak and Mustafa Bahsi. Some properties of circulant matrices with Ducci sequences. Linear Algebra and its Applications, 542(??):557–568, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305347.

Schulz:2015:TCS

[SBB15] F. Schulz, R. Brits, and G. Braatvedt. Trace characterizations and socle identifications in Banach algebras. Linear Algebra and its Applications, 472(??):151– 166, May 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000191.

Sburlati:2017:CPD

[Sbu17] Giovanni Sburlati. Congruences for permanents and determinants of circulants. Linear Algebra and its Applications, 531(??):257–267, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303671.

Simmons:2017:SLQ

[SCD17] David E. Simmons, Justin P. Coon, and Animesh Datta. Symmetric Laplacians, quantum density matrices and their von Neumann entropy. Linear Algebra and its Applications, 532(??):534–549, November 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304020.

Schrijver:2015:TTR

[Sch15] Alexander Schrijver. On traces of tensor representations of diagrams. Linear Algebra and its Applications, 476(??):28–41, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001445.

Schweitzer:2016:MCE

[Sch16] Marcel Schweitzer. Monotone convergence of the extended Krylov subspace method for Laplace–Stieltjes functions of Hermitian positive definite matrices. Linear Algebra and its Applications, 507(??):486–498, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302580.

Schmale:2017:ACS

[Sch17] Wiland Schmale. Algebraic condition for the singularity of certain Toeplitz pencils. Linear Algebra and its Applications, 533(??):386– 396, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304482.

Shao:2016:SPP

[SdJY^{+}16] Meiyue Shao, Felipe H. da Jornada, Chao Yang, Jack Deslippe, and Steven G. Louie. Structure preserving parallel algorithms for solving the Bethe–Salpeter eigenvalue problem. Linear Algebra and its Applications, 488(??):148–167, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005637.

Sharma:2017:NIP

[SDK17] R. Sharma, P. Devi, and R. Kumari. A note on inequalities for positive linear maps. Linear Algebra and its Applications, 528(??): 113–123, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302087.

Seeger:2018:CEA

[See18] Alberto Seeger. Complementarity eigenvalue analysis of connected graphs. Linear Algebra and its Applications, 543(??):205–225, April 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306997.

Seigal:2018:GDR

[Sei18] Anna Seigal. Gram determinants of real binary tensors. Linear Algebra and its Applications, 544 (??):350–369, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300260.

Sergeev:2015:ESC

[Ser15] Sergei Sergeev. Extremals of the supereigenvector cone in max algebra: a combinatorial description. Linear Algebra and its Applications, 479(??):106–117, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001895.

Serre:2016:NCS

[Ser16] Denis Serre. Non-commutative standard polynomials applied to matrices. Linear Algebra and its Applications, 490(??):202–223, February 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006667.

Saez:2018:GCP

[SFW18] Meritxell Sáez, Elisenda Feliu, and Carsten Wiuf. Graphical criteria for positive solutions to linear systems. Linear Algebra and its Applications, 552(??): 166–193, September 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302088.

Shao:2016:SIS

[SG16] Yanling Shao and Yubin Gao. The scrambling index set of primitive minimally strong digraphs. Linear Algebra and its Applications, 500(??):1–14, July 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001671.

Sui:2017:FNM

[SG17a] Xuefang Sui and Paolo Gondolo. Factorizations into normal matrices in indefinite inner product spaces. Linear Algebra and its Applications, 516(??):143–166, March 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305560.

Sui:2017:NPD

[SG17b] Xuefang Sui and Paolo Gondolo. A new polar decomposition in a scalar product space. Linear Algebra and its Applications, 516 (??):126–142, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305559.

Sang:2016:LPQ

[SGH16] Man Sang, Suogang Gao, and Bo Hou. Leonard pairs and quantum algebra U_{q}(sl_{2}). Linear Algebra and its Applications, 510(??): 346–360, December 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630386X.

Sepitka:2015:RSN

[SH15a] Peter Sepitka and Roman Simon Hilscher. Recessive solutions for nonoscillatory discrete symplectic systems. Linear Algebra and its Applications, 469(??):243–275, March 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007642.

Slawski:2015:EPD

[SH15b] Martin Slawski and Matthias Hein. Estimation of positive definite M-matrices and structure learning for attractive Gaussian Markov random fields. Linear Algebra and its Applications, 473 (??):145–179, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514002444.

Sharma:2016:SSC

[Sha16] Uday Bhaskar Sharma. Simultaneous similarity classes of commuting matrices over a finite field. Linear Algebra and its Applications, 501 (??):48–97, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300015.

Shchesnovich:2016:PTC

[Shc16] Valery S. Shchesnovich. The permanent-on-top conjecture is false. Linear Algebra and its Applications, 490(??):196–201, February 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

Shemesh:2016:STR

[She16] Dan Shemesh. A simultaneous triangularization result. Linear Algebra and its Applications, 498(??):394– 398, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005716.

Shebrawi:2017:NRI

[She17] Khalid Shebrawi. Numerical radius inequalities for certain 2 × 2 operator matrices II. Linear Algebra and its Applications, 523 (??):1–12, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300964.

Shitov:2015:BFS

[Shi15]  Yaroslav Shitov.  Bases

of a free semimodule are small. Linear Algebra and its Applications, 466(??): 38–40, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006491.

Shimabukuro:2016:MAA

[Shi16a] Osamu Shimabukuro. Modular adjacency algebras of dual polar schemes. Linear Algebra and its Applications, 489(??):238–252, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006072.

Shinzawa:2016:UDP

[Shi16b] N. Shinzawa. Ultra discrete permanent and the consistency of max plus linear equations. Linear Algebra and its Applications, 506(??):445–477, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301811.

Shitov:2016:CTL

[Shi16c] Yaroslav Shitov. A counterexample on tropical linear spaces. Linear Algebra and its Applications, 511(??): 19–21, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303871.

Shitov:2016:EOS

[Shi16d] Yaroslav Shitov. Extending orthogonal subsets of semimodules. Linear Algebra and its Applications, 508(??): 225–233, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302798.

Shitov:2016:NSR

[Shi16e] Yaroslav Shitov. A note on square roots of nonnegative matrices. Linear Algebra and its Applications, 497 (??):62–65, May 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001245.

Shitov:2016:CPS

[Shi16f] Yaroslav Shitov. On the characteristic polynomial of a supertropical adjoint matrix. Linear Algebra and its Applications, 499(??):26– 30, June 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516001634.

Shitov:2016:SSZ

[Shi16g] Yaroslav Shitov. Sums of square-zero endomorphisms of a free module. Linear Algebra and its Applications, 507(??):191–197, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302373.

Shitov:2017:COI

[Shi17a] Yaroslav Shitov. A counterexample on outer inverses in semigroups. Linear Algebra and its Applications, 513(??):120–121, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304566.

Shitov:2017:TRS

[Shi17b] Yaroslav Shitov. Three remarks on spectra of zerononzero patterns. Linear Algebra and its Applications, 534(??):73–77, December 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304846.

Shitov:2018:CFM

[Shi18a] Yaroslav Shitov. Commuting functions of matrices over topological fields. Linear Algebra and its Applications, 553(??):362–364, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302398.

Shitov:2018:DSM

[Shi18b] Yaroslav Shitov. On the determinant of a sparse 0– 1 matrix. Linear Algebra and its Applications, 554 (??):49–50, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302593.

Shitov:2018:RLT

[Shi18c] Yaroslav Shitov. On the rank of a Latin tensor. Linear Algebra and its Applications, 544(??):299–305, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518300259.

Shitov:2019:POR

[Shi19] Yaroslav Shitov. A partial order that is not rank/ trace complete. Linear Algebra and its Applications, 582(??):99–102, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303349.

Simon:2017:UPT

[Sim17] Barry Simon. Unitaries permuting two orthogonal projections. Linear Algebra and its Applications, 528(??): 436–441, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301945.

Simson:2018:CSC

[Sim18] Daniel Simson. A Coxeter spectral classification of positive edge-bipartite graphs I. Dynkin types B_{\}, C_{\}, F_{△}, G_{∈}, E_{}, E_{}, E_{∀}. Linear Algebra and its Applications, 557(??):105–133, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518303318.

Simson:2019:SCS

[Sim19] Daniel Simson. Symbolic computation of strong Gram congruences for Cox-regular positive edge-bipartite graphs with loops. Linear Algebra and its Applications, 573 (??):90–143, July 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300916.

Sinchuk:2017:NKK

[Sin17] Sergey Sinchuk. A note on K_{2} of Kac–Moody groups over a field. Linear Algebra and its Applications, 534(??):174–180, December 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

Shi:2016:POP

[SJD16] Weijuan Shi, Guoxing Ji, and Hongke Du. Pairs of orthogonal projections with a fixed difference. Linear Algebra and its Applications, 489(??):288–297, January 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006151.

Soto:2015:NPM

[SJS15] Ricardo L. Soto, Ana I. Julio, and Mario Salas. Nonnegative persymmetric matrices with prescribed elementary divisors. Linear Algebra and its Applications, 483(??):139–157, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003511.

Schwetlick:2017:BNM

[SK17a] Hubert Schwetlick and Ute Kandler. Block Newton Method and Block Rayleigh Quotient Iteration for computing invariant subspaces of general complex matrices. Linear Algebra and its Applications, 526(??):60– 94, August 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301933.

Sharma:2017:PLM

[SK17b] R. Sharma and R. Kumari. Positive linear maps and perturbation bounds of matrices. Linear Algebra and its Applications, 528(??): 62–75, September 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001440.

Sanca:2018:FPT

[SKC18] Ernest Sanca, Vladimir R. Kostić, and Ljiljana Cvetković. Fractional pseudospectra and their localizations. Linear Algebra and its Applications, 559(??):244–269, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304245.

Skripka:2017:PSS

[Skr17] Anna Skripka. On positivity of spectral shift functions. Linear Algebra and its Applications, 523(??): 118–130, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301271.

Skulj:2017:PBD

[Sku17]  Damjan Skulj.  Pertur-

bation bounds and degree of imprecision for uniquely convergent imprecise Markov chains. Linear Algebra and its Applications, 533(??):336–356, November 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304421.

Sarkans:2016:SHO

[SL16] E. Sarkans and H. Logemann. Stability of higherorder discrete-time Lur’e systems. Linear Algebra and its Applications, 506(??): 183–211, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516301744.

Shen:2017:ACT

[SLY17] W. P. Shen, C. Li, and J. C. Yao. Approximate Cayley transform methods for inverse eigenvalue problems and convergence analysis. Linear Algebra and its Applications, 523(??): 187–219, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301313.

Shaked-Monderer:2016:SGW

[SM16] Naomi Shaked-Monderer. SPN graphs: When copositive = SPN. Linear Algebra and its Applications, 509(??):82–113, November 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516302804.

See corrigendum [SM18].

Shaked-Monderer:2018:CSS

[SM18] Naomi Shaked-Monderer. Corrigendum to “SPN graphs: When copositive = SPN” [Linear Algebra Appl. 509 (2016) 82–113]. Linear Algebra and its Applications, 541(??):285–286, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306894.

See [SM16].

Smaga:2015:UOD

[Sma15] Lukasz Smaga. Uniquely Eoptimal designs with n ≡ 2 (mod 4) correlated observations. Linear Algebra and its Applications, 473 (??):297–315, May 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514005667.

Shaked-Monderer:2016:SCG

[SMBDK16] Naomi Shaked-Monderer, Abraham Berman, Mirjam Dür, and M. Rajesh Kannan. SPN completable graphs. Linear Algebra and its Applications, 498 (??):58–73, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006909.

Sababheh:2019:NIG

[SMH19] Mohammed Sababheh, Hamid Reza Moradi, and Zahra Heydarbeygi. New inequalities for the generalized Karcher mean. Linear Algebra and its Applications, 580(??): 184–199, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302757.

Santos:2017:JMC

[SNDM17] Carla Santos, Célia Nunes, Cristina Dias, and João Tiago Mexia. Joining models with commutative orthogonal block structure. Linear Algebra and its Applications, 517(??):235–245, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306085.

Sau:2016:PSA

[SNP16] Nguyen H. Sau, P. Niamsup, and Vu N. Phat. Positivity and stability analysis for linear implicit difference delay equations. Linear Algebra and its Applications, 510(??):25–41, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516303391.

Simsek:2016:GHR

[SÖ16] Hakan Simsek and Mustafa Özdemir. Generating hyperbolical rotation matrix for a given hyperboloid. Linear Algebra and its Applications, 496(??):221–245, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516000537.

Sodomaco:2018:PES

[Sod18] Luca Sodomaco. The product of the eigenvalues of a symmetric tensor. Linear Algebra and its Applications, 554(??):224–248, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302738.

Soltan:2018:PSC

[Sol18a] Valeriu Soltan. Polarity and separation of cones. Linear Algebra and its Applications, 538(??):212–224, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305980.

Solymosi:2018:SNM

[Sol18b] József Solymosi. The sum of nonsingular matrices is often nonsingular. Linear Algebra and its Applications, 552(??):159–165, September 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302131.

Soltan:2019:MTC

[Sol19] Valeriu Soltan. Moreautype characterizations of polar cones. Linear Algebra and its Applications, 567 (??):45–62, April 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519300199.

Somodi:2017:IZF

[Som17] Marius Somodi. On the Ihara zeta function and resistance distance-based indices. Linear Algebra and its Applications, 513(??): 201–209, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304591.

Sothanaphan:2017:DBM

[Sot17] Nat Sothanaphan. Determinants of block matrices with noncommuting blocks. Linear Algebra and its Applications, 512(??):202–218, January 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304645.

Serre:2016:GBL

[SP16] François Serre and Markus Püschel. Generalizing block LU factorization: a lowerupper-lower block triangular decomposition with minimal off-diagonal ranks. Linear Algebra and its Applications, 509(??):114–142, November 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302920.

Sain:2019:NIP

[SPBB19] Debmalya Sain, Kallol Paul, Pintu Bhunia, and Santanu Bag. On the numerical index of polyhedral Banach spaces. Linear Algebra and its Applications, 577(??):121–133, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951930182X.

Sain:2015:ONA

[SPH15] Debmalya Sain, Kallol Paul, and Sourav Hait. Operator norm attainment and Birkhoff–James orthogonality. Linear Algebra and its Applications, 476(??):85– 97, July 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515001469.

Sanchez:2018:ESI

[SR18] Víctor Cabezas Sánchez and Fernando Pablos Romo. Explicit solutions of infinite systems of linear equations from reflexive generalized inverses of finite potent endomorphisms. Linear Algebra and its Applications, 559(??):125–144, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304312.

Sanguanwong:2019:EGM

[SR19] Ratsiri Sanguanwong and Kijti Rodtes. The equality of generalized matrix functions on the set of all symmetric matrices. Linear Algebra and its Applications, 565 (??):65–81, March 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830572X.

Srinivasan:2017:NHR

[Sri17] Murali K. Srinivasan. Note on hook representations of the symmetric group. Linear Algebra and its Applications, 515(??):6–10, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516305213.

Saluev:2015:GPC

[SS15a] Tigran Saluev and Iskander Sitdikov. Generic properties and a criterion of an operator norm. Linear Algebra and its Applications, 485(??):1–20, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379515004346.

Stampach:2015:SFS

[SS15b] F. Stampach and P. Stovícek. Special functions and spectrum of Jacobi matrices. Linear Algebra and its Applications, 464(??):38–61, January 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379513004230.

Sun:2015:ESS

[SS15c] Huaqing Sun and Yuming Shi. On essential spectra of singular linear Hamiltonian systems. Linear Algebra and its Applications, 469(??): 204–229, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007654.

Simic:2016:PRS

[SS16] Slobodan K. Simić and Zoran Stanić. Polynomial reconstruction of signed graphs. Linear Algebra and its Applications, 501 (??):390–408, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516300477.

Schrijver:2017:SAP

[SS17a] Alexander Schrijver and Bart Sevenster. The Strong Arnold Property for 4connected flat graphs. Linear Algebra and its Applications, 522(??):153–160, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951730068X.

Szanto:2017:SSS

[SS17b] Csaba Szántó and István Szöllősi. A short solution to the subpencil problem involving only column minimal indices. Linear Algebra and its Applications, 517 (??):99–119, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951630581X.

Smoktunowicz:2018:STS

[SS18] Agata Smoktunowicz and Alicja Smoktunowicz. Settheoretic solutions of the Yang–Baxter equation and new classes of R-matrices. Linear Algebra and its Applications, 546(??):86–114, June 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518300533.

Solhjem:2019:SMP

[SS19] Sara Solhjem and Jessica Striker. Sign matrix polytopes from Young tableaux. Linear Algebra and its Applications, 574(??):84–122, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Stor:2015:FSE

[SSB15a] N. Jakovcević Stor, I. Slapnicar, and J. L. Barlow. Forward stable eigenvalue decomposition of rank-one modifications of diagonal matrices. Linear Algebra and its Applications, 487(??): 301–315, December 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005406.

Stor:2015:AED

[SSB15b] Nevena Jakovcević Stor, Ivan Slapnicar, and Jesse L. Barlow. Accurate eigenvalue decomposition of real symmetric arrowhead matrices and applications. Linear Algebra and its Applications, 464(??):62–89, January 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379513006265.

Salinelli:2016:EES

[SSCS16] E. Salinelli, S. Serra-Capizzano, and D. Sesana. Eigenvalue–eigenvector structure of Schoenmakers–Coffey matrices via Toeplitz technology and applications. Linear Algebra and its Applications, 491(??):138–160, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001810.

Silva:2016:CTE

[SSF16] Teresa M. Silva, Luís Silva, and Sara Fernandes. Convergence time to equilibrium distributions of autonomous and periodic nonautonomous graphs. Linear Algebra and its Applications, 488(??):199–215, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005650.

Soleymani:2015:HFI

[SSH15] F. Soleymani, Predrag S. Stanimirović, and F. Khaksar Haghani. On hyperpower family of iterations for computing outer inverses possessing high efficiencies. Linear Algebra and its Applications, 484(??):477–495, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004103.

Shi:2015:NBE

[SSS15] Minjia Shi, Keisuke Shiromoto, and Patrick Solé. A note on a basic exact sequence for the Lee and Euclidean weights of linear codes over F_{ℓ}. Linear Algebra and its Applications, 475 (??):151–153, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000622.

Sebestyen:2015:EPO

[SST15] Zoltán Sebestyén, Zsolt Szúcs, and Zsigmond Tarcsay. Extensions of positive operators and functionals. Linear Algebra and its Applications, 472 (??):54–80, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000579.

Sikiric:2017:RFC

[SSV17] Mathieu Dutour Sikirić, Achill Schürmann, and Frank Vallentin. Rational factorizations of completely positive matrices. Linear Algebra and its Applications, 523(??):46–51, June 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517300939.

Stanic:2019:BLE

[Sta19] Zoran Stanić. Bounding the largest eigenvalue of signed graphs. Linear Algebra and its Applications, 573(??):80–89, July 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301120.

Stefanovski:2015:RAB

[Ste15] Jovan Stefanovski. A reformulation of augmented basic interpolation problem and an application to H_{∞} control. Linear Algebra and its Applications, 485(??):103– 123, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004322.

Ster:2018:EMZ

[Ste18] Janez Ster. On expressing matrices over Z_{2} as the sum of an idempotent and a nilpotent. Linear Algebra and its Applications, 544 (??):339–349, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518300223.

Strapasson:2018:NSL

[Str18] João Eloir Strapasson. A note on suborthogonal lattices. Linear Algebra and its Applications, 543(??):31– 41, April 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306869.

Stuart:2017:SPS

[Stu17] Jeffrey L. Stuart. Sign kpotent sign patterns and ray k-potent ray patterns that admit k-potence. Linear Algebra and its Applications, 527(??):163–183, August 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730174X.

Smith:2018:DSL

[STW18] Derek A. Smith, Lorenzo Traldi, and William Watkins. Duality and the signed Laplacian matrix of a graph. Linear Algebra and its Applications, 548(??):1–18, July 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518300892.

Smith:2019:NIS

[STW19] Derek A. Smith, Lorenzo Traldi, and William Watkins. A note on 2-isomorphisms and the signed Laplacian matrix of a graph. Linear Algebra and its Applications, 563(??):277–286, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305251.

Szokol:2015:PPG

[STZ15] Patrícia Szokol, Ming-Cheng Tsai, and Jun Zhang. Preserving problems of geodesic-affine maps and related topics on positive definite matrices. Linear Algebra and its Applications, 483(??):293–308, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003626.

Schulze:2018:DDS

[SUBG18] Philipp Schulze, Benjamin Unger, Christopher Beattie, and Serkan Gugercin. Data-driven structured realization. Linear Algebra and its Applications, 537(??): 250–286, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305645.

Sugie:2017:NSO

[Sug17] Jitsuro Sugie. Nonoscillation of second-order linear difference systems with varying coefficients. Linear Algebra and its Applications, 531(??):22–37, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303312.

Solus:2016:EPS

[SUY16] Liam Solus, Caroline Uhler, and Ruriko Yoshida. Extremal positive semidefinite matrices whose sparsity pattern is given by graphs without K_{5} minors. Linear Algebra and its Applications, 509(??):247–275, November 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303263.

Stehlik:2017:SCD

[SV17] Petr Stehlík and Petr Vanek. Sign-changing diagonal perturbations of Laplacian matrices of graphs. Linear Algebra and its Applications, 531(??):64–82, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303439.

Saito:2015:TSP

[SW15] Naoki Saito and Ernest Woei. Tree simplification and the ‘plateaux’ phenomenon of graph Laplacian eigenvalues. Linear Algebra and its Applications, 481(??):263–279, September 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002931.

Sirotkin:2016:AIE

[SW16] Gleb Sirotkin and Ben Wallis. Almost-invariant and essentially-invariant halfspaces. Linear Algebra and its Applications, 507(??): 399–413, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302427.

Somphotphisut:2018:CTZ

[SW18a] Somchai Somphotphisut and Keng Wiboonton. A characterization of trace zero bisymmetric nonnegative 5× 5 matrices. Linear Algebra and its Applications, 540 (??):1–10, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306353.

Somphotphisut:2018:BNI

[SW18b] Somchai Somphotphisut and Keng Wiboonton. On the bisymmetric nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 543(??):42– 71, April 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306870.

Shen:2019:SIM

[SWB19] Jihong Shen, Yue Wang, and Changjiang Bu. Some inequalities on the minimum H-eigenvalue of the Fan product of Z-tensors. Linear Algebra and its Applications, 579(??):55–71, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S002437951930240X.

Szymanski:2018:CJN

[SWZ18] Konrad Szymański, Stephan Weis, and Karol Zyczkowski. Classification of joint numerical ranges of three Hermitian matrices of size three. Linear Algebra and its Applications, 545(??): 148–173, May 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306456.

Stosic:2016:PIC

[SXD16] Marko Stosić, João Xavier, and Marija Dodig. Projection on the intersection of convex sets. Linear Algebra and its Applications, 509(??):191–205, November 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302956.

Shimabukuro:2015:MAA

[SY15] Osamu Shimabukuro and Masayoshi Yoshikawa. Modular adjacency algebras of Grassmann graphs. Linear Algebra and its Applications, 466(??):208–217, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951400682X.

Saburov:2016:CCS

[SY16a] Mansoor Saburov and Nur Atikah Yusof. Counterexamples to the conjecture on stationary probability vectors of the second-order Markov chains. Linear Algebra and its Applications, 507(??): 153–157, October 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302294.

Shao:2016:SPT

[SY16b] Jiayu Shao and Lihua You. On some properties of three different types of triangular blocked tensors. Linear Algebra and its Applications, 511(??):110–140, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303901.

Shao:2017:SPL

[SY17] Jiayu Shao and Xiying Yuan. Some properties of the Laplace and normalized Laplace spectra of uniform hypergraphs. Linear Algebra and its Applications, 531(??):98–117, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303397.

Simson:2017:IAL

[SZ17] Daniel Simson and Katarzyna Zajac. Inflation algorithm for loop-free non-negative edge-bipartite graphs of corank at least two. Linear Algebra and its Applications, 524(??):109–152, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517301209.

Skripka:2019:SUP

[SZ19] Anna Skripka and Maxim Zinchenko. Stability and uniqueness properties of Taylor approximations of matrix functions. Linear Algebra and its Applications, 582(??):218–236, December 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303313.

Shao:2023:NPB

[SZ23] Meiyue Shao and Sizhe Zhang. A note on “Perturbation bounds for Williamson’s symplectic normal form”. Linear Algebra and its Applications, 677(??):194–199, November 15, 2023. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379523002987.

See [IGW17].

Sznt:2016:CDU

[Szá16] András Szántó. Complementary decompositions and unextendible mutually unbiased bases. Linear Algebra and its Applications, 496 (??):392–406, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001014.

Sun:2019:SPG

[SZB19] Lizhu Sun, Jiang Zhou, and Changjiang Bu. Spectral properties of general hypergraphs. Linear Algebra and its Applications, 561(??):

187–203, January 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304701.

Stanimirovic:2016:NNA

[SZW16] Predrag S. Stanimirović, Ivan S. Zivković, and Yimin Wei. Neural network approach to computing outer inverses based on the full rank representation. Linear Algebra and its Applications, 501(??):344–362, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516300465.

Tabata:2016:LBI

[Tab16] Ryo Tabata. Limiting behavior of immanants of certain correlation matrix. Linear Algebra and its Applications, 510(??):230–245, December 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303640.

Tan:2016:FSF

[Tan16a] Yi-Jia Tan. Free sets and free subsemimodules in a semimodule. Linear Algebra and its Applications, 496 (??):527–548, May 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000975.

Tanaka:2016:STP

[Tan16b] Ryotaro Tanaka. The solution of Tingley’s problem for the operator norm unit sphere of complex n × n matrices. Linear Algebra and its Applications, 494 (??):274–285, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000239.

Tanimoto:2019:EM

[Tan19] Ryuji Tanimoto. Exponential matrices. Linear Algebra and its Applications, 572 (??):153–251, July 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300941.

Tarazaga:2018:SSS

[Tar18] Pablo Tarazaga. On the structure of the set of symmetric matrices with the Perron–Frobenius property. Linear Algebra and its Applications, 549(??):219– 232, July 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301253.

Turek:2015:HUM

[TC15] Ondrej Turek and Taksu Cheon. Hermitian unitary matrices with modular permutation symmetry. Linear Algebra and its Applications, 469(??): 569–593, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008015.

Tudisco:2015:CPN

[TCD15] Francesco Tudisco, Valerio Cardinali, and Carmine Di Fiore. On complex power nonnegative matrices. Linear Algebra and its Applications, 471(??): 449–468, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514008246.

Trefois:2015:ZFN

[TD15] Maguy Trefois and Jean-Charles Delvenne. Zero forcing number, constrained matchings and strong structural controllability. Linear Algebra and its Applications, 484(??):199–218, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500378X.

Terwilliger:2015:LRT

[Ter15] Paul Terwilliger. Loweringraising triples and U_{q}(sl_{2}). Linear Algebra and its Applications, 486(??):1–172, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004735.

Terwilliger:2017:LTR

[Ter17a] Paul Terwilliger. Leonard triples of q-Racah type and their pseudo intertwiners. Linear Algebra and its Applications, 515(??):145–174, February 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305456.

Terwilliger:2017:OAP

[Ter17b] Paul Terwilliger. The q-Onsager algebra and the positive part of U_{q}( ˆ∫↕_{2}). Linear Algebra and its Applications, 521(??):19–56, May 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379517300563.

Terwilliger:2018:PWM

[Ter18] Paul Terwilliger. A poset Φ_{n} whose maximal chains are in bijection with the n × n alternating sign matrices. Linear Algebra and its Applications, 554(??): 79–85, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302647.

Tan:2019:PST

[TFC19] Ying-Ying Tan, Keqin Feng, and Xiwang Cao. Perfect state transfer on abelian Cayley graphs. Linear Algebra and its Applications, 563(??):331–352, February 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518305342.

Turek:2019:GCH

[TG19] Ondrej Turek and Dardo Goyeneche. A generalization of circulant Hadamard and conference matrices. Linear Algebra and its Applications, 569(??):241–265, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379519300424.

Tahaei:2016:CRB

[TH16a] Maedeh S. Tahaei and Seyed Naser Hashemi. The coefficients of the reduced Bartholdi zeta function. Linear Algebra and its Applications, 509(??):1–18, November 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951630266X.

Tam:2016:NSR

[TH16b] Bit-Shun Tam and Peng-Ruei Huang. Nonnegative square roots of matrices. Linear Algebra and its Applications, 498(??):404– 440, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515006746.

Tam:2017:NGG

[TH17] Bit-Shun Tam and Tsu-Hsien Huang. Nullities of graphs with given order, matching number and cyclomatic number revisited. Linear Algebra and its Applications, 535(??):105–140, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305128.

Teng:2015:CRA

[TLL15] Zhongming Teng, Linzhang Lu, and Ren-Cang Li. Cluster-robust accuracy bounds for Ritz subspaces. Linear Algebra and its Applications, 480(??):11–26, September 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002542.

Takemura:2016:TTD

[TNK16] Kazuo Takemura, Atsushi Nagai, and Yoshinori Kametaka. Two types of discrete Sobolev inequalities on a weighted Toeplitz graph. Linear Algebra and its Applications, 507(??):344–355, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302464.

Tokcan:2017:WRB

[Tok17]  Neriman Tokcan.  On

the Waring rank of binary forms. Linear Algebra and its Applications, 524 (??):250–262, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301593.

Tomon:2015:CNR

[Tom15] István Tomon. On the chromatic number of regular graphs of matrix algebras. Linear Algebra and its Applications, 475(??): 154–162, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001378.

Tavakolipour:2018:TET

[TS18] Hanieh Tavakolipour and Fatemeh Shakeri. On tropical eigenvalues of tridiagonal Toeplitz matrices. Linear Algebra and its Applications, 539(??):198–218, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306377.

Tsatsomeros:2016:GMP

[Tsa16a] M. J. Tsatsomeros. Geometric mapping properties of semipositive matrices. Linear Algebra and its Applications, 498(??):349– 359, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004267.

Tsatsomeros:2016:BRM

[Tsa16b] Michael Tsatsomeros. Book review: Matrices: Algebra, Analysis and Applications, Shmuel Friedland. World Scientific Publishing Co., New Jersey (2016), ISBN: 978-981-4667-96-8. Linear Algebra and its Applications, 508(??):301–303, November 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303251.

Topcu:2016:SCP

[TSH16] Hatice Topcu, Sezer Sorgun, and Willem H. Haemers. On the spectral characterization of pineapple graphs. Linear Algebra and its Applications, 507(??):267–273, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630235X.

Tsukerman:2015:TST

[Tsu15] Emmanuel Tsukerman. Tropical spectral theory of tensors. Linear Algebra and its Applications, 481(??):94– 106, September 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002815.

Turnsek:2017:ROS

[Tur17] Aleksej Turnsek. A remark on orthogonality and symmetry of operators in B(H). Linear Algebra and its Applications, 535(??):141– 150, December 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730513X.

Trefois:2015:BFM

[TVD15] Maguy Trefois, Paul Van Dooren, and Jean-Charles Delvenne. Binary factorizations of the matrix of all ones. Linear Algebra and its Applications, 468 (??):63–79, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514000299.

Tao:2016:GCS

[TW16] J. Tao and Guoqiang Wang. A generalization of the Craig–Sakamoto theorem to Euclidean Jordan algebras. Linear Algebra and its Applications, 493(??):134–145, March 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500720X.

Tian:2017:GWD

[TW17] Fenglei Tian and Dein Wong. Graphs whose distance matrix has at most three negative eigenvalues. Linear Algebra and its Applications, 530(??):470–484, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Tian:2018:RBM

[TW18a] Fenglei Tian and Dein Wong. Relation between the matching number and the second largest distance Laplacian eigenvalue of a graph. Linear Algebra and its Applications, 558(??): 174–185, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304075.

Townsend:2018:SVM

[TW18b] Alex Townsend and Heather Wilber. On the singular values of matrices with high displacement rank. Linear Algebra and its Applications, 548(??):19–41, July 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301009.

Tsatsomeros:2019:SM

[TW19] Michael J. Tsatsomeros and Megan Wendler. Semimonotone matrices. Linear Algebra and its Applications, 578(??):207–224, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302113.

Tian:2016:LBD

[TWM16] Fenglei Tian, Dein Wong, and Xiaobin Ma. Lower bounds of distance Laplacian spectral radii of n-vertex graphs in terms of matching number. Linear Algebra and its Applications, 506(??): 579–587, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302452.

Tian:2015:PFC

[TWR15] Fenglei Tian, Dein Wong, and Jianling Rou. Proof for four conjectures about the distance Laplacian and distance signless Laplacian eigenvalues of a graph. Linear Algebra and its Applications, 471(??):10–20, April 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379514008180.

Tang:2018:WEL

[TXZ18] Gaohua Tang, Guoli Xia, and Yiqiang Zhou. When is every linear transformation a sum of an idempotent one and a locally nilpotent one? Linear Algebra and its Applications, 543 (??):226–233, April 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Tam:2015:KFR

[TY15] Tin-Yau Tam and Wen Yan. On Ky Fan’s result on eigenvalues and real singular values of a matrix. Linear Algebra and its Applications, 468(??):3–17, March 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379513006095.

Tyaglov:2017:SIP

[Tya17]  Mikhail Tyaglov.  Self-

interlacing polynomials. Linear Algebra and its Applications, 535(??):12–34, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304962.

Tygert:2019:RAD

[Tyg19] Mark Tygert. Regressionaware decompositions. Linear Algebra and its Applications, 565(??):208–224, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305780.

Trombetti:2019:NAG

[TZ19] Rocco Trombetti and Yue Zhou. Nuclei and automorphism groups of generalized twisted Gabidulin codes. Linear Algebra and its Applications, 575(??):1– 26, August 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301624.

Tang:2017:MTI

[TZK17] Yongsheng Tang, Shixin Zhu, and Xiaoshan Kai. MacWilliams type identities on the Lee and Euclidean weights for linear codes over Z_{ℓ}. Linear Algebra and its Applications, 516 (??):82–92, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305602.

Uchiyama:2015:OIG

[Uch15] Mitsuru Uchiyama. Operator inequalities: From a general theorem to concrete inequalities. Linear Algebra and its Applications, 465(??):161–175, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006065.

Urure:2018:IIU

[UG18] Ronald Ismael Quispe Urure and Dimas José Gonçalves. Identities with involution for 2 × 2 upper triangular matrices algebra over a finite field. Linear Algebra and its Applications, 544 (??):223–253, May 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518300132.

Khan:2015:SRC

[uIKF15] Murad ul Islam Khan and Yi-Zheng Fan. On the spectral radius of a class of non-odd-bipartite even uniform hypergraphs. Linear Algebra and its Applications, 480(??):93–106, September 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002207.

Ulukok:2019:MIP

[Ulu19]  Zübeyde Ulukök.  More

inequalities for positive semidefinite 2 × 2 block matrices and their blocks. Linear Algebra and its Applications, 572(??):51–67, July 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S002437951930103X.

Usevich:2016:ALS

[UM16] Konstantin Usevich and Ivan Markovsky. Adjusted least squares fitting of algebraic hypersurfaces. Linear Algebra and its Applications, 502(??):243–274, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004371.

Urschel:2018:NDG

[Urs18] John C. Urschel. Nodal decompositions of graphs. Linear Algebra and its Applications, 539(??):60–71, February 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306298.

Uchiyama:2015:SDO

[UUG15] Mitsuru Uchiyama, Atsushi Uchiyama, and Mariko Giga. Superadditivity and derivative of operator functions. Linear Algebra and its Applications, 465(??):401–411, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514005916.

Udagawa:2015:FOM

[UWYY15] Yoichi Udagawa, Shuhei Wada, Takeaki Yamazaki, and Masahiro Yanagida. On a family of operator means involving the power difference means. Linear Algebra and its Applications, 485(??):124–131, November 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004140.

Udagawa:2017:SPW

[UYY17] Yoichi Udagawa, Takeaki Yamazaki, and Masahiro Yanagida. Some properties of weighted operator means and characterizations of interpolational means. Linear Algebra and its Applications, 517(??): 217–234, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306061.

Uhlig:2019:TVM

[UZ19] Frank Uhlig and Yunong Zhang. Time-varying matrix eigenanalyses via Zhang Neural Networks and lookahead finite difference equations. Linear Algebra and its Applications, 580(??): 417–435, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302800.

VanBarel:2016:DRF

[Van16a] Marc Van Barel. Designing rational filter functions for solving eigenvalue problems by contour integration. Linear Algebra and its Applications, 502(??): 346–365, August 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003481.

VandeVoorde:2016:DSF

[Van16b] Geertrui Van de Voorde. Desarguesian spreads and field reduction for elements of the semilinear group. Linear Algebra and its Applications, 507(??):96–120, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516302130.

Vannieuwenhoven:2017:CNT

[Van17] Nick Vannieuwenhoven. Condition numbers for the tensor rank decomposition. Linear Algebra and its Applications, 535(??):35–86, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304901.

Vassiliou:2015:PNH

[Vas15] P.-C. G. Vassiliou. On the periodicity of nonhomogeneous Markov chains and systems. Linear Algebra and its Applications, 471 (??):654–684, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000464.

Van-Brunt:2018:IDN

[VB18] Alexander Van-Brunt. Infinitely divisible nonnegative matrices, M-matrices, and the embedding problem for finite state stationary Markov chains. Linear Algebra and its Applications, 541(??):163–176, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306468.

Vos:2019:BTU

[VB19] Alexis De Vos and Stijn De Baerdemacker. The Birkhoff theorem for unitary matrices of prime-power dimension. Linear Algebra and its Applications, 578 (??):27–52, October 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302095.

Dooren:2018:RPM

[VD18] Paul Van Dooren and Froilán M. Dopico. Robustness and perturbations of minimal bases. Linear Algebra and its Applications, 542(??):246–281, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302999.

vanDam:2016:NBM

[vDS16] E. R. van Dam and R. Sotirov. New bounds for the max-k-cut and chromatic number of a graph. Linear Algebra and its Applications, 488(??): 216–234, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005704.

Vecharynski:2016:GSB

[Vec16] Eugene Vecharynski. A generalization of Saad’s bound on harmonic Ritz vectors of Hermitian matrices. Linear Algebra and its Applications, 494(??):219–235, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S002437951600015X.

Vanegas:2017:NLS

[VF17] E. O. Quintero Vanegas and Juan C. Gutierrez Fernandez. Nilpotent linear spaces and Albert’s Problem. Linear Algebra and its Applications, 518(??):57– 78, April 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306152.

Vieira:2018:ABS

[VG18] Evilson Vieira and Ronaldo Garcia. Asymptotic behavior of the shape of planar polygons by linear flows. Linear Algebra and its Applications, 557(??):508–528, November 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303951.

Vandaele:2017:LEC

[VGG17] Arnaud Vandaele, Nicolas Gillis, and François Glineur. On the linear extension complexity of regular ngons. Linear Algebra and its Applications, 521(??): 217–239, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306127.

Veerman:2018:SCL

[VHB18] J. J. P. Veerman, David K. Hammond, and Pablo E. Baldivieso. Spectra of certain large tridiagonal matrices. Linear Algebra and its Applications, 548(??):123– 147, July 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301071.

Vijayakumar:2015:RGR

[Vij15] G. R. Vijayakumar. A result on generating root systems by linearly independent subsets. Linear Algebra and its Applications, 478(??): 228–240, August 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515002104.

Virosztek:2016:QDP

[Vir16] Dániel Virosztek. Quantum f-divergence preserving maps on positive semidefinite operators acting on finite dimensional Hilbert spaces. Linear Algebra and its Applications, 501(??): 242–253, July 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300404.

Virosztek:2019:JCQ

[Vir19] Dániel Virosztek. Jointly convex quantum Jensen divergences. Linear Algebra and its Applications, 576(??):67–78, September 1, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518301046.

Vecharynski:2016:PSD

[VK16] Eugene Vecharynski and Andrew Knyazev. Preconditioned steepest descentlike methods for symmetric indefinite systems. Linear Algebra and its Applications, 511(??):274–295, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304104.

Veerman:2019:DCW

[VK19] J. J. P. Veerman and Ewan Kummel. Diffusion and consensus on weakly connected directed graphs. Linear Algebra and its Applications, 578(??):184–206, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S002437951930223X.

Vera-Lopez:2015:CCC

[VLGS15] A. Vera-López and M. A. García-Sánchez. Counting conjugacy classes for unitriangular matrices. Linear Algebra and its Applications, 486(??):275–294, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004826.

Volcic:2017:DNR

[Vol17] Jurij Volcic. On domains of noncommutative rational functions. Linear Algebra and its Applications, 516 (??):69–81, March 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516305572.

Van:2017:MRC

[VQ17] Hieu Ha Van and Rachel Quinlan. On the maximum rank of completions of entry pattern matrices. Linear Algebra and its Applications, 525(??):1–19, July 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301362.

Van:2019:ANE

[VQ19] Hieu Ha Van and Rachel Quinlan. Almost-nonsingular entry pattern matrices. Linear Algebra and its Applications, 578(??):334–355, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

Varghese:2018:DEC

[VSV18] Anu Varghese, Wasin So, and A. Vijayakumar. Distance energy change of complete bipartite graph due to edge deletion. Linear Algebra and its Applications, 553(??):211–222, September 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302374.

VanBarel:2018:PES

[VT18] Marc Van Barel and Françoise Tisseur. Polynomial eigenvalue solver based on tropically scaled Lagrange linearization. Linear Algebra and its Applications, 542 (??):186–208, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302628.

VanderMeulen:2018:BPR

[VV18] Kevin N. Vander Meulen and Trevor Vanderwoerd. Bounds on polynomial roots using intercyclic companion matrices. Linear Algebra and its Applications, 539(??): 94–116, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306286.

vanWyk:2017:LSI

[vWZ17] Leon van Wyk and Michal Ziembowski. Lie solvability and the identity [x_{1}, y_{1}][x_{2}, y_{2}] ⋯ [x_{q}, y_{q}] = 0 in certain matrix algebras. Linear Algebra and its Applications, 533(??):235– 257, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304500.

Wada:2018:WDA

[Wad18] Shuhei Wada. When does Ando–Hiai inequality hold? Linear Algebra and its Applications, 540(??):234–243, March 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306584.

Wang:2015:NAZ

[Wan15a] Long Wang. A note on automorphisms of the zerodivisor graph of upper triangular matrices. Linear Algebra and its Applications, 465(??):214–220, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006326.

Wang:2015:FID

[Wan15b] Yu Wang. Functional identities of degree 2 in arbitrary triangular rings. Linear Algebra and its Applications, 479(??):171–184, August 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002566.

Wang:2016:CED

[Wan16a] Hongxing Wang. Core-EP decomposition and its applications. Linear Algebra and its Applications, 508(??): 289–300, November 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516303354.

Wang:2016:ICP

[Wan16b] Wei Wang. Identifiability of covariance parameters in linear mixed effects models. Linear Algebra and its Applications, 506(??): 603–613, October 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302397.

Wang:2016:CCT

[Wan16c] Yu Wang. Commuting (centralizing) traces and Lie (triple) isomorphisms on triangular algebras revisited. Linear Algebra and its Applications, 488(??): 45–70, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005662.

Wang:2019:LAQ

[Wan19] Wei Wang. On the linear algebra in the quaternionic pluripotential theory. Linear Algebra and its Applications, 562(??): 223–241, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304981.

Warrington:2017:OBT

[War17] Gregory S. Warrington. Orthogonal bases for transportation polytopes applied to latin squares, magic squares and Sudoku boards. Linear Algebra and its Applications, 531(??):285–304, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303701.

Wu:2015:COT

[WC15] Sheng-Jhih Wu and Moody T. Chu. Constructing optimal transition matrix for Markov chain Monte Carlo.

Linear Algebra and its Applications, 487(??):184–202, December 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005315.

Wang:2017:ONI

[WDFS17] Miaoyan Wang, Khanh Dao Duc, Jonathan Fischer, and Yun S. Song. Operator norm inequalities between tensor unfoldings on the partition lattice. Linear Algebra and its Applications, 520(??):44–66, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300459.

Wang:2017:KCE

[WDV17] Xiangrong Wang, Johan L. A. Dubbeldam, and Piet Van Mieghem. Kemeny’s constant and the effective graph resistance. Linear Algebra and its Applications, 535(??):231–244, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305256.

Wei:2015:IEP

[Wei15] Ying Wei. Inverse eigenvalue problem of Jacobi matrix with mixed data. Linear Algebra and its Applications, 466(??):102–116, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006168.

Wei:2017:APG

[Wei17] Yimin Wei. Acute perturbation of the group inverse. Linear Algebra and its Applications, 534(??): 135–157, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304858.

Wang:2019:UBS

[WF19] Long Wang and Xianwen Fang. Upper bound of skew energy of an oriented graph in terms of its skew rank. Linear Algebra and its Applications, 569(??): 195–205, May 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519300448.

Wen:2019:IOF

[WFS19] Shilin Wen, Junsheng Fang, and Rui Shi. On irreducible operators in factor von Neumann algebras. Linear Algebra and its Applications, 565(??):239–243, March 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305779.

Wang:2015:ASB

[WG15] Rui Wang and You Gao. Association schemes based on the subspaces of type (2, 1, 0) in singular symplectic space over finite fields. Linear Algebra and its Applications, 483(??):249–267, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515003559.

Wang:2018:SLE

[WH18] Dijian Wang and Yaoping Hou. On the sum of Laplacian eigenvalues of a signed graph. Linear Algebra and its Applications, 555(??): 39–52, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302891.

Wang:2019:MUP

[WH19a] Ruidong Wang and Xujian Huang. The Mazur– Ulam property for twodimensional somewhere-flat spaces. Linear Algebra and its Applications, 562(??): 55–62, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304658.

Wu:2019:CFC

[WH19b] Senlin Wu and Chan He. Covering functionals of convex polytopes. Linear Algebra and its Applications, 577(??):53–68, September 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519301806.

Wu:2017:SAP

[WHC17] Xiufeng Wu, Junjie Huang, and Alatancang Chen. Selfadjoint perturbations of spectra for upper triangular operator matrices. Linear Algebra and its Applications, 531(??):1–21, October 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303051.

Wang:2017:LTE

[WHG17] Yan Wang, Bo Hou, and Suogang Gao. Leonard triples extended from a given totally almost bipartite Leonard pair of Bannai/Itô type. Linear Algebra and its Applications, 519 (??):111–135, April 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306383.

Wimmer:2016:CIS

[Wim16] Harald K. Wimmer. Contour integral solutions of Sylvester-type matrix equations. Linear Algebra and its Applications, 493(??): 537–543, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000021.

Winkler:2016:PCB

[Win16] Joab R. Winkler. Polynomial computations for blind image deconvolution. Linear Algebra and its Applications, 502(??):77–103, August 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006059.

Wang:2015:MUI

[WL15] Xue-Feng Wang and Ren-Cang Li. Monotonicity of unitarily invariant norms. Linear Algebra and its Applications, 466(??):254–266, February 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514006910.

Wang:2016:POB

[WL16] Hongxing Wang and Xiaoji Liu. Partial orders based on core-nilpotent decomposition. Linear Algebra and its Applications, 488(??): 235–248, January 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951500573X.

Wei:2018:RGL

[WLH18] Xiaohan Wei, Qing Ling, and Zhu Han. Robust group lasso: Model and recoverability. Linear Algebra and its Applications, 557(??):134–173, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379518303185.

Wang:2019:IEB

[WLL19] Zhifeng Wang, Chaoqian Li, and Yaotang Li. Infimum of error bounds for linear complementarity problems of Σ-SDD and Σ_{1}-SSD matrices. Linear Algebra and its Applications, 581(??):285– 303, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303076.

Wang:2016:PEN

[WLP16] Xuefeng Wang, Chi-Kwong Li, and Yiu-Tung Poon. Perturbing eigenvalues of nonnegative matrices. Linear Algebra and its Applications, 498(??):3–20, June 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514002262.

Wang:2019:CPR

[WLWZ19] Guangfu Wang, Shuchao Li, Wei Wei, and Siqi Zhang. On the characteristic polynomials and H-ranks of the weighted mixed graphs. Linear Algebra and its Applications, 581(??):383–404, November 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519303143.

Wang:2017:BGE

[WM17a] Long Wang and Xiaobin Ma. Bounds of graph energy in terms of vertex cover number. Linear Algebra and its Applications, 517(??): 207–216, March 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306048.

Wang:2017:RGS

[WM17b] Wei Wang and Lihuan Mao. A remark on the generalized spectral characterization of the disjoint union of graphs. Linear Algebra and its Applications, 518 (??):1–13, April 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516306139.

Wong:2015:NMU

[WMC15] Dein Wong, Xiaobin Ma, and Li Chen. Nonlinear mappings on upper triangular matrices derivable at zero point. Linear Algebra and its Applications, 483(??): 236–248, October 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003596.

Wickman:2017:DGP

[WO17] Clare Wickman and Kasso Okoudjou. Duality and geodesics for probabilistic frames. Linear Algebra and its Applications, 532(??): 198–221, November 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303348.

Wojcik:2015:OPS

[Wój15]  Paweł Wójcik.  Opera-

tors preserving sesquilinear form. Linear Algebra and its Applications, 469(??): 531–538, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007988.

Wojcik:2019:CLS

[Wój19] Paweł Wójcik. Characterization of linear similarities through functional equation and mappings preserving orthogonalities. Linear Algebra and its Applications, 579(??):206–216, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302411.

Wolf:2018:EGF

[Wol18] Reinhard Wolf. Estimating the gap of finite metric spaces of strict p-negative type. Linear Algebra and its Applications, 556(??): 171–199, November 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303239.

Wu:2017:CEJ

[WP17] Gang Wu and Hong-Kui Pang. On the correction equation of the Jacobi– Davidson method. Linear Algebra and its Applications, 522(??):51–70, June 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300836.

Wang:2019:CGG

[WQH19] Wei Wang, Lihong Qiu, and Yulin Hu. Cospectral graphs, GM-switching and regular rational orthogonal matrices of level p. Linear Algebra and its Applications, 563(??):154–177, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305159.

Wang:2018:ISS

[WT18] Cailu Wang and Yuegang Tao. Interval strong solutions of interval systems of max-plus linear equations. Linear Algebra and its Applications, 537(??): 148–159, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305700.

Wang:2018:CPL

[WW18] Zhijun Wang and Dein Wong. The characteristic polynomial of lexicographic product of graphs. Linear Algebra and its Applications, 541(??):177–184, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517306675.

Wong:2018:LBG

[WWC18] Dein Wong, Xinlei Wang, and Rui Chu. Lower bounds of graph energy in terms of matching number. Linear Algebra and its Applications, 549(??):276– 286, July 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518301502.

Wang:2018:SGC

[WWT18] Xinlei Wang, Dein Wong, and Fenglei Tian. Signed graphs with cut points whose positive inertia indexes are two. Linear Algebra and its Applications, 539(??): 14–27, February 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305384.

Xu:2019:IEP

[wXjMZL19] Wei wei Xu, Li juan Ma, Lei Zhu, and Hao Liu. On interval estimates of perturbations of generalized eigenvalues for diagonalizable pairs. Linear Algebra and its Applications, 562(??): 15–43, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304415.

Xu:2014:MGM

[wXL14] Wei wei Xu and Hao Liu. A modified general modulus-based matrix splitting method for linear complementarity problems of Hmatrices. Linear Algebra and its Applications, 458(??): 626–637, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514003887.

See corrigendum [wXZ19].

Xu:2019:CSM

[wXZ19] Wei wei Xu and Lei Zhu. Corrigendum to “A modified general modulus-based matrix splitting method for linear complementarity problems of H-matrices” [Linear Algebra Appl. 458 (2014) 626–637]. Linear Algebra and its Applications, 561(??):204–206, January 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/

science/article/pii/S0024379518304191.

See [wXL14].

Wang:2015:CNN

[WYL15] Shaoxin Wang, Hu Yang, and Hanyu Li. Condition numbers for the nonlinear matrix equation and their statistical estimation. Linear Algebra and its Applications, 482(??):221–240, October 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500364X.

Wang:2019:RGP

[WYL19] Jie Wang, Xiying Yuan, and Lele Liu. Regular graphs with a prescribed complete multipartite graph as a star complement. Linear Algebra and its Applications, 579(??):302–319, October 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302563.

Wang:2017:MGH

[WYLW17] Yi Wang, Bo-Jun Yuan, Shuang-Dong Li, and Chong-Jun Wang. Mixed graphs with H-rank 3. Linear Algebra and its Applications, 524(??):22–34, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301386.

Wu:2015:PSC

[WZ15] Tingzeng Wu and Heping Zhang. Per-spectral characterizations of graphs with extremal per-nullity. Linear Algebra and its Applications, 484(??):13–26, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003717.

Wang:2017:PEA

[WZ17a] Xing Tao Wang and Lei Zhang. Partial eigenvalue assignment with time delay in high order system using the receptance. Linear Algebra and its Applications, 523 (??):335–345, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517301350.

Wei:2017:ICM

[WZ17b] Yangjiang Wei and Yi Ming Zou. Inverses of Cartan matrices of Lie algebras and Lie superalgebras. Linear Algebra and its Applications, 521 (??):283–298, May 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517300654.

Wang:2019:RAR

[WZ19] Weiping Wang and Chenlu Zhang. Riordan arrays and related polynomial sequences. Linear Algebra and its Applications, 580(??): 262–291, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519302605.

Wen:2015:LSL

[WZL15] Qin Wen, Qin Zhao, and Huiqing Liu. The least signless Laplacian eigenvalue of non-bipartite graphs with given stability number. Linear Algebra and its Applications, 476(??):148– 158, July 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007010.

Wang:2017:EBA

[WZL17] Wei-Guo Wang, Lei-Hong Zhang, and Ren-Cang Li. Error bounds for approximate deflating subspaces for linear response eigenvalue problems. Linear Algebra and its Applications, 528(??):273–289, September 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516303603.

Wang:2019:LBC

[WZW19] Yan Wang, LianZhu Zhang, and ShouLiu Wei. Length bounds for cycle bases of graphs. Linear Algebra and its Applications, 560(??): 146–154, January 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518304737.

Wang:2019:SLC

[WZZ19] Wen-Huan Wang, Lei Zhong, and Lian-Jiang Zheng. The signless Laplacian coefficients and the incidence energy of the graphs without even cycles. Linear Algebra and its Applications, 563(??):476–493, February 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518305135.

Xie:2018:SNI

[XCJ18] Ze-Jia Xie, Che-Man Cheng, and Xiao-Qing Jin. Some norm inequalities for commutators of contracted tensor products. Linear Algebra and its Applications, 540 (??):244–256, March 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306614.

Xu:2018:CTC

[XCQ18] Changqing Xu, Zhibing Chen, and Liqun Qi. On {0, 1} CP tensors and CP pseudographs. Linear Algebra and its Applications, 557(??):287–306, November 15, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379518303732.

Xu:2017:NMR

[XF17] Qingxiang Xu and Xiaochun Fang. A note on majorization and range inclusion of adjointable operators on Hilbert C^{∗}-modules. Linear Algebra and its Applications, 516(??):118–125, March 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305511.

Xu:2017:LMP

[XFZD17] Jinli Xu, Ajda Fosner, Baodong Zheng, and Yuting Ding. Linear maps preserving r-potents of tensor products of matrices. Linear Algebra and its Applications, 520(??):67–76, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300447.

Xiao:2018:IRP

[XL18] Chuanfu Xiao and Hanyu Li. Improved random perturbation intervals of symmetric eigenvalue problem. Linear Algebra and its Applications, 541(??):1– 12, March 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517306547.

Xue:2018:SRG

[XLLS18] Jie Xue, Huiqiu Lin, Shuting Liu, and Jinlong Shu. On the A_{α}-spectral radius of a graph. Linear Algebra and its Applications, 550(??): 105–120, August 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518301484.

Xu:2016:CPT

[XLQC16] Changqing Xu, Ziyan Luo, Liqun Qi, and Zhibing Chen. {0, 1} completely positive tensors and multihypergraphs. Linear Algebra and its Applications, 510(??):110–123, December 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516303536.

Xu:2017:PSC

[XS17] Guixin Xu and Yuming Shi. Perturbations of spectra of closed subspaces in Banach spaces. Linear Algebra and its Applications, 531(??): 547–574, October 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303762.

Xue:2018:PCD

[XS18] Jie Xue and Jinlong Shu. Proof of a conjecture on the distance Laplacian spectral radius of graphs. Linear Algebra and its Applications, 540(??):84–94, March 1, 2018. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517306365.

Xu:2017:MPM

[XSW17] Qingxiang Xu, Chuanning Song, and Guorong Wang. Multiplicative perturbations of matrices and the generalized triple reverse order law for the Moore–Penrose inverse. Linear Algebra and its Applications, 530(??): 366–383, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517303142.

Xi:2019:DSL

[XSW19a] Weige Xi, Wasin So, and Ligong Wang. The (distance) signless Laplacian spectral radius of digraphs with given arc connectivity. Linear Algebra and its Applications, 581(??):85– 111, November 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302915.

Xi:2019:GDM

[XSW19b] Weige Xi, Wasin So, and Ligong Wang. The generalized distance matrix of digraphs. Linear Algebra and its Applications, 577(??):270–286, September 15, 2019. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379519302022.

Xu:2016:HTV

[Xu16] Changqing Xu. Hankel tensors, Vandermonde tensors and their positivities. Linear Algebra and its Applications, 491(??):56–72, February 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001184.

Xu:2017:PDS

[Xu17a] Bangteng Xu. Pseudo-direct sums and wreath products of loose-coherent algebras with applications to coherent configurations. Linear Algebra and its Applications, 530(??):202–219, October 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517302975.

Xu:2017:IUB

[Xu17b] Xuefeng Xu. An improved upper bound for the number of distinct eigenvalues of a matrix after perturbation. Linear Algebra and its Applications, 523(??): 109–117, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301234.

Xu:2015:ACC

[XvdBvdLS15] Genjiu Xu, René van den Brink, Gerard van der Laan, and Hao Sun. Associated consistency characterization of two linear values for TU games by matrix approach. Linear Algebra and its Applications, 471 (??):224–240, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000282.

Xie:2016:SFG

[XW16] Pengpeng Xie and Yimin Wei. The stability of formulae of the Gohberg– Semencul–Trench type for Moore–Penrose and group inverses of Toeplitz matrices. Linear Algebra and its Applications, 498(??):117– 135, June 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515000580.

Xiao:2018:FTL

[XWD18] Peng Xiao, Ligong Wang, and Yanfei Du. The first two largest spectral radii of uniform supertrees with given diameter. Linear Algebra and its Applications, 536(??):103–119, January 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305335.

Xiao:2017:MSR

[XWL17] Peng Xiao, Ligong Wang, and Yong Lu. The maximum spectral radii of uniform supertrees with given degree sequences. Linear Algebra and its Applications, 523 (??):33–45, June 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300952.

Xu:2012:TLC

[XX12] Chuang Xu and Runzhang Xu. Tripotency of a linear combination of two involutory matrices and a tripotent matrix that mutually commute. Linear Algebra and its Applications, 437(9):2091– 2109, November 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379512004399.

See corrigendum [Kis15].

Xie:2016:TSM

[XXZ16] Shui-Lian Xie, Hong-Ru Xu, and Jin-Ping Zeng. Twostep modulus-based matrix splitting iteration method for a class of nonlinear complementarity problems. Linear Algebra and its Applications, 494(??):1–10, April 1, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516000045.

Yamagishi:2015:PHF

[Yam15] Masakazu Yamagishi. Periodic harmonic functions on lattices and Chebyshev polynomials. Linear Algebra and its Applications, 476(??):1–15, July 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515001482.

Yang:2016:UTM

[Yan16] Yang Yang. Upper triangular matrices and Billiard Arrays. Linear Algebra and its Applications, 493(??): 508–536, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515007600.

Yang:2018:FDI

[Yan18] Yang Yang. Finite-dimensional irreducible □_{q}-modules and their Drinfel’d polynomials. Linear Algebra and its Applications, 537(??): 160–190, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517305712.

Yaskov:2016:CLE

[Yas16] Pavel Yaskov. Controlling the least eigenvalue of a random Gram matrix. Linear Algebra and its Applications, 504(??):108–123, September 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

Yuan:2017:CGS

[YCL17] Xiying Yuan, Hongyan Chen, and Lele Liu. On the characterization of graphs by star complements. Linear Algebra and its Applications, 533(??):491–506, November 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304676.

Yan:2019:CSP

[YdB19] Dan Yan and Michiel de Bondt. The classification of some polynomial maps with nilpotent Jacobians. Linear Algebra and its Applications, 565(??): 287–308, March 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518305834.

Yang:2018:UAC

[YDHX18] Sheng-Liang Yang, Yan-Ni Dong, Tian-Xiao He, and Yan-Xue Xu. A unified approach for the Catalan matrices by using Riordan arrays. Linear Algebra and its Applications, 558(??): 25–43, December 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830377X.

Yang:2018:HRA

[YDYY18] Sheng-Liang Yang, Yan-Ni Dong, Lin Yang, and Juan Yin. Half of a Riordan array and restricted lattice paths. Linear Algebra and its Applications, 537 (??):1–11, January 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951730561X.

Ye:2017:NCM

[Ye17] Ke Ye. New classes of matrix decompositions. Linear Algebra and its Applications, 514(??):47–81, February 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304979.

Yuan:2015:NTL

[YG15] Xi-Ying Yuan and Ji-Ming Guo. Note on the k th Laplacian eigenvalues of trees with perfect matchings. Linear Algebra and its Applications, 483(??):115–127, October 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003493.

Yuan:2015:FRS

[YHY15] Pingzhi Yuan, Zilong He, and Lihua You. Further results and some open problems on the primitive degree of nonnegative tensors. Linear Algebra and its Applications, 480(??): 72–92, September 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515002244.

Yin:2018:CLT

[Yin18] Yanqing Yin. Central limit theorem for linear spectral statistics of large dimensional quaternion sample covariance matrices. Linear Algebra and its Applications, 554(??):275–315, October 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302714.

Yin:2019:HFA

[Yin19] Guojian Yin. A harmonic FEAST algorithm for non-Hermitian generalized eigenvalue problems. Linear Algebra and its Applications, 578 (??):75–94, October 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519301946.

Yang:2017:SJM

[YL17] Chaojun Yang and Fangyan Lu. Surjective Jordan maps and Jordan triple maps. Linear Algebra and its Applications, 535(??):1–11, December 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517304986.

Yao:2015:TPH

[YLB15] Hongmei Yao, Lei Liu, and Changjiang Bu. Tensors product and hyperdeterminant of boundary formats product. Linear Algebra and its Applications, 483 (??):1–20, October 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003419.

Yang:2016:APL

[YLT16] Yu Yang, Denny H. Leung, and Wai-Shing Tang. All 2-positive linear maps from M_{3}(C) to M_{3}(C) are decomposable. Linear Algebra and its Applications, 503(??): 233–247, August 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300738.

Yu:2019:TPK

[YMZ19] Ying-Ying Yu, Hui Ma, and Chun-Gang Zhu. Total positivity of a kind of generalized Toric-Bernstein basis. Linear Algebra and its Applications, 579(??): 449–462, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302642.

Yordzhev:2017:CFS

[Yor17]  Krasimir Yordzhev.  On

the cardinality of a factor set of binary matrices. Linear Algebra and its Applications, 534(??): 122–134, December 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304937.

Youssef:2015:EBF

[You15] Pierre Youssef. Extracting a basis with fixed block inside a matrix. Linear Algebra and its Applications, 469 (??):28–38, March 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514007502.

Yuan:2016:PCL

[YQS16] Xiying Yuan, Liqun Qi, and Jiayu Shao. The proof of a conjecture on largest Laplacian and signless Laplacian H-eigenvalues of uniform hypergraphs. Linear Algebra and its Applications, 490(??):18–30, February 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006084.

Shu:2017:NFS

[ySpW17] Qian yu Shu and Xue ping Wang. Note on the free sets and free subsemimodules in semimodules. Linear Algebra and its Applications, 520 (??):125–133, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517300289.

Yuan:2016:OSU

[YSS16] Xiying Yuan, Jiayu Shao, and Haiying Shan. Ordering of some uniform supertrees with larger spectral radii. Linear Algebra and its Applications, 495 (??):206–222, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000343.

Yuan:2017:EPS

[YSZ17] Yaping Yuan, Jiong Sun, and Anton Zettl. Eigenvalues of periodic Sturm Liouville problems. Linear Algebra and its Applications, 517(??):148–166, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305614.

Yan:2016:PMN

[YT16] Dan Yan and Guoping Tang. Polynomial maps with nilpotent Jacobians in dimension three. Linear Algebra and its Applications, 489(??): 298–323, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951500614X.

Yuan:2014:MIF

[Yua14] Xiying Yuan. Maxima of the Q-index: forbidden odd cycles. Linear Algebra and its Applications, 458(??): 207–216, October 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514003796.

See corrigendum [Yua15].

Yuan:2015:CSQ

[Yua15]  Xiying Yuan.  Corrigen-

dum to “Maxima of the Qindex: forbidden odd cycles” [Linear Algebra Appl. 458 (2014) 207–216]. Linear Algebra and its Applications, 465(??):426–429, January 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006119.

See [Yua14].

Yuan:2016:PPL

[Yua16]  Xiying Yuan.  Proof of

a problem on Laplacian eigenvalues of trees. Linear Algebra and its Applications, 503(??):180–189, August 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516300519.

Yura:2015:HDS

[Yur15]  Fumitaka Yura.  Hankel

determinant solution for elliptic sequence. Linear Algebra and its Applications, 484(??):27–45, November 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003699.

Yuster:2017:EEW

[Yus17] Raphael Yuster. The effect of edge weights on clique weights. Linear Algebra and its Applications, 515(??): 298–312, February 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305626.

Yu:2016:DIA

[YWD16] Anqi Yu, Xiunan Wang, and Chunyuan Deng. On the Drazin inverse of an anti-triangular block matrix. Linear Algebra and its Applications, 489(??): 274–287, January 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006096.

Ye:2017:GIM

[YYMK17] Dong Ye, Yujun Yang, Bholanath Mandal, and Douglas J. Klein. Graph invertibility and median eigenvalues. Linear Algebra and its Applications, 513(??): 304–323, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304931.

Yu:2019:SUH

[YYQ19] Guihai Yu, Xiying Yuan, and Hui Qu. Signed kuniform hypergraphs and tensors. Linear Algebra and its Applications, 580(??): 1–13, November 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302630.

You:2019:SEQ

[YYSX19] Lihua You, Man Yang, Wasin So, and Weige Xi. On the spectrum of an equitable quotient matrix and its application. Linear Algebra and its Applications, 577(??):21–40, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930165X.

Yuan:2015:SUB

[YZL15] Xiying Yuan, Man Zhang, and Mei Lu. Some upper bounds on the eigenvalues of uniform hypergraphs. Linear Algebra and its Applications, 484(??): 540–549, November 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515003766.

Yao:2019:SVI

[YZL^{+}19] Hongmei Yao, Can Zhang, Lei Liu, Jiang Zhou, and Changjiang Bu. Singular value inclusion sets of rectangular tensors. Linear Algebra and its Applications, 576(??):181–199, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Yuan:2018:GPS

[YZLC18] Xiying Yuan, Qingqing Zhao, Lele Liu, and Hongyan Chen. On graphs with prescribed star complements. Linear Algebra and its Applications, 559(??):80–94, December 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304282.

Yu:2018:CSN

[YZZ18] Baomin Yu, Xianzhong Zhao, and Lingli Zeng. A congruence on the semiring of normal tropical matrices. Linear Algebra and its Applications, 555(??): 321–335, October 15, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303112.

Zajac:2019:SLF

[Zaj19]  Katarzyna Zajac.  On

the structure of loop-free non-negative edge-bipartite graphs. Linear Algebra and its Applications, 579(??): 262–283, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930254X.

Zamani:2016:OVP

[Zam16] Ali Zamani. The operatorvalued parallelism. Linear Algebra and its Applications, 505(??):282–295, September 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516301537.

Zamani:2019:NRI

[Zam19] Ali Zamani. A -numerical radius inequalities for semi-Hilbertian space operators. Linear Algebra and its Applications, 578(??):159–183, October 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519302216.

Zaspel:2016:SCM

[Zas16]  Peter Zaspel.  Subspace

correction methods in algebraic multi-level frames. Linear Algebra and its Applications, 488(??):505–521, January 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515005418.

Zhang:2015:UGS

[ZC15] Hui Zhang and Lizhi Cheng. On uniqueness guarantees of solution in convex regularized linear inverse problems. Linear Algebra and its Applications, 486(??): 475–483, December 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515005169.

Zhang:2015:MPB

[ZD15] Huan Zhang and Peter M. Dower. A max-plus based fundamental solution for a class of discrete time linear regulator problems. Linear Algebra and its Applications, 471(??):693–729, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000373.

Zhou:2016:DMT

[ZD16a] Hui Zhou and Qi Ding. The distance matrix of a tree with weights on its arcs. Linear Algebra and its Applications, 511(??):365– 377, December 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951630427X.

Zhou:2016:PDM

[ZD16b] Hui Zhou and Qi Ding. The product distance matrix of a tree with matrix weights on its arcs. Linear Algebra and its Applications, 499 (??):90–98, June 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516001725.

Zemel:2016:GRG

[Zem16] Shaul Zemel. Generalized Riordan groups and operators on polynomials. Linear Algebra and its Applications, 494(??):286–308, April 1, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516000240.

Zhou:2015:MPM

[ZF15] Anwa Zhou and Jinyan Fan. Monotonically positive matrices. Linear Algebra and its Applications, 485(??):467– 479, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004553.

Zhang:2018:LQE

[ZG18] Rong Zhang and Shu-Guang Guo. On the least Qeigenvalue of a non-bipartite Hamiltonian graph. Linear Algebra and its Applications, 538(??):89–102, February 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517305918.

Zhou:2016:SOG

[ZGW16] Kai Zhou, Zhenhua Gu, and Zhe-Xian Wan. Subconstituents of orthogonal graphs of characteristic 2. Linear Algebra and its Applications, 507(??):251–266, October 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302117.

Zhang:2015:MPP

[ZH15] Wen Zhang and Jinchuan Hou. Maps preserving peripheral spectrum of generalized products of operators. Linear Algebra and its Applications, 468(??):87– 106, March 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379514000706.

Zhang:2017:EDS

[ZH17] Xuefeng Zhang and Wenchao Huang. Extension of diagonal stability and stabilization for continuous-time fractional positive linear systems. Linear Algebra and its Applications, 530(??): 500–517, October 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517303816.

Zhang:2015:MTV

[Zha15a] Fuzhen Zhang. Book review: Matrix Theory, vol. 147, Xingzhi Zhan, in: Grad. Stud. Math.. American Mathematical Society (2013), ISBN: 978-0-8218-9491-0. Linear Algebra and its Applications, 471(??):1– 9, April 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007964.

Zhang:2015:EMR

[Zha15b] Pingping Zhang. Extension of Matic’s results. Linear Algebra and its Applications, 486(??):328–334, December 1, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379515004917.

Zhao:2015:PIM

[Zha15c] Chenggui Zhao. Parameterinvariant models for load balancing on heterogeneous networks. Linear Algebra and its Applications, 471 (??):369–382, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000312.

Zhao:2017:LLC

[Zha17] Wenhua Zhao. The LNED and LFED conjectures for algebraic algebras. Linear Algebra and its Applications, 534(??):181–194, December 1, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517304925.

Zhang:2018:CPC

[Zha18a] Qinghong Zhang. Completely positive cones: Are they facially exposed? Linear Algebra and its Applications, 558(??):195–204, December 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304117.

Zhang:2018:IAO

[Zha18b] Ye Zhang. An identity for (δ, ϵ)-approximately orthogonality preserving mappings. Linear Algebra and its Applications, 554(??): 358–370, October 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379518302854.

Zhang:2018:SSN

[Zha18c] Yun Zhang. Some spectral norm inequalities on Hadamard products of nonnegative matrices. Linear Algebra and its Applications, 556(??):162–170, November 1, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830329X.

Zhang:2019:SIR

[Zha19a] Pingping Zhang. On some inequalities related to positive block matrices. Linear Algebra and its Applications, 576(??):258–267, September 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951830497X.

Zhang:2019:IRI

[Zha19b]  Yun Zhang.  Interpolat-

ing the Rotfel’d inequality for unitarily invariant norms. Linear Algebra and its Applications, 574(??):60– 66, August 1, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301302.

Zhan:2016:MWS

[ZHC16] Xu-Zhou Zhan, Yong-Jian Hu, and Gong-Ning Chen. On maximal weight solutions of two matricial moment problems in the nondegenerate case. Linear Algebra and its Applications, 506(??):413–444, October 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516302002.

Zheng:2019:MMD

[Zhe19]  De-Yin Zheng.  Matrix

methods for determinants of Pascal-like matrices. Linear Algebra and its Applications, 577(??):94–113, September 15, 2019. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379519301818.

Zhou:2017:IDM

[Zho17] Hui Zhou. The inverse of the distance matrix of a distance well-defined graph. Linear Algebra and its Applications, 517(??):11–29, March 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516305870.

Zhu:2019:NEC

[Zhu19] Hao Zhu. A note on eigenvalues of a class of singular continuous and discrete linear Hamiltonian systems. Linear Algebra and its Applications, 570(??):46–60, June 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S002437951930059X.

Zimmermann:2015:CNA

[Zim15] R. Zimmermann. On the condition number anomaly of Gaussian correlation matrices. Linear Algebra and its Applications, 466(??): 512–526, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007174.

Zizler:2016:NSC

[Ziz16]  Peter Zizler.  On non-

stationary cyclic convolution and combination. Linear Algebra and its Applications, 499(??):31–42, June 15, 2016. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516001609.

Zhang:2017:LDI

[ZK17] Xian Zhang and Mykola Khrypchenko. Lie derivations of incidence algebras.

Linear Algebra and its Applications, 513(??):69–83, January 15, 2017. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304840.

Zhang:2017:SUH

[ZKSB17] Wei Zhang, Liying Kang, Erfang Shan, and Yanqin Bai. The spectra of uniform hypertrees. Linear Algebra and its Applications, 533(??): 84–94, November 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S002437951730438X.

Zhang:2015:CSL

[ZL15a] Minjie Zhang and Shuchao Li. Corrigendum to “On the signless Laplacian spectra of k-trees” [Linear Algebra Appl. 467 (2015) 136– 148]. Linear Algebra and its Applications, 485(??):527– 530, November 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515004437.

See [ZL15b].

Zhang:2015:SLS

[ZL15b] Minjie Zhang and Shuchao Li. On the signless Laplacian spectra of k-trees. Linear Algebra and its Applications, 467(??):136–148, February 15, 2015. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514007344.

See corrigendum [ZL15a].

Zhai:2015:SCE

[ZLG15] Mingqing Zhai, Huiqiu Lin, and Shicai Gong. Spectral conditions for the existence of specified paths and cycles in graphs. Linear Algebra and its Applications, 471 (??):21–27, April 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514008027.

Zhang:2017:CDI

[ZLG17] Minjie Zhang, Shuchao Li, and Ivan Gutman. Connectivity, diameter, independence number and the distance spectral radius of graphs. Linear Algebra and its Applications, 529(??):30– 50, September 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379517302392.

Zhang:2017:SPS

[ZLKB17] Wei Zhang, Lele Liu, Liying Kang, and Yanqin Bai. Some properties of the spectral radius for general hypergraphs. Linear Algebra and its Applications, 513(??): 103–119, January 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516304657.

Zhao:2018:DZT

[ZLLL18] Jianxing Zhao, Qilong Liu, Chaoqian Li, and Yaotang Li. Dashnic–Zusmanovich type matrices: a new subclass of nonsingular Hmatrices. Linear Algebra and its Applications, 552(??): 277–287, September 1, 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518302222.

Zheng:2016:NML

[ZLQ16] Hua Zheng, Wen Li, and Wei Qu. A non-modulus linear method for solving the linear complementarity problem. Linear Algebra and its Applications, 495 (??):38–50, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516000355.

Zhang:2017:WCG

[ZLS17] Guangliang Zhang, Muhuo Liu, and Haiying Shan. Which Q-cospectral graphs have same degree sequences.

Linear Algebra and its Applications, 520(??):274–285, May 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S0024379516306073.

Zhong:2016:FRM

[ZM16] Jin Zhong and Chao Ma. Further results on the minimum rank of regular classes of (0, 1)-matrices. Linear Algebra and its Applications, 511(??):447–459, December 15, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516304232.

Zhang:2019:EGI

[ZMT19] Daochang Zhang, Dijana Mosić, and Tin-Yau Tam. On the existence of group inverses of Peirce corner matrices. Linear Algebra and its Applications, 582(??): 482–498, December 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519303271.

Zou:2017:AGM

[Zou17] Limin Zou. An arithmeticgeometric mean inequality for singular values and its applications. Linear Algebra and its Applications, 528(??):25–32, September 1, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379516000185.

Zou:2019:UAG

[Zou19]  Limin Zou.  Unification

of the arithmetic-geometric mean and Hölder inequalities for unitarily invariant norms. Linear Algebra and its Applications, 562(??): 154–162, February 1, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518304713.

Zhang:2017:DOM

[ZS17] Teng Zhang and Amit Singer. Disentangling orthogonal matrices. Linear Algebra and its Applications, 524(??):159–181, July 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517301520.

Song:2015:CGN

[zSqST15] Ya zhi Song, Xiao qiu Song, and Bit-Shun Tam. A characterization of graphs G with nullity |V (G)| − 2m + 2c(G). Linear Algebra and its Applications, 465(??): 363–375, January 15, 2015.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379514006314.

Zhang:2019:MFS

[ZSQZ19] Liming Zhang, Zhihua Sun, Xiaoying Qu, and Xiqiang Zhao. Matrix factorization of the n × n shift Bell matrix. Linear Algebra and its Applications, 577(??):214– 239, September 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379519301958.

Zhou:2016:SCM

[ZSWB16] Jiang Zhou, Lizhu Sun, Yuanpeng Wei, and Changjiang Bu. Some characterizations of M-tensors via digraphs. Linear Algebra and its Applications, 495(??): 190–198, April 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0024379516000562.

Zuevsky:2018:ATE

[Zue18] Alexander Zuevsky. Affine Toda equations and solutions in the homogeneous grading. Linear Algebra and its Applications, 542 (??):149–161, ???? 2018. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856

Zuiddam:2017:NGB

[Zui17] Jeroen Zuiddam. A note on the gap between rank and border rank. Linear Algebra and its Applications, 525 (??):33–44, July 15, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379517301775.

Zusmanovich:2017:NSG

[Zus17a] Pasha Zusmanovich. Nonsemigroup gradings of associative algebras. Linear Algebra and its Applications, 523(??):52–58, June 15, 2017. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379517300940.

Zusmanovich:2017:SEM

[Zus17b] Pasha Zusmanovich. Special and exceptional mock-Lie algebras. Linear Algebra and its Applications, 518 (??):79–96, April 1, 2017. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379516306280.

Zhang:2016:MLE

[ZW16] Haixia Zhang and Yi Wang. On the maximum Laplacian Estrada index of graphs. Linear Algebra and its Applications, 492(??):68–77, March 1, 2016. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515006680.

Zhou:2018:UBN

[ZWS18] Qi Zhou, Dein Wong, and Dongqin Sun. An upper bound of the nullity of a graph in terms of order and maximum degree. Linear Algebra and its Applications, 555(??):314–320, October 15, 2018. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379518303094.

Zhang:2019:COG

[ZXW19] Yuanshuai Zhang, Feng Xu, and Dein Wong. Characterization of oriented graphs of rank 2. Linear Algebra and its Applications, 579(??): 136–147, October 15, 2019. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0024379519302307.

Zhu:2015:GIF

[ZZC15] Huihui Zhu, Xiaoxiang Zhang, and Jianlong Chen. Generalized inverses of a factorization in a ring with involution. Linear Algebra and its Applications, 472 (??):142–150, May 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0024379515000543.

Zhang:2015:NCP

[ZZR15] Li-Li Zhang, Yun-Peng Zhang, and Zhi-Ru Ren. New convergence proofs of modulus-based synchronous multisplitting iteration methods for linear complementarity problems. Linear Algebra and its Applications, 481(??):83–93, September 15, 2015. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0024379515002803.